### v1

# spacing=nonfrench

Abstract
Let $\Gamma$ be a bounded Jordan curve and $\Omega_{i},\Omega_{e}$ its two complementary components. For $1<p<\infty,\,s\in(0,1)$ we define $\mathcal{B}_{p,p}^{s}(\Omega_{i,e})$ as the set of functions $f:\Gamma\to\mathbb{C}$ having harmonic extension $u$ respectively in $\Omega_{i}$ and $\Omega_{e}$ such that
$\iint_{\Omega_{i,e}}|\nabla u(z)|^{p}d(z,\Gamma)^{(1-s)p-1}dxdy<+\infty.$
If $\Gamma$ is further assumed to be rectifiable we define $B_{p,p}^{s}(\Gamma)$ as the space of functions $f\in L^{p}(\Gamma)$ such that
$\iint_{\Gamma\times\Gamma}\frac{|f(z)-f(\zeta)|^{p}}{|z-\zeta|^{1+ps}}|dz||d\zeta|<+\infty.$
When $\Gamma$ is the unit circle these three spaces coincide with the homogeneous fractional Besov-Sobolev space. For a general rectifiable curve these spaces need not coincide and our first goal is to investigate the cases of equality: while the chord-arc property is the necessary and sufficient condition for equality in the classical case of $s=1/p,\,p\geq 2$, this is no longer the case for general $s\in(0,1)$. We show however that equality holds for radial-Lipschitz curves. In the general (possibly non-rectifiable) case we study boundary values of functions in $\mathcal{B}_{p,p}^{s}(\Omega_{i,e})$ and give conditions for equality of these trace-spaces that we then call $\mathcal{B}^{s}_{p,p}(\Gamma)$. Using Plemelj-Calderón property we further identify $\mathcal{B}^{s}_{p,p}(\Gamma)$ with the space of restrictions of a weighted Sobolev space of the plane. Finally we re-interpretate some of our results as the "almost"-Dirichlet principle in the spirit of Maz’ya.
Keywords: Cauchy integral, fractional Sobolev spaces, chord-arc curves, Lipschitz curves, Dirichlet spaces
2020 MSC: Primary 42B20, 46E35; Secondary 31A05, 30H35
1 Introduction
1.1 The Dirichlet Principle
In all this work we will be considering a bounded Jordan curve $\Gamma$ and call $\Omega_{i},\Omega_{e}$ respectively the interior and exterior connected components of its complement in the sphere. If $f$ is a continuous function on $\Gamma$ and $F$ a continuous extension of $f$ to $\overline{\Omega}_{i,e}$ which is furthermore assumed to be $C^{1}$ in $\Omega_{i,e}$ we call (2)(2-)energy the quantity
$\mathcal{E}_{i,e}(F)=\iint_{\Omega_{i,e}}|\nabla F(z)|^{2}dxdy\leq+\infty.$
The Dirichlet principle asserts that among all the extensions of $f$ the one with the lowest energy is the harmonic one. Notice that the (2)(2-)energy of the harmonic extension of a continuous function may be infinite, in which case every extension has infinite (2)(2-)energy. When $\Gamma=\mathbb{T}$, the unit circle, we define $\mathcal{D}(\mathbb{D}_{i})$, the Dirichlet space over the unit disk $\Omega_{i}=\mathbb{D}_{i}$ as the space of harmonic functions in $\mathbb{D}_{i}$ with finite (2)(2-)energy. This space coincides with the space of Poisson integrals of $L^{2}(\mathbb{T})$-functions such that
$\sum\limits_{n\in\mathbb{Z}}|n||\hat{f}(n)|^{2}<\infty.$
As Douglas observed in [14] it also coincides with the space of Poisson integrals of $L^{2}(\mathbb{T})$-functions such that
$\iint_{\mathbb{T}\times\mathbb{T}}\frac{|f(z)-f(\zeta)|^{2}}{|z-\zeta|^{2}}|dz||d\zeta|<\infty.$
The map $z\mapsto 1/\bar{z}$ is a reflection about the unit circle which is bi-Lipschitz from $\{1/2<|z|<2\}$ onto itself. It follows that if $u_{i}\in\mathcal{D}(\mathbb{D}_{i})$ is the Poisson integral of $f\in L^{2}(\mathbb{T})$ then the Poisson extension $u_{e}$ of $f$ in $\mathbb{D}_{e}=\overline{\mathbb{C}}\setminus\overline{\mathbb{D}}_{i}$ also has a finite energy. We may thus view the Dirichlet space as the subset of $L^{2}(\mathbb{T})$ of functions whose harmonic extension to $\overline{\mathbb{C}}\setminus\mathbb{T}$ has finite energy.
The Dirichlet space is a special case of homogeneous fractional Sobolev spaces. More precisely, if $0<s<1$, we define $H^{s}(\mathbb{T})$ as the subspace of $L^{2}(\mathbb{T})$ of functions $f$ such that
$\|f\|^{2}_{H^{s}}=\sum\limits_{n\in\mathbb{Z}}|n|^{2s}|\hat{f}(n)|^{2}<\infty.$
It is not difficult to see that the above quantity, the square of the $H^{s}$-semi-norm, is equivalent (in particular, in the case of $s=1/2$ it is equal, up to a constant multiplicative factor,) to two other:
  1. 1.
    The (square of the) Littlewood-Paley semi-norm
    $\iint_{\mathbb{D}_{i}}|\nabla u_{i}(z)|^{2}(1-|z|)^{1-2s}dxdy,$
  2. 2.
    The (square of the) Douglas norm
    $\iint_{\mathbb{T}\times\mathbb{T}}\frac{|f(z)-f(\zeta)|^{2}}{|z-\zeta|^{1+2s}}|dz||d\zeta|.$
The mention of Littlewood-Paley comes from the fact that for $s=0$ the Littlewood-Paley semi-norm is exactly the $L^{2}$-norm of $\mathsf{g}(f)$, the Littlewood-Paley function associated to $f$ [37].
In terms of Besov spaces,
$H^{s}(\mathbb{T})=B_{2,2}^{s}(\mathbb{T}),$
and that this space also coincides with $(I-\Delta)^{-s/2}(L^{2}(\mathbb{T}))$, namely the space of Bessel potentials of order $s$.
We can further generalize these spaces by replacing $2$ with $p>1$ and define $B_{p,p}^{s}(\mathbb{T})$ as the set of $L^{p}(\mathbb{T})$-functions such that either
  1. 1.
    (Littlewood-Paley)
    $\iint_{\mathbb{D}_{i}}|\nabla u_{i}(z)|^{p}(1-|z|)^{(1-s)p-1}dxdy<\infty,$
  2. 2.
    $\iint_{\mathbb{T}\times\mathbb{T}}\frac{|f(z)-f(\zeta)|^{p}}{|z-\zeta|^{1+ps}}|dz||d\zeta|<\infty.$
It has to be noticed that if $p\neq 2$,
$B_{p,p}^{s}(\mathbb{T})\neq(I-\Delta)^{-s/2}(L^{p}(\mathbb{T})),$
so that, to avoid confusions, we will always use the Besov terminology at least for $p\neq 2$.
The main goal of this paper is to extend these notions to more general Jordan curves. The first restriction we want to impose on the Jordan curve $\Gamma$ is that we want it to share with $\mathbb{T}$ the reflection property, that is we want the existence of a reflection across $\Gamma$ which is bi-Lipschitz in a neighborhood of $\Gamma$. The curves satisfying this property are well-known: they are the quasicircles, i.e., the quasiconformal images of circles [1]. Geometrically, the Jordan curve $\Gamma$ is a quasicircle if and only if there exists $C>1$ such that
$\forall z_{1},z_{2}\in\Gamma,\quad\min(\mathrm{diam}(\gamma_{1}),\mathrm{diam}(\gamma_{2}))\leq C|z_{1}-z_{2}|$
where $\gamma_{j},j=1,2$, are the two subarcs of $\Gamma$ with endpoints $z_{1},z_{2}$.
A quasicircle need not be smooth, and the Hausdorff dimension of a quasicircle can take any value in $[1,2)$. As an example, the Von Koch snowflake curve is a quasicircle without any tangent and whose Hausdorff dimension, which coincides with its Minkowski dimension, is $\log 4/\log 3$.
From now on, all the Jordan curves that we consider will be assumed to be quasicircles.
1.2 Statement of the Main Theorems
Let $\Gamma$ be a quasicircle and $\Omega_{i,e}$ the inner and outer connected components of $\overline{\mathbb{C}}\setminus\Gamma$ as above. Suppose $\Omega$ is one of these two components and correspondingly $\mathbb{D}$ is one of $\mathbb{D}_{i}$ and $\mathbb{D}_{e}$. We define
$\mathcal{B}_{p,p}^{s}(\Omega)=\{u\;\mathrm{harmonic\;in}\;\Omega:\;\|u\|_{\mathcal{B}_{p,p}^{s}(\Omega)}^{p}:=\iint_{\Omega}|\nabla u(z)|^{p}d(z,\Gamma)^{(1-s)p-1}dxdy<\infty\}.$
We define for functions in these spaces and for some values of p,sp,\,s their boundary values on $\Gamma$, which will allow us to identify $\mathcal{B}_{p,p}^{s}(\Omega)$ with a space of functions defined on $\Gamma$ that we denote by $\mathcal{B}_{p,p}^{s}(\Omega\!\to\!\Gamma)$.
In order to define this space we distinguish three cases, namely, s>1/p,s=1/p,s<1/ps>1/p,\,s=1/p,\,s<1/p.
  1. 1.
    Case $s>1/p$: we will see that in this case $\mathcal{B}_{p,p}^{s}(\Omega)\subset\Lambda^{\alpha}(\overline{\Omega})$, the Hölder space with exponent $\alpha=s-1/p$. The boundary value of a function in $\mathcal{B}_{p,p}^{s}(\Omega)$ is then just its restriction to $\Gamma$.
  2. 2.
    Case $s=1/p$: we use here the remarkable fact that these spaces are conformally invariant. In particular, if $f\in\mathcal{B}_{p,p}^{s}(\Omega)$, then $f\circ\varphi\in\mathcal{B}_{p,p}^{s}(\mathbb{D})$ where $\varphi$ is a Riemann mapping from $\mathbb{D}$ onto $\Omega$, and this isomorphism $T$ is almost isometric in the sense that, due to the Koebe distortion theorem,
    $\forall f\in\mathcal{B}_{p,p}^{1/p}(\Omega),\quad\frac{1}{4}\|f\|_{\mathcal{B}_{p,p}^{1/p}(\Omega)}\leq\|Tf\|_{\mathcal{B}_{p,p}^{1/p}(\mathbb{D})}\leq 4\|f\|_{\mathcal{B}_{p,p}^{1/p}(\Omega)}.$
    Since this latter space is included in the harmonic Hardy space $h^{p}(\mathbb{D})$ we conclude that $f$ has a well-defined boundary value in $L^{p}(\omega,\Gamma)$, where $\omega$ is the harmonic measure for $\Omega$.
  3. 3.
    Case $s<1/p$: we proceed here differently. Let $C_{c}^{\infty}(\mathbb{C})|_{\Gamma}=\{f|_{\Gamma}:\;f\in C_{c}^{\infty}(\mathbb{C})\}$, $C_{c}^{\infty}(\mathbb{C})$-functions being $C^{\infty}$-functions with compact support in $\mathbb{C}$. Let $\mathcal{D}_{p}^{s}(\Gamma)$ be the set of functions in $C_{c}^{\infty}(\mathbb{C})|_{\Gamma}$ such that their harmonic extensions to $\Omega_{i,e}$ belongs to $\mathcal{B}_{p,p}^{s}(\Omega_{i,e})$, and we denote the set of such extensions by $\mathcal{D}_{p}^{s}(\Omega_{i,e})$. We define the space of boundary values $\mathcal{B}_{p,p}^{s}(\Omega_{i,e}\!\to\!\Gamma)$ as the abstract closure of this space, i.e., equivalence classes of Cauchy sequences in $\mathcal{D}_{p}^{s}(\Gamma)$, equipped with the $\mathcal{B}_{p,p}^{s}(\Omega_{i,e})$-norm. Notice that this procedure only defines boundary values for the closures of $\mathcal{D}_{p}^{s}(\Gamma)$ in $\mathcal{B}_{p,p}^{s}(\Omega_{i,e})$, which may not be the full space. However, the space $\mathcal{B}_{p,p}^{s}(\Omega_{i,e})$ in this case is quite problematic, and the closure of $\mathcal{D}_{p}^{s}(\Omega_{i,e})$ in $\mathcal{B}_{p,p}^{s}(\Omega_{i,e})$ is actually a more robust variant of this space. .
For all these cases we define
$\mathcal{B}_{p,p}^{s}(\Gamma)=\mathcal{B}_{p,p}^{s}(\Omega_{i}\to\Gamma)\cap\mathcal{B}_{p,p}^{s}(\Omega_{e}\to\Gamma),$
equipped with the natural norm. One of our main tasks will be to examine when these three spaces coincide. The case 2 above is particularly interesting especially since the two harmonic measures may be mutually singular. This drawback has been ruled by [36] and [5] by exhibiting a change of variable realizing a "transmission" operator between the spaces on both sides which happens to be an isomorphism, thus showing the equality of the three spaces for all quasicircles.
When $\Gamma$ is furthermore assumed to be rectifiable we define, following Douglas [14],
$B_{p,p}^{s}(\Gamma)=\{f\in L^{p}(\Gamma):\;\|f\|_{B_{p,p}^{s}(\Gamma)}^{p}:=\iint_{\Gamma\times\Gamma}\frac{|f(z)-f(\zeta)|^{p}}{|z-\zeta|^{1+ps}}|dz||d\zeta|<\infty\}.$
Here and in what follows, $|dz|$ denotes the arc-length measure. If $\Gamma$ is the unit circle the three spaces $\mathcal{B}_{p,p}^{s}(\Omega_{i}\to\Gamma),\mathcal{B}_{p,p}^{s}(\Omega_{e}\to\Gamma),B_{p,p}^{s}(\Gamma)$ coincide (see p.151 in [38]), but they have no reason to coincide in general. Finding cases of equality is one of our goals. In order to state the first result in this direction, we need some definitions:
A curve $\Gamma$ is said to be chord-arc (or $K$-chord-arc) if it is the bi-Lipschitz image of a circle. These curves are geometrically characterized by the fact that they are rectifiable and that
$\exists K>1:\quad\forall z_{1},z_{2}\in\Gamma,\,\min(\mathrm{length}(\gamma_{1}),\mathrm{length}(\gamma_{2}))\leq K|z_{1}-z_{2}|,$
where $\gamma_{j},\,j=1,2$, are the two subarcs of $\Gamma$ with endpoints $z_{1},z_{2}$. Notice that chord-arc curves are in particular rectifiable quasicircles (but the converse is not true).
We will also need the notion of radial-Lipschitz curves. These are the curves given in polar coordinates by the equation $z(\theta)=r(\theta)e^{i\theta}$ where $r:\mathbb{R}\to(0,+\infty)$ is $2\pi$-periodic and Lipschitz continuous. The radial-Lipschitz curves are in particular chord-arc.
In [44] and [45] the present authors have proven
Theorem 1.1.
Let $p>1$. If $\Gamma$ is chord-arc then
$\mathcal{B}_{p,p}^{1/p}(\Omega_{i}\to\Gamma)=\mathcal{B}_{p,p}^{1/p}(\Omega_{e}\to\Gamma)=B_{p,p}^{1/p}(\Gamma).$ (1.1)
Conversely, if $\Gamma$ is a rectifiable quasicircle such that (1.1) holds and $p\geq 2$ then $\Gamma$ is chord-arc.
The methods we used for the proof of this theorem depend heavily on the special features of the case $s=1/p$, namely, conformal invariance and the existence of a transmission operator.
In this paper our main results in this direction are Theorems 5.5, 5.7, 5.8 combined with the claims in appendix that loosely speaking say that
Theorem 1.2.
If $\Gamma$ is a radial-Lipschitz curve and $p=2$, $s\in(0,1)$ then
$\mathcal{B}_{p,p}^{s}(\Omega_{i}\!\to\!\Gamma)=\mathcal{B}_{p,p}^{s}(\Omega_{e}\!\to\!\Gamma)=B_{p,p}^{s}(\Gamma).$ (1.2)
This can be generalized to any $p\in(1,\infty)$ under a slightly different assumption related to $p$.
We will also prove a result in this regard for general quasicircles. Before stating it we need to introduce some notions:
Astala [3] has introduced the notion of "Minkowski content" $h(\Gamma)$: the proper definition will be recalled below but let us just mention here that $h(\Gamma)$ is the Minkowski dimension if $\Gamma$ is a self-similar fractal set (as is the Koch snowflake curve for example).
The Muckenhoupt weights $A_{p}(\mathbb{C}),\,p>1$, are the weights $\omega$ such that the Hardy-Littlewood maximal function is bounded on $L^{p}(\omega,\mathbb{C})$. Let $\omega$ be such a weight. We define the weighted Sobolev space $W^{1,p}(\omega,\mathbb{C})$ as being the space of tempered distributions $f$ such that $f$ is locally integrable and, in the sense of distributions, $\nabla f\in L^{p}(\omega,\mathbb{C})$.
Theorem 1.3 (see Theorem 3.16).
If $\Gamma$ is a quasicircle and
$\frac{1}{p}<1-\frac{1}{p}\leq s<\frac{p+K_{*}-h(\Gamma)}{K_{*}p}$
where $K_{*}\leq 2$ is a geometric quantity depending only on $\Gamma$ to be defined in section $3$ then
$\mathcal{B}_{p,p}^{s}(\Omega_{i}\to\Gamma)=\mathcal{B}_{p,p}^{s}(\Omega_{e}\to\Gamma)=W^{1,p}(\omega,\mathbb{C})|_{\Gamma}$ (1.3)
where $\omega(z)=d(z,\Gamma)^{(1-s)p-1}$.
Theorem 1.3 allows us to improve theorem 1.2 for certain values of $p,s$:
Corollary 1.4.
If $\Gamma$ is chord-arc and $p>2,s\in[1-\frac{1}{p},\frac{p+K_{*}-1}{K_{*}p})$ then (1.3) holds.
Another corollary is an "almost-Dirichlet principle" generalizing, in dimension $2$ a result of Maz’ya simplified by Mironescu-Russ ([29], [30]). We say that the quasidisk $\Omega$ satisfies the almost-Dirichlet principle for $p,s$ if any continuous function $f:\Gamma\to\mathbb{C}$ having a continuous extension $F:\overline{\Omega}\to\mathbb{C}$ which is in $C^{1}(\Omega)$ and satisfies
$\iint_{\Omega}|\nabla F(z)|^{p}d(z,\Gamma)^{(1-s)p-1}dxdy<\infty$
is in $\mathcal{B}_{p,p}^{s}(\Omega\to\Gamma)$, or, in other words, is such that
$\iint_{\Omega}|\nabla u(z)|^{p}d(z,\Gamma)^{(1-s)p-1}dxdy<\infty,$
where $u$ is the harmonic extension of $f$ in $\Omega$.
Corollary 1.5.
Under the assumption of Theorem 1.3 the almost-Dirichlet principle holds for $\mathcal{B}_{p,p}^{s}(\Omega)$.
One of the tools for proving Theorem 1.3 will be the use of Plemelj-Calderón property which is of independent interest.
1.3 Plemelj-Calderón Problem
It is an old problem, given a bounded Jordan curve $\Gamma$ in the plane and a (complex valued) function $f$ defined on $\Gamma$, to find two holomorphic functions $G_{i},G_{e}$ defined on the interior-connected and the exterior-connected components of $\Gamma$ such that
$f=G_{i}|_{\Gamma}+G_{e}|_{\Gamma},$
the boundary traces being defined in some sense.
This problem was actually raised and solved by Sokhotski in 1873 before being rediscovered by Plemelj as a main ingredient of his attempt to solve Hilbert’s 22th problem in 1908 ([34]). Assume first that $\Gamma$ is smooth and $f$ is a $C^{1}$-function on $\Gamma$. The idea is to use the Cauchy integral and define
$F(z)=\frac{1}{2\pi i}\int_{\Gamma}\frac{f(\zeta)}{\zeta-z}d\zeta,\qquad z\in\mathbb{C}\setminus\Gamma.$
This defines a function which is holomorphic outside the curve $\Gamma$. Moreover, this function has boundary values on $\Gamma$ from inside and outside, and one has Plemelj formula on $\Gamma$:
$F_{i}|_{\Gamma}=Tf+1/2f,$
$F_{e}|_{\Gamma}=Tf-1/2f$
where the Cauchy integral
$Tf(z)=\frac{1}{2\pi i}\text{p.v.}\int_{\Gamma}\frac{f(\zeta)}{\zeta-z}d\zeta=\frac{1}{2\pi i}\lim\limits_{\varepsilon\to 0}\int_{|\zeta-z|>\varepsilon}\frac{f(\zeta)}{\zeta-z}d\zeta.$
The problem is then solved with $G_{i}=F_{i}$ and $G_{e}=-F_{e}$.
Of course the result remains true for much more general curves and functions. Subsequent generalizations relax the smoothness requirements on the curve $\Gamma$ and the function $f$: In order for Plemelj formula to hold true one indeed only needs $\Gamma$ to be rectifiable and $f\in L^{1}(\Gamma)$. The modern approach of this problem is then to solve the problem in a given space: Calderón ([6]) asked for instance for which rectifiable curves it is true that for any $f\in L^{2}(\Gamma)$ Plemelj functions $F_{i}|_{\Gamma}$ and $F_{e}|_{\Gamma}$ are also in $L^{2}(\Gamma)$? Equivalently, when is the operator $T$ bounded on $f\in L^{2}(\Gamma)$? This problem has been solved by Coifman-McIntosh-Meyer ([10]) for Lipschitz curves (Calderón had previously solved it for curves with small Lipschitz bound ([6])) and then the last word was given by David ([11]) who proved that $T$ is bounded on $f\in L^{2}(\Gamma)$ (or $L^{p}(\Gamma)$, $1<p<\infty$) if and only if $\Gamma$ is Ahlfors-regular, meaning that the length of the part of $\Gamma$ lying in a disk of radius $r$ is no bigger than $Cr$ for some $C$ independent of the chosen disk.
Now Plemelj-Calderón problem may be addressed for curves that are not necessarily rectifiable. An example has been given in [48] where Plemelj-Calderón problem for Hölder classes has been raised and solved in some particular cases. The idea is to replace the Cauchy integral by the identity
$f(z)=-\frac{1}{\pi}\iint_{\mathbb{C}}\frac{\bar{\partial}f(\zeta)}{\zeta-z}d\xi d\eta$
which is true for any test function $f\in C_{c}^{\infty}(\mathbb{C})$, and from which it follows that, in the case of rectifiable boundary,
$F_{i}(z)$ $=-\frac{1}{\pi}\iint_{\Omega_{e}}\frac{\bar{\partial}f(\zeta)}{\zeta-z}d\xi d\eta,\qquad z\in\Omega_{i};$
$F_{e}(z)$ $=\frac{1}{\pi}\iint_{\Omega_{i}}\frac{\bar{\partial}f(\zeta)}{\zeta-z}d\xi d\eta,\qquad z\in\Omega_{e}$
where $\Omega_{i,e}$ stand for the interior and exterior components of $\Gamma$. Later Astala ([3]) obtained more precise results for the Hölder classes: these will be discussed in section 3 where we will draw the parallel with fractional Sobolev spaces.
In this paper we address the following Plemelj-Calderón problem: given $p>1,\,s\in(0,1)$, can one write a function $f\in\mathcal{B}_{p,p}^{s}(\Gamma)$ as a sum of two functions $F_{i,e}$ that are holomorphic in $\Omega_{i,e}$ and whose boundary values belong to $\mathcal{B}_{p,p}^{s}(\Gamma)$?
The main result of the present paper in this direction is Theorems 2.3, 2.4, 3.5 and Corollaries 3.8, 3.17. A priori version of these results is as follows.
Theorem 1.6.
Let $\Gamma$ be a quasicircle , $p>1,\,s\in(0,1)$. Then a priori Plemelj-Calderón property holds for $C_{c}^{\infty}(\mathbb{C})|_{\Gamma}$, if
$(p-1)(h(\Gamma)-1)<sp<p+1-h(\Gamma).$
2 Critical Besov Spaces on Quasicircles
Recall that a quasidisk is the image of the unit disk $\mathbb{D}_{i}$ by a quasiconformal homeomorphism of the plane, and its boundary curve is called a quasicircle. An homeomorphism $\Phi$ of the plane is called quasiconformal if its gradient in the sense of distribution is in $L^{2}_{loc}(\mathbb{C})$ and if in addition there exists a constant $k<1$ such that
$\frac{\partial\Phi}{\partial\bar{z}}=\mu(z)\frac{\partial\Phi}{\partial z}$
for a function $\mu\in L^{\infty}(\mathbb{C})$ with $\|\mu\|_{\infty}\leq k$. If we put $K=\frac{1+k}{1-k}$ then $K$ represents the maximal eccentricity of infinitesimal images of circles by the quasiconformal mapping.
Any quasicircle $\Gamma$ admits a quasiconformal reflection across $\Gamma$ which is bi-Lipschitz around $\Gamma$. Here, by a quasiconformal reflection we mean an anti-quasiconformal self-homeomorphism $R$ of $\overline{\mathbb{C}}$ such that $R\circ R=\text{id}$ on $\overline{\mathbb{C}}$ and $R=\text{id}$ when restricted to $\Gamma$. Precisely, let $\Phi$ be a quasiconformal homeomorphism of $\overline{\mathbb{C}}$ we then have $R(z)=\Phi\circ(1/\overline{\Phi^{-1}(z)})$ is a quasiconformal reflection with respect to the quasicircle $\Gamma:=\partial\Phi(\mathbb{D}_{i})$.
Let $\Omega_{i}$ and $\Omega_{e}$ be domains bounded by a quasicircle $\Gamma$. If $\varphi_{i}$ and $\varphi_{e}$ are conformal isomorphisms from $\mathbb{D}_{i}$ onto $\Omega_{i}$ and from $\mathbb{D}_{e}$ onto $\Omega_{e}$, respectively, then they can be extended to a homeomorphism between the closures such that $h:=\varphi_{e}^{-1}\circ\varphi_{i}$ is a homeomorphism of $\mathbb{T}$ which is quasisymmetric in the sense that there exists a constant $C>0$ such that
$C^{-1}\leq\frac{|h(e^{i(t+\alpha)})-h(e^{it})|}{|h(e^{it})-h(e^{i(t-\alpha)})|}\leq C$
for every $t\in\mathbb{R}$ and $\pi/2<\alpha\leq\pi/2$. Moreover, every quasisymmetry arises in this way from some quasicircle. For these results on quasiconformal theory, we refer to [1] for details.
Let $\Gamma$ be a Jordan curve and $p>1$. We recall that the spaces $\mathcal{B}_{p,p}^{1/p}(\Omega_{i,e})$ , called the $p$-critical Besov spaces, have the remarkable property of being conformally invariant: if $\Omega,\Omega^{\prime}$ are two Jordan domains and $\varphi:\Omega\to\Omega^{\prime}$ is a biholomorphism then the map $f\mapsto f\circ\varphi$ is a quasi-isometry between $\mathcal{B}_{p,p}^{1/p}(\Omega^{\prime})$ and $\mathcal{B}_{p,p}^{1/p}(\Omega)$. In order to be as much as possible self-contained, here is a sketch of proof: by the change of variable $\zeta=\varphi(z)$ we have
$\iint_{\Omega^{\prime}}|\nabla u(\zeta)|^{p}d(\zeta,\Gamma^{\prime})^{p-2}dxdy=\iint_{\Omega}|\nabla u(\varphi(z))|^{p}d(\varphi(z),\Gamma^{\prime})^{p-2}|\varphi^{\prime}(z)|^{2}dxdy.$
Now the Koebe distortion theorem [1] implies that
$\forall z\in\Omega,\quad\frac{1}{4}\leq\frac{d(\varphi(z),\Gamma^{\prime})}{d(z,\Gamma)|\varphi^{\prime}(z)|}\leq 4,$ (2.1)
and the result follows.
Applying this result for $\Omega^{\prime}=\mathbb{D}$, $\partial\mathbb{D}$ being the unit circle $\mathbb{T}$, we see that all the $p$-critical Besov spaces over $\Omega$ are quasi-isometric to $\mathcal{B}_{p,p}^{1/p}(\mathbb{D})$. It is a classical result [37] that the space $\mathcal{C}_{p}(\mathbb{D})$ consisting of functions in $\mathcal{B}_{p,p}^{1/p}(\mathbb{D})$ that have a continuous extension to $\overline{\mathbb{D}}$ is dense in $\mathcal{B}_{p,p}^{1/p}(\mathbb{D})$. This can also be derived from that the set of polynomials is dense in the Bergman space $A^{p}_{p-2}$, the set of analytic functions $g$ in $\mathbb{D}$ such that
$\iint_{\mathbb{D}}|g(z)|^{p}|1-|z|^{2}|^{p-2}dxdy<\infty.$
Since a bi-holomorphic homeomorphism between two Jordan domains extends to a homeomorphism between the closures the same property will hold for the analogous spaces $\mathcal{C}_{p}(\Omega)$.
The next step is to prove that $\mathcal{C}_{p}(\Omega_{i})|_{\Gamma}=\mathcal{C}_{p}(\Omega_{e})|_{\Gamma}$ for any quasicircle $\Gamma$. We first prove the result for $p=2$.
Let $u_{i}\in\mathcal{C}_{2}(\Omega_{i})$ and $f=u_{i}|_{\Gamma}$. Recall that there exists a quasiconformal reflection $R$ across $\Gamma$. Put $v(z)=u_{i}(R(z)),z\in\Omega_{e}$. By the bi-Lipschitz property of $R$ around $\Gamma$ we have that
$\iint_{\Omega_{e}}|\nabla v(z)|^{2}dxdy<\infty$
(stricly speaking, there is a technical problem at $\infty$, which is settled in the appendix). We invoke then the Dirichlet principle which implies that
$\iint_{\Omega_{e}}|\nabla u_{e}(z)|^{2}dxdy\leq\iint_{\Omega_{e}}|\nabla v(z)|^{2}dxdy<+\infty,$
where $u_{e}:=P_{e}(f)$ stands for the harmonic extension of $f$ in $\Omega_{e}$. That is, we have that the transmission operator
$\mathcal{C}_{2}(\Omega_{i})\ni u_{i}\mapsto u_{e}\in\mathcal{C}_{2}(\Omega_{e})$
is bounded. Actually, it is a bounded isomorphism since the boundedness of the inverse operator can be seen by exchanging the roles of $\Omega_{i}$ and $\Omega_{e}$, which proves the result.
If $p>1$ and $0<s<1$ (except the case of $p=2$ and $s=1/2$) we loose the Dirichlet principle and use instead the almost-Dirichlet principle. Recall that the almost-Dirichlet principle holds for $\mathcal{B}_{p,p}^{s}(\Omega)$ if for any $f$ defined on $\Gamma$, $u$ denotes its harmonic extension in $\Omega$ and $v$ is any continuous extension of $f$ to $\Omega$ which is in $C^{1}(\Omega)$, we have
$\iint_{\Omega}|\nabla u(z)|^{p}d(z,\Gamma)^{(1-s)p-1}dxdy\leq C\iint_{\Omega}|\nabla v(z)|^{p}d(z,\Gamma)^{(1-s)p-1}dxdy$
for some constant $C$ depending only on $\Omega,\,p,\,s$. It was proved in [30] that the almost-Dirichlet principle holds for $\mathcal{B}_{p,p}^{s}(\mathbb{D}_{i,e})$ for all p>1, 0<s<1p>1,\,0<s<1.
Lemma 2.1.
Let $\Gamma$ be a quasicircle and $1<p<\infty$. The almost-Dirichlet principle holds for the space $\mathcal{B}_{p,p}^{1/p}(\Omega_{i,e})$.
Proof.
The almost-Dirichlet principle in $\mathbb{D}_{i}$ applied to $s=1/p$, reads
$\iint_{\mathbb{D}_{i}}|\nabla u(z)|^{p}(1-|z|)^{p-2}dxdy\leq C\iint_{\mathbb{D}_{i}}|\nabla v(z)|^{p}(1-|z|)^{p-2}dxdy,$ (2.2)
with the same notations as before: $u$ is the harmonic function with the boundary value $u|_{\Gamma}=f$ and $v$ is any $C^{1}$ extension of $f$ to $\mathbb{D}_{i}$. Let $\varphi$ be a conformal mapping from $\mathbb{D}_{i}$ onto $\Omega_{i}$. By the conformal invariance of the space $W^{1,p}(\omega,\mathbb{D}_{i})$ where $\omega(z)=(1-|z|)^{p-2}$ and using the change of variables, we have
$\iint_{\Omega_{i}}|\nabla(u\circ\varphi^{-1})|^{p}d(z,\,\Gamma)^{p-2}dxdy\leq C\iint_{\Omega_{i}}|\nabla(v\circ\varphi^{-1})|^{p}d(z,\,\Gamma)^{p-2}dxdy.$ (2.3)
Noting that $u\circ\varphi^{-1}$ is harmonic we have proved the almost-Dirichlet principle for the space $\mathcal{B}_{p,p}^{1/p}(\Omega_{i}).$ The assertion for the space $\mathcal{B}_{p,p}^{1/p}(\Omega_{e})$ can be treated similarly. ∎
As a consequence, if $f$ is a continuous function on $\Gamma$, its harmonic extension to $\Omega_{i}$ is in $\mathcal{C}_{p}(\Omega_{i})$ if and only if its harmonic extension to $\Omega_{e}$ is in $\mathcal{C}_{p}(\Omega_{e})$. Then, $C_{p}(\Omega_{i}\!\to\!\Gamma)=C_{p}(\Omega_{e}\!\to\!\Gamma)$ so that we may define $\mathcal{C}_{p}(\Gamma):=\mathcal{C}_{p}(\Omega_{i,e}\!\to\!\Gamma)$.
Let $\varphi_{i,e}$ denote the Riemann mappings from $\mathbb{D}_{i,e}$ onto $\Omega_{i,e}$ as before. By the conformal invariance, we have that $f\circ\varphi_{i,e}\in\mathcal{C}_{p}(\mathbb{T})$ and $f\circ\varphi_{e}=f\circ\varphi_{i}\circ h$ with $h:=\varphi_{i}^{-1}\circ\varphi_{e}$ being a circle homeomorphism called conformal welding. As we have recalled above this homeomorphism is a quasisymmetry: we deduce from this that $V_{h}:g\mapsto g\circ h$ is an isomorphism of $\mathcal{C}_{p}(\mathbb{T})$ which extends to an isomorphism of $\mathcal{B}_{p,p}^{1/p}(\mathbb{T})$ with $V_{h}^{-1}=V_{h^{-1}}$. Now if one consider any quasisymmetry $h$ of the unit circle we know [1] that $h$ is the conformal welding of some quasicircle. We deduce from this geometric approach that $V_{h}$ is an isomorphism of $\mathcal{B}_{p,p}^{1/p}(\mathbb{T})$ for any quasisymmetry $h$. Conversely, Nag-Sullivan [33] for the case $p=2$ and Bourdaud [5] for the general case have proven that if $h$ is a circle homeomorphism and if $V_{h}$ is an isomorphism of $\mathcal{B}_{p,p}^{1/p}(\mathbb{T})$ then $h$ must be a quasisymmetry. Vodop’Yanov[42], and later Bourdaud [5] moreover proved that if $s\neq 1/p$ then the circle homeomorphisms $h$ such that $V_{h}$ operates on $\mathcal{B}_{p,p}^{s}(\mathbb{T})$ isomorphically are the bi-Lipschitz ones which form a proper subgroup of the group of quasisymmetries.
2.1 Boundary Values of Functions $\mathcal{B}_{p,p}^{1/p}(\Omega)$
When dealing with a function $u$ in $\mathcal{B}_{p,p}^{1/p}(\Omega_{i})$ that is continuous up to the boundary the boundary values are obvious: it is just the restriction to the boundary, and moreover we have just seen that if we take the harmonic extension of this boundary value in $\Omega_{e}$ we get a function $v\in\mathcal{B}_{p,p}^{1/p}(\Omega_{e})$ with a norm equivalent to the one we started with. Now let us consider a general function $u\in\mathcal{B}_{p,p}^{1/p}(\Omega_{i})$: as we have seen, this function is the limit of a sequence $(u_{n})\in\mathcal{C}_{p}(\Omega_{i})$ and we put $f_{n}$ to be the boundary value of $u_{n}$. The harmonic extension to $\Omega_{e}$ of $f_{n}$ is then a Cauchy sequence in $\mathcal{B}_{p,p}^{1/p}(\Omega_{e})$, thus convergent to some $v\in\mathcal{B}_{p,p}^{1/p}(\Omega_{e})$ and we may regard the couple $(u,v)$ as an abstract version of the boundary value of $u$, and also $v$.
If we want to be more concrete, we use the Riemann maps $\varphi_{i,e}$: $u\circ\varphi_{i}\in\mathcal{B}_{p,p}^{1/p}(\mathbb{D}_{i})\subset h^{p}(\mathbb{D})$, the classical harmonic Hardy space of the disk, so it has radial boundary values almost everywhere on the circle defining a function $b_{i}\in L^{p}(\mathbb{T})$. Notice that this fact translates in $\Omega_{i}$ by saying that $u$ has limits along $\omega_{i}$-almost every internal ray, where $\omega_{i}$ stands for harmonic measure in $\Omega_{i}$. Notice that this limit $b_{i}$ characterizes $u$. We may do the same thing with $v$ and define an analogous function $b_{e}\in L^{p}(\mathbb{T})$. When $u\in\mathcal{C}_{p}(\Omega_{i})$ we have seen that $b_{e}=V_{h}(b_{i})$ where $h$ is the conformal welding of $\Gamma$. Unfortunately it does not make sense in $L^{p}(\mathbb{T})$ since $h$ need not be absolutely continuous with respect to the Lebesgue measure, a fact which transfers geometrically to the fact that $\omega_{i}$ and $\omega_{e}$ need not be mutually absolutely continuous. Nevertheless, the relation $b_{e}=V_{h}(b_{i})$ remains true if we interpret $V_{h}$ on $\mathcal{B}_{p,p}^{1/p}(\mathbb{T})$ as the completion of $V_{h}|_{\mathcal{C}_{p}^{1/p}(\mathbb{T})}$.
In the case of $p=2$, Schippers-Staubach [36] has given an even more concrete description of the boundary values: we outline their argument. On the one hand, for any $u\in\mathcal{B}_{2,2}^{1/2}(\Omega_{i})$ we have $u\circ\varphi_{i}\in\mathcal{B}_{2,2}^{1/2}(\mathbb{D}_{i})$ as before. In this case, $u\circ\varphi_{i}$ has radial boundary values everywhere on the circle $\mathbb{T}$ except on a Borel set $F_{1}$ of Logarithmic capacity zero, defining a function $b_{i}\in L^{2}(\mathbb{T})$. On the other hand, let $h:=\varphi_{i}^{-1}\circ\varphi_{e}$, the conformal welding with respect to the quasicircle $\Gamma$, that is a quasisymmetry (so is $h^{-1}$) as we pointed out above. Since a quasisymmetry takes a Borel set of Logarithmic capacity zero to a Borel set of Logarithmic capacity zero (see [2]) we see that $F_{2}:=h^{-1}(F_{1})$ is also a Borel set of Logarithmic capacity zero. Then $b_{e}=b_{i}\circ\varphi_{i}^{-1}\circ\varphi_{e}=V_{h}(b_{i})$ is well-defined on $\mathbb{T}\setminus F_{2}$ and belongs to $L^{2}(\mathbb{T})$. It is known that the union $F_{1}\cup F_{2}$ is also of Logarithmic capacity zero, and in particular, Lebesgue measure zero since a set of Logarithmic capacity zero has zero Lebesgue measure. Put $P_{e}(b_{e})$ to be the harmonic extension in $\mathbb{D}_{e}$ of $b_{e}$, and $v:=P_{e}(b_{e})\circ\varphi_{e}^{-1}$ is the harmonic extension in $\Omega_{e}$ of $b_{i}\circ\varphi_{i}^{-1}$. We can see that $u$ and $v$ have the same boundary values along every internal ray on $\Gamma$, except on a set whose harmonic measure is zero with respect to both $\Omega_{i}$ and $\Omega_{e}$. Moreover, since the spaces $\mathcal{B}_{p,p}^{1/p}(\Omega_{i,e})$ increase with $p$ the same argument also holds for the case of $1<p<2$.
2.2 Plemelj-Calderón Problem for $\mathcal{B}_{2,2}^{1/2}(\Gamma)$
We can now address Plemelj-Calderón problem for $\mathcal{B}_{2,2}^{1/2}(\Gamma)$. The argument is broken into two steps. In the first we consider the function in the space $D(\Gamma)$, the restriction space to $\Gamma$ of $C_{c}^{\infty}(\mathbb{C})$. The second step consists in an approximation process.
For any test function $f$, it holds that
$f(z)=-\frac{1}{\pi}\iint_{\mathbb{C}}\frac{\bar{\partial}f(\zeta)}{\zeta-z}d\xi d\eta,\qquad z\in\mathbb{C}.$ (2.4)
Indeed,
$\frac{1}{\pi}\iint_{\mathbb{C}}\frac{\bar{\partial}f(\zeta)}{z-\zeta}d\xi d\eta=\bar{\partial}f(z)\ast\left(\frac{1}{\pi z}\right)=\bar{\partial}\left(\frac{1}{\pi z}\right)\ast f(z)=\delta_{0}\ast f(z)=f(z)$
where $\ast$ stands for convolution, and $\delta_{0}$ denotes the Dirac measure at 0. Define
$\widehat{F_{i}}(z)=-\frac{1}{\pi}\iint_{\Omega_{e}}\frac{\bar{\partial}f(\zeta)}{\zeta-z}d\xi d\eta,\qquad z\in\Omega_{i},$ (2.5)
$\widehat{F_{e}}(z)=\frac{1}{\pi}\iint_{\Omega_{i}}\frac{\bar{\partial}f(\zeta)}{\zeta-z}d\xi d\eta,\qquad z\in\Omega_{e}.$ (2.6)
It is easily seen that $\widehat{F_{i}}$ and $\widehat{F_{e}}$ are holomorphic functions in $\Omega_{i}$ and $\Omega_{e}$, respectively with $\widehat{F_{e}}(z)=O(\frac{1}{|z|})$ at $\infty$. Furthermore, it can be shown that both $\widehat{F_{i}}$ and $\widehat{F_{e}}$ define continuous functions at each $z\in\mathbb{C}$. It follows from (2.4) that on $\Gamma$,
$f=\widehat{F_{i}}-\widehat{F_{e}}$. (2.7)
Here, we have used that the quasicircle $\Gamma$ has zero area with respect to the two-dimensional Lebesgue measure. Taking derivative we have that
$\widehat{F_{i}}^{\prime}(z)=-\frac{1}{\pi}\iint_{\Omega_{e}}\frac{\bar{\partial}f(\zeta)}{(\zeta-z)^{2}}d\xi d\eta=B(\bar{\partial}f\chi_{\Omega_{e}})(z).$
Here, $\chi_{\Omega_{e}}$ is the characteristic function of the domain $\Omega_{e}$ and $B$ denotes the Beurling operator, i.e., the convolution with $\frac{1}{\pi}\text{p.v.}\frac{1}{z^{2}}$ which is an isometry of $L^{2}(\mathbb{C})$. Then,
$\|\widehat{F_{i}}\|_{\mathcal{B}_{2,2}^{1/2}(\Omega_{i})}$ $=\|\widehat{F_{i}}^{\prime}\|_{L^{2}(\Omega_{i})}=\|B(\bar{\partial}f\chi_{\Omega_{e}})\|_{L^{2}(\Omega_{i})}$ (2.8)
$\leq\|B(\bar{\partial}f\chi_{\Omega_{e}})\|_{L^{2}(\mathbb{C})}\leq\|\bar{\partial}f\|_{L^{2}(\Omega_{e})}\leq\|\bar{\partial}f\|_{L^{2}(\mathbb{C})}.$
Similarly, it follows from (2.6) that
$\|\widehat{F_{e}}\|_{\mathcal{B}_{2,2}^{1/2}(\Omega_{e})}\leq\|\bar{\partial}f\|_{L^{2}(\mathbb{C})}.$ (2.9)
Lemma 2.2.
For any $f\in C_{c}^{\infty}(\mathbb{C})$, the integrals in (2.5) and (2.6) depend only on the boundary values $f$ on $\Gamma$, not on the specific extensions belonging to $B_{2,2}^{1/2}(\Omega_{i,e})$ of $f$ to respectively $\Omega_{i}$ and $\Omega_{e}$. In particular, $\widehat{F_{i}}$ and $\widehat{F_{e}}$ defined in respectively (2.5) and (2.6) equal to
$F_{i}(z)=-\frac{1}{\pi}\iint_{\Omega_{e}}\frac{\bar{\partial}u_{e}(\zeta)}{\zeta-z}d\xi d\eta,\qquad z\in\Omega_{i},$ (2.10)
$F_{e}(z)=\frac{1}{\pi}\iint_{\Omega_{i}}\frac{\bar{\partial}u_{i}(\zeta)}{\zeta-z}d\xi d\eta,\qquad z\in\Omega_{e}$ (2.11)
so that
$f=F_{i}-F_{e}$ (2.12)
on $\Gamma$. Here, $u_{i}$ and $u_{e}$ are harmonic extensions of $f$ to respectively $\Omega_{i}$ and $\Omega_{e}$.
Proof.
Note that $u_{i}$ and $u_{e}$ are continuous on $\overline{\Omega}_{i}$ and $\overline{\Omega}_{e}$, respectively since the boundary function $f$ is continuous on $\Gamma$. By the Dirichlet principle, $u_{i,e}\in\mathcal{B}_{2,2}^{1/2}(\Omega_{i,e})$. Using the similar computation we can see that the holomorphic functions $F_{i}\in\mathcal{B}_{2,2}^{1/2}(\Omega_{i})$ and $F_{e}\in\mathcal{B}_{2,2}^{1/2}(\Omega_{e})$.
By the theorem of Gol’dshtein-Latfullin-Vodop’yanov [18] (see also [22]) $u_{i}$ can be extended to $U_{i}\in W^{1,2}(\mathbb{C})$, defining $U_{i}=u_{i}\circ R$ on $\Omega_{e}$ where $R$ is a quasiconformal reflection across $\Gamma$. There exists a sequence $(U_{i,n})$ of $C_{c}^{\infty}(\mathbb{C})$ converging to $U_{i}$ in $W^{1,2}(\mathbb{C})$ (see e.g. Ch.11 in [26]). Denote $U_{i,n}|_{\Gamma}$ by $f_{i,n}$. By what proceeds, $f_{i,n}$ may be written as $\widehat{F_{i,n}}-\widehat{F_{e,n}}$ on $\Gamma$. Here,
$\widehat{F_{i,n}}(z)=-\frac{1}{\pi}\iint_{\Omega_{e}}\frac{\bar{\partial}U_{i,n}(\zeta)}{\zeta-z}d\xi d\eta,\qquad z\in\Omega_{i}$
and
$\widehat{F_{e,n}}(z)=\frac{1}{\pi}\iint_{\Omega_{i}}\frac{\bar{\partial}U_{i,n}(\zeta)}{\zeta-z}d\xi d\eta,\qquad z\in\Omega_{e}.$
Notice that $\widehat{F_{i,n}}$ and $\widehat{F_{e,n}}$ are holomorphic in $\Omega_{i},\,\Omega_{e}$ respectively, continuous in $\mathbb{C}$, and $F_{e,n}(z)=O(\frac{1}{|z|})$ as $z\to\infty$. Similar to the reasoning in (2.8) we have that
$\begin{split}\|\widehat{F_{e,n}}-F_{e}\|_{\mathcal{B}_{2,2}^{1/2}(\Omega_{e})}&=\|(\widehat{F_{e,n}}-F_{e})^{\prime}\|_{L^{2}(\Omega_{e})}\leq\|\bar{\partial}U_{i,n}-\bar{\partial}u_{i})\|_{L^{2}(\Omega_{i})}\\ &\leq\|\nabla(U_{i,n}-u_{i})\|_{L^{2}(\Omega_{i})}\leq\|U_{i,n}-U_{i}\|_{W^{1,2}(\mathbb{C})}\to 0\end{split}$
[…]
$\|\nabla(P_{i}(f_{i,n})-u_{i})\|_{L^{2}(\Omega_{i})}\leq\|\nabla(U_{i,n}-u_{i})\|_{L^{2}(\Omega_{i})}\to 0$
where $P_{i}(f_{i,n})$ denotes the harmonic extension in $\Omega_{i}$ of $f_{i,n}$. We thus have that $f_{i,n}(z)\to f(z)$ at each point $z\in\Gamma$.
[…]
$f=G_{i}-F_{e}.$ (2.13)
Let us now consider $u_{e}$ the harmonic extension of $f$ in $\Omega_{e}$ and $U_{e}$ an extension of $u_{e}$ belonging to $W^{1,2}(\mathbb{C})$. If we take a sequence $(U_{e,n})$ of $C_{c}^{\infty}(\mathbb{C})$ converging to $U_{e}$ we can prove, as before, that
$f=F_{i}-G_{e}.$ (2.14)
By uniqueness of the decomposition, proved in [27] mainly based on an identity in [39], we see that these two decomposition (2.13),(2.14) are same as (2.7) so that $\widehat{F_{i}}=F_{i}$,  $\widehat{F_{e}}=F_{e}$ and $f=F_{i}-F_{e}$ on $\Gamma$. ∎
Concerning the above proof, we give two side remarks. Firstly, using a part of the argument in the proof of Lemma 2.2, we can see that $C_{c}^{\infty}(\mathbb{C})|_{\Gamma}$ is a dense subset of $\mathcal{B}_{2,2}^{1/2}(\Gamma)$. Precisely, for any $f\in\mathcal{B}_{2,2}^{1/2}(\Gamma)$, its harmonic extension $u_{i}$ in $\Omega_{i}$ can be extended to $U_{i}\in W^{1,2}(\mathbb{C})$. There exists a sequence $(U_{i,n})$ of $C_{c}^{\infty}(\mathbb{C})$ converging to $U_{i}$ in $W^{1,2}(\mathbb{C})$. Using the Dirichlet principle, we may conclude the assertion.
Secondly, we explain that if the Jordan curve $\Gamma$ is rectifiable then the assertion in Lemma 2.2 can be derived easily from the generalized Green’s formula (see p.150 in [28]). It says that let $f$ have $L^{1}$-derivatives in the domain $G$ and if $\overline{\Omega}\subset G$ is a Jordan domain with rectifiable boundary $\Gamma=\partial\Omega$ then
$\iint_{\Omega}f_{\bar{z}}dxdy=-\frac{i}{2}\int_{\Gamma}fdz.$
Let $f\in C_{c}^{\infty}(\mathbb{C})$ and $\Gamma$ be a rectifiable Jordan curve and let $u_{e}$ be the harmonic extension of $f$, restricted on $\Gamma$, to $\Omega_{e}$, that is continuous on $\Omega_{e}\cup\Gamma$. By the Dirichlet principle, $u_{e}\in W^{1,2}(\Omega_{e})$. As we noticed before, $u_{e}$ can be extended to $U_{e}\in W^{1,2}(\mathbb{C})$ using the quasiconformal reflection across $\Gamma$. Note that $f-U_{e}$ has $L^{2}$-derivatives in $\mathbb{C}$. By the generalized Green’s formula,
$\iint_{\Omega_{e}}\frac{\bar{\partial}(f-u_{e})(\zeta)}{\zeta-z}d\xi d\eta=-\frac{i}{2}\int_{\Gamma}\frac{(f-u_{e})|_{\Gamma}(\zeta)}{\zeta-z}d\zeta,\quad z\in\Omega_{i}$
which equals to 0 since $(f-u_{e})|_{\Gamma}=0$ on $\Gamma$. Thus, we see that
$\iint_{\Omega_{e}}\frac{\bar{\partial}f(\zeta)}{\zeta-z}d\xi d\eta=\iint_{\Omega_{e}}\frac{\bar{\partial}u_{e}(\zeta)}{\zeta-z}d\xi d\eta$
which implies that the integral does not depend on the specific extension provided that $\Gamma$ is rectifiable.
Theorem 2.3.
Let $\Gamma$ be a quasicircle. Any function $f\in\mathcal{B}_{2,2}^{1/2}(\Gamma)$ admits a unique decomposition $f=F_{i}-F_{e}$ on $\Gamma$, see (2.13), with $F_{i,e}$ being holomorphic, $F_{i,e}\in\mathcal{B}_{2,2}^{1/2}(\Omega_{i,e})$ and
$\|F_{i,e}\|_{\mathcal{B}_{2,2}^{1/2}(\Omega_{i,e})}\leq C\|f\|_{\mathcal{B}_{2,2}^{1/2}(\Gamma)}$
for every $f\in\mathcal{B}_{2,2}^{1/2}(\Gamma)$ where the constant $C$ depends only on $\Gamma$.
Proof.
Let now $f$ be any function of $\mathcal{B}_{2,2}^{1/2}(\Gamma)$. It can be shown that the functions $F_{i}$ and $F_{e}$, defined as (2.10) and (2.11), are still well-defined in this general case, holomorphic in the domains of definition and continuous in $\mathbb{C}$. By the same reasoning as above, we have
$\begin{split}&\|F_{i}\|_{\mathcal{B}_{2,2}^{1/2}(\Omega_{i})}\leq\|u_{e}\|_{\mathcal{B}_{2,2}^{1/2}(\Omega_{e})}\simeq\|f\|_{\mathcal{B}_{2,2}^{1/2}(\Gamma)},\\ &\|F_{e}\|_{\mathcal{B}_{2,2}^{1/2}(\Omega_{e})}\leq\|u_{i}\|_{\mathcal{B}_{2,2}^{1/2}(\Omega_{i})}\simeq\|f\|_{\mathcal{B}_{2,2}^{1/2}(\Gamma)}.\end{split}$
Here and in what follows, the notation $A\simeq B$" means there exists a constant $C$ such that $A/C\leq B\leq CA$. In the above the implicit constant depends only on $\Gamma$.
Recall that $C_{c}^{\infty}(\mathbb{C})|_{\Gamma}$ is a dense subset of $\mathcal{B}_{2,2}^{1/2}(\Gamma)$. Using (2.13) it is not hard to show $f=F_{i}|_{\Gamma}-F_{e}|_{\Gamma}$ by an approximation process. The details are left to the reader. ∎
2.3 Plemelj-Calderón Problem for $\mathcal{B}_{p,p}^{1-1/p}(\Gamma)$
The space $\mathcal{B}_{2,2}^{1/2}(\Gamma)$ has two remarkable properties: it is conformally invariant and its definition does not involve the distance to the boundary. As a corollary, as we have seen, Plemelj-Calderón property holds for this space for all quasicircles. For $p\neq 2$ these properties split: the conformality property remains for the spaces $\mathcal{B}_{p,p}^{1/p}(\Gamma),\,p>1$ but in its definition a distance term $d(z,\Gamma)^{p-2}$ appears. This case will be discussed in the next section, together with the general case of $s\in(0,1)$ with a distance term $d(z,\Gamma)^{(1-s)p-1}$. The other property transfers to $\mathcal{B}_{p,p}^{1-1/p}(\Gamma)$. Notice that $1-1/p=1/p^{\prime}$ where $p^{\prime}$ is the coefficient conjugate to $p$.
Define $\widetilde{\mathcal{B}}_{p,p}^{1-1/p}(\Gamma)$ as the closure of $C_{c}^{\infty}(\mathbb{C})|_{\Gamma}\cap\mathcal{B}_{p,p}^{1-1/p}(\Omega_{i}\to\Gamma)\cap\mathcal{B}_{p,p}^{1-1/p}(\Omega_{e}\to\Gamma)$ in $\mathcal{B}_{p,p}^{1-1/p}(\Gamma)$. We have the following
Theorem 2.4.
Let $\Gamma$ be a quasicircle and $p>2$. Every $f\in\widetilde{\mathcal{B}}_{p,p}^{1-1/p}(\Gamma)$ admits a unique decomposition $f=F_{i}-F_{e}$ with $F_{i,e}$ being holomorphic, $F_{i,e}\in\mathcal{B}_{p,p}^{1-1/p}(\Omega_{i,e})$ and
$\|F_{i,e}\|_{\mathcal{B}_{p,p}^{1-1/p}(\Omega_{i,e})}\leq C\|f\|_{\mathcal{B}_{p,p}^{1-1/p}(\Gamma)}$
for every $f\in\widetilde{\mathcal{B}}_{p,p}^{1-1/p}(\Gamma)$ where the constant $C$ depends only on $\Gamma$ and $p$.
Before we proceed to the proof of this theorem, one must properly define what we mean by $\mathcal{B}_{p,p}^{1-1/p}(\Gamma)$. It happens that functions in $\mathcal{B}_{p,p}^{1-1/p}(\Omega_{i,e})$ are Hölder continuous in the closure of $\Omega_{i,e}$ with exponent $\alpha=1-2/p$. This may be seen by two different ways. Firstly, s=11/p>1/ps=1-1/p>1/p, we will prove later that functions in $\mathcal{B}_{p,p}^{s}(\Omega_{i,e})$ are $\alpha$-Hölder continuous in the closure of $\Omega_{i,e}$ with $\alpha=s-1/p$. The other way to prove the same thing is to appeal to Jones’s theorem: a function in $\mathcal{B}_{p,p}^{1-1/p}(\Omega_{i,e})$ can be extended to a function in $W^{1,p}(\mathbb{C})$ even though clearly one cannot extend in this case by using quasiconformal reflection argument, used in the case of $p=2$ as above (see [22], also see [8]). The Morrey’s inequality states that $W^{1,p}(\mathbb{C})\subset\Lambda^{\alpha}(\mathbb{C})$, $p>2$ (note that this inclusion relation holds in $\mathbb{R}^{n}$ for $p>n$) for $\alpha=1-2/p$, the space of $\alpha$-Hölder continuous functions in $\mathbb{C}$. Knowing this fact one may re-define the space $\mathcal{B}_{p,p}^{1-1/p}(\Gamma)$ as the space of $\Lambda^{1-2/p}$-functions on $\Gamma$ whose harmonic extensions to $\Omega_{i,e}$ belong to $\mathcal{B}_{p,p}^{1-1/p}(\Omega_{i,e})$, that is,
$\mathcal{B}_{p,p}^{1-1/p}(\Gamma)=\Lambda^{1-2/p}(\Gamma)\cap\mathcal{B}_{p,p}^{1-1/p}(\Omega_{i}\to\Gamma)\cap\mathcal{B}_{p,p}^{1-1/p}(\Omega_{e}\to\Gamma).$
Proof.
Thanks to Jones’ theorem the proof is exactly the same as for $p=2$ because the Beurling transform is bounded on $L^{p}(\mathbb{C})$ for all $p>1$. Notice that the argument of uniqueness used in the case $p=2$ goes through here because $L^{p}(\Omega_{i})\subset L^{2}(\Omega_{i})$ for $p>2$. ∎
3 Fractional Sobolev Spaces on Quasidisks
3.1 Plemelj-Calderón Property for General $\mathcal{B}_{p,p}^{s}(\Gamma)$
As mentioned in the proceeding section, in order to deal with Plemelj-Calderón property for general $p>1,s\in(0,1)$ we need to deal with boundedness of the Beurling transform on $L^{p}(\omega,\,\mathbb{C})$ where $\omega$ is a weight. By a weight we mean a nonnegative locally integrable function on $\mathbb{C}$. Since the Beurling transform is a Calderón-Zygmund operator, we know [9] that this property holds if and only if $\omega$ belongs to the so-called class $A_{p}$ whose definition we now recall.
Definition 3.1.
A weight $\omega:\mathbb{C}\to\mathbb{R}^{+}$ is said to satisfy the $A_{1}$ condition on $\mathbb{C}$ if there exists $C>0$ such that for any disk $D$ of the plane,
$\frac{1}{|D|}\iint_{D}\omega(z)dxdy\leq C\omega(z)$
for almost all $z\in D$; while $\omega$ is said to satisfy the $A_{\infty}$ condition if
$\frac{1}{|D|}\iint_{D}\omega(z)dxdy\leq C\exp\left(\frac{1}{|D|}\iint_{D}\log\omega(z)dxdy\right).$
For any $p>1$, the $A_{p}$ class is the set of weights $\omega:\mathbb{C}\to\mathbb{R}^{+}$ such that there exists $C>1$ such that for every disk $D$ of the plane,
$\frac{1}{|D|}\iint_{D}\omega(z)dxdy\left(\frac{1}{|D|}\iint_{D}w(z)^{-\frac{1}{p-1}}dxdy\right)^{p-1}\leq C.$ (3.1)
For future use let us notice that $A_{p}$ weights on $\mathbb{T}$ are defined similarly. It is known that for all $p>1$, $A_{1}\subset A_{p}\subset A_{\infty}$, $\omega\in A_{p}$ is equivalent to $\omega^{-1/(p-1)}\in A_{p^{\prime}}$ where $1/p+1/p^{\prime}=1$, and $A_{p_{1}}\subset A_{p_{2}}$ for $1<p_{1}<p_{2}$. By Hölder inequality the left part of the inequality (3.1) is always $\geq 1$: one can thus interpret this class as verifying some kind of reverse Hölder inequality. If the weight $\omega$ has $A_{p}$ then any Calderon-Zygmund operator is bounded on the weighted space $L^{p}(\omega,\mathbb{C})$ (see [9]).
If $f\in C_{c}^{\infty}(\mathbb{C})$ we recall that
$f(z)$ $=-\frac{1}{\pi}\iint_{\mathbb{C}}\frac{\bar{\partial}f(\zeta)}{\zeta-z}d\xi d\eta$
$=F_{i}(z)-F_{e}(z)$
where
$F_{i}(z)=-\frac{1}{\pi}\iint_{\Omega_{e}}\frac{\bar{\partial}f(\zeta)}{\zeta-z}d\xi d\eta,\quad F_{e}(z)=\frac{1}{\pi}\iint_{\Omega_{i}}\frac{\bar{\partial}f(\zeta)}{\zeta-z}d\xi d\eta.$
Notice $F_{i}^{\prime}(z)=B(\bar{\partial}f\chi_{\Omega_{e}})(z)=B(\bar{\partial}u_{e}\chi_{\Omega_{e}})(z)$, $F_{e}^{\prime}(z)=-B(\bar{\partial}f\chi_{\Omega_{i}})(z)=-B(\bar{\partial}u_{i}\chi_{\Omega_{i}})(z)$ where $u_{i,e}$ are the harmonic extensions of $f|_{\Gamma}$ to $\Omega_{i,e}$, we have that a sufficient condition on $\Gamma$ that implies $F_{i,e}\in\mathcal{B}_{p,p}^{s}(\Omega_{i,e})$ is that the weight $\omega(z)=d(z,\Gamma)^{(1-s)p-1}$ satisfies the $A_{p}$ condition. Precisely, under this condition we have
$\|F_{i}\|_{\mathcal{B}_{p,p}^{s}(\Omega_{i})}\leq C\|u_{e}\|_{\mathcal{B}_{p,p}^{s}(\Omega_{e})}\leq C\|f\|_{\mathcal{B}_{p,p}^{s}(\Gamma)},$
$\|F_{e}\|_{\mathcal{B}_{p,p}^{s}(\Omega_{e})}\leq C\|u_{i}\|_{\mathcal{B}_{p,p}^{s}(\Omega_{i})}\leq C\|f\|_{\mathcal{B}_{p,p}^{s}(\Gamma)}$
where $C$ is an absolute constant.
The dependence of the condition that $d(z,\Gamma)^{(1-s)p-1}$ having $A_{p}$ on the geometry of the curve has been studied by Astala [3]. In order to state his results we first need to introduce some notions.
For a compact set $E\subset\mathbb{C}$ and $0<\delta\leq 2$ let
$M_{\delta}(E,t)=\frac{|E+D(0,t)|}{t^{2-\delta}}.$
Here, $E+D(0,t)$, by the definition $\{e+z:\;e\in E,\;z\in D(0,t)\}$, is a set of points at a distance less than $t$ from $E$ which is called Minkowski sausage. We then define a kind of Minkowski content by
$h_{\delta}(E)=\sup\limits_{0<t\leq\mathrm{diam}(E)}{M_{\delta}(E,t)}.$
Definition 3.2.
We say that a Jordan curve $\Gamma$ is $\delta$-regular if there exists $C>0$ such that for every disk $D(z,R)\subset\mathbb{C}$,
$h_{\delta}(\Gamma\cap D(z,R))\leq CR^{\delta}.$
It is known that $1$-regularity is equivalent to Ahlfors-regularity.
It has been proved that for any quasicircle $\Gamma$ there exists $\delta<2$ such that $\Gamma$ is $\delta$-regular. We may thus have, for a quasicrcle $\Gamma$,
$h(\Gamma)=\inf\{\delta:\Gamma\,\mathrm{is}\,\delta\text{-}\mathrm{regular}\}\in[1,2)$
and we may say the degree of regularity of $\Gamma$ is $h(\Gamma)$. For more information on $h(\Gamma)$, and its relation with the Hausdorff dimension $\text{dim}_{H}(\Gamma)$ of $\Gamma$, see [3].
Definition 3.3.
A compact subset $E$ of the complex plane is said to be porous if there exists $c\in(0,1)$ and $r_{0}>1$ such that for every $z\in\mathbb{C}$ and $0<r\leq r_{0}$, the disk $D(z,r)$ contains a disk of radius $cr$ which does not intersect $E$.
Theorem 3.4 ([3]).
Let $\alpha\in(0,1)$ and $\Gamma$ be a porous Jordan curve. We have, for any $1<p<\infty$,
$d(z,\Gamma)^{\alpha-1}\in A_{p}\Leftrightarrow d(z,\Gamma)^{\alpha-1}\in A_{1}\Leftrightarrow\alpha>h(\Gamma)-1.$ (3.2)
Define $\widetilde{\mathcal{B}}_{p,p}^{s}(\Gamma)$ to be the closure of
$C_{c}^{\infty}(\mathbb{C})|_{\Gamma}\cap\mathcal{B}_{p,p}^{s}(\Omega_{i}\to\Gamma)\cap\mathcal{B}_{p,p}^{s}(\Omega_{e}\to\Gamma)$
in $\mathcal{B}_{p,p}^{s}(\Gamma).$ Notice that every quasicircle is porous (see [41]). We may now state the principal theorem of this section:
Figure 1: Domain formed by points $(h(\Gamma),s)$: case p>2 and case p<2
Theorem 3.5.
Let $\Gamma$ be a quasicircle and $1<p<\infty,\;0<s<1$. If the point $(h(\Gamma),s)$ locates in the shadowed region (see Figure 1); that is,
$(h(\Gamma)-1)\frac{p-1}{p}<s<\frac{p+1-h(\Gamma)}{p},$ (3.3)
then any $f\in\widetilde{\mathcal{B}}_{p,p}^{s}(\Gamma)$ has a decomposition $f=F_{i}-F_{e}$ with $F_{i,e}$ being holomorphic, $F_{i,e}\in\mathcal{B}_{p,p}^{s}(\Omega_{i,e})$ and
$\|F_{i,e}\|_{\mathcal{B}_{p,p}^{s}(\Omega_{i,e})}\leq C\|f\|_{\mathcal{B}_{p,p}^{s}(\Gamma)}$ (3.4)
for every $f\in\widetilde{\mathcal{B}}_{p,p}^{s}(\Gamma)$ where the constant $C$ depends only on $\Gamma$ and p,sp,\;s.
Proof.
Following the discussion above, it is clear that (3.4) holds if it holds that the weight $d(z,\Gamma)^{(1-s)p-1}\in A_{p}$. Assume first that $s>\frac{p-1}{p}$ so that (1s)p<1(1-s)p<1. By Theorem 3.4, $d(z,\Gamma)^{(1-s)p-1}\in A_{p}\Leftrightarrow(1-s)p>h(\Gamma)-1\Leftrightarrow s<\frac{p+1-h(\Gamma)}{p}$.
Assume now $s<\frac{p-1}{p}$. Then, $d(z,\Gamma)^{(1-s)p-1}\in A_{p}\Leftrightarrow d(z,\Gamma)^{\frac{1-(1-s)p}{p-1}}\in A_{p^{\prime}}$. Here, $\frac{1-(1-s)p}{p-1}=\frac{sp}{p-1}-1$ with $\frac{sp}{p-1}<1$. By Theorem 3.4 again we have thus that $d(z,\Gamma)^{(1-s)p-1}\in A_{p}\Leftrightarrow s>\frac{p-1}{p}(h-1)$. The decomposition can be proved similar to that of Theorem 2.3. The case of s=11/ps=1-1/p has been addressed in the last section. ∎
We define $\mathcal{A}$ to be the admissible set involved in Theorem 3.5, that is,
$\mathcal{A}=\{(h(\Gamma),p,s)\in[1,2)\times(1,+\infty)\times(0,1):(h(\Gamma)-1)(p-1)<sp<p+1-h(\Gamma)\},$
so that if $\Gamma$ is a quasicircle and $p>1,\,s\in(0,1)$ such that $(h(\Gamma),p,s)\in\mathcal{A}$ then Plemelj-Calderón property holds for $\widetilde{\mathcal{B}}_{p,p}^{s}(\Gamma)$. Figure 1 shows slices of $\mathcal{A}$ first with $p>1$ fixed and Figure 2 with $h(\Gamma)$ given.
Figure 2: Domain formed by points $(p,s)$: case h>3/2 and case h<3/2
We end this part with a comment on the triangular domain above the shadowed region in Figure 1. Let the curve $\Gamma$ be a self-similar isotropic fractal which satisfies the open set condition. Its fractal dimension $d(\Gamma)$ equals to $h(\Gamma)$ in this case (see [32]). For instance, the Koch snowflake curve, we mentioned before, is a such quasicircle with $d(\Gamma)=\log 4/\log 3$. Denote by $N(\varepsilon)$ the minimal number of disks of radius $\varepsilon$ necessary to cover $\Gamma$. Then,
$d(\Gamma)=\lim_{\varepsilon\to 0}\frac{\log N(\varepsilon)}{\log\frac{1}{\varepsilon}}.$
Taking $\varepsilon=2^{-(n+1)}$ with sufficiently large $n$ one has the difference between the number $N(\varepsilon)$ and $2^{(n+1)d(\Gamma)}$ approaches 0 as $n$ increases to $\infty$. By dividing the domain $\Omega_{i}$ into pieces:
$\{z\in\Omega_{i}:\;d(z,\Gamma)\geq 2^{-1}\},\quad\{z\in\Omega_{i}:\;2^{-(j+1)}\leq d(z,\Gamma)<2^{-j}\},\quad j=1,2,3,\cdots$
and then computing directly one can see that the integral $\iint_{\Omega_{i}}d(z,\Gamma)^{(1-s)p-1}dxdy$ converges if and only if the series $\sum 2^{-n(p-sp+1-h(\Gamma))}$ converges, namely, $s<(p+1-h(\Gamma))/p$. Consequently, assuming $s\geq(p+1-h(\Gamma))/p$ one can see that in this case even the simplest function $u(z)=z$ is not in the space $\mathcal{B}^{s}_{p,p}(\Omega_{i})$.
3.2 Plemelj-Calderón Property for $\mathcal{B}_{p,p}^{1/p}(\Gamma)$
Recall that the dense subspace $\mathcal{C}_{p}(\Omega_{i,e})$ of $\mathcal{B}_{p,p}^{1/p}(\Omega_{i,e})$ has well-defined boundary values so that $\mathcal{C}_{p}(\Omega_{i}\!\to\!\Gamma)=\mathcal{C}_{p}(\Omega_{e}\!\to\!\Gamma)$, this space of boundary values called $\mathcal{C}_{p}(\Gamma)$. We may then define in an abstract way boundary values $\mathcal{B}_{p,p}^{1/p}(\Gamma)$ of functions in $\mathcal{B}_{p,p}^{1/p}(\Omega_{i,e})$ as equivalence classes of Cauchy sequences in $\mathcal{C}_{p}(\Gamma)$.
We may notice that the dense subspace defined in this way is not very tractable since we do not know the conformal mapping explicitly. The following result gives an explicit dense subspace of $\mathcal{B}_{p,p}^{1/p}(\Gamma)$.
Theorem 3.6.
Let $\Gamma$ be a quasicircle. If $h(\Gamma)<p<\frac{h(\Gamma)}{h(\Gamma)-1}$, then $C_{c}^{\infty}(\mathbb{C})|_{\Gamma}$ is dense in $\mathcal{B}_{p,p}^{1/p}(\Gamma)$.
Remark 3.7.
Under the assumption of Theorem 3.6, when $s<1/p$ the only thing that we can say is that $\mathcal{B}_{p,p}^{s}(\Gamma)$ contains $\mathcal{B}_{p,p}^{1/p}(\Gamma)$ and thus also $C_{c}^{\infty}(\mathbb{C})|_{\Gamma}$.
Proof of Theorem 3.6.
The assumption on $p$ is exactly the condition for $(h(\Gamma),p,\frac{1}{p})\in\mathcal{A}$. If it is satisfied we then have that $\omega(z)=d(z,\Gamma)^{p-2}$ is an $A_{p}$ weight, in particular, integrable locally in $\mathbb{C}$. Suppose $g\in C_{c}^{\infty}(\mathbb{C})$. Then,
$\iint_{\Omega_{i,e}}|\nabla g|^{p}d(z,\Gamma)^{p-2}dxdy\leq\iint_{\mathbb{C}}|\nabla g|^{p}d(z,\Gamma)^{p-2}dxdy<\infty.$
By that and the almost-Dirichlet principle (Lemma 2.1) one may show $g|_{\Gamma}\in\mathcal{B}_{p,p}^{1/p}(\Gamma)$.
For a general $f\in\mathcal{B}_{p,p}^{1/p}(\Gamma)$, we will use the argument on the boundary values of $\mathcal{B}_{p,p}^{1/p}(\Omega_{i,e})$ in the first paragraph of section 2.1. Let $u$ be the harmonic extension to $\Omega_{i}$ of $f$. Noting that $\omega\in A_{p}$ we see that $u$ has an extension $U\in W^{1,p}(\omega,\;\mathbb{C})$ to the whole plane (see [8]), and there exists a sequence $U_{n}$ in $C_{c}^{\infty}(\mathbb{C})$ such that (see Lemma 3.13 and also [8])
$\|U_{n}-u\|_{W^{1,p}(\omega,\;\Omega_{i})}\leq\|U_{n}-U\|_{W^{1,p}(\omega,\;\mathbb{C})}\to 0,\qquad n\to\infty.$ (3.5)
By the almost-Dirichlet principle (Lemma 2.1), (3.5) implies
$\|P_{i}(U_{n}|_{\Gamma})-u\|_{\mathcal{B}_{p,p}^{1/p}(\Omega_{i})}\to 0.$
The notations $P_{i}$ here and $P_{e}$ below denote the harmonic extensions in respectively $\Omega_{i}$ and $\Omega_{e}$ of functions defined on $\Gamma$. (3.5) also implies that the sequence $(U_{n})$ is a Cauchy sequence in $W^{1,p}(\omega,\Omega_{e})$, and then, using the almost-Dirichlet principle again, the sequence $(P_{e}(U_{n}|_{\Gamma}))$ is also a Cauchy sequence in $W^{1,p}(\omega,\Omega_{e})$, thus convergent to some $v\in\mathcal{B}_{p,p}^{1/p}(\Omega_{e})$.
Consequently,
$\|U_{n}|_{\Gamma}-f\|_{\mathcal{B}_{p,p}^{1/p}(\Gamma)}^{p}=\|P_{i}(U_{n}|_{\Gamma})-u\|_{\mathcal{B}_{p,p}^{1/p}(\Omega_{i})}^{p}+\|P_{e}(U_{n}|_{\Gamma})-v\|_{\mathcal{B}_{p,p}^{1/p}(\Omega_{e})}^{p}\to 0.$
This completes the proof of Theorem 3.6. ∎
As a corollary of Theorem 3.5 and Theorem 3.6 we immediately have the following.
Corollary 3.8.
Let $\Gamma$ be a quasicircle. If $h(\Gamma)<p<\frac{h(\Gamma)}{h(\Gamma)-1}$ then any function $f\in\mathcal{B}_{p,p}^{1/p}(\Gamma)$ admits the decomposition $f=F_{i}-F_{e}$ on $\Gamma$ with $F_{i}$ and $F_{e}$ being holomorphic, $F_{i,e}\in\mathcal{B}_{p,p}^{1/p}(\Omega_{i,e})$ and
$\|F_{i,e}\|_{\mathcal{B}_{p,p}^{1/p}(\Omega_{i,e})}\leq C\|f\|_{\mathcal{B}_{p,p}^{1/p}(\Gamma)}$
for every $f\in\mathcal{B}_{p,p}^{1/p}(\Gamma)$ where the constant $C$ depends on $\Gamma$ and $p$.
Remark 3.9.
We already know that this decomposition is unique if $p=2$. Since the spaces $\mathcal{B}_{p,p}^{1/p}(\mathbb{D})$ increase with $p$ we still have uniqueness for $h(\Gamma)<p\leq 2$.
3.3 Boundary Values of Functions in $\mathcal{B}_{p,p}^{s}(\Omega)$
Let $\Gamma$ be a quasicircle and $\Omega$ one of its complementary domains $\Omega_{i,e}$ as above. The boundary values of the critical case $s=1/p$ have been treated in the last section. As announced in the introduction, we will define now what we mean by boundary values of functions in $\mathcal{B}_{p,p}^{s}(\Omega),\,p>1,\,s\in(0,1)$. As we shall see the cases $s<1/p$ and $s>1/p$ will show to be very different. We start with the case of $s>1/p$.
Proposition 3.10.
Let $\Omega$ be a (bounded or unbounded) domain bounded by a bounded Jordan curve $\Gamma$. For any $u\in\mathcal{B}_{p,p}^{s}(\Omega)$ with $s\in(1/p,1)$ there exists a constant $C>0$ such that
$|\nabla u(z)|\leq Cd(z,\Gamma)^{\frac{-1-(1-s)p}{p}},\qquad z\in\Omega.$
Proof.
Let $z\in\Omega$ and $D$ is the disk $D(z,d(z,\Gamma)/2)$. By the mean value property of harmonic functions one may write
$\nabla u(z)=\frac{1}{|D|}\iint_{D}\nabla u(\zeta)d\xi d\eta.$
(Here and in the sequel we write $|D|$ for the Lebesgue measure of $D$.) so that, by Hölder inequality,
$|\nabla u(z)|^{p}$ $\leq\frac{4}{\pi d(z,\Gamma)^{2}}\iint_{D}|\nabla u(\zeta)|^{p}d\xi d\eta$
$\leq Cd(z,\Gamma)^{-1-(1-s)p}\iint_{D}|\nabla u(\zeta)|^{p}d(\zeta,\Gamma)^{(1-s)p-1}d\xi d\eta$
$\leq C\|u\|_{\mathcal{B}_{p,p}^{s}(\Omega)}^{p}d(z,\Gamma)^{-1-(1-s)p}.$
It follows that there exists $C>0$ such that $|\nabla u(z)|\leq Cd(z,\Gamma)^{\alpha-1}$ with $\alpha=s-1/p$. When $\Gamma$ is a quasicircle we may apply the following result due to Gehring-Martio ([17]).
Proposition 3.11.
Let $\Omega$ be a (bounded or unbounded) domain bounded by a quasicircle $\Gamma$ and $\alpha\in(0,1]$. The following are equivalent:
  1. (1)
    $\forall z\in\Omega,\;|\nabla u(z)|\leq Cd(z,\Gamma)^{\alpha-1}$ for some constant $C>0$;
  2. (2)
    $u\in\Lambda^{\alpha}(\overline{\Omega})$ .
Here, $\Lambda^{\alpha}(\overline{\Omega})$ stands for the space of Hölder functions of order $\alpha$.
If now $s\in(1/p,1)$ then $\alpha=s-1/p\in(0,1)$ and Gehring-Martio theorem implies that functions in $\mathcal{B}_{p,p}^{s}(\Omega)$ are continuous on $\overline{\Omega}$ , so that boundary values of functions in $\mathcal{B}_{p,p}^{s}(\Omega_{i,e})$ are well-defined and they characterize $u$, as being the unique harmonic extension of this continuous boundary value. We may now define the space $\mathcal{B}_{p,p}^{s}(\Gamma)$ for this range $s\in(1/p,1)$ as being the space of (s1/p)(s-1/p)-Hölder functions $f$ on $\Gamma$ whose harmonic extension $u_{i,e}$ to $\Omega_{i,e}$ belongs to $\mathcal{B}_{p,p}^{s}(\Omega_{i,e})$, and the space $\mathcal{B}_{p,p}^{s}(\Gamma)$ is assigned the natural norm $\|\cdot\|_{\mathcal{B}_{p,p}^{s}(\Gamma)}$ so that $\|f\|^{p}_{\mathcal{B}_{p,p}^{s}(\Gamma)}=\|u_{i}\|^{p}_{\mathcal{B}_{p,p}^{s}(\Omega_{i})}+\|u_{e}\|^{p}_{\mathcal{B}_{p,p}^{s}(\Omega_{e})}$.
3.3.1 Trace Mappings
As we have already seen, in the case of $s>1/p$ the spaces $\mathcal{B}_{p,p}^{s}(\Omega_{i,e})$ have well-defined boundary values and, with the terminology adopted in the introduction, the trace mapping
$\mathcal{B}_{p,p}^{s}(\Omega_{i,e})\to\mathcal{B}_{p,p}^{s}(\Omega_{i,e}\to\Gamma)$
is an isomorphism. The first purpose of this part is to determine conditions that imply
$\mathcal{B}_{p,p}^{s}(\Omega_{i}\to\Gamma)=\mathcal{B}_{p,p}^{s}(\Omega_{e}\to\Gamma).$
We already know the result for $s=1/p$ and $h(\Gamma)<p<h(\Gamma)/(1-h(\Gamma))$ (but in this case the boundary value has a different meaning). Now the space $\Lambda^{s}(\Omega)$ introduced above in connection with Gehring-Martio theorem coincides with the Besov space $\mathcal{B}_{\infty,\infty}^{s}(\Omega)$. A theorem of Hinkkanen [20] (see also [48]) states that suppose the curve $\Gamma$ is a Jordan curve and if $f\in\Lambda^{s}(\Gamma)$ with $s<1/2$ then the harmonic extension of $f$ to $\Omega_{i,e}$ is in $\mathcal{B}_{\infty,\infty}^{s}(\Omega)$ with norm bounded by that of $f$ and the index $1/2$ is critical. Hinkkanen’s result is true for all Jordan curves, and it has been improved to $s<1/K$ in [48] for quasicircles such that their Riemann mapping has a $K$-quasiconformal extension with $1\leq K\leq 2$. We thus define
$K_{*}:=\min{(K,2)}.$
We may now state the main result of this part.
Theorem 3.12.
Let $\Gamma,p,s$ be such that
$h(\Gamma)<p\leq\infty,\;1/p\leq s\;\mathrm{and}\;s<\frac{p+K_{*}-h(\Gamma)}{K_{*}p}$ (3.6)
then
$\mathcal{B}_{p,p}^{s}(\Omega_{i}\to\Gamma)=\mathcal{B}_{p,p}^{s}(\Omega_{e}\to\Gamma).$ (3.7)
Proof.
Let $P_{e}$, as before, be the operator which takes a function on $\Gamma$ to its harmonic extension in $\Omega_{e}$. As we have recalled that if
$h(\Gamma)<q<\frac{h(\Gamma)}{1-h(\Gamma)},\,\;\;0<\sigma<1/K_{*}$
then
$P_{e}:\,\mathcal{B}_{q,q}^{1/q}(\Omega_{i}\to\Gamma)$ $\to\mathcal{B}_{q,q}^{1/q}(\Omega_{e}),$
$P_{e}:\,\mathcal{B}_{\infty,\infty}^{\sigma}(\Omega_{i}\to\Gamma)$ $\to\mathcal{B}_{\infty,\infty}^{\sigma}(\Omega_{e}),$
are both bounded. We want to interpolate these results. Using complex interpolation, we have, for $0<\theta<1$,
$\mathcal{B}_{q,q}^{1/q}(\Omega),\mathcal{B}_{\infty,\infty}^{\sigma}(\Omega)]_{\theta}=\mathcal{B}_{p,p}^{s}(\Omega),$
with
$s$ $=\frac{1-\theta}{q}+\theta\sigma,$
$\frac{1}{p}$ $=\frac{1-\theta}{q}.$
We must decide what the couples $(p,s)$ are that may be "caught" in this way in view of the constraints: the relations above are equivalent to
$\sigma$ $=\frac{s-\frac{1}{p}}{\theta},$
$q$ $=(1-\theta)p,$
and the constraints are
$0<\frac{s-\frac{1}{p}}{\theta}$ $<1/K_{*},$
$h(\Gamma)<(1-\theta)p$ $<\frac{h(\Gamma)}{h(\Gamma)-1}.$
That correspond to $\theta$ being in the intersection of two intervals
$(K_{*}(s-\frac{1}{p}),1),\;(1-\frac{h(\Gamma)}{p(h(\Gamma)-1)},1-\frac{h(\Gamma)}{p}).$
We see that in order to have a solution we must have $s>1/p,p>h(\Gamma)$, and $K_{*}(s-\frac{1}{p})<1-\frac{h(\Gamma)}{p}$, which is the condition of the statement of the theorem. ∎
3.3.2 Weighted Sobolev Spaces
Let $\omega$ be an $A_{p}$ weight defined in $\mathbb{C}$, for some $p>1$. The weighted Sobolev space $W^{1,p}(\omega,\mathbb{C})$ is the space of tempered distributions $T$ such that $\nabla T$, taken in the sense of distributions, is a function in $L^{p}(\omega,\mathbb{C})$. We will use these spaces for $\omega(z)=d(z,\Gamma)^{(1-s)p-1}\in A_{p}$ provided that $(h(\Gamma),p,s)\in\mathcal{A}$.
For the usual technical reasons we need to know that $C_{c}^{\infty}(\mathbb{C})$-functions are dense in $W^{1,p}(\omega,\mathbb{C})$.
Lemma 3.13.
If $\omega\in A_{p}$ then the space $C_{c}^{\infty}(\mathbb{C})$ is dense in $W^{1,p}(\omega,\mathbb{C})$.
Proof.
Let us consider $f\in W^{1,p}(\omega,\mathbb{C})$: without loss of generally we may assume that $f$ has compact support. Let then $(\varphi_{\varepsilon})$ be a $C_{c}^{\infty}(\mathbb{C})$-approximation of identity, so that $f_{\varepsilon}=f*\varphi_{\varepsilon}\in C_{c}^{\infty}(\mathbb{C})$. Now
$\|f-f_{\varepsilon}\|_{W^{1,p}(\omega,\;\mathbb{C})}^{p}=\iint_{\mathbb{C}}|\nabla f(z)-\nabla f_{\varepsilon}(z)|^{p}\omega(z)dxdy$
which converges to 0 by the Lebesgue dominated convergence theorem. Indeed $|\nabla f_{\varepsilon}|$ is controlled by the Hardy-Littlewood maximal function of $\nabla f$ and this maximal function is in $L^{p}(\omega,\mathbb{C})$ from the fact that $\omega\in A_{p}$. ∎
In the classical case $\omega\equiv 1$ we recall that if $p>2$ then $W^{1,p}(\mathbb{C})\subset\Lambda^{\alpha}(\mathbb{C})$ with $\alpha=1-2/p$. In the weighted case we do not know this embedding result in general but we have two particular analogues.
Proposition 3.14.
The following two embedding assertions hold:
  • If $\omega\in A_{1}$ then every function in $W^{1,p}(\omega,\mathbb{C})$, $p>2$, is continuous;
  • If $\omega=d(z,\Gamma)^{(1-s)p-1}$ is only assumed to be in $A_{q}$ for some $q>1$ then the same result is true if $p>\frac{6-2h(\Gamma)}{2-h(\Gamma)}$.
Remark 3.15.
When $\omega(z)=d(z,\Gamma)^{(1-s)p-1}$ it is known that for any $q>1$, $\omega\in A_{q}\Leftrightarrow\omega\in A_{1}$ if 0<(1s)p<10<(1-s)p<1, i.e., s>11/ps>1-1/p.
Proof of Proposition 3.14.
Suppose first that $\omega\in A_{1}$. Choose $q$ so that $2<q<p$ and let $B$ be an open ball in $\mathbb{C}$ . By Hölder inequality, we get
$\int_{B}|\nabla f(x)|^{q}dxdy\leq\left(\int_{B}|\nabla f(x)|^{p}\omega(x)dxdy\right)^{\frac{q}{p}}\left(\int_{B}\omega(x)^{-\frac{q}{p-q}}dxdy\right)^{1-\frac{q}{p}}<\infty$
because $f\in W^{1,p}(\omega,\mathbb{C})$, and $\omega\in A_{1}\subset A_{p/q}$, which is equivalent to $\omega^{-q/(p-q)}\in A_{p/(p-q)}$, and in particular implies that $\omega^{-q/(p-q)}$ is locally integrable.
For the second case let us recall first that if $B$ is a ball containing $\Gamma$ then
$\iint_{B}d(z,\Gamma)^{\beta}dxdy<\infty$
if $\beta>h(\Gamma)-2$. By the previous computation it follows that a sufficient condition for the validity of the theorem is
$\frac{2}{p-2}>h(\Gamma)-2$
which is equivalent to the condition stated. ∎
We return to Theorem 3.12. It states that under certain hypothesis on $(h(\Gamma),p,s)$ we have
$\mathcal{B}_{p,p}^{s}(\Omega_{i}\to\Gamma)=\mathcal{B}_{p,p}^{s}(\Omega_{e}\to\Gamma).$
Let us call this space of boundary values $\mathcal{B}_{p,p}^{s}(\Gamma)$.
Theorem 3.16.
With the following assumption
$\frac{1}{p}<1-\frac{1}{p}\leq s<\frac{p+K_{*}-h(\Gamma)}{K_{*}p}$ (3.8)
we have that the elements of $W^{1,p}(\omega,\mathbb{C})$ are continuous functions in $\mathbb{C}$ and that
$\mathcal{B}_{p,p}^{s}(\Gamma)=W^{1,p}(\omega,\mathbb{C})|_{\Gamma},$ (3.9)
where $\omega(z)=d(z,\Gamma)^{(1-s)p-1}$ and $W^{1,p}(\omega,\mathbb{C})|_{\Gamma}=\{f|_{\Gamma}:\,f\in W^{1,p}(\omega,\mathbb{C})\}.$
Before we start the proof, let us comment on this statement. First of all, the assumption made implies that $\omega\in A_{p}$ and that
$\mathcal{B}_{p,p}^{s}(\Omega_{i}\to\Gamma)=\mathcal{B}_{p,p}^{s}(\Omega_{e}\to\Gamma)\;(=\mathcal{B}_{p,p}^{s}(\Gamma)).$
Moreover the inequality 1/p<11/p1/p<1-1/p is equivalent to $p>2$ and the inequality s>11/ps>1-1/p further implies, as we have noticed above, that $\omega\in A_{1}$, from which it follows that the elements of $W^{1,p}(\omega,\mathbb{C})$ are continuous functions (Proposition 3.14). Finally, the inequality
$1-\frac{1}{p}<\frac{p+K_{*}-h(\Gamma)}{K_{*}p}$
is equivalent to $p<\frac{2K_{*}-h(\Gamma)}{K_{*}-1}$.
Proof of Theorem 3.16.
We first show that $\mathcal{B}_{p,p}^{s}(\Gamma)\subset W^{1,p}(\omega,\mathbb{C})|\Gamma$. Let $f\in\mathcal{B}_{p,p}^{s}(\Gamma)$ and let $u_{i}$ the harmonic extension of $f$ to $\Omega_{i}$ . By hypothesis $u_{i}\in W^{1,p}(\omega,\Omega_{i})$ and a theorem of [8] asserts that since $\Gamma$ is a quasicircle and $\omega\in A_{p}$, $u_{i}$ can be extended to a function of $W^{1,p}(\omega,\mathbb{C})$.
In order to prove that $W^{1,p}(\omega,\mathbb{C})|_{\Gamma}\subset\mathcal{B}_{p,p}^{s}(\Gamma)$, let us first consider $f\in C_{c}^{\infty}(\mathbb{C})$. Recall that
$F_{i}(z)=-\frac{1}{\pi}\iint_{\Omega_{e}}\frac{\bar{\partial}f(\zeta)}{\zeta-z}d\xi d\eta,\quad F_{e}(z)=\frac{1}{\pi}\iint_{\Omega_{i}}\frac{\bar{\partial}f(\zeta)}{\zeta-z}d\xi d\eta$
are holomorphic respectively in $\Omega_{i}$ and $\Omega_{e}$, and the fact that $\omega\in A_{p}$ implies that $F_{i,e}\in\mathcal{B}_{p,p}^{s}(\Omega_{i,e}\to\Gamma)$, which is equal to $\mathcal{B}_{p,p}^{s}(\Gamma)$ under the assumption of Theorem 3.12. Using $f|_{\Gamma}=F_{i}-F_{e}$ we may see $f|_{\Gamma}\in\mathcal{B}_{p,p}^{s}(\Gamma)$.
Let now $f\in W^{1,p}(\omega,\mathbb{C})$. To prove $f|_{\Gamma}\in\mathcal{B}_{p,p}^{s}(\Gamma)$ we use a standard approximation argument by Lemma 3.13. There is a sequence $(f_{n})\in C_{c}^{\infty}(\mathbb{C})\cap W^{1,p}(\omega,\mathbb{C})$ converging to $f$ for the $W^{1,p}(\omega,\mathbb{C})$ norm. By the preceding discussion we can write $f_{n}|_{\Gamma}=(F_{n})_{i}-(F_{n})_{e}$, with $(F_{n})_{i,e}\in\mathcal{B}_{p,p}^{s}(\Omega_{i,e})$ and moreover $(F_{n})_{i}$ and $(F_{n})_{e}$ are converging sequences in $\mathcal{B}_{p,p}^{s}(\Omega_{i,e})$. Since $\mathcal{B}_{p,p}^{s}(\Omega_{i}\to\Gamma)=\mathcal{B}_{p,p}^{s}(\Omega_{e}\to\Gamma)$, $f|_{\Gamma}\in\mathcal{B}_{p,p}^{s}(\Gamma)$ follows. ∎
Corollary 3.17.
Under the assumption of Theorem 3.16, the space $C_{c}^{\infty}(\mathbb{C})|_{\Gamma}$ is dense in $\mathcal{B}_{p,p}^{s}(\Gamma)$, and thus the conclusions in Theorem 3.5 hold for all $f\in\mathcal{B}_{p,p}^{s}(\Gamma)$.
We end this section by a question. Is it true that within the hypothesis of Theorem 3.16 if $f\in\mathcal{B}_{p,p}^{s}(\Gamma)$ then its extension to the whole plane by the harmonic extension on both sides belongs to $W^{1,p}(\omega,\mathbb{C})$? We do not know the answer to this question in general but only for the cases with improved regularity (bigger $s$):
Proposition 3.18.
Let $\Gamma$ be a quasicircle and let $f$ be a function which is $\alpha$-Hölder continuous in $\mathbb{C}$ with $\alpha>h(\Gamma)-1$ and of class $C^{1}$ in $\mathbb{C}\backslash\Gamma$. Then if the function $F(z)=\nabla f(z),\,z\in\mathbb{C}\backslash\Gamma$ (thus defined almost everywhere) is in $L^{1}_{loc}(\mathbb{C})$, we have that, in the sense of distributions, $\nabla f=F$.
Proof.
Let $\varphi$ be a $C_{c}^{\infty}$-function in $\mathbb{R}$ with support included in [1,1][-1,1]. As usual we define the approximation of identity $\varphi_{\varepsilon}(x)=\frac{1}{\varepsilon}\varphi(\frac{x}{\varepsilon})$. From this we can define the approximation of identity in two variables
$\psi_{\varepsilon}(z)=\psi_{\varepsilon}(x+iy)=\varphi_{\varepsilon}(x)\varphi_{\varepsilon}(y).$
We may assume without loss of generality that $f$ has compact support, the function $f*\psi_{\varepsilon}$ is then a $C^{\infty}$-function with compact support and it converges uniformly in $\mathbb{C}$ towards $f$. From this we deduce that
$\iint_{\mathbb{C}}\nabla(f*\psi_{\varepsilon})(z)h(z)dxdy=-\iint_{\mathbb{c}}f*\psi_{\varepsilon}(z)\nabla h(z)dxdy$
for any $h\in C_{c}^{\infty}(\mathbb{C})$, and the right-hand side converges to $\iint_{\mathbb{C}}f(z)\nabla h(z)dxdy$ as $\varepsilon$ converges to 0.
We split the left-hand side into two parts:
Let $\Gamma_{\varepsilon}$ be the set of points at distance less than $2\varepsilon$ from $\Gamma$. Then
$\iint_{\mathbb{C}\backslash\Gamma_{\varepsilon}}\nabla(f*\psi_{\varepsilon})(z)h(z)dxdy=\iint_{\mathbb{C}\backslash\Gamma_{\varepsilon}}(\nabla F)*\psi_{\varepsilon}(z)h(z)dxdy$
which converges as $\varepsilon$ goes to 0 towards $\iint_{\mathbb{C}}\nabla F(z)h(z)dxdy$. In order to finish the proof of the proposition, it suffices now to prove that
$\iint_{\Gamma_{\varepsilon}}\frac{\partial(f*\psi_{\varepsilon})(z)}{\partial x}dxdy$
converges to 0, the same fact for $y$ being the same. But
$\frac{\partial(f*\psi_{\varepsilon})}{\partial x}(z)=f*\frac{\partial(\psi_{\varepsilon})}{\partial x}(z)$
which is also equal to
$\int_{[-\varepsilon,\varepsilon]}\varphi_{\varepsilon}(t_{2})\left(\int_{[-\varepsilon,\varepsilon]}(f(x-t_{1},y-t_{2})-f(x,y-t_{2}))\varphi_{\varepsilon}^{\prime}(t_{1})dt_{1}\right)dt_{2},$
because $\int_{[-\varepsilon,\varepsilon]}\varphi_{\varepsilon}^{\prime}(t_{1})dt_{1}=0$. Now $|f(x-t_{1},y-t_{2})-f(x,y-t_{2})|\leq C\varepsilon^{\alpha}$ and $|\varphi_{\varepsilon}^{\prime}(t_{1})|\leq C\varepsilon^{-2}$. Since we integrate on an interval of length $2\varepsilon$, the absolute value of the whole integral is bounded by $C\varepsilon^{\alpha-1}$.
Finally,
$\iint_{\Gamma_{\varepsilon}}\left|f*\frac{\partial(\psi_{\varepsilon})}{\partial x}(z)\right|dxdy\leq|\Gamma_{\varepsilon}|\varepsilon^{\alpha-1}.$
But we know that $|K\Gamma_{\varepsilon}|\varepsilon^{h(\Gamma)-2}\leq C$; we conclude that
$|\Gamma_{\varepsilon}|\varepsilon^{\alpha-1}\leq C\varepsilon^{\alpha+1-h(\Gamma)},$
and the result follows. ∎
Based on Proposition 3.18, a sufficient condition for a positive answer to the question is the hypothesis of Theorem 3.16 and $s>\frac{1}{p}+(h(\Gamma)-1)$. To be precise, the function $f\in\mathcal{B}_{p,p}^{s}(\Gamma)$, defined to be $u_{i,e}$ in $\Omega_{i,e}$, satisfies the assumption of Proposition 3.18, and thus belongs to $L^{1}_{\text{loc}}(\mathbb{C})$ in the sense of distribution, so that $f\in W^{1,p}(\omega,\mathbb{C})$. But for the condition $s>\frac{1}{p}+(h(\Gamma)-1)$ to be compatible with (3.8) some elementary arithmetic left to the reader leads to the following sufficient conditions:
$p>2,\quad\frac{1}{p}+h(\Gamma)-1<s<\frac{p+K_{*}-h(\Gamma)}{pK_{*}}.$
Notice that in the case of a chord-arc curve ($h(\Gamma)=1$) there are admissible values of $s$ for all $p>2$. In general, for having amissible values of $s$ one needs
$h(\Gamma)<\frac{K_{*}+1}{K_{*}+\frac{1}{p}}$
and for this inequality to hold for any $p>2$ and $K_{*}\leq 2$, $h(\Gamma)<\frac{6}{5}$ is necessary and sufficient.
3.3.3 Almost-Dirichlet Principle
We start by recalling the classical Dirichlet principle. If $\Omega$ is a Jordan domain with $\partial\Omega=\Gamma$ and $f$ is a continuous function on $\Gamma$ having a continuous extension $F:\overline{\Omega}\to\mathbb{C}$ which is in $C^{1}(\Omega)$. then, if $u$ stands for the harmonic extension of $f$ to $\Omega$, we have
$\iint_{\Omega}|\nabla u(z)|^{2}dxdy\leq\iint_{\Omega}|\nabla F(z)|^{2}dxdy.$
Now let $\Gamma$ be a quasicircle and $p>1,\,s\in(0,1)$: we will say that $\mathcal{B}_{p,p}^{s}(\Omega)$ satisfies an almost-Dirichlet principle if there exists $C>0$ such that
$\iint_{\Omega}|\nabla u(z)|^{p}d(z,\Gamma)^{(1-s)p-1}dxdy\leq C\iint_{\Omega}|\nabla F(z)|^{p}d(z,\Gamma)^{(1-s)p-1}dxdy.$
Maz’ya [29] has proven the almost-Dirichlet principle for the unit disk for all $p>1,s\in(0,1)$ and his proof has been simplified by [30]. Theorem 3.16 shows that if $\Gamma$ satisfies the hypothesis of this theorem, then the almost-Dirichlet principle holds.
3.3.4 The Case of pp–Dirichlet Energy
In this part we highlight the case of $p$-Dirichlet energy corresponding to $\omega\equiv 1$, i.e., s=11/ps=1-1/p.
Theorem 3.19.
If $\Gamma$ is a quasicircle then for all $p\in\left(2,\frac{2K_{*}-h(\Gamma)}{K_{*}-1}\right)$, there exists a constant $C=C(p,\Gamma,s)$ such that for $v\in C_{c}^{\infty}(\mathbb{C})$ ,
$\iint_{\Omega}|\nabla u(z)|^{p}dxdy\leq C\iint_{\Omega}|\nabla v(z)|^{p}dxdy,$ (3.10)
where $u$ stands for the harmonic extension of $v|_{\Gamma}$ in $\Omega$.
The proof is just by putting s=11/ps=1-1/p in Theorem 3.16 and taking care of the constraints: details are left to the reader. In particular, if $\Gamma$ is chord-arc then (3.10) holds for $p\in[2,3)$, and we have the following extension if $K<2$:
Corollary 3.20.
If $\Gamma$ is chord-arc with $K<2$ then the almost-Dirichlet principle holds for $2\leq p<\frac{2K-1}{K-1}.$
Notice that the corollary implies that for any $p>2$ the almost-Dirichlet principle is valid if $K$ is close enough to $1$.
4 Fractional Sobolev Spaces on Chord-arc Domains
4.1 A Review of Classical Facts about Fractional Sobolev Spaces on Plane Domains.
4.1.1 Fractional Sobolev Spaces on Plane Domains
We have already dealt with the Sobolev space $W^{1,p}(\mathbb{C})$, $1<p<\infty$ consisting of tempered distributions $f\in\mathcal{S}^{\prime}(\mathbb{C})$ such that $f$ is locally integrable and $\nabla f$, taken in the sense of distributions, is in $L^{p}(\mathbb{C})$. The semi-norm
$\|f\|_{W^{1,p}(\mathbb{C})}=\left(\iint_{\mathbb{C}}|\nabla f(z)|^{p}dxdy\right)^{\frac{1}{p}}$
equips $W^{1,p}(\mathbb{C})$ with a structure of Banach space modulo constants called homogeneous Sobolev space in the literature. The inhomogeneous Sobolev space is given by $\mathbb{W}^{1,p}(\mathbb{C})=W^{1,p}(\mathbb{C})\cap L^{p}(\mathbb{C})$. It is a Banach space equipped with the norm
$\|f\|_{\mathbb{W}^{1,p}(\mathbb{C})}=\left(\|f\|_{W^{1,p}(\mathbb{C})}^{p}+\|f\|_{L^{p}(\mathbb{C})}^{p}\right)^{\frac{1}{p}}.$
Let $\Omega$ be an open set in $\mathbb{C}$. $W^{1,p}(\Omega)$ and $\mathbb{W}^{1,p}(\Omega)$ should be defined similarly. We point out that if the open set $\Omega$ supports classical Poincaré inequality
$\|f\|_{L^{p}(\Omega)}\leq C\|f\|_{W^{1,p}(\Omega)}$
for every $f\in C_{c}^{\infty}(\Omega)$ where the constant $C$ depends only on $\Omega$ and $p$ then both spaces are known to coincide. For instance, the bounded chord-arc domain is a such domain. The bounded quasidisk also supports Poincaré inequality since it is an extension domain for Sobolev spaces $W^{1,p}(\mathbb{C})$ (see [22, 8]) so that one can apply Poincaré inequality in a bigger sufficiently regular domain which, combined with the extension theorem, leads to the conclusion.
For $0<s<1$ we define the fractional Besov-Sobolev space $B_{p,p}^{s}(\mathbb{C})$ as the space of $L^{p}(\mathbb{C})$-functions $f:\mathbb{C}\to\mathbb{C}$ such that
$\int_{\mathbb{C}}\int_{\mathbb{C}}\frac{|f(z)-f(\zeta)|^{p}}{|z-\zeta|^{2+sp}}dzd\zeta<\infty.$
We remark that suppose $s\geq 1$ and $\Omega$ is a connected open set in $\mathbb{C}$ then any measurable function $f:\Omega\to\mathbb{C}$ such that the integral of the above integrand $|f(z)-f(\zeta)|^{p}/|z-\zeta|^{2+sp}$ over $\Omega\times\Omega$ converges is actually constant. These spaces may be viewed as intermediate spaces between $W^{0,p}(\mathbb{C})$, i.e., $L^{p}(\mathbb{C})$ and $\mathbb{W}^{1,p}(\mathbb{C})$. More precisely, they may be obtained as real-interpolate spaces between these two spaces: $L^{p}(\mathbb{C}),\mathbb{W}^{1,p}(\mathbb{C})]_{s,p}=B^{s}_{p,p}(\mathbb{C})$, see [40]. We call them Besov-Sobolev spaces because if one uses complex interpolation instead we obtain the spaces $H_{p}^{s}(\mathbb{C})$ consisting of Bessel potential of $L^{p}$-functions: $L^{p}(\mathbb{C}),\mathbb{W}^{1,p}(\mathbb{C})]_{s}=H^{s}_{p}(\mathbb{C})$. and $B_{p,p}^{s}(\mathbb{C})=H_{p}^{s}(\mathbb{C})$ only if $p=2$ (if $p\neq 2$ these two spaces are not even isomorphic).
For further use we extend the definition of Besov-Sobolev spaces to $s\in(1,2)$ by saying that $f\in B_{p,p}^{s}(\mathbb{C})$ with $s\in(1,2)$ if $f\in B_{p,p}^{s-1}(\mathbb{C})$ and $\nabla f\in B_{p,p}^{s-1}(\mathbb{C})$. Similarly, it can be further extended to other non-integer orders $s>0$. It is known that the space $C_{c}^{\infty}(\mathbb{C})$ is dense in $B_{p,p}^{s}(\mathbb{C})$ for non-integer orders $s>0$ and $1<p<\infty$.
Let us now consider a chord-arc sub-domain of the plane. We recall that a chord-arc domain is the image of a disk by a bi-Lipschitz homeomorphism of the plane. The name comes from the fact that a Jordan domain $\Omega$ is chord-arc if and only if $\Gamma=\partial\Omega$ is a chord-arc curve.
We define $B_{p,p}^{s}(\Omega)$ simply as the space of restrictions to $\Omega$ of functions in $B_{p,p}^{s}(\mathbb{C})$. Such a function is in particular such that
$\int_{\Omega}\int_{\Omega}\frac{|f(z)-f(\zeta)|^{p}}{|z-\zeta|^{2+sp}}dzd\zeta<\infty$
and functions satisfying this last property can be extended to a function in $B_{p,p}^{s}(\mathbb{C})$.
4.1.2 Fractional Sobolev Spaces on Chord-arc Curves
If $\Omega$ is a (bounded or unbounded) chord-arc domain with boundary $\Gamma$ we may define for $0<s<1$, $p>1$,
$B_{p,p}^{s}(\Gamma)=\{f\in L^{p}(\Gamma):\,\iint_{\Gamma\times\Gamma}\frac{|f(z)-f(\zeta)|^{p}}{|z-\zeta|^{1+ps}}|dz||d\zeta|<\infty\}.$
We define the trace operator $\gamma$ on $C_{c}^{\infty}(\mathbb{C})$ by $\gamma(u)=u|_{\Gamma}$.
Theorem 4.1 (p.182 in [24]).
Let $s\in(1/p,1+1/p)$. The trace operator $\gamma$ can be extended to a bounded linear operator from $B_{p,p}^{s}(\Omega)$ onto $B_{p,p}^{s-1/p}(\Gamma)$, the order having $1/p$-loss, whose kernel is the closure of $C_{c}^{\infty}(\Omega)$ in $B_{p,p}^{s}(\Omega)$.
4.2 Plemelj-Calderón Problem on Chord-arc Domains
We now come back to Plemelj-Calderón problem. By a deep theorem of David [11] using the Lipschitz result [10], the Cauchy integral operator $T$ is bounded on $L^{p}(\Gamma)$ (i.e., $B_{p,p}^{0}(\Gamma)$) for every $p\in(1+\infty)$ (see section 1.3). Using it we will show the boundedness of the operator $T$ on $B_{p,p}^{1}(\Gamma)$.
Theorem 4.2.
The operator $T$ is bounded on the Sobolev space $B_{p,p}^{1}(\Gamma)$, and moreover, the operator norm on $B_{p,p}^{1}(\Gamma)$ equals to that on $L^{p}(\Gamma)$.
Before we proceed to the proof of this theorem, one must properly define what we mean by $B^{1}_{p,p}(\Gamma)$ when $\Gamma$ is a chord-arc curve (or $K$-chord-arc curve for specifying its chord-arc constant). Without loss of generality we suppose $\text{length}(\Gamma)=2\pi$. Let $\varphi:\,\mathbb{D}\to\Omega$ be the Riemann map from the unit disk $\mathbb{D}$ onto the interior domain $\Omega$ of $\Gamma$, which can be extended to a homeomorphism of the closures such that $\varphi$ restricted on $\mathbb{T}$ is a quasisymmetry (see [35]). Let $t\mapsto z(t)$ denote the arc-length parametrization of $\Gamma$ and $\lambda(e^{it})=z(t)$ so that $\varphi=\lambda\circ h$ where $h$ is an absolutely continuous homeomorphism of $\mathbb{T}$ with $|h^{\prime}|=|\varphi^{\prime}|$. It is easy to see $\lambda$ is a bi-Lipschitz homeomorphism from $\mathbb{T}$ onto $\Gamma$. Indeed, for any $t_{1},t_{2}\in[0,2\pi)$,
$K^{-1}|e^{it_{1}}-e^{it_{2}}|\leq|\lambda(e^{it_{1}})-\lambda(e^{it_{2}})|\leq\pi/2|e^{it_{1}}-e^{it_{2}}|.$
We say $f\in L^{p}(\Gamma)$ belongs to $B_{p,p}^{1}(\Gamma)$ if $f\circ\lambda$ is an anti-derivative of a function $h\in L^{p}(\mathbb{T})$. We define the function $f^{\prime}$ by $f^{\prime}\circ\lambda\lambda^{\prime}=h$. Then $f^{\prime}\in L^{p}(\Gamma)$. The space $B_{p,p}^{1}(\Gamma)$ is endowed with the natural norm $\|f\|_{B_{p,p}^{1}(\Gamma)}=\|f^{\prime}\|_{L^{p}(\Gamma)}$.
To prepare for the proof we introduce the following definition. Set $\Gamma_{r}=\tau(|z|=r)$ to be "circular curves", where $\tau$ is a Riemann map that takes $\mathbb{D}$ onto $\Omega$. The Hardy space $E^{p}(\Omega)$ of index $p$ on the chord-arc domain $\Omega$ consists of holomorphic functions $F$ on $\Omega$ with a finite norm
$\|F\|_{p}=\Bigg(\frac{1}{2\pi}\sup_{r}\int_{\Gamma_{r}}|F(w)|^{p}|dw|\Bigg)^{1/p}.$
Notice that $\tau$ is a homeomorphism of the closures $\overline{\mathbb{D}}$ onto $\Omega\cup\Gamma$, and absolutely continuous on $\mathbb{T}$ since $\Gamma$ is rectifiable.
Proof of Theorem 4.2.
Let $f\in B_{p,p}^{1}(\Gamma)$. By David’s theorem we may write
$f^{\prime}=G_{i}+G_{e}$ (4.1)
on $\Gamma$. Here, $G_{i}$ and $G_{e}$ belong to the Hardy spaces of index $p$ on $\Omega_{i}$ and $\Omega_{e}$, respectively, and then $G_{i}$ and $G_{e}$ have non-tangential boundary limits almost everywhere on $\Gamma$ with respect to the arc-length measure, denoted still by $G_{i}$ and $G_{e}$, so that $\|G_{i}\|_{L^{p}(\Gamma)}$ and $\|G_{e}\|_{L^{p}(\Gamma)}$ are both controlled from above by $\|f\|_{B_{p,p}^{1}(\Gamma)}$.
Recall that $G_{i}\in E^{p}(\Omega_{i})\subset E^{1}(\Omega_{i})$, so that $G_{i}\circ\tau\tau^{\prime}\in E^{1}(\mathbb{D}_{i})$. Let now $\Phi_{i}$ be an anti-derivative of $G_{i}$ on $\Omega_{i}$. Then we see $(\Phi_{i}\circ\tau)^{\prime}=G_{i}\circ\tau\tau^{\prime}\in E^{1}(\mathbb{D}_{i})$. By a result of Hardy-Littlewood (see e.g. p.89 in [16]), we conclude that $\Phi_{i}\circ\tau$ is continuous on $\overline{\mathbb{D}}_{i}$ and absolutely continuous on $\mathbb{T}$ such that $(\Phi_{i}\circ\tau)^{\prime}(\zeta)=\lim_{r\to 1}(\Phi_{i}\circ\tau)^{\prime}(r\zeta)$ almost everywhere on $\mathbb{T}$, and thus $(\Phi_{i}\circ\tau)^{\prime}(\zeta)=G_{i}\circ\tau(\zeta)\tau^{\prime}(\zeta)$ almost everywhere on $\mathbb{T}$. Using it, let us define $\Phi_{i}^{\prime}=G_{i}$ on $\Gamma$, and we can similarly define $\Phi_{e}^{\prime}=G_{e}$ on $\Gamma$. Combined with (4.1) that leads to, by adjusting the constants,
$f=\Phi_{i}+\Phi_{e}$
on $\Gamma$ with the norm of $\Phi_{i,e}$ in $B^{1}_{p,p}(\Gamma)$ controlled from above by $\|f\|_{B_{p,p}^{1}(\Gamma)}$. By Plemelj formula we conclude that the operator $T$ is bounded on $B_{p,p}^{1}(\Gamma)$ with respect to the norm $\|\cdot\|_{B_{p,p}^{1}(\Gamma)}$, and moreover, we have $\|T\|_{B_{p,p}^{1}(\Gamma)\to B_{p,p}^{1}(\Gamma)}=\|T\|_{L^{p}(\Gamma)\to L^{p}(\Gamma)}$. ∎
Now we get ready to show the main result of this section: the operator $T$ is bounded on $B_{p,p}^{s}(\Gamma)$ for $0<s<1$. Because of the isomorphism between $B_{p,p}^{s}(\Gamma)$ and $B_{p,p}^{s}(\mathbb{T})$, this boils down to proving that the operator with kernel
$\text{p.v.}\frac{\lambda^{\prime}(\zeta)}{\lambda(\zeta)-\lambda(\xi)}$
is bounded on $B_{p,p}^{s}(\mathbb{T})$. Set $g=f\circ\lambda$. This operator is more precisely defined by
$\widetilde{T}g(\xi)=\frac{1}{2\pi i}\text{p.v.}\int_{\mathbb{T}}\frac{g(\zeta)\lambda^{\prime}(\zeta)}{\lambda(\zeta)-\lambda(\xi)}d\zeta.$
Let us introduce Calderón’s interpolation theorem.
Theorem 4.3 (p.38 in [40], Ch.6 in [4] and also [7]).
Let $s_{0},s_{1}\in[0,1]$. If $s_{0}\neq s_{1}$ then we have, in terms of complex interpolation theory of Banach spaces,
$B_{p,p}^{s}(\mathbb{T})=[B_{p,p}^{s_{0}}(\mathbb{T}),B_{p,p}^{s_{1}}(\mathbb{T})]_{\theta}$
where $s=(1-\theta)s_{0}+\theta s_{1}$ and the exponent $\theta\in(0,1)$.
The following functorial property of complex interpolation is a basic assertion in interpolation theory. Let $\{A_{0},A_{1}\}$ and $\{B_{0},B_{1}\}$ be two interpolation couples of Banach spaces and let $L$ be a linear operator mapping from $A_{0}+A_{1}$ into $B_{0}+B_{1}$ such that its restriction to $A_{j}$ is a linear and bounded operator from $A_{j}$ into $B_{j}$ with norm $M_{j}$, where $j=0,1$. Then the restriction of $L$ to $A_{0},A_{1}]_{\theta}$, $0<\theta<1$ is a linear and bounded operator from $A_{0},A_{1}]_{\theta}$ into $B_{0},B_{1}]_{\theta}$ with norm $M_{\theta}$. Further, it is known that the complex interpolation method is an exact interpolation functor of exponent $\theta$ (see e.g. p.88 in [4]) meaning that
$M_{\theta}\leq M_{0}^{1-\theta}M_{1}^{\theta}.$ (4.2)
It now suffices to invoke Calderón’s interpolation theorem to conclude the following.
Theorem 4.4.
For $0<s<1$ and $1<p<\infty$, the operator $T$ is bounded on $B_{p,p}^{s}(\Gamma)$.
Proof.
By taking $s_{0}=0$ and $s_{1}=1$ in Theorem 4.3, we have that
$B_{p,p}^{s}(\mathbb{T})=[L^{p}(\mathbb{T}),B_{p,p}^{1}(\mathbb{T})]_{s}.$ (4.3)
By David’s theorem and Theorem 4.2, we conclude that the operator $\widetilde{T}$ is bounded on $B_{p,p}^{s}(\mathbb{T})$ with the operator norm $\|\widetilde{T}\|_{B_{p,p}^{s}(\mathbb{T})\to B_{p,p}^{s}(\mathbb{T})}=\|\widetilde{T}\|_{L^{p}(\mathbb{T})\to L^{p}(\mathbb{T})}$, and equivalently, the operator $T$ is bounded on $B_{p,p}^{s}(\Gamma)$ with the operator norm $\|T\|_{B_{p,p}^{s}(\Gamma)\to H^{s}(\Gamma)}=\|T\|_{L^{p}(\Gamma)\to L^{p}(\Gamma)}$. ∎
It follows from Plemelj formula that every function $f\in B_{p,p}^{s}(\Gamma)$ may be written uniquely as $f=\Phi_{i}+\Phi_{e}$ with $\|\Phi_{i,e}\|_{B_{p,p}^{s}(\Gamma)}\leq C\|f\|_{B_{p,p}^{s}(\Gamma)}$ for some constant $C$, $\Phi_{i,e}$ being boundary values of holomorphic functions in $\Omega_{i}$ and $\Omega_{e}$, respectively. The uniqueness of decomposition for $0<s\leq 1$ just follows from the case of $s=0$.
4.2.1 On a Theorem of Murai
Clearly, Theorem 4.4 is also valid for chord-arc curves $\Gamma$ passing through $\infty$, and in particular for Lipschitz curves $\Gamma$ (i.e., there is a Lipschitz function $A:\mathbb{R}\to\mathbb{R}$ whose graph is $\Gamma$). A theorem by Murai [31] states that for a Lipschitz curve $\Gamma$ with Lipschitz norm $M$, $\|T\|_{L^{p}(\Gamma)\to L^{p}(\Gamma)}$ is no more than $C(1+M)^{3/2}$ (see also [12]), where $C$ is a universal constant. If we plug this information in the preceding theorem we obtain the same bound for $\|T\|_{B_{p,p}^{1-1/p}(\Gamma)\to B_{p,p}^{1-1/p}(\Gamma)}$. But the norm $\|T\|_{B_{p,p}^{1-1/p}(\Gamma)\to B_{p,p}^{1-1/p}(\Gamma)}$ depends only on the $L^{p}(\mathbb{C})$-boundedness of the Beurling transform (see section 2), so that it is actually independent of $M$. In order to get better estimates of $\|T\|_{B_{p,p}^{s}(\Gamma)\to B_{p,p}^{s}(\Gamma)}$ we thus use Calderón’s interpolation theorem between $L^{p}(\Gamma)$ and $B_{p,p}^{1-1/p}(\Gamma)$ for 0<s<11/p0<s<1-1/p and between $B_{p,p}^{1-1/p}(\Gamma)$ and $B_{p,p}^{1}(\Gamma)$ for 11/p<s<11-1/p<s<1. We use more precisely that
$B_{p,p}^{s}(\Gamma)$ $=[L^{p}(\Gamma),B_{p,p}^{1-1/p}(\Gamma)]_{\frac{ps}{p-1}},\;\;\;\;\;\,0<s<1-1/p;$
$B_{p,p}^{s}(\Gamma)$ $=[B_{p,p}^{1-1/p}(\Gamma),B_{p,p}^{1}(\Gamma)]_{ps+1-p},\;\;1-1/p<s<1.$
Using (4.2), we get to
Theorem 4.5.
If $\Gamma$ is a Lipschitz curve with Lipschitz constant $M$ then we have, for $0<s<1$,
$\|T\|_{B_{p,p}^{s}(\Gamma)\to B_{p,p}^{s}(\Gamma)}\leq C(1+M)^{\frac{3}{2}|1-\frac{ps}{p-1}|}$
where $C$ depends only on $p$ .
5 Douglas versus Littlewood-Paley: the Chord-arc Case
5.1 The Conjugate Operator
Let $\Gamma$ be a chord-arc curve of length $2\pi$, and $1<p<\infty$, $0<s<1$. We have seen two ways of generalizing the fractional Sobolev spaces from $\mathbb{R}$ (or from $\mathbb{T}$) to $\Gamma$. The first one, which can be called the Douglas way is
$B_{p,p}^{s}(\Gamma)=\bigg\{f\in L^{p}(\Gamma):\iint_{\Gamma\times\Gamma}\frac{|f(z)-f(\zeta)|^{p}}{|z-\zeta|^{1+ps}}|dz||d\zeta|<\infty\bigg\}.$
The second one, via Littlewood-Paley theory, is
$\mathcal{B}_{p,p}^{s}(\Gamma)=\bigg\{f\in L^{p}(\Gamma):\iint_{\Omega_{i,e}}|\nabla u_{i,e}(z)|^{p}d(z,\Gamma)^{(1-s)p-1}dxdy<\infty\bigg\}.$
Here, $u_{i,e}$ stands for the harmonic extension of $f$ to $\Omega_{i,e}$. We have seen that Plemelj-Calderón problem is solvable in some sense on both spaces $B_{p,p}^{s}(\Gamma)$ and $\mathcal{B}_{p,p}^{s}(\Gamma)$ so that the natural question arises of whether these spaces coincide.
Assume that $\Omega$ is bounded and $\mathbb{D}$ is the unit disk. Let $z_{0}$ be a point in $\Omega$. If $u$ is harmonic in $\Omega$ it is well known that there exists a unique harmonic function $\tilde{u}$ in $\Omega$ such that $\tilde{u}(z_{0})=0$ and $u+i\tilde{u}$ is holomorphic in $\Omega$. In the case of $\Omega=\mathbb{D}$, $\tilde{u}$ is the harmonic (Poisson) extension of $H(f)$ where $H$ is the Hilbert transform:
$H(f)(e^{i\theta})=\lim_{\epsilon\to 0}\frac{1}{2\pi}\int_{|\theta-\varphi|>\epsilon}\cot\Big(\frac{\theta-\varphi}{2}\Big)f(e^{i\varphi})d\varphi.$
By Cauchy-Riemann equations, $|\nabla\tilde{u}|=|\nabla u|$, so that $\mathcal{B}_{p,p}^{s}(\Omega)$ is conjugate-invariant. We denote the subspace of $\mathcal{B}_{p,p}^{s}(\Omega)$ consisting of holomorphic functions by $\mathcal{HB}_{p,p}^{s}(\Omega)$, i.e., the set of $u+i\tilde{u}$ with $u,\tilde{u}\in\mathcal{B}_{p,p}^{s}(\Omega)$.
A necessary condition for $B_{p,p}^{s}(\Gamma)=\mathcal{B}_{p,p}^{s}(\Gamma)$ to hold is thus that $B_{p,p}^{s}(\Gamma)$ is stable by conjugation. Let $\varphi:\,\mathbb{D}\to\Omega$ be the Riemann map with $\varphi(0)=z_{0}$. Recall that $t\mapsto z(t)$ denotes the arc-length parametrization of $\Gamma$ and $\lambda(e^{it})=z(t)$ so that $\varphi=\lambda\circ h$ where $h$ is an absolutely continuous homeomorphism of $\mathbb{T}$ with $|h^{\prime}|=|\varphi^{\prime}|$. Recall that $\lambda$ is a bi-Lipschitz homeomorphism from $\mathbb{T}$ onto $\Gamma$. A simple computation shows that if $f\in B_{p,p}^{s}(\Gamma)$; that is $g=f\circ\lambda\in B_{p,p}^{s}(\mathbb{T})$, and $u$ is its harmonic extension to $\Omega$, then $\tilde{u}$ is the harmonic extension of $\tilde{f}=\tilde{g}\circ\lambda^{-1}$. Here,
$\tilde{g}=V_{h}^{-1}\circ H\circ V_{h}(g),$
where $V_{h}(g)=g\circ h$.
It is known (see [9], p.247 in [16]) that $V_{h}^{-1}HV_{h}$ is bounded on $L^{p}(\mathbb{T})$ if and only if $|h^{\prime}|=|\varphi^{\prime}|$ belongs to $A_{p}$ on $\mathbb{T}$. Based on this comment on the limiting case $s=0$, we now state a theorem about the general case $0\leq s\leq 1$.
Theorem 5.1.
For $p>1,\,0\leq s\leq 1$, suppose $\Gamma$ is such that
  • $|\varphi^{\prime}|\in A_{p}$ if $1<p\leq 2$;
  • $|\varphi^{\prime}|\in A_{p^{\prime}}$$(1/p^{\prime}+1/p=1)$ if $p>2$.
Then, $B_{p,p}^{s}(\Gamma)$ is stable by conjugation, and moreover, the conjugate operator $f\mapsto\tilde{f}$ on $B_{p,p}^{s}(\Gamma)$ is bounded.
Before going to the proof, let us comment on the hypothesis of Theorem 5.1. If $\omega$ is an $A_{\infty}$ weight let us define $\delta(\omega)=\inf\{q>1:\,\omega\in A_{q}\}$. It is known [9] that if $\omega\in A_{q}$ for some $q>1$ then $\delta(\omega)<q$ (see p.254 in [16]). Let us then define $\delta=\delta(|\varphi^{\prime}|)$. Noting that the hypothesis imply that $|\varphi^{\prime}|\in A_{2}$, we see that the hypothesis of the theorem are equivalent to
$\delta<q<\frac{\delta}{\delta-1}.$
On the other hand, it is known (see [21]) that $\Gamma$ being chord-arc implies $|\varphi^{\prime}|$ having $A_{\infty}$ but there are examples of chord-arc curves such that $|\varphi^{\prime}|\notin A_{2}$ (see [23]), we thus know the chord-arc condition is not sufficient for Theorem 5.1 to hold.
Proof.
We have just seen that $V_{h}^{-1}HV_{h}$ is bounded on $L^{p}(\mathbb{T})$ if and only if $|\varphi^{\prime}|\in A_{p}$ for any $1<p<\infty$. Let now $f\in B_{p,p}^{1}(\Gamma)$ so that $g=f\circ\lambda\in B_{p,p}^{1}(\mathbb{T})$, the set of $g\in L^{p}(\mathbb{T})$ such that $g^{\prime}\in L^{p}(\mathbb{T})$. We have $(g\circ h)^{\prime}=g^{\prime}\circ hh^{\prime}\in L^{p}(|h^{\prime}|^{1-p})$. Recall that a weight $\omega\in A_{p^{\prime}}$ if and only if $\omega^{-\frac{1}{p^{\prime}-1}}\in A_{p}$. By that, $|h^{\prime}|^{1-p}\in A_{p}$ is equivalent to $|h^{\prime}|=|\varphi^{\prime}|\in A_{p^{\prime}}$. Then, we have $H((g\circ h)^{\prime})\in L^{p}(|h^{\prime}|^{1-p})$ with a norm bounded by the $L^{p}(|h^{\prime}|^{1-p})$-norm of $(g\circ h)^{\prime}$ (see [9]); that implies $(\tilde{g})^{\prime}\in L^{p}(\mathbb{T})$ with $\tilde{g}=V_{h}^{-1}HV_{h}(g)$, and thus $\tilde{f}=\tilde{g}\circ\lambda^{-1}\in B_{p,p}^{1}(\Gamma)$ such that $\|\tilde{f}\|_{B_{p,p}^{1}(\Gamma)}$ bounded by $\|f\|_{B_{p,p}^{1}(\Gamma)}$. Since the weight $A_{p}$ increases with $p$ and from the following complex interpolation theorem (see Theorem 4.3):
$L^{p}(\mathbb{T}),B_{p,p}^{1}(\mathbb{T})]_{s}=B^{s}_{p,p}(\mathbb{T})$ (5.1)
the theorem now follows. ∎
Specially, a natural condition on the curve $\Gamma$ implying that $|\varphi^{\prime}|\in A_{2}$ is the radial-Lipschitz condition. Indeed, if $\Gamma$ is a radial-Lipschitz curve it can be shown that $|\varphi^{\prime}|$ satisfies the Helson-Szegö condition (see section 5.3 for a specific proof): $\log|\varphi^{\prime}|=u+Hv$ with $u\in L^{\infty},\;v\in L^{\infty}$ and $\|v\|_{\infty}<\pi/2$, which is equivalent to $|\varphi^{\prime}|\in A_{2}$ as follows from the Helson-Szegö theorem ([16]). We may now state:
Corollary 5.2.
For $0\leq s\leq 1$, if $\Gamma$ is such that $|\varphi^{\prime}|\in A_{2}$ then $B_{2,2}^{s}(\Gamma)$ is stable by conjugation, and moreover, the conjugate operator $f\mapsto\tilde{f}$ on $B_{2,2}^{s}(\Gamma)$ is bounded. The conclusion holds in particular for radial-Lipschitz curves.
For $0\leq s\leq 1$, $1<p<\infty$, we denote the subspace of $B_{p,p}^{s}(\Gamma)$ consisting of functions $f+i\tilde{f}$ such that $f,\;\tilde{f}\in B_{p,p}^{s}(\Gamma)$ by $\text{HB}_{p,p}^{s}(\Gamma)$.
5.2 The Operator $V_{s,p}$ on $\mathcal{B}^{s}_{p,p}(\Omega)$
Recall that $\mathcal{B}^{s}_{p,p}(\Omega)$, $0\leq s<1$ and $1<p<\infty$, is the set of harmonic functions $u$ on the domain $\Omega$ bounded by the Jordan curve $\Gamma$ such that its norm $\|u\|_{\mathcal{B}^{s}_{p,p}(\Omega)}<\infty$. Suppose $\Gamma$ is chord-arc which in particular is rectifiable. It is known that the non-tangential boundary values $f$ of $u$ exist almost everywhere on $\Gamma$ with respect to the arc-length measure and $u$ can be recovered from $f$ using the harmonic extension. The space of boundary functions of $\mathcal{B}^{s}_{p,p}(\Omega)$ is denoted by $\mathcal{B}_{p,p}^{s}(\Omega\!\to\!\Gamma)$ so that $\mathcal{B}_{p,p}^{s}(\Gamma)=\mathcal{B}_{p,p}^{s}(\Omega_{i}\!\to\!\Gamma)\cap\mathcal{B}_{p,p}^{s}(\Omega_{e}\!\to\!\Gamma)$. We may identify $\mathcal{B}_{p,p}^{s}(\Omega)$ with $\mathcal{B}_{p,p}^{s}(\Omega\!\to\!\Gamma)$ for convenience, and will switch between the two freely. Similarly, for $0\leq s\leq 1$ and $1<p<\infty$ we also identify $B_{p,p}^{s}(\Gamma)$ with the set $B_{p,p}^{s}(\Omega)$ of its harmonic extensions in $\Omega$.
Let $\Omega,\,\Omega^{\prime}$ be two Jordan domains containing 0 and $\varphi$ a holomorphic diffeomorphism from $\Omega$ onto $\Omega^{\prime}$ fixing 0. Let $f\in\mathcal{HB}_{p,p}^{s}(\Omega^{\prime})$. Using the change of variable $\zeta=\varphi(z)$ we have
$\iint_{\Omega^{\prime}}d(\zeta,\Gamma^{\prime})^{(1-s)p-1}|f^{\prime}(\zeta)|^{p}d\xi d\eta=\iint_{\Omega}d(\varphi(z),\Gamma^{\prime})^{(1-s)p-1}|(f\circ\varphi)^{\prime}(z)|^{p}|\varphi^{\prime}(z)|^{2-p}dxdy.$
By the Koebe distortion theorem (2.1) there exists a universal constant $C>1$ such that
$C^{-|(1-s)p-1|}\iint_{\Omega}d(z,\Gamma)^{(1-s)p-1}|V_{s,p}(f)^{\prime}(z)|^{p}dxdy$
$\leq$ $\iint_{\Omega^{\prime}}d(\zeta,\Gamma^{\prime})^{(1-s)p-1}|f^{\prime}(\zeta)|^{p}d\xi d\eta$ (5.2)
$\leq$ $C^{|(1-s)p-1|}\iint_{\Omega}d(z,\Gamma)^{(1-s)p-1}|V_{s,p}(f)^{\prime}(z)|^{p}dxdy,$
where $V_{s,p}$ is the operator defined by
$V_{s,p}(f)(z)=\int_{0}^{z}(f\circ\varphi)^{\prime}(u)\varphi^{\prime}(u)^{1/p-s}du.$
In other words, the operator $V_{s,p}$ is a bounded isomorphism between $\mathcal{HB}_{p,p}^{s}(\Omega)$ and $\mathcal{HB}_{p,p}^{s}(\Omega^{\prime})$ with the operator norm
$\|V_{s,p}\|,\,\|V_{s,p}^{-1}\|\leq C^{|(1-s)-1/p|}.$
Notice that for s=11/ps=1-1/p, this operator is nothing but the composition by $\varphi$ and that, in this case, $V_{1-1/p,p}$ is an isometry.
For convenience of later use, when $p=2$, the operator $V_{s,p}$ will be simplified to be $V_{s}$.
5.2.1 The Operator $V_{s}$ on $\mathcal{B}^{s}_{2,2}(\Omega)$
In order to understand better the operator $V_{s}$ let us rewrite it by using an integration by parts:
$V_{s}(f)(z)$ $=f\circ\varphi(z)\varphi^{\prime}(z)^{1/2-s}-(1/2-s)\int_{0}^{z}f\circ\varphi(u)\varphi^{\prime}(u)^{1/2-s}\frac{\varphi^{\prime\prime}(u)}{\varphi^{\prime}(u)}du$
$=T_{s}f(z)-(1/2-s)S(T_{s}(f))(z),$
where
$T_{s}f(z)=f\circ\varphi(z)\varphi^{\prime}(z)^{1/2-s}$
and
$Sg(z)=\int_{0}^{z}g(u)\frac{\varphi^{\prime\prime}(u)}{\varphi^{\prime}(u)}du.$
Let us now specialize to $s=0$ and $(\Omega,\Omega^{\prime})=(\mathbb{D},\Omega)$. Recall that the Hardy space $E^{2}(\Omega)$ is the space of holomorphic functions $f:\,\Omega\to\mathbb{C}$ such that $T_{0}(f)\in E^{2}(\mathbb{D})$, the classical Hardy space of the unit disk. We say that a function $g\in{\rm BMOA}(\mathbb{D})$ if $g\in E^{2}(\mathbb{D})$ and if in addition the boundary values of $g$ on $\mathbb{T}$ is of bounded mean oscillation (abbr. BMO) in the sense that
$\sup_{I\subset\mathbb{T}}\frac{1}{|I|}\int_{I}|g(z)-g_{I}||dz|<\infty,$
where the supremum is taken over all sub-arcs $I$ of $\mathbb{T}$ and $g_{I}$ denotes the integral mean of $g$ over $I$. It is known that $\log\varphi^{\prime}\in{\rm BMOA}(\mathbb{D})$ if $\varphi(\mathbb{D})$ is chord-arc, but not vice versa. By Fefferman-Stein, $\log\varphi^{\prime}\in{\rm BMOA}(\mathbb{D})$ if and only if
$d\mu=\bigg|\frac{\varphi^{\prime\prime}(z)}{\varphi^{\prime}(z)}\bigg|^{2}(1-|z|)dxdy$
is a Carleson measure in $\mathbb{D}$ (see e.g. Ch.VI in [16]).
Theorem 5.3.
Let $\Omega$ be a Jordan domain containing 0 and $\varphi$ the Riemann mapping from $\mathbb{D}$ onto $\Omega$ fixing 0. The following statements hold:
  1. (1)
    If $\log\varphi^{\prime}\in{\rm BMOA}(\mathbb{D})$ then $E^{2}(\Omega)\subset\mathcal{HB}_{2,2}^{0}(\Omega)$;
  2. (2)
    If $\Omega$ is a chord-arc domain then $\mathcal{HB}_{2,2}^{0}(\Omega)\subset E^{2}(\Omega)$.
Moreover, the inclusions are continuous with respect to the norms $\|\cdot\|_{E^{2}(\Omega)}$ and $\|\cdot\|_{\mathcal{B}_{2,2}^{0}(\Omega)}$.
Proof.
Suppose $\log\varphi^{\prime}\in{\rm BMOA}(\mathbb{D})$; that is, $\mu$ is a Carleson measure in $\mathbb{D}$. Let $f\in E^{2}(\Omega)$ so that $T_{0}f\in E^{2}(\mathbb{D})$. Then by Carleson (see e.g. Theorem 3.9 in [16])
$\iint_{\mathbb{D}}|T_{0}(f)|^{2}d\mu\leq C\|T_{0}f\|_{E^{2}(\mathbb{D})}^{2}$
where $C$ is a constant depending only on the Carleson norm of $\mu$. Since $E^{2}(\mathbb{D})=\mathcal{HB}_{2,2}^{0}(\mathbb{D})$ we have
$\iint_{\mathbb{D}}|(T_{0}f)^{\prime}|^{2}(1-|z|)dxdy<\infty.$
Finally,
$\iint_{\mathbb{D}}|(V_{0}f)^{\prime}(z)|^{2}(1-|z|)dxdy$ $=\iint_{\mathbb{D}}\Big|(T_{0}f)^{\prime}(z)-\frac{1}{2}T_{0}f(z)\cdot\frac{\varphi^{\prime\prime}(z)}{\varphi^{\prime}(z)}\Big|^{2}(1-|z|)dxdy$
$\leq 2\iint_{\mathbb{D}}|(T_{0}f)^{\prime}|^{2}(1-|z|)dxdy+\frac{1}{2}\iint_{\mathbb{D}}|T_{0}(f)|^{2}d\mu$
$<\infty.$
Combined with (5.2), this completes the proof of statement $(1)$.
For the proof of statement $(2)$ we will need the definition of the (Littlewood-Paley) $\mathsf{g}$-function of a holomorphic function $f$ in $\mathbb{D}$:
$\mathsf{g}(f)(e^{i\theta})=\left(\int_{0}^{1}(1-r)|f^{\prime}(re^{i\theta})|^{2}dr\right)^{1/2}.$
Suppose now that $f\in\mathcal{HB}_{2,2}^{0}(\Omega)$. Then $V_{0}(f)\in\mathcal{HB}_{2,2}^{0}(\mathbb{D})$, which in turn implies that
$\iint_{\mathbb{D}}(1-|u|)|(f\circ\varphi)^{\prime}(u)|^{2}|\varphi^{\prime}(u)|dudv<\infty,$
or, in other words, that $\mathsf{g}(f\circ\varphi)\in L^{2}(\mathbb{T},|\varphi^{\prime}|d\theta)$. Since $\Gamma$ is assumed to be chord-arc, we have that $|\varphi^{\prime}|$ has the weight $A_{\infty}$ on $\mathbb{T}$. We can then apply a theorem of Gundy-Wheeden [19] (see also [21]) which implies that $\mathsf{g}(f\circ\varphi)\in L^{2}(\mathbb{T},|\varphi^{\prime}|d\theta)$ if and only if the non-tangential maximal function $\mathsf{n}(f\circ\varphi)$ of $f\circ\varphi$ is in $L^{2}(\mathbb{T},|\varphi^{\prime}|d\theta)$. On the other hand, for any "circular curves" $\Gamma_{r}$ it holds that (see p.233 in [21])
$\int_{\Gamma_{r}}|f(\zeta)|^{2}d\sigma(\zeta)\leq C\|\mathsf{n}(f\circ\varphi)\|^{2}_{L^{2}(\mathbb{T},|\varphi^{\prime}|d\theta)},$
which implies $f\in E^{2}(\Omega)$. ∎
Notice that for this theorem we do not need $|\varphi^{\prime}|$ having $A_{2}$ on $\mathbb{T}$. This condition is nevertheless necessary for the following corollary to hold by Corollary 5.2 since if it is not attached then the (real) space of real parts of $E^{2}(\Omega\!\to\!\Gamma)$-functions is a proper subspace of the real space $L_{\mathbb{R}}^{2}(\Gamma,d\sigma)$
Corollary 5.4.
Let $\Omega$ be a chord-arc domain bounded by $\Gamma$. If $|\varphi^{\prime}|\in A_{2}$ then $B_{2,2}^{0}(\Gamma)=\mathcal{B}_{2,2}^{0}(\Omega\!\to\!\Gamma)$.
5.3 The Equality $\mathcal{B}_{2,2}^{s}(\Omega\!\to\!\Gamma)=B_{2,2}^{s}(\Gamma)$
The aim of this part is to prove the following theorem that is one of the main theorems of this paper.
Theorem 5.5.
Let $\Omega$ be a bounded chord-arc domain bounded by $\Gamma$ with $0\in\Omega$, and $\varphi$ its Riemann mapping fixing 0. If $\Gamma$ is such that
$\log|\varphi^{\prime}|=b_{1}+Hb_{2},\qquad b_{1},\;Hb_{1},\;b_{2}\in L^{\infty}(\mathbb{T})\;\text{and}\;\|b_{2}\|_{\infty}<\frac{\pi}{2},$ (5.3)
then, for $0\leq s\leq 1$, $\mathcal{B}_{2,2}^{s}(\Omega\!\to\!\Gamma)=B_{2,2}^{s}(\Gamma)$ with comparable norms. The conclusions hold in particular for radial-Lipschitz domains.
Before going further, let us proceed from understanding why the conditions (5.3) are assumed. The Helson-Szegö theorem (see Ch.IV in [16]) says that $\log|\varphi^{\prime}|=b_{1}+Hb_{2}$ with $b_{1},b_{2}\in L^{\infty}(\mathbb{T})$ and $\|b_{2}\|_{\infty}<\pi/2$ if and only if $|\varphi^{\prime}|\in A_{2}$ on $\mathbb{T}$, which is only used to guarantee that $B_{2,2}^{s}(\Gamma)$ is stable by conjugation for all $0\leq s\leq 1$ (see Corollary 5.2). For the proof of the corresponding analytic case: $\mathcal{HB}_{2,2}^{s}(\Omega\!\to\!\Gamma)=\text{HB}_{2,2}^{s}(\Gamma)$, this condition can thus be ignored. Note that $\text{Arg}\varphi^{\prime}=H\log|\varphi^{\prime}|=Hb_{1}-b_{2}$. The extra assumption that $Hb_{1}\in L^{\infty}(\mathbb{T})$ is used to derive that $\text{Arg}\varphi^{\prime}$ is bounded, geometrically which means that the boundary curve cannot spiral too much. For instance, for the double spiral curve, $\text{Arg}\varphi^{\prime}$ keeps increasing and goes to infinity.
Let $r(\theta):\mathbb{R}\to\mathbb{(}0,+\infty)$ be a $2\pi$-periodic continuous function. Then the curve $\Gamma=\{r(\theta)e^{i\theta}:\;\theta\in[0,2\pi)\}$ is called a starlike Jordan curve with respect to 0 and the domain $\Omega$ bounded by $\Gamma$ is called a starlike domain. Recall that if moreover $r(\theta)$ is a Lipschitz function, i.e., there exists a positive constant $M$ such that $\|r^{\prime}\|_{\infty}\leq M$ then the curve $\Gamma$ is called a radial-Lipschitz curve and $\Omega$ is called a radial-Lipschitz domain. Suppose $\varphi$ is a Riemann mapping from $\mathbb{D}$ onto $\Omega$ fixing 0. It is known that $\Omega$ is starlike with respect to 0 if and only if
$\varphi^{\prime}(0)\neq 0\;\text{and}\;\text{Arg}\frac{z\varphi^{\prime}(z)}{\varphi(z)}\in\left(-\frac{\pi}{2}\alpha,\frac{\pi}{2}\alpha\right),$ (5.4)
where $\alpha=1$. The condition $\varphi^{\prime}(0)\neq 0$ implies that $\log(\varphi(z)/z)$ is continuous on $\overline{\mathbb{D}}$ so that $\log|\varphi(z)/z|\in L^{\infty}$, and $\text{Arg}(\varphi(z)/z)\in L^{\infty}$. The condition (5.4), combined with this, shows that $\text{Arg}\varphi^{\prime}\in L^{\infty}$. A simple computation shows that $\Omega$ is a radial-Lipschitz domain if and only if the condition (5.4) holds where the constant $\alpha<1$ depends on $M$, from which it follows that $\Omega$ is a chord-arc domain (see p.172 in [35]). Notice that
$\log|\varphi^{\prime}(z)|=\log\left|\frac{\varphi(z)}{z}\right|+\log\left|\frac{z\varphi^{\prime}(z)}{\varphi(z)}\right|=\log\left|\frac{\varphi(z)}{z}\right|-H\left(\text{Arg}\frac{z\varphi^{\prime}(z)}{\varphi(z)}\right),$
and
$H\left(\log\left|\frac{\varphi(z)}{z}\right|\right)=\text{Arg}\frac{\varphi(z)}{z}.$
We have that a radial-Lipschitz curve $\Gamma$ satisfies the condition (5.3). Consequently, the radial-Lipschitz domain has all properties of Theorem 5.5.
The following abstract interpolation Theorem of Voigt [43] is the main tool in our proof of Theorem 5.5:
Theorem 5.6.
Assume that $(X_{0},X_{1})$, $(Y_{0},Y_{1})$ are two interpolation pairs of Banach spaces, and assume that $\check{X}$ is a dense subspace of $(X_{0}\cap X_{1},\|\cdot\|_{X_{0}\cap X_{1}})$. Denote the strip $S:=\{z=s+it\in\mathbb{C}:\;0\leq s\leq 1\}$ and the interpolation spaces $X_{s}:=[X_{0},X_{1}]_{s}$, $Y_{s}:=[Y_{0},Y_{1}]_{s}$. Let $(T_{z};\;z\in S)$ be a family of linear mappings $T_{z}:\check{X}\to Y_{0}+Y_{1}$ with the following properties:
  1. (i)
    For all $x\in\check{X}$ the function $T_{(\cdot)}x:S\to Y_{0}+Y_{1}$ is continuous, bounded on $S$, and analytic in $\stackrel{{\scriptstyle\circ}}{{S}}$;
  2. (ii)
    for $j=0,1$, $x\in\check{X}$, the function $\mathbb{R}\ni t\mapsto T_{j+it}x\in Y_{j}$ is continuous, and
    $M_{j}:=\sup\{\|T_{j+it}x\|_{Y_{j}};t\in\mathbb{R},x\in\check{X},\|x\|_{X_{j}}\leq 1\}<\infty.$
Then, for all $s\in[0,1]$, $T_{s}(\check{X})\subset Y_{s}$,
$\|T_{s}x\|_{Y_{s}}\leq M_{0}^{1-s}M_{1}^{s}\|x\|_{X_{s}}\quad for\;all\;x\in\check{X}.$
We now start the proof of Theorem 5.5. Clearly, we only need to show $\mathcal{HB}_{2,2}^{s}(\Omega\!\to\!\Gamma)=\text{HB}_{2,2}^{s}(\Gamma)$ with comparable norms. Recall that Bergman spaces with standard weights $A^{p}_{\alpha}$ are defined as the sets of holomorphic functions $g$ on $\mathbb{D}$ such that $g\in L^{p}(\mathbb{D},(1-|z|^{2})^{\alpha}dxdy)$, i.e.,
$\|g\|_{A^{p}_{\alpha}}^{p}=\iint_{\mathbb{D}}|g(z)|^{p}(1-|z|^{2})^{\alpha}dxdy<\infty$
where $p>0$ and $\alpha>-1$. Here, the assumption that $\alpha>-1$ is essential because the space $L^{p}(\mathbb{D},(1-|z|^{2})^{\alpha}dxdy)$ does not contain any holomorphic function other than 0 when $\alpha\leq-1$. The definition of $A^{p}_{\alpha}$ is extended to the case where $\alpha$ is any real number; that is consistent with the traditional definition when $\alpha>-1$ (see Theorem 13 in [47]). From the definition it can be seen that if $\alpha<\beta$ then the strict inclusion $A^{p}_{\alpha}\subset A^{p}_{\beta}$ holds.
For our purpose we in particular mention that the space $A^{2}_{-1}$ is the classical Hardy space $E^{2}(\mathbb{D})$, i.e., that is the set of holomorphic functions $g$ in $\mathbb{D}$ such that
$\int_{\mathbb{D}}|g^{\prime}(z)|^{2}(1-|z|^{2})dxdy<\infty.$
We also need to use the case of $p=2$ and $\alpha=1-2s$ where $0\leq s<1$, namely,
$A^{2}_{1-2s}=\{g\;\text{analytic in}\;\mathbb{D}:\iint_{\mathbb{D}}|g(z)|^{2}(1-|z|^{2})^{1-2s}dxdy<\infty\}.$
This is a closed linear subspace of the Hilbert space $L^{2}(\mathbb{D},(1-|z|^{2})^{1-2s}dxdy)$. For these spaces, the following interpolation theorem (see Theorem 36 in [47]) holds true:
$A^{2}_{1},A^{2}_{-1}]_{s}=A^{2}_{1-2s},\qquad 0<s<1$ (5.5)
with equivalent norms.
Suppose $\varphi$ is a conformal map from $\mathbb{D}$ onto $\Omega$ fixing 0. For any $f\in\mathcal{HB}_{2,2}^{s}(\Omega)$, $0\leq s<1$, we have seen from (5.2) that
$\iint_{\mathbb{D}}|V_{s}(f)^{\prime}(z)|^{2}(1-|z|)^{1-2s}dxdy\simeq\iint_{\Omega}|f^{\prime}(\zeta)|^{2}d(\zeta,\Gamma)^{1-2s}d\xi d\eta$
where the implicit constants depend only on $s$, and
$V_{s}(f)^{\prime}(z)=(f\circ\varphi)^{\prime}(z)\varphi^{\prime}(z)^{1/2-s}.$
From it we see that
$f\in\mathcal{HB}_{2,2}^{s}(\Omega)\;\Leftrightarrow\;V_{s}(f)^{\prime}\in A^{2}_{1-2s}$ (5.6)
with comparable norms. Note that any $g\in A^{2}_{1-2s}$ can be written in the form $V_{s}(f)^{\prime}$ by picking $f(z)=\int_{0}^{z}g\circ\varphi^{-1}(u)(\varphi^{-1})^{\prime}(u)^{3/2-s}du$.
Motivated by equivalence (5.6), we define $\mathcal{HB}_{2,2}^{1}(\Omega)$ as the set of analytic functions $f$ on $\Omega$ satisfying $V_{1}(f)^{\prime}\in A^{2}_{-1}$. It is easy to see that $f$ is an anti-derivative of a function in $E^{2}(\Omega)$. To be more precise, recall that $A^{2}_{-1}=E^{2}(\mathbb{D})$ we have
$\int_{\Gamma}|f^{\prime}(\zeta)|^{2}|d\zeta|=\int_{\mathbb{T}}|f^{\prime}\circ\varphi(z)\varphi^{\prime}(z)^{1/2}|^{2}|dz|=\int_{\mathbb{T}}|V_{1}(f)^{\prime}(z)|^{2}|dz|<\infty.$
By this definition, $\mathcal{HB}_{2,2}^{1}(\Omega\!\to\!\Gamma)=\text{HB}_{2,2}^{1}(\Gamma)$, and we also have $\mathcal{HB}_{2,2}^{0}(\Omega\!\to\!\Gamma)=\text{HB}_{2,2}^{0}(\Gamma)$ with equivalent norms by Theorem 5.3. Combining this with (5.6), we can see that this boils down to prove
$f\in\text{HB}_{2,2}^{s}(\Omega)\;\Leftrightarrow\;V_{s}(f)^{\prime}\in A^{2}_{1-2s}$ (5.7)
with comparable norms for $0<s<1$. For this purpose, we will use Theorem 5.6 in the following.
Recall that $(A^{2}_{1},A^{2}_{-1})$ is an interpolation pair and $A^{2}_{1},A^{2}_{-1}]_{s}=A^{2}_{1-2s},\;0<s<1$. Notice that the boundary curve $\Gamma$ of $\Omega$ is a chord-arc curve which is the image of $\mathbb{T}$ under the bi-Lipschitz mapping $\lambda$, we then conclude by (4.3) that $(\text{HB}_{2,2}^{0}(\Gamma),\text{HB}_{2,2}^{1}(\Gamma))$ is an interpolation pair and $\text{HB}_{2,2}^{0}(\Gamma),\text{HB}_{2,2}^{1}(\Gamma)]_{s}=\text{HB}_{2,2}^{s}(\Gamma)$. Note that $\Omega$ is a chord-arc domain, in particular a Smirnov domain, that is equivalent to that the Hardy space $E^{2}(\Omega)$ (i.e., $\text{HB}_{2,2}^{0}(\Omega)$) coincides with the $L^{2}(\Gamma)$ closure of the set $\check{B}$ of polynomials on $\Gamma$ (see Theorem 10.6 in [15]). Since $\text{HB}_{2,2}^{1}(\Gamma)$ is the set of functions $f\in E^{2}(\Omega)$ such that $f^{\prime}\in E^{2}(\Omega)$, the set $\check{B}$ is also a dense subspace of $\text{HB}_{2,2}^{1}(\Gamma)$. By the inclusion relation $\text{HB}_{2,2}^{1}(\Gamma)\subset\text{HB}_{2,2}^{0}(\Gamma)$, we have that $\text{HB}_{2,2}^{0}(\Gamma)\cap\text{HB}_{2,2}^{1}(\Gamma)=\text{HB}_{2,2}^{1}(\Gamma)$.
Define a family of linear mappings $(T_{z};\;z\in S)$ on $\check{B}$ by
$T_{z}(f):=e^{cz^{2}}V_{z}(f)=e^{cz^{2}}f^{\prime}\circ\varphi(\varphi^{\prime})^{\frac{3}{2}-z}.$
Here, $c$ is a constant which will be determined later. Now we will prove that for each $z\in S$, the linear mapping $T_{z}$ sends $\check{B}$ to $A^{2}_{1}$, that is also the sum space $A^{2}_{1}+A^{2}_{-1}$, and satisfies the properties (i) and (ii) of Theorem 5.6.
First, we prove that for each $z=s+it\in S$, each polynomial $f\in\check{B}$, the integral
$\iint_{\mathbb{D}}|T_{z}(f)(\zeta)|^{2}(1-|\zeta|)d\xi d\eta$ (5.8)
converges. From simple computation it follows that $|e^{cz^{2}}|^{2}=e^{2c(s^{2}-t^{2})}$, $|(\varphi^{\prime})^{\frac{3}{2}-z}|^{2}=e^{2t\text{Arg}\varphi^{\prime}}|\varphi^{\prime}|^{3-2s}$ and $|f^{\prime}\circ\varphi|^{2}<C_{1}$ since the diameter of $\Omega$ is finite. Here, the constant $C_{1}$ depends on $f$ and $\Omega$. Then,
$\iint_{\mathbb{D}}|T_{z}(f)(\zeta)|^{2}(1-|\zeta|)d\xi d\eta\leq C_{1}e^{2c}\left(e^{-ct^{2}}e^{t\|\text{Arg}\varphi^{\prime}\|_{\infty}}\right)^{2}\iint_{\mathbb{D}}|\varphi^{\prime}|^{2(1-s)}|\varphi^{\prime}(\zeta)|(1-|\zeta|)d\xi d\eta$
Set $A:=\|\text{Arg}\varphi^{\prime}\|_{\infty}$ and pick $c\geq A^{2}/4$ we then have that $e^{-ct^{2}}e^{tA}\leq e$. By the Koebe distortion theorem (2.1), $|\varphi^{\prime}(\zeta)|(1-|\zeta|)$ is comparable to $d(\varphi(\zeta),\;\Gamma)$, that is bounded due to the boundedness of the diameter of $\Omega$. By Hölder inequality,
$\iint_{\mathbb{D}}|\varphi^{\prime}|^{2(1-s)}d\xi d\eta\leq\pi^{s}\left(\iint_{\mathbb{D}}|\varphi^{\prime}|^{2}d\xi d\eta\right)^{1-s}<\infty.$
We can then conclude that the integral (5.8) converges, and moreover it has an upper-bound depending only on $\Omega$ and given $f$. That implies that the linear mapping $T_{z}$ sends $\check{B}$ to $A^{2}_{1}$, and for any $f\in\check{B}$ the function $T_{(\cdot)}f:\;S\to A^{2}_{1}$ is bounded. In particular,
$\begin{split}\|T_{it}f\|_{A^{2}_{1}}^{2}&=\iint_{\mathbb{D}}|T_{it}(f)(\zeta)|^{2}(1-|\zeta|^{2})d\xi d\eta\\ &\leq e^{2c+3}\iint_{\mathbb{D}}|f^{\prime}\circ\varphi|^{2}|\varphi^{\prime}|^{2}|\varphi^{\prime}|(1-|\zeta|)d\xi d\eta\\ &\leq 4e^{2c+3}\iint_{\Omega}|f^{\prime}(z)|^{2}d(z,\Gamma)dxdy,\end{split}$
that is comparable to $\|f\|^{2}_{B_{2,2}^{0}(\Omega)}$ by Theorem 5.3. Thus,
$M_{0}:=\sup\{\|T_{it}f\|_{A^{2}_{1}};\;t\in\mathbb{R},f\in\check{B},\|f\|_{B_{2,2}^{0}(\Omega)}\leq 1\}<\infty.$
Next, for any $t\in\mathbb{R}$, any $f\in\check{B}$ with $\|f\|_{B_{2,2}^{1}(\Omega)}\leq 1$ the norm $\|T_{1+it}(f)\|_{A^{2}_{-1}}$, by Theorem 5.3, is comparable to $\int_{\mathbb{T}}|T_{1+it}(f)(\zeta)|^{2}|d\zeta|$. Noting
$|T_{1+it}(f)(\zeta)|^{2}=|f^{\prime}\circ\varphi(\zeta)|^{2}|\varphi^{\prime}(\zeta)|e^{2c}\left(e^{-ct^{2}}e^{t\text{Arg}\varphi^{\prime}(\zeta)}\right)^{2},\qquad\zeta\in\mathbb{T},$
we then have
$\int_{\mathbb{T}}|T_{1+it}(f)(\zeta)|^{2}|d\zeta|\leq e^{2(1+c)}\int_{\Gamma}|f^{\prime}(\zeta)|^{2}|d\zeta|\leq e^{2(1+c)}$
Thus,
$M_{1}:=\sup\{\|T_{1+it}f\|_{A^{2}_{-1}};\;t\in\mathbb{R},f\in\check{B},\|f\|_{B_{2,2}^{1}(\Omega)}\leq 1\}<\infty.$
By the continuity of parametrized integral, all functions $S\ni z\mapsto T_{z}f\in A^{2}_{1}$,   $\mathbb{R}\ni t\mapsto T_{it}f\in A^{2}_{1}$, and $\mathbb{R}\ni t\mapsto T_{1+it}f\in A^{2}_{-1}$ are continuous.
Finally, to prove that for any $f\in\check{B}$, $\stackrel{{\scriptstyle\circ}}{{S}}\ni z\mapsto T_{z}f\in A^{2}_{1}$ is a holomorphic map, we use a general result about the infinite dimensional holomorphy (see p.206 in [25]). It says that suppose $X$ and $Y$ are Banach spaces over the complex numbers and $U$ is a domain of $X$ then a continuous function $f:U\to Y$ is holomorphic if there exists a total subset $A$ of the dual $Y^{*}$ such that for every $\alpha\in A$ the function $\alpha\circ f:U\to\mathbb{C}$ is holomorphic. Here, a subset $A$ of $Y^{*}$ is total if $\alpha(y)=0$ for all $\alpha\in A$ implies that $y=0$.
It is known that suppose $\alpha>-1$ and $0<p<\infty$ the inequality
$(1-|\zeta|^{2})^{\frac{2+\alpha}{p}}|g(\zeta)|\leq C\|g\|_{A^{p}_{\alpha}}$ (5.9)
holds for some constant $C$ depending only on $\alpha$ and any $\zeta\in\mathbb{D}$ (see Theorem 2.1 in [46], p.73 in [47]). For each $\zeta\in\mathbb{D}$, define $l_{\zeta}(g)=g(\zeta)$ for $g\in A^{2}_{1}$. By taking $p=2,\alpha=1$ in (5.9), we see that $\|l_{\zeta}\|\leq C(1-|\zeta|^{2})^{-3/2}$, which implies that $l_{\zeta}\in(A^{2}_{1})^{*}$. Set $A:=\{l_{\zeta};\;\zeta\in\mathbb{D}\}$. Clearly, $A$ is a total subset of $(A^{2}_{1})^{*}$. Now for each $l_{\zeta}\in A$, one can see that the function
$\stackrel{{\scriptstyle\circ}}{{S}}\ni z\mapsto l_{\zeta}\circ T_{z}(f)=e^{cz^{2}}f^{\prime}\circ\varphi(\zeta)(\varphi^{\prime})^{3/2-z}(\zeta)\in\mathbb{C}$
is holomorphic.
We can then invoke Theorem 5.6, and see that for all $s\in(0,1)$, $T_{s}(\check{B})\subset A^{2}_{1-2s}$, and moreover,
$\|T_{s}(f)\|_{A^{2}_{1-2s}}\leq M_{0}^{1-s}M_{1}^{s}\|f\|_{B_{2,2}^{s}(\Omega)},\qquad\text{for\; all}\;f\in\check{B}.$ (5.10)
Since $\check{B}$ is dense in $\text{HB}_{2,2}^{0}(\Gamma)\cap\text{HB}_{2,2}^{1}(\Gamma)$ and $\text{HB}_{2,2}^{0}(\Gamma)\cap\text{HB}_{2,2}^{1}(\Gamma)$ is dense in $\text{HB}_{2,2}^{s}(\Gamma)$ for all $s\in[0,1]$ (see [43]), it follows immediately that $\check{B}$ is dense in $\text{HB}_{2,2}^{s}(\Gamma)$. Thus, (5.10) holds true for all $f\in\text{HB}_{2,2}^{s}(\Gamma)$. The bounded inverse theorem implies the inverse of the operator $T_{s}$ is also bounded. This completes the proof of Theorem 5.5.
5.4 The Equality $\mathcal{B}_{p,p}^{s}(\Omega\!\to\!\Gamma)=B_{p,p}^{s}(\Gamma)$
The aim of this part is to prove the following
Theorem 5.7.
Let $\Omega$ be a bounded chord-arc domain bounded by $\Gamma$ with $0\in\Omega$ and $\varphi$ its Riemann mapping fixing 0. Let $0<s<1$, $1<p<\infty$. If $\Gamma$ is such that $\text{Arg}\,\varphi^{\prime}\in L^{\infty}$ and either
  • when $p\geq 2$, $|\varphi^{\prime}|\in A_{p^{\prime}}$$(1/p+1/p^{\prime}=1)$;
  • when $p<2$, $|\varphi^{\prime}|\in A_{p}$
then it holds that $\mathcal{B}_{p,p}^{s}(\Omega\!\to\!\Gamma)=B_{p,p}^{s}(\Gamma)$ with comparable norms.
We recall that if $\delta$ stands for the infinimum of the set of $q$ such that $|\varphi^{\prime}|\in A_{q}$ then the conditions in the theorem are equivalent to
$\delta<q<\frac{\delta}{\delta-1}.$
In particular, $\delta<2$ so that $|\varphi^{\prime}|\in A_{2}$. That implies that $B_{p,p}^{s}(\Gamma)$ is stable by conjugation by Theorem 5.1. Together with $\text{Arg}\,\varphi^{\prime}\in L^{\infty}$, that also implies that $\mathcal{B}_{2,2}^{s}(\Omega\!\to\!\Gamma)=B_{2,2}^{s}(\Gamma)$, $0\leq s\leq 1$, with comparable norms by Theorem 5.5. These two results will be used in the proof of Theorem 5.7.
On the other hand, for the proof of Theorem 5.7 we also need to use the claim: $\mathcal{B}_{p,p}^{1/p}(\Omega\!\to\!\Gamma)=B_{p,p}^{1/p}(\Gamma)$ with comparable norms for $1<p<\infty$ provided that the curve $\Gamma$ is chord-arc. That was proved in our recent paper [45] using the conformal invariance of the space $\mathcal{B}_{p,p}^{1/p}(\Omega)$. We outline its proof here: suppose $f\in B_{p,p}^{1/p}(\Gamma)$ and its harmonic extension in $\Omega$ is still denoted by $f$ then the assertion follows from the following reasoning:
$\|f\|_{B_{p,p}^{1/p}(\Gamma)}\simeq\|f\circ\lambda\|_{B_{p,p}^{1/p}(\mathbb{T})}\simeq\|f\circ\varphi\|_{B_{p,p}^{1/p}(\mathbb{T})}\simeq\|f\circ\varphi\|_{\mathcal{B}_{p,p}^{1/p}(\mathbb{D})}=\|f\|_{\mathcal{B}_{p,p}^{1/p}(\Omega)},$
where the implicit constants depend only on $\Omega$ and $p$. The above steps in order are due to $(1)$ $\lambda$ is a bi-Lipschitz homeomorphism of $\mathbb{T}$ onto $\Gamma$; $(2)$ $\lambda^{-1}\circ\varphi$ is a quasisymmetry; $(3)$ when $\Gamma=\mathbb{T}$ this is well-known as we mentioned before; $(4)$ the space $\mathcal{B}_{p,p}^{1/p}(\Omega)$ is conformal invariant.
Recall that $\mathcal{B}_{p,p}^{s}(\Omega)$ is conjugate invariant. We have seen, under the assumption of Theorem 5.7, that $B_{p,p}^{s}(\Gamma)$ is also conjugate invariant, in order to prove Theorem 5.7 we thus only need to prove the holomorphic case:
Theorem 5.8.
Let $\Omega$ be a bounded chord-arc domain bounded by $\Gamma$ with $0\in\Omega$ and $\varphi$ its Riemann mapping fixing 0. Let $0<s<1$, $1<p<\infty$. If $\Gamma$ is such that $\text{Arg}\,\varphi^{\prime}\in L^{\infty}$ then it holds that
$\mathcal{HB}_{p,p}^{s}(\Omega\!\to\!\Gamma)=\text{HB}_{p,p}^{s}(\Gamma)$ (5.11)
with comparable norms. The conclusion holds in particular for radial-Lipschitz domains.
Proof.
For any $f\in\mathcal{HB}_{p,p}^{s}(\Omega)$, $0<s<1$, $1<p<\infty$ we have seen from (5.2) that
$\iint_{\mathbb{D}}|V_{s,p}(f)^{\prime}(z)|^{p}(1-|z|)^{(1-s)p-1}dxdy\simeq\iint_{\Omega}|f^{\prime}(\zeta)|^{p}d(\zeta,\Gamma)^{(1-s)p-1}d\xi d\eta$
where the implicit constant depends only on $s$ and $p$, and
$V_{s,p}(f)^{\prime}(z)=(f\circ\varphi)^{\prime}(z)\varphi^{\prime}(z)^{1/p-s}.$
From it we see that
$f\in\mathcal{HB}_{p,p}^{s}(\Omega)\;\Leftrightarrow\;V_{s,p}(f)^{\prime}\in A^{p}_{(1-s)p-1}$ (5.12)
with comparable norms. Noting that any $g\in A^{p}_{(1-s)p-1}$ can be written in the form $V_{s,p}(f)^{\prime}$ by choosing $f(z)=\int_{0}^{z}g\circ\varphi^{-1}(u)(\varphi^{-1})^{\prime}(u)^{1+1/p-s}du$, we have that the operator $\mathcal{HB}_{p,p}^{s}(\Omega)\ni f\mapsto V_{s,p}(f)^{\prime}\in A^{p}_{(1-s)p-1}$ is a bounded isomorphism. This boils down to prove
$f\in\text{HB}_{p,p}^{s}(\Omega)\;\Leftrightarrow\;V_{s,p}(f)^{\prime}\in A^{p}_{(1-s)p-1}$ (5.13)
with comparable norms. For this purpose, we will use Theorem 5.6 in the following.
We have proved that for any $1<p<\infty$, $\mathcal{B}_{p,p}^{1/p}(\Omega\!\to\!\Gamma)=B_{p,p}^{1/p}(\Gamma)$ if $\Gamma$ is a chord-arc curve; and for any $0<s<1$, $\mathcal{B}_{2,2}^{s}(\Omega\!\to\!\Gamma)=B_{2,2}^{s}(\Gamma)$ if moreover $\Gamma$ satisfies that $|\varphi^{\prime}|\in A_{2}$ and $\text{Arg}\varphi^{\prime}\in L^{\infty}$, thus the equivalence (5.13) has already been known to hold for $s=1/p$, $1<p<\infty$, and also holds for $0<s<1$, $p=2$.
Recall that $(A^{p}_{p-2},A^{2}_{1-2s})$ is an interpolation pair and for $0\leq\theta\leq 1$,
$A^{p}_{p-2},A^{2}_{1-2s}]_{\theta}=A^{q}_{(1-\alpha)q-1}.$ (5.14)
We have also seen that $(\text{HB}_{p,p}^{1/p}(\Gamma),\text{HB}_{2,2}^{s}(\Gamma))$ is an interpolation pair and
$\text{HB}_{p,p}^{1/p}(\Gamma),\text{HB}_{2,2}^{s}(\Gamma)]_{\theta}=\text{HB}_{q,q}^{\alpha}(\Gamma)$ (5.15)
provided that $\Gamma$ is a chord-arc curve. Here,
$\begin{split}\alpha=\alpha(\theta)&:=(1-\theta)/p+\theta s,\\ q=q(\theta)&:=2p/(2(1-\theta)+\theta p).\end{split}$ (5.16)
Some elementary arithmetic leads to that for any given $\alpha\in(0,1)$ and $q\in(1,\infty)$, there exist $p\in(1,\infty)$, $s\in(0,\,\min(1,5/2-2/p))$ and $\theta\in[0,1]$ such that (5.16) holds. Recall that in the proof of Theorem 5.5 we have shown that $\check{B}$, the set of polynomials on $\Gamma$, is dense in $\text{HB}_{2,2}^{s}(\Gamma)$. By using the similar argument, we can see that $\check{B}$ is also dense in $\text{HB}_{p,p}^{1/p}(\Gamma)$ and thus $\check{B}$ is a dense subspace of the intersection space $\text{HB}_{p,p}^{1/p}(\Gamma)\cap\text{HB}_{2,2}^{s}(\Gamma)$.
Define a family of linear mappings $(\widetilde{T}_{z};\;z=\theta+it\in S)$, $S$ standing for a vertical strip with $0\leq\theta\leq 1$ as before, on $\check{B}$ by
$\widetilde{T}_{z}(f):=e^{cz^{2}}V_{\alpha(z),q(z)}(f)^{\prime}=e^{cz^{2}}f^{\prime}\circ\varphi(\varphi^{\prime})^{\frac{1}{q(z)}+1-\alpha(z)}=e^{cz^{2}}f^{\prime}\circ\varphi(\varphi^{\prime})^{(\frac{1}{2}-s)z+1}.$
Here, we still choose the constant $c\geq A^{2}/4$ as in the proof of Theorem 5.5 where $A:=\|\text{Arg}\varphi^{\prime}\|_{\infty}$, so that $e^{-ct^{2}}e^{tA}\leq e$. We will prove that for each $z=\theta+it\in S$, the linear mapping $\widetilde{T}_{z}$ sends $\check{B}$ to $A^{p}_{p-2}+A^{2}_{1-2s}$, and the family $(\widetilde{T}_{z})$ satisfies the properties (i) and (ii) of Theorem 5.6.
We first estimate $|\widetilde{T}_{z}(f)(\zeta)|^{p}$ for each $z=\theta+it\in S$ and each $f\in\check{B}$. It follows from simple computation that
$|e^{cz^{2}}|^{p}=e^{pc(\theta^{2}-t^{2})}\leq e^{pc}e^{-ct^{2}p},$
and
$|(\varphi^{\prime})^{(1/2-s)z+1}|^{p}=e^{t(s-1/2)p\text{Arg}\varphi^{\prime}}\left(|\varphi^{\prime}|^{1+(1/2-s)\theta}\right)^{p}\leq e^{|t|pA}\left(|\varphi^{\prime}|^{1+(1/2-s)\theta}\right)^{p}.$
There exists a constant $C_{1}$ such that $|f^{\prime}\circ\varphi|^{p}<C_{1}$ since the diameter of the domain $\Omega$ with 0 as an interior point is finite. Then,
$|\widetilde{T}_{z}(f)(\zeta)|^{p}\leq C_{1}e^{pc}\left(e^{-ct^{2}+|t|A}\right)^{p}\left(|\varphi^{\prime}|^{1+(1/2-s)\theta}\right)^{p}\leq C_{1}e^{pc}e^{p}\left(|\varphi^{\prime}|^{1+(1/2-s)\theta}\right)^{p}.$ (5.17)
To show the linear mapping $\widetilde{T}_{z}$ sends $\check{B}$ to $A^{p}_{p-2}+A^{2}_{1-2s}$, we split into the following three cases:
Case $1$. when $s\leq 1/2$
$\iint_{\mathbb{D}}|\widetilde{T}_{z}(f)(\zeta)|^{2}(1-|\zeta|)^{1-2s}d\xi d\eta$ $\leq C_{1}e^{2c}e^{2}\iint_{\mathbb{D}}\left(|\varphi^{\prime}|^{1+(1/2-s)\theta}\right)^{2}(1-|\zeta|)^{1-2s}d\xi d\eta$
$\leq C_{1}e^{2c}e^{2}\iint_{\mathbb{D}}|\varphi^{\prime}|^{2}|\varphi^{\prime}|^{(1-2s)\theta}(1-|\zeta|)^{(1-2s)\theta}d\xi d\eta$
which is bounded by a constant depending only on $s$ and $\Omega$.
Case $2$. when $s>1/2$ and $p\geq 2$
$\iint_{\mathbb{D}}|\widetilde{T}_{z}(f)(\zeta)|^{p}(1$- $|\zeta|)^{p-2}d\xi d\eta\leq C_{1}e^{pc}e^{p}\iint_{\mathbb{D}}\left(|\varphi^{\prime}|^{1+(1/2-s)\theta}\right)^{p}(1-|\zeta|)^{p-2}d\xi d\eta$
$\leq$ $C_{1}e^{pc}e^{p}\iint_{\mathbb{D}}\left((|\varphi^{\prime}|(1-|\zeta|))^{1+(1/2-s)\theta}\right)^{p-2}|\varphi^{\prime}|^{2+(1-2s)\theta}d\xi d\eta$
$\leq$ $C_{1}e^{pc}e^{p}\text{diam}(\Gamma)^{p-2}\iint_{\mathbb{D}}|\varphi^{\prime}|^{2+(1-2s)\theta}d\xi d\eta$
which, using Hölder inequality, is bounded by a constant depending only on $p$ and $\Omega$.
Case $3$. When $1/2<s<\min(1,5/2-2/p)$ and $1<p<2$, it suffices to show that for all $\theta\in(0,1)$,
$\iint_{\mathbb{D}}|\varphi^{\prime}(z)|^{2+p\theta(\frac{1}{2}-s)}d(\varphi(z),\Gamma)^{p-2}dxdy<\infty.$
First of all, since $s>1/2$ and $p<2$ we have that
$1<2+p\theta(\frac{1}{2}-s)<2.$
By Hölder inequality,
$\begin{split}&\frac{1}{\pi}\iint_{\mathbb{D}}|\varphi^{\prime}(z)|^{2+p\theta(\frac{1}{2}-s))}d(\varphi(z),\Gamma)^{p-2}dxdy\\ \leq&\left(\frac{1}{\pi}\iint_{\mathbb{D}}|\varphi^{\prime}(z)|^{2}d(\varphi(z),\Gamma)^{\frac{2(p-2)}{2+p\theta(\frac{1}{2}-s)}}dxdy\right)^{\frac{2+p\theta(\frac{1}{2}-s)}{2}},\end{split}$
and the last integral is equal, by the change of variable $\zeta=\varphi(z)$, to
$\iint_{\Omega}d(\zeta,\Gamma)^{\frac{2(p-2)}{2+p\theta(\frac{1}{2}-s)}}d\xi d\eta.$
The domain being Lipschitz, this integral is finite if and only if
$\frac{2(p-2)}{2+p\theta(\frac{1}{2}-s)}>-1,$
or, equivalently, if and only if
$0<\theta<2\frac{1-\frac{1}{p}}{s-\frac{1}{2}}.$
A sufficient condition for the finiteness will thus be that
$2\frac{1-\frac{1}{p}}{s-\frac{1}{2}}>1\Leftrightarrow s<\frac{5}{2}-\frac{2}{p}.$
Consequently, when $1<p<\infty,\,0<s<\min(1,5/2-2/p)$, the linear ma??pping $\widetilde{T}_{z}$ sends $\check{B}$ to $A^{p}_{p-2}+A^{2}_{1-2s}$, and for all $f\in\check{B}$, the mapping $\widetilde{T}_{(\cdot)}(f):S\to A^{p}_{p-2}+A^{2}_{1-2s}$ is bounded.
Taking $\theta=0$ in (5.17) we can see
$\iint_{\mathbb{D}}|\widetilde{T}_{it}(f)|^{p}(1-|z|^{2})^{p-2}dxdy\leq C(p,c)\iint_{\Omega}|f^{\prime}(\zeta)|^{p}d(\zeta,\Gamma)^{p-2}d\xi d\eta$
where the constant $C(p,c)$ depends only on $p$ and $c$. By that we get
$M_{0}$ $:=\sup\{\|\widetilde{T}_{it}f\|_{A^{p}_{p-2}};\;t\in\mathbb{R},f\in\check{B},\|f\|_{B_{p,p}^{1/p}(\Gamma)}\leq 1\}<\infty.$
Taking $\theta=1$ in (5.17) we can see
$\iint_{\mathbb{D}}|\widetilde{T}_{1+it}(f)|^{2}(1-|z|^{2})^{1-2s}dxdy\leq C(c,s)\iint_{\Omega}|f^{\prime}(\zeta)|^{2}d(\zeta,\Gamma)^{1-2s}d\xi d\eta$
where $C(s,c)$ is a constant depending only on $s$ and $c$. From this it follows that
$M_{1}$ $:=\sup\{\|\widetilde{T}_{1+it}f\|_{A^{2}_{1-2s}};\;t\in\mathbb{R},f\in\check{B},\|f\|_{B_{2,2}^{s}(\Gamma)}\leq 1\}<\infty.$
By the continuity of parametrized integral, all functions $S\ni z\mapsto\widetilde{T}_{z}f\in A^{p}_{p-2}+A^{2}_{1-2s}$,   $\mathbb{R}\ni t\mapsto\widetilde{T}_{it}f\in A^{p}_{p-2}$, and $\mathbb{R}\ni t\mapsto\widetilde{T}_{1+it}f\in A^{2}_{1-2s}$ are continuous. Finally, for any $f\in\check{B}$, the proof of holomorphy of the mapping $\stackrel{{\scriptstyle\circ}}{{S}}\ni z\mapsto\widetilde{T}_{z}f\in A^{p}_{p-2}+A^{2}_{1-2s}$ is similar to that in the proof of Theorem 5.5. We omit the details here. This completes the proof of Theorem 5.8. ∎
5.5 Some Final Remarks
In this part, we let $0<s<1$ and $1<p<\infty$, and suppose $\Gamma$ satisfies the assumption of Theorem 5.7, namely $\Gamma$ is a bounded chord-arc curve with 0 in its interior domain $\Omega$, and $\Gamma$ is such that $\text{Arg}\,\varphi^{\prime}\in L^{\infty}$ and $|\varphi^{\prime}|\in A_{q}$ with $q=\min{(p,p^{\prime})}$. Here, $1/p+1/p^{\prime}=1$ and $\varphi_{i}^{\prime}$ is a Riemann mapping from the unit disk $\mathbb{D}$ onto $\Omega$ keeping 0 fixed. Recall that when $p=2$ we have shown that the radial-Lipschitz curve satisfies this assumption.
Since $h(\Gamma)=1$ one may see that $(h(\Gamma),p,s)$ is always in the admissible set $\mathcal{A}$. It follows from Theorem 6.2 and Theorem 5.7 (or Theorem 5.5 for the case of $p=2$) that in this case $\mathcal{B}_{p,p}^{s}(\Omega_{i}\!\to\!\Gamma)=\mathcal{B}_{p,p}^{s}(\Omega_{e}\!\to\!\Gamma)$ with comparable norms. Recall that we explored earlier that this equality also holds under a different assumption where there are some restrictions on $p$ and $s$ in Theorem 3.12.
If $p>2$, $s\geq 1-1/p$ we then have that
  • (i)
    $\mathcal{B}_{p,p}^{s}(\Gamma)=W^{1,p}(\omega,\mathbb{C})|_{\Gamma}$ where $\omega(z)=d(z,\Gamma)^{(1-s)p-1}$, i.e., the space $\mathcal{B}_{p,p}^{s}(\Gamma)$ is the trace space of $W^{1,p}(\omega,\mathbb{C})$ on $\Gamma$;  (by Theorem 3.16)
  • (ii)
    The space $C_{c}^{\infty}(\mathbb{C})|_{\Gamma}$ is dense in $\mathcal{B}_{p,p}^{s}(\Gamma)$, and thus Plemelj-Calderón property holds for any function in $\mathcal{B}_{p,p}^{s}(\Gamma)$.  (by Corollary 3.17 and Theorem 3.5)
  • (iii)
    for any $f\in\mathcal{B}_{p,p}^{s}(\Gamma)$, its extension to $\mathbb{C}$ by the harmonic extensions on both connected components $\Omega_{i,e}$ of $\overline{\mathbb{C}}\setminus\Gamma$ belongs to $W^{1,p}(\omega,\mathbb{C})$;  (by section 3.3.2)
  • (iv)
    Any function in $\mathcal{B}_{p,p}^{s}(\Omega_{i,e})$ satisfies the almost-Dirichlet principle.  (by section 3.2.3)
The above statements are claims: given $p>2$ and thus the class of curves $\Gamma$ for which claims are true depends on $p$ (i,e., $|\varphi^{\prime}|\in A_{p^{\prime}}$). If we reverse the point of view and start with a curve $\Gamma$ satisfying $\text{Arg}\,\varphi^{\prime}\in L^{\infty}$ and $|\varphi^{\prime}|\in A_{2}$ then the claims are true for $p\in(\delta,\frac{\delta}{\delta-1})$ where $\delta$ is the infinimum of the set of $p^{\prime}s$ such that $|\varphi^{\prime}|\in A_{p}$. By the theory of $A_{\infty}$ weights, $\delta<2$.
6 Appendix
Let $\Gamma$ be a Jordan curve with $\infty\notin\Gamma$, and $\Omega_{i,e}$ the connected components of $\overline{\mathbb{C}}\setminus\Gamma$ as above. In all this work we have treated $\Omega_{i}$ and $\Omega_{e}$ similarly: in doing so we have overlooked the fact that $\Omega_{e}$ is unbounded, thus requiring an extra argument. The purpose of this appendix is to fill this gap.
Without loss of generality, suppose that $\Gamma$ is included in the unit disk $\mathbb{D}$ with 0 being an interior point. The map $z\mapsto 1/z$ transforms $\Omega_{e}$ into a bounded domain $\Omega^{\prime}_{i}$ with boundary $\Gamma^{\prime}=1/\Gamma$; $\Omega_{i}^{\prime}$ contains 0 which corresponds to $\infty$ in $\Omega_{e}$. We say $u$ is harmonic in the domain $\Omega_{e}$ with $\infty$ being an interior point if $u(1/z)$ is harmonic in the bounded domain $\Omega_{i}^{\prime}$. Let $p>1$ and $s\in(0,1)$ as usual and consider $u\in\mathcal{B}_{p,p}^{s}(\Omega_{e})$. Let $v:\Omega_{i}^{\prime}\to\mathbb{C}$ being defined by $v(z)=u(1/z)$. By a change of the variable $z=1/w$ we conclude by the Koebe distortion theorem (2.1) that
$\iint_{\Omega_{e}}|\nabla u(w)|^{p}d(w,\Gamma)^{(1-s)p-1}dudv\simeq\iint_{\Omega_{i}^{\prime}}|\nabla v(z)|^{p}d(z,\Gamma^{\prime})^{(1-s)p-1}|z|^{2(sp-1)}dxdy$ (6.1)
where the implicit constant is a universal one. We have the following
Proposition 6.1.
If $u\in\mathcal{B}_{p,p}^{s}(\Omega_{e})$ then $v\in\mathcal{B}_{p,p}^{s}(\Omega_{i}^{\prime}).$
Proof.
Let $r<1/4$ and $R_{1},R_{2}$ be such that $\frac{1}{2}+r<R_{1}<R_{2}<1-r$. For any $\zeta$ such that $|\zeta|<r$ the annulus $A_{\zeta}=\{R_{1}<|z-\zeta|<R_{2}\}$ is included in the annulus $\{\frac{1}{2}<|z|<1\}$. By the mean value property for harmonic functions we have, for $|\zeta|<r$,
$|\nabla v(\zeta)|=\frac{1}{\pi R_{i}^{2}}\iint_{|z-\zeta|<R_{i}}\nabla v(z)dxdy,\quad i=1,2$
so that
$|\nabla v(\zeta)|\leq\frac{1}{\pi(R_{2}^{2}-R_{1}^{2})}\iint_{\{1/2<|z|<1\}}|\nabla v(z)|dxdy.$
Noting (6.1) we can conclude by using Hölder inequality that $|\nabla v|$ is uniformly bounded in the disk $|\zeta|<r$ by a constant depending only on $\Gamma$ multiplied by the norm of $u$ in $\mathcal{B}_{p,p}^{s}(\Omega_{e})$. This estimate implies that
$\iint_{|z|<r}|\nabla v(z)|^{p}d(z,\Gamma^{\prime})^{(1-s)p-1}dxdy<\infty.$ (6.2)
In order to control the rest of the integral giving the norm of $v$:
$\iint_{\Omega^{\prime}_{i}\setminus(|z|<r)}|\nabla v(z)|^{p}d(z,\Gamma^{\prime})^{(1-s)p-1}dxdy,$ (6.3)
we first observe that the map $w=1/z$ is bi-Lipschitz in the domain $\Omega^{\prime}_{i}\setminus(|z|<r)$ and then that there exists a constant $C$ depending only on $\Gamma$ such that
$\frac{1}{C|w|}d(w,\Gamma)\leq d(1/w,\Gamma^{\prime})\leq\frac{C}{|w|}d(w,\Gamma).$
The details are left to the reader. ∎
Theorem 6.2.
If $\mathcal{B}_{p,p}^{s}(\Omega_{i}\!\to\!\Gamma)=B_{p,p}^{s}(\Gamma)$, and the identity operator from $\mathcal{B}_{p,p}^{s}(\Omega_{i}\!\to\!\Gamma)$ onto $B_{p,p}^{s}(\Gamma)$ is a bounded isomorphism, then the same assertions hold for the unbounded component $\Omega_{e}$ of $\overline{\mathbb{C}}\setminus\Gamma$, and in particular, the transmission operator from $\mathcal{B}_{p,p}^{s}(\Omega_{i})$ onto $\mathcal{B}_{p,p}^{s}(\Omega_{e})$,
$\mathcal{B}_{p,p}^{s}(\Omega_{i})\to\mathcal{B}_{p,p}^{s}(\Omega_{i}\!\to\!\Gamma)\to\mathcal{B}_{p,p}^{s}(\Omega_{e}),$
is a bounded isomorphism.
Proof.
Let us use the above proposition to prove that $\mathcal{B}_{p,p}^{s}(\Omega_{e}\!\to\!\Gamma)=B_{p,p}^{s}(\Gamma)$. If $u\in\mathcal{B}_{p,p}^{s}(\Omega_{e})$ then $v\in\mathcal{B}_{p,p}^{s}(\Omega^{\prime}_{i})$ so that $v\in\mathcal{B}_{p,p}^{s}(\Omega_{i}^{\prime}\!\to\!\Gamma^{\prime})=B_{p,p}^{s}(\Gamma^{\prime})$ from which it follows easily that $u\in B_{p,p}^{s}(\Gamma)$. ∎
Statements and Declarations. The authors declare that there is no conflict of interest, and there is no data associated to this work.
Acknowledgments. This work is supported by the National Natural Science Foundation of China (Grant No. 12271218 and 12571083) and the University of Orléans. The authors also warmly thank the Poincaré Institute in Paris for its hospitality through the framework of the program "Research in Paris". Special thanks also to Gérard Bourdaud for his help and for interesting discussions.

### v2

@@
 Let $\Gamma$ be a bounded Jordan curve and $\Omega_{i},\Omega_{e}$ its two complementary components.
-For $1<p<\infty,\,s\in(0,1)$ we define $\mathcal{B}_{p,p}^{s}(\Omega_{i,e})$ as the set of functions $f:\Gamma\to\mathbb{C}$ having harmonic extension $u$ respectively in $\Omega_{i}$ and $\Omega_{e}$ such that
+For $p\in(1,\infty),\,s\in(0,1)$ we define the two spaces $\mathcal{B}_{p,p}^{s}(\Omega_{i,e})$ as the set of harmonic functions $u$ respectively in $\Omega_{i}$ and $\Omega_{e}$ such that
 $\iint_{\Omega_{i,e}}|\nabla u(z)|^{p}d(z,\Gamma)^{(1-s)p-1}dxdy<+\infty.$
-If $\Gamma$ is further assumed to be rectifiable we define $B_{p,p}^{s}(\Gamma)$ as the space of functions $f\in L^{p}(\Gamma)$ such that
+When it is possible to identify these spaces with spaces of functions on the boundary (trace spaces), we address the question of their equality.
+When $\Gamma$ is the unit circle, these two spaces coincide with homogeneous fractional Besov-Sobolev spaces and the framework of quasicircles appears to be an appropriate generalization.
+In this framework, we study the boundedness of the Plemelj-Calderón operator and apply the results to show that for some values of $p,s$, if the two spaces coincide, they are restrictions to $\Gamma$ of some weighted Sobolev space.
+If $\Gamma$ is further assumed to be rectifiable, we define $B_{p,p}^{s}(\Gamma)$ as the space of functions $f\in L^{p}(\Gamma)$ such that
 $\iint_{\Gamma\times\Gamma}\frac{|f(z)-f(\zeta)|^{p}}{|z-\zeta|^{1+ps}}|dz||d\zeta|<+\infty.$
-When $\Gamma$ is the unit circle these three spaces coincide with the homogeneous fractional Besov-Sobolev space.
-For a general rectifiable curve these spaces need not coincide and our first goal is to investigate the cases of equality: while the chord-arc property is the necessary and sufficient condition for equality in the classical case of $s=1/p,\,p\geq 2$, this is no longer the case for general $s\in(0,1)$.
-We show however that equality holds for radial-Lipschitz curves.
-In the general (possibly non-rectifiable) case we study boundary values of functions in $\mathcal{B}_{p,p}^{s}(\Omega_{i,e})$ and give conditions for equality of these trace-spaces that we then call $\mathcal{B}^{s}_{p,p}(\Gamma)$.
-Using Plemelj-Calderón property we further identify $\mathcal{B}^{s}_{p,p}(\Gamma)$ with the space of restrictions of a weighted Sobolev space of the plane.
-Finally we re-interpretate some of our results as the "almost"-Dirichlet principle in the spirit of Maz’ya.
-Keywords: Cauchy integral, fractional Sobolev spaces, chord-arc curves, Lipschitz curves, Dirichlet spaces
+Again, these spaces coincide with the homogeneous fractional Besov-Sobolev spaces for the unit circle. While the chord-arc property is the necessary and sufficient condition for the equality
+$\mathcal{B}_{p,p}^{s}(\Omega_{i})=\mathcal{B}_{p,p}^{s}(\Omega_{e})=B_{p,p}^{s}(\Gamma)$
+in the case of $s=1/p,\,p\geq 2$, this is no longer the case for general $s\in(0,1)$.
+However, we show that equality holds for radial-Lipschitz curves.
+Finally, we re-interpretate some of our results as some "almost"-Dirichlet principle in the spirit of Maz’ya.
+Keywords: Cauchy integral, fractional Sobolev spaces, Besov spaces, quasicircles, chord-arc curves, Lipschitz curves, Dirichlet spaces
 2020 MSC: Primary 42B20, 46E35; Secondary 31A05, 30H35
 1 Introduction
-1.1 The Dirichlet Principle
-In all this work we will be considering a bounded Jordan curve $\Gamma$ and call $\Omega_{i},\Omega_{e}$ respectively the interior and exterior connected components of its complement in the sphere.
-If $f$ is a continuous function on $\Gamma$ and $F$ a continuous extension of $f$ to $\overline{\Omega}_{i,e}$ which is furthermore assumed to be $C^{1}$ in $\Omega_{i,e}$ we call (2)(2-)energy the quantity
+1.1 Dirichlet principle
+In all this work, we will consider a bounded Jordan curve $\Gamma$ and call $\Omega_{i},\Omega_{e}$ the interior and exterior connected components of its complement in the sphere.
+If $f$ is a continuous function on $\Gamma$ and $F$ a continuous extension of $f$ to $\overline{\Omega}_{i,e}$, which is furthermore assumed to be $C^{1}$ in $\Omega_{i,e}$, we call (2)(2-)energy the quantity
 $\mathcal{E}_{i,e}(F)=\iint_{\Omega_{i,e}}|\nabla F(z)|^{2}dxdy\leq+\infty.$
-The Dirichlet principle asserts that among all the extensions of $f$ the one with the lowest energy is the harmonic one.
+The Dirichlet principle states that among all the extensions of $f$ the one with the lowest energy is the harmonic one.
 Notice that the (2)(2-)energy of the harmonic extension of a continuous function may be infinite, in which case every extension has infinite (2)(2-)energy.
-When $\Gamma=\mathbb{T}$, the unit circle, we define $\mathcal{D}(\mathbb{D}_{i})$, the Dirichlet space over the unit disk $\Omega_{i}=\mathbb{D}_{i}$ as the space of harmonic functions in $\mathbb{D}_{i}$ with finite (2)(2-)energy.
+When $\Gamma=\mathbb{T}$, the unit circle, we define $\mathcal{D}(\mathbb{D}_{i})$, the Dirichlet space on the unit disk $\Omega_{i}=\mathbb{D}_{i}$ as the space of harmonic functions in $\mathbb{D}_{i}$ with finite (2)(2-)energy.
 This space coincides with the space of Poisson integrals of $L^{2}(\mathbb{T})$-functions such that
@@
 $\iint_{\mathbb{T}\times\mathbb{T}}\frac{|f(z)-f(\zeta)|^{2}}{|z-\zeta|^{2}}|dz||d\zeta|<\infty.$
-The map $z\mapsto 1/\bar{z}$ is a reflection about the unit circle which is bi-Lipschitz from $\{1/2<|z|<2\}$ onto itself.
+The map $z\mapsto 1/\bar{z}$ is a reflection about the unit circle that is bi-Lipschitz from $\{1/2<|z|<2\}$ onto itself.
 It follows that if $u_{i}\in\mathcal{D}(\mathbb{D}_{i})$ is the Poisson integral of $f\in L^{2}(\mathbb{T})$ then the Poisson extension $u_{e}$ of $f$ in $\mathbb{D}_{e}=\overline{\mathbb{C}}\setminus\overline{\mathbb{D}}_{i}$ also has a finite energy.
-We may thus view the Dirichlet space as the subset of $L^{2}(\mathbb{T})$ of functions whose harmonic extension to $\overline{\mathbb{C}}\setminus\mathbb{T}$ has finite energy.
 The Dirichlet space is a special case of homogeneous fractional Sobolev spaces.
@@
 $\|f\|^{2}_{H^{s}}=\sum\limits_{n\in\mathbb{Z}}|n|^{2s}|\hat{f}(n)|^{2}<\infty.$
-It is not difficult to see that the above quantity, the square of the $H^{s}$-semi-norm, is equivalent (in particular, in the case of $s=1/2$ it is equal, up to a constant multiplicative factor,) to two other:
+It is not difficult to see that the above quantity, the square of the $H^{s}$-semi-norm, is equivalent (in particular, in the case of $s=1/2$ it is equal, up to a constant multiplicative factor) to two other:
   1. 1.
@@
     $\iint_{\mathbb{T}\times\mathbb{T}}\frac{|f(z)-f(\zeta)|^{2}}{|z-\zeta|^{1+2s}}|dz||d\zeta|.$
-The mention of Littlewood-Paley comes from the fact that for $s=0$ the Littlewood-Paley semi-norm is exactly the $L^{2}$-norm of $\mathsf{g}(f)$, the Littlewood-Paley function associated to $f$ [37].
+The mention of Littlewood-Paley comes from the fact that for $s=0$ the Littlewood-Paley semi-norm is exactly the $L^{2}$-norm of $\mathsf{g}(f)$, the Littlewood-Paley function associated with $f$ (see [44]).
 In terms of Besov spaces,
 $H^{s}(\mathbb{T})=B_{2,2}^{s}(\mathbb{T}),$
-and that this space also coincides with $(I-\Delta)^{-s/2}(L^{2}(\mathbb{T}))$, namely the space of Bessel potentials of order $s$.
+and this space also coincides with $(I-\Delta)^{-s/2}(L^{2}(\mathbb{T}))$, that is, the space of Bessel potentials of order $s$.
 We can further generalize these spaces by replacing $2$ with $p>1$ and define $B_{p,p}^{s}(\mathbb{T})$ as the set of $L^{p}(\mathbb{T})$-functions such that either
@@
     $\iint_{\mathbb{T}\times\mathbb{T}}\frac{|f(z)-f(\zeta)|^{p}}{|z-\zeta|^{1+ps}}|dz||d\zeta|<\infty.$
-It has to be noticed that if $p\neq 2$,
+It should be noted that if $p\neq 2$,
 $B_{p,p}^{s}(\mathbb{T})\neq(I-\Delta)^{-s/2}(L^{p}(\mathbb{T})),$
-so that, to avoid confusions, we will always use the Besov terminology at least for $p\neq 2$.
+so that, to avoid confusion, we will use the Besov terminology.
+The spaces $B_{p,p}^{s}(\mathbb{T})$, $s\in(0,1)$, $p\in(1,\infty)$, are "intermediate" spaces between two Sobolev spaces, so they are also called fractional Sobolev spaces in the literature.
+Throughout this paper, we call them (along with their generalizations to the more general Jordan curves) fractional Besov-Sobolev spaces.
 The main goal of this paper is to extend these notions to more general Jordan curves.
-The first restriction we want to impose on the Jordan curve $\Gamma$ is that we want it to share with $\mathbb{T}$ the reflection property, that is we want the existence of a reflection across $\Gamma$ which is bi-Lipschitz in a neighborhood of $\Gamma$.
+The first restriction on the Jordan curve $\Gamma$ is that we want it to share with $\mathbb{T}$ the reflection property, that is, the existence of a reflection across $\Gamma$ which is bi-Lipschitz in a neighborhood of $\Gamma$.
 The curves satisfying this property are well-known: they are the quasicircles, i.e., the quasiconformal images of circles [1].
@@
 where $\gamma_{j},j=1,2$, are the two subarcs of $\Gamma$ with endpoints $z_{1},z_{2}$.
-A quasicircle need not be smooth, and the Hausdorff dimension of a quasicircle can take any value in $[1,2)$.
-As an example, the Von Koch snowflake curve is a quasicircle without any tangent and whose Hausdorff dimension, which coincides with its Minkowski dimension, is $\log 4/\log 3$.
+A quasicircle does not necessarily need to be smooth, and the Hausdorff dimension of a quasicircle can take any value in $[1,2)$.
+As an example, the Von Koch snowflake curve is a quasicircle without any tangent, and its Hausdorff dimension, which coincides with its Minkowski dimension, is $\log 4/\log 3$.
 From now on, all the Jordan curves that we consider will be assumed to be quasicircles.
-1.2 Statement of the Main Theorems
+1.2 Statement of main theorems
 Let $\Gamma$ be a quasicircle and $\Omega_{i,e}$ the inner and outer connected components of $\overline{\mathbb{C}}\setminus\Gamma$ as above.
-Suppose $\Omega$ is one of these two components and correspondingly $\mathbb{D}$ is one of $\mathbb{D}_{i}$ and $\mathbb{D}_{e}$.
-We define
-$\mathcal{B}_{p,p}^{s}(\Omega)=\{u\;\mathrm{harmonic\;in}\;\Omega:\;\|u\|_{\mathcal{B}_{p,p}^{s}(\Omega)}^{p}:=\iint_{\Omega}|\nabla u(z)|^{p}d(z,\Gamma)^{(1-s)p-1}dxdy<\infty\}.$
-We define for functions in these spaces and for some values of p,sp,\,s their boundary values on $\Gamma$, which will allow us to identify $\mathcal{B}_{p,p}^{s}(\Omega)$ with a space of functions defined on $\Gamma$ that we denote by $\mathcal{B}_{p,p}^{s}(\Omega\!\to\!\Gamma)$.
-In order to define this space we distinguish three cases, namely, s>1/p,s=1/p,s<1/ps>1/p,\,s=1/p,\,s<1/p.
+Suppose $\Omega$ is one of these two components and, correspondingly, $\mathbb{D}$ is one of $\mathbb{D}_{i}$ and $\mathbb{D}_{e}$.
+We define, for $p>1$ and $0<s<1$,
+$\mathcal{B}_{p,p}^{s}(\Omega)=\{u\;\mathrm{harmonic\;in}\;\Omega:\;\|u\|_{\mathcal{B}_{p,p}^{s}(\Omega)}^{p}=\iint_{\Omega}|\nabla u(z)|^{p}d(z,\Gamma)^{(1-s)p-1}dxdy<\infty\}.$
+We now define for functions in these spaces and for some values of p,sp,\,s their boundary values on $\Gamma$, which will allow us to identify $\mathcal{B}_{p,p}^{s}(\Omega)$ with a space of functions defined on $\Gamma$ that we denote by $\mathcal{B}_{p,p}^{s}(\Omega\!\to\!\Gamma)$.
+In order to define this space, we distinguish three cases, namely s>1/p,s=1/p,s<1/ps>1/p,\,s=1/p,\,s<1/p.
   1. 1.
@@
     $\forall f\in\mathcal{B}_{p,p}^{1/p}(\Omega),\quad\frac{1}{4}\|f\|_{\mathcal{B}_{p,p}^{1/p}(\Omega)}\leq\|Tf\|_{\mathcal{B}_{p,p}^{1/p}(\mathbb{D})}\leq 4\|f\|_{\mathcal{B}_{p,p}^{1/p}(\Omega)}.$
-    Since this latter space is included in the harmonic Hardy space $h^{p}(\mathbb{D})$ we conclude that $f$ has a well-defined boundary value in $L^{p}(\omega,\Gamma)$, where $\omega$ is the harmonic measure for $\Omega$.
+    Since the latter space is included in the harmonic Hardy space $h^{p}(\mathbb{D})$, we conclude that $f$ has a well-defined boundary value in $L^{p}(\omega,\Gamma)$, where $\omega$ is the harmonic measure for $\Omega$.
   3. 3.
@@
 Let $C_{c}^{\infty}(\mathbb{C})|_{\Gamma}=\{f|_{\Gamma}:\;f\in C_{c}^{\infty}(\mathbb{C})\}$, $C_{c}^{\infty}(\mathbb{C})$-functions being $C^{\infty}$-functions with compact support in $\mathbb{C}$.
-Let $\mathcal{D}_{p}^{s}(\Gamma)$ be the set of functions in $C_{c}^{\infty}(\mathbb{C})|_{\Gamma}$ such that their harmonic extensions to $\Omega_{i,e}$ belongs to $\mathcal{B}_{p,p}^{s}(\Omega_{i,e})$, and we denote the set of such extensions by $\mathcal{D}_{p}^{s}(\Omega_{i,e})$.
-We define the space of boundary values $\mathcal{B}_{p,p}^{s}(\Omega_{i,e}\!\to\!\Gamma)$ as the abstract closure of this space, i.e., equivalence classes of Cauchy sequences in $\mathcal{D}_{p}^{s}(\Gamma)$, equipped with the $\mathcal{B}_{p,p}^{s}(\Omega_{i,e})$-norm.
-Notice that this procedure only defines boundary values for the closures of $\mathcal{D}_{p}^{s}(\Gamma)$ in $\mathcal{B}_{p,p}^{s}(\Omega_{i,e})$, which may not be the full space.
+Let $\mathcal{D}_{p}^{s}(\Gamma)$ be the set of functions in $C_{c}^{\infty}(\mathbb{C})|_{\Gamma}$ such that their harmonic extensions to $\Omega_{i,e}$ belong to $\mathcal{B}_{p,p}^{s}(\Omega_{i,e})$, and we denote the set of such extensions by $\mathcal{D}_{p}^{s}(\Omega_{i,e})$.
+We define the space of boundary values $\mathcal{B}_{p,p}^{s}(\Omega_{i,e}\!\to\!\Gamma)$ as the abstract closure of this space, i.e., the equivalence classes of Cauchy sequences in $\mathcal{D}_{p}^{s}(\Gamma)$, equipped with the $\mathcal{B}_{p,p}^{s}(\Omega_{i,e})$-norm.
+Notice that this procedure only defines boundary values for the closures of $\mathcal{D}_{p}^{s}(\Omega_{i,e})$ in $\mathcal{B}_{p,p}^{s}(\Omega_{i,e})$, which may not be the full space.
 However, the space $\mathcal{B}_{p,p}^{s}(\Omega_{i,e})$ in this case is quite problematic, and the closure of $\mathcal{D}_{p}^{s}(\Omega_{i,e})$ in $\mathcal{B}_{p,p}^{s}(\Omega_{i,e})$ is actually a more robust variant of this space.
-.
-For all these cases we define
+For all these cases, we define
 $\mathcal{B}_{p,p}^{s}(\Gamma)=\mathcal{B}_{p,p}^{s}(\Omega_{i}\to\Gamma)\cap\mathcal{B}_{p,p}^{s}(\Omega_{e}\to\Gamma),$
 equipped with the natural norm.
-One of our main tasks will be to examine when these three spaces coincide.
-The case 2 above is particularly interesting especially since the two harmonic measures may be mutually singular.
-This drawback has been ruled by [36] and [5] by exhibiting a change of variable realizing a "transmission" operator between the spaces on both sides which happens to be an isomorphism, thus showing the equality of the three spaces for all quasicircles.
-When $\Gamma$ is furthermore assumed to be rectifiable we define, following Douglas [14],
-$B_{p,p}^{s}(\Gamma)=\{f\in L^{p}(\Gamma):\;\|f\|_{B_{p,p}^{s}(\Gamma)}^{p}:=\iint_{\Gamma\times\Gamma}\frac{|f(z)-f(\zeta)|^{p}}{|z-\zeta|^{1+ps}}|dz||d\zeta|<\infty\}.$
+The case 2 above is particularly interesting, especially since the two harmonic measures may be mutually singular.
+This drawback has been ruled out by [40] and [5] by exhibiting a change of variables realizing a "transmission" operator between the spaces on both sides which happens to be an isomorphism, thus showing the equality of the three spaces for all quasicircles.
+When $\Gamma$ is furthermore assumed to be rectifiable, we define, following Douglas [14],
+$B_{p,p}^{s}(\Gamma)=\{f\in L^{p}(\Gamma):\;\|f\|_{B_{p,p}^{s}(\Gamma)}^{p}=\iint_{\Gamma\times\Gamma}\frac{|f(z)-f(\zeta)|^{p}}{|z-\zeta|^{1+ps}}|dz||d\zeta|<\infty\}.$
 Here and in what follows, $|dz|$ denotes the arc-length measure.
-If $\Gamma$ is the unit circle the three spaces $\mathcal{B}_{p,p}^{s}(\Omega_{i}\to\Gamma),\mathcal{B}_{p,p}^{s}(\Omega_{e}\to\Gamma),B_{p,p}^{s}(\Gamma)$ coincide (see p.151 in [38]), but they have no reason to coincide in general.
+If $\Gamma$ is the unit circle, the three spaces $\mathcal{B}_{p,p}^{s}(\Omega_{i}\to\Gamma),\mathcal{B}_{p,p}^{s}(\Omega_{e}\to\Gamma),B_{p,p}^{s}(\Gamma)$ coincide (see p.151 in [45]), but they have no reason to coincide in general.
 Finding cases of equality is one of our goals.
@@
 $\exists K>1:\quad\forall z_{1},z_{2}\in\Gamma,\,\min(\mathrm{length}(\gamma_{1}),\mathrm{length}(\gamma_{2}))\leq K|z_{1}-z_{2}|,$
-where $\gamma_{j},\,j=1,2$, are the two subarcs of $\Gamma$ with endpoints $z_{1},z_{2}$. Notice that chord-arc curves are in particular rectifiable quasicircles (but the converse is not true).
+where $\gamma_{j},\,j=1,2$, are the two subarcs of $\Gamma$ with endpoints $z_{1},z_{2}$. Notice that chord-arc curves are, in particular, rectifiable quasicircles (but the converse is not true).
 We will also need the notion of radial-Lipschitz curves.
-These are the curves given in polar coordinates by the equation $z(\theta)=r(\theta)e^{i\theta}$ where $r:\mathbb{R}\to(0,+\infty)$ is $2\pi$-periodic and Lipschitz continuous.
+These are the curves given in polar coordinates by the equation $z(\theta)=r(\theta)e^{i\theta}$ where $r:\mathbb{R}\to(0,+\infty)$ is $2\pi$-periodic and Lipschitz continuous (The norm of Lipschitz curve will be defined in Section 5.).
 The radial-Lipschitz curves are in particular chord-arc.
-In [44] and [45] the present authors have proven
 Theorem 1.1.
-Let $p>1$. If $\Gamma$ is chord-arc then
+Let $p>1$. If $\Gamma$ is a chord-arc, then
 $\mathcal{B}_{p,p}^{1/p}(\Omega_{i}\to\Gamma)=\mathcal{B}_{p,p}^{1/p}(\Omega_{e}\to\Gamma)=B_{p,p}^{1/p}(\Gamma).$ (1.1)
 Conversely, if $\Gamma$ is a rectifiable quasicircle such that (1.1) holds and $p\geq 2$ then $\Gamma$ is chord-arc.
-The methods we used for the proof of this theorem depend heavily on the special features of the case $s=1/p$, namely, conformal invariance and the existence of a transmission operator.
-In this paper our main results in this direction are Theorems 5.5, 5.7, 5.8 combined with the claims in appendix that loosely speaking say that
-Theorem 1.2.
-If $\Gamma$ is a radial-Lipschitz curve and $p=2$, $s\in(0,1)$ then
+The second statement was first proved in [50] and [51] by the present authors.
+For the proof of the first statement, we refer to [27, 29, 28] (see also [41, 40, 43]).
+The special features of the case $s=1/p$ lie in (and the proof of this theorem depends primarily on) conformal invariance of spaces $\mathcal{B}_{p,p}^{1/p}(\Omega_{i,e})$ and the existence of a transmission operator.
+In Section 5 we will prove our main results in this direction (see Theorems 5.5 and 5.7) that loosely speaking say that
+Theorem 1.2 (see Corollary 5.10).
+If $\Gamma$ is a radial-Lipschitz curve with norm $M$ then
 $\mathcal{B}_{p,p}^{s}(\Omega_{i}\!\to\!\Gamma)=\mathcal{B}_{p,p}^{s}(\Omega_{e}\!\to\!\Gamma)=B_{p,p}^{s}(\Gamma).$ (1.2)
-This can be generalized to any $p\in(1,\infty)$ under a slightly different assumption related to $p$.
-We will also prove a result in this regard for general quasicircles. Before stating it we need to introduce some notions:
-Astala [3] has introduced the notion of "Minkowski content" $h(\Gamma)$: the proper definition will be recalled below but let us just mention here that $h(\Gamma)$ is the Minkowski dimension if $\Gamma$ is a self-similar fractal set (as is the Koch snowflake curve for example).
-The Muckenhoupt weights $A_{p}(\mathbb{C}),\,p>1$, are the weights $\omega$ such that the Hardy-Littlewood maximal function is bounded on $L^{p}(\omega,\mathbb{C})$.
+for any $s\in(0,1)$ and $p\in\left(1+\frac{2\arctan M}{\pi},1+\frac{\pi}{2\arctan M}\right)$.
+In particular, this interval converges to $(1,\infty)$ as $M\to 0$ and to a singleton $\{2\}$ as $M\to\infty$.
+In Section 3, we will also prove a result in this regard for general quasicircles. Before stating it, we need to introduce some notions.
+Astala [3] has introduced the notion of "Minkowski content" $h(\Gamma)$: the proper definition will be recalled below but let us just mention here that $h(\Gamma)$ is the Minkowski dimension if $\Gamma$ is a self-similar fractal set (as is the Von Koch snowflake curve, for example), and that chord-arc curves are exactly quasicircles with $h(\Gamma)=1$.
+The Muckenhoupt weights $A_{p},\,p>1$, in $\mathbb{C}$ are the weights $\omega$ such that the Hardy-Littlewood maximal function is bounded in $L^{p}(\omega,\mathbb{C})$.
 Let $\omega$ be such a weight.
-We define the weighted Sobolev space $W^{1,p}(\omega,\mathbb{C})$ as being the space of tempered distributions $f$ such that $f$ is locally integrable and, in the sense of distributions, $\nabla f\in L^{p}(\omega,\mathbb{C})$.
+We define the weighted Sobolev space $W^{1,p}(\omega,\mathbb{C})$ as the space of tempered distributions $f$ such that $f$ is locally integrable and, in the sense of distributions, $\nabla f\in L^{p}(\omega,\mathbb{C})$.
 Theorem 1.3 (see Theorem 3.5).
-If $\Gamma$ is a quasicircle and
-$\frac{1}{p}<1-\frac{1}{p}\leq s<\frac{p+K_{*}-h(\Gamma)}{K_{*}p}$
-where $K_{*}\leq 2$ is a geometric quantity depending only on $\Gamma$ to be defined in section $3$ then
-$\mathcal{B}_{p,p}^{s}(\Omega_{i}\to\Gamma)=\mathcal{B}_{p,p}^{s}(\Omega_{e}\to\Gamma)=W^{1,p}(\omega,\mathbb{C})|_{\Gamma}$ (1.3)
-where $\omega(z)=d(z,\Gamma)^{(1-s)p-1}$.
-Theorem 1.3 allows us to improve theorem 1.2 for certain values of $p,s$:
-Corollary 1.4.
-If $\Gamma$ is chord-arc and $p>2,s\in[1-\frac{1}{p},\frac{p+K_{*}-1}{K_{*}p})$ then (1.3) holds.
-Another corollary is an "almost-Dirichlet principle" generalizing, in dimension $2$ a result of Maz’ya simplified by Mironescu-Russ ([29], [30]).
+Let $\Gamma$ be a quasicircle and $\omega(z)=d(z,\Gamma)^{(1-s)p-1}$. If
+$p>2,\quad\frac{h(\Gamma)}{2}-\frac{h(\Gamma)-1}{p}<s<1-\frac{h(\Gamma)-1}{p}$ (1.3)
+(the latter reduces to $1/2<s<1$ when $\Gamma$ is chord-arc) and
+$\mathcal{B}_{p,p}^{s}(\Omega_{i}\to\Gamma)=\mathcal{B}_{p,p}^{s}(\Omega_{e}\to\Gamma),$ (1.4)
+then
+$\mathcal{B}_{p,p}^{s}(\Omega_{i}\to\Gamma)=\mathcal{B}_{p,p}^{s}(\Omega_{e}\to\Gamma)=W^{1,p}(\omega,\mathbb{C})|_{\Gamma}.$ (1.5)
+We remark here that in this theorem the condition (1.3) is used to guarantee that $W^{1,p}(\omega,\mathbb{C})$ consists of continuous functions so that its restriction on $\Gamma$ makes sense, and the condition (1.3) can also ensure that $\omega$ is an $A_{p}$-weight.
+Examples of pairs $(p,s)$ such that equation (1.4) holds are $(p,1/p)$ and $(\infty,s)$ with $s<1/2$.
+If one could interpolate between these special spaces, one could obtain much more spaces satisfying (1.4).
+We will investigate this possibility in our forthcoming paper.
+Another class of quasicircles satisfying (1.4) is given by Theorem 1.2: it follows that in this case
+$\mathcal{B}_{p,p}^{s}(\Omega_{i}\to\Gamma)=\mathcal{B}_{p,p}^{s}(\Omega_{e}\to\Gamma)=B_{p,p}^{s}(\Gamma)=W^{1,p}(\omega,\mathbb{C})|_{\Gamma}.$
+A corollary of this study is an "almost-Dirichlet principle" generalizing, in dimension $2$ a result of Maz’ya simplified by Mironescu-Russ ([32], [33]).
 We say that the quasidisk $\Omega$ satisfies the almost-Dirichlet principle for $p,s$ if any continuous function $f:\Gamma\to\mathbb{C}$ having a continuous extension $F:\overline{\Omega}\to\mathbb{C}$ which is in $C^{1}(\Omega)$ and satisfies
@@
 where $u$ is the harmonic extension of $f$ in $\Omega$.
-Corollary 1.5.
-Under the assumption of Theorem 1.3 the almost-Dirichlet principle holds for $\mathcal{B}_{p,p}^{s}(\Omega)$.
-One of the tools for proving Theorem 1.3 will be the use of Plemelj-Calderón property which is of independent interest.
-1.3 Plemelj-Calderón Problem
-It is an old problem, given a bounded Jordan curve $\Gamma$ in the plane and a (complex valued) function $f$ defined on $\Gamma$, to find two holomorphic functions $G_{i},G_{e}$ defined on the interior-connected and the exterior-connected components of $\Gamma$ such that
-$f=G_{i}|_{\Gamma}+G_{e}|_{\Gamma},$
-the boundary traces being defined in some sense.
+Corollary 1.4 (see Corollary 5.12).
+If $\Gamma$ is a radial-Lipschitz curve with norm $M$, then $\mathcal{B}_{p,p}^{s}(\Omega)$ has the almost-Dirichlet property if
+$\frac{1}{2}<s<1\quad\text{and}\quad 2<p<1+\frac{\pi}{2\arctan M},$
+that converges to $(2,\infty)$ as $M\to 0$ and to a singleton $\{2\}$ as $M\to\infty$.
+One of the tools for proving Theorem 1.3 will be Plemelj-Calderón property, which is of independent interest.
+1.3 Plemelj-Calderón problem: statement of two more main theorems
+It is an old problem, given a bounded Jordan curve $\Gamma$ in the plane and a (complex valued) function $f$ defined on $\Gamma$, to find two holomorphic functions $G_{i},G_{e}$ defined in the interior and the exterior-connected components $\Omega_{i,e}$ of $\Gamma$ such that
+$f=G_{i}|_{\Gamma}+G_{e}|_{\Gamma}.$
+Here, the boundary traces are defined in some sense.
 This problem was actually raised and solved by Sokhotski in 1873 before being rediscovered by Plemelj as a main ingredient of his attempt to solve Hilbert’s 22th problem in 1908 ([38]).
@@
 $F(z)=\frac{1}{2\pi i}\int_{\Gamma}\frac{f(\zeta)}{\zeta-z}d\zeta,\qquad z\in\mathbb{C}\setminus\Gamma.$
-This defines a function which is holomorphic outside the curve $\Gamma$. Moreover, this function has boundary values on $\Gamma$ from inside and outside, and one has Plemelj formula on $\Gamma$:
+This defines a function that is holomorphic outside the curve $\Gamma$. Moreover, this function has boundary values on $\Gamma$ from inside and outside, and one has Plemelj formulae on $\Gamma$:
 $F_{i}|_{\Gamma}=Tf+1/2f,$
 $F_{e}|_{\Gamma}=Tf-1/2f$
-where the Cauchy integral
+where $F_{i}=F|_{\Omega_{i}}$, $F_{e}=F|_{\Omega_{e}}$ and the Cauchy integral
 $Tf(z)=\frac{1}{2\pi i}\text{p.v.}\int_{\Gamma}\frac{f(\zeta)}{\zeta-z}d\zeta=\frac{1}{2\pi i}\lim\limits_{\varepsilon\to 0}\int_{|\zeta-z|>\varepsilon}\frac{f(\zeta)}{\zeta-z}d\zeta.$
 The problem is then solved with $G_{i}=F_{i}$ and $G_{e}=-F_{e}$.
-Of course the result remains true for much more general curves and functions.
-Subsequent generalizations relax the smoothness requirements on the curve $\Gamma$ and the function $f$: In order for Plemelj formula to hold true one indeed only needs $\Gamma$ to be rectifiable and $f\in L^{1}(\Gamma)$.
-The modern approach of this problem is then to solve the problem in a given space: Calderón ([6]) asked for instance for which rectifiable curves it is true that for any $f\in L^{2}(\Gamma)$ Plemelj functions $F_{i}|_{\Gamma}$ and $F_{e}|_{\Gamma}$ are also in $L^{2}(\Gamma)$?
+Of course, the result remains true for much more general curves and functions.
+Subsequent generalizations relax the smoothness requirements on the curve $\Gamma$ and the function $f$: In order for Plemelj formulae to hold, one indeed only needs $\Gamma$ to be rectifiable and $f\in L^{1}(\Gamma)$.
+The modern approach of this problem is then to solve the problem in a given space: Calderón ([6]) asked, for instance, for which rectifiable curves it is true that for any $f\in L^{2}(\Gamma)$ Plemelj functions $F_{i}|_{\Gamma}$ and $F_{e}|_{\Gamma}$ are also in $L^{2}(\Gamma)$?
 Equivalently, when is the operator $T$ bounded on $f\in L^{2}(\Gamma)$?
-This problem has been solved by Coifman-McIntosh-Meyer ([10]) for Lipschitz curves (Calderón had previously solved it for curves with small Lipschitz bound ([6])) and then the last word was given by David ([11]) who proved that $T$ is bounded on $f\in L^{2}(\Gamma)$ (or $L^{p}(\Gamma)$, $1<p<\infty$) if and only if $\Gamma$ is Ahlfors-regular, meaning that the length of the part of $\Gamma$ lying in a disk of radius $r$ is no bigger than $Cr$ for some $C$ independent of the chosen disk.
+This problem has been solved by Coifman-McIntosh-Meyer ([10]) for Lipschitz curves (Calderón had previously solved it for curves with small Lipschitz bound ([6])) and then the last word was given by David ([11]) (using the Lipschitz result [10]) who proved that $T$ is bounded on $f\in L^{2}(\Gamma)$ (or $L^{p}(\Gamma)$, $1<p<\infty$) if and only if $\Gamma$ is Ahlfors-regular, meaning that the length of the part of $\Gamma$ lying in a disk of radius $r$ is no bigger than $Cr$ for some $C$ independent of the chosen disk.
 Now Plemelj-Calderón problem may be addressed for curves that are not necessarily rectifiable.
 An example has been given in [54] where Plemelj-Calderón problem for Hölder classes has been raised and solved in some particular cases.
-The idea is to replace the Cauchy integral by the identity
+The idea is to replace the Cauchy integral with the identity
 $f(z)=-\frac{1}{\pi}\iint_{\mathbb{C}}\frac{\bar{\partial}f(\zeta)}{\zeta-z}d\xi d\eta$
-which is true for any test function $f\in C_{c}^{\infty}(\mathbb{C})$, and from which it follows that, in the case of rectifiable boundary,
-$F_{i}(z)$ $=-\frac{1}{\pi}\iint_{\Omega_{e}}\frac{\bar{\partial}f(\zeta)}{\zeta-z}d\xi d\eta,\qquad z\in\Omega_{i};$
-$F_{e}(z)$ $=\frac{1}{\pi}\iint_{\Omega_{i}}\frac{\bar{\partial}f(\zeta)}{\zeta-z}d\xi d\eta,\qquad z\in\Omega_{e}$
-where $\Omega_{i,e}$ stand for the interior and exterior components of $\Gamma$.
-Later Astala ([3]) obtained more precise results for the Hölder classes: these will be discussed in section 3 where we will draw the parallel with fractional Sobolev spaces.
-In this paper we address the following Plemelj-Calderón problem: given $p>1,\,s\in(0,1)$, can one write a function $f\in\mathcal{B}_{p,p}^{s}(\Gamma)$ as a sum of two functions $F_{i,e}$ that are holomorphic in $\Omega_{i,e}$ and whose boundary values belong to $\mathcal{B}_{p,p}^{s}(\Gamma)$?
-The main result of the present paper in this direction is Theorems 2.3, 2.4, 3.5 and Corollaries 3.8, 3.17. A priori version of these results is as follows.
+which is true for any test function $f\in C_{c}^{\infty}(\mathbb{C})$, and from which it follows that, in the case of a rectifiable boundary,
+$F_{i}(z)$ $=-\frac{1}{\pi}\iint_{\Omega_{e}}\frac{\bar{\partial}f(\zeta)}{\zeta-z}d\xi d\eta,\qquad z\in\Omega_{i},$
+$F_{e}(z)$ $=\frac{1}{\pi}\iint_{\Omega_{i}}\frac{\bar{\partial}f(\zeta)}{\zeta-z}d\xi d\eta,\qquad z\in\Omega_{e}.$
+Later, Astala ([3]) obtained more precise results for the Hölder classes: these will be discussed in Section 2 where we will draw a parallel with fractional Besov-Sobolev spaces $\mathcal{B}_{p,p}^{s}(\Gamma)$.
+Recently, Plemelj-Calderón problem for critical Besov spaces (the case of $s=1/p$) has been treated by Liu-Shen [28], and by Matsuzaki [31].
+Schippers-Staubach has further discussed this issue in a broader sense; see [42] and references therein.
+In Section 2, we address the following Plemelj-Calderón problem: given $p>1,\,s\in(0,1)$, can one write a function $f\in\mathcal{B}_{p,p}^{s}(\Gamma)$ as a sum of two functions $F_{i,e}$ that are holomorphic in $\Omega_{i,e}$ and whose boundary values belong to $\mathcal{B}_{p,p}^{s}(\Gamma)$?
+The main result of the present paper in this direction is Theorems 2.3, 2.4, 2.9 and Corollaries 2.13, 3.6 (each focusing on different $(p,s)$-values).
+The a priori version of these results is as follows.
 Theorem 1.5.
-Let $\Gamma$ be a quasicircle , $p>1,\,s\in(0,1)$. Then a priori Plemelj-Calderón property holds for $C_{c}^{\infty}(\mathbb{C})|_{\Gamma}$, if
-$(p-1)(h(\Gamma)-1)<sp<p+1-h(\Gamma).$
-2 Critical Besov Spaces on Quasicircles
-Recall that a quasidisk is the image of the unit disk $\mathbb{D}_{i}$ by a quasiconformal homeomorphism of the plane, and its boundary curve is called a quasicircle.
-An homeomorphism $\Phi$ of the plane is called quasiconformal if its gradient in the sense of distribution is in $L^{2}_{loc}(\mathbb{C})$ and if in addition there exists a constant $k<1$ such that
+Suppose that $\Gamma$ is a quasicircle. For $p>1,\,s\in(0,1)$, the a priori Plemelj-Calderón property holds for $C_{c}^{\infty}(\mathbb{C})|_{\Gamma}\cap\mathcal{B}_{p,p}^{s}(\Gamma)$ if
+$(p-1)(h(\Gamma)-1)<sp<p+1-h(\Gamma)$
+that reduces to $0<s<1$ when $\Gamma$ is chord-arc.
+Assume that the curve $\Gamma$ is rectifiable so that the space $B_{p,p}^{s}(\Gamma)$ is well-defined.
+We will see that Plemelj-Calderón problem on $B_{p,p}^{s}(\Gamma)$ is much easier to handle: In Section 4, we will show the following:
+Theorem 1.6 (see Theorem 4.4).
+Suppose that the curve $\Gamma$ is chord-arc. For $0<s<1$ and $1<p<\infty$, the Cauchy integral operator $T$ is bounded on $B_{p,p}^{s}(\Gamma)$.
+2 Plemelj-Calderón problem for $\mathcal{B}_{p,p}^{s}(\Gamma)$ on quasicircles: the non-rectifiable case
+The main goal of this section is to address Plemelj-Calderón Problem for the space $\mathcal{B}_{p,p}^{s}(\Gamma)$, $p>1$ and $0<s<1$, on quasicircles $\Gamma$.
+We now introduce several notions closely related to quasicircles which will be used later.
+Recall that a quasidisk is the image of the unit disk $\mathbb{D}_{i}$ under a quasiconformal homeomorphism of the plane, and its boundary curve is called a quasicircle.
+An homeomorphism $\Phi$ of the plane is called quasiconformal if its gradient in the sense of distribution is in $L^{2}_{loc}(\mathbb{C})$ and if, in addition, there exists a constant $k<1$ such that
 $\frac{\partial\Phi}{\partial\bar{z}}=\mu(z)\frac{\partial\Phi}{\partial z}$
@@
 If we put $K=\frac{1+k}{1-k}$ then $K$ represents the maximal eccentricity of infinitesimal images of circles by the quasiconformal mapping.
-Any quasicircle $\Gamma$ admits a quasiconformal reflection across $\Gamma$ which is bi-Lipschitz around $\Gamma$.
+Any quasicircle $\Gamma$ admits a quasiconformal reflection across $\Gamma$ that is bi-Lipschitz around $\Gamma$.
 Here, by a quasiconformal reflection we mean an anti-quasiconformal self-homeomorphism $R$ of $\overline{\mathbb{C}}$ such that $R\circ R=\text{id}$ on $\overline{\mathbb{C}}$ and $R=\text{id}$ when restricted to $\Gamma$.
-Precisely, let $\Phi$ be a quasiconformal homeomorphism of $\overline{\mathbb{C}}$ we then have $R(z)=\Phi\circ(1/\overline{\Phi^{-1}(z)})$ is a quasiconformal reflection with respect to the quasicircle $\Gamma:=\partial\Phi(\mathbb{D}_{i})$.
+Precisely, let $\Phi$ be a quasiconformal homeomorphism of $\overline{\mathbb{C}}$ we then have $R(z)=\Phi\circ(1/\overline{\Phi^{-1}(z)})$ is a quasiconformal reflection with respect to the quasicircle $\Gamma=\partial\Phi(\mathbb{D}_{i})$.
 Let $\Omega_{i}$ and $\Omega_{e}$ be domains bounded by a quasicircle $\Gamma$.
-If $\varphi_{i}$ and $\varphi_{e}$ are conformal isomorphisms from $\mathbb{D}_{i}$ onto $\Omega_{i}$ and from $\mathbb{D}_{e}$ onto $\Omega_{e}$, respectively, then they can be extended to a homeomorphism between the closures such that $h:=\varphi_{e}^{-1}\circ\varphi_{i}$ is a homeomorphism of $\mathbb{T}$ which is quasisymmetric in the sense that there exists a constant $C>0$ such that
+If $\varphi_{i}$ and $\varphi_{e}$ are conformal isomorphisms from $\mathbb{D}_{i}$ onto $\Omega_{i}$ and from $\mathbb{D}_{e}$ onto $\Omega_{e}$, respectively, then they can be extended to a homeomorphism between the closures such that $h=\varphi_{e}^{-1}\circ\varphi_{i}$ is a homeomorphism of $\mathbb{T}$ that is quasisymmetric in the sense that there exists a constant $C>0$ such that
 $C^{-1}\leq\frac{|h(e^{i(t+\alpha)})-h(e^{it})|}{|h(e^{it})-h(e^{i(t-\alpha)})|}\leq C$
 for every $t\in\mathbb{R}$ and $\pi/2<\alpha\leq\pi/2$. Moreover, every quasisymmetry arises in this way from some quasicircle. For these results on quasiconformal theory, we refer to [1] for details.
+2.1 Boundary values of functions in a dense subspace of $\mathcal{B}_{p,p}^{1/p}(\Omega)$
 Let $\Gamma$ be a Jordan curve and $p>1$.
-We recall that the spaces $\mathcal{B}_{p,p}^{1/p}(\Omega_{i,e})$ , called the $p$-critical Besov spaces, have the remarkable property of being conformally invariant: if $\Omega,\Omega^{\prime}$ are two Jordan domains and $\varphi:\Omega\to\Omega^{\prime}$ is a biholomorphism then the map $f\mapsto f\circ\varphi$ is a quasi-isometry between $\mathcal{B}_{p,p}^{1/p}(\Omega^{\prime})$ and $\mathcal{B}_{p,p}^{1/p}(\Omega)$.
-In order to be as much as possible self-contained, here is a sketch of proof: by the change of variable $\zeta=\varphi(z)$ we have
+We recall that the spaces $\mathcal{B}_{p,p}^{1/p}(\Omega_{i,e})$ , called the $p$-critical Besov spaces, have the remarkable property of being conformally invariant: if $\Omega,\Omega^{\prime}$ are two Jordan domains and $\varphi:\Omega\to\Omega^{\prime}$ is a biholomorphism, then the map $u\mapsto u\circ\varphi$ is a quasi-isometry between $\mathcal{B}_{p,p}^{1/p}(\Omega^{\prime})$ and $\mathcal{B}_{p,p}^{1/p}(\Omega)$.
+In order to be as self-contained as possible, here is a sketch of the proof: by the change of variables $\zeta=\varphi(z)$ we have
 $\iint_{\Omega^{\prime}}|\nabla u(\zeta)|^{p}d(\zeta,\Gamma^{\prime})^{p-2}dxdy=\iint_{\Omega}|\nabla u(\varphi(z))|^{p}d(\varphi(z),\Gamma^{\prime})^{p-2}|\varphi^{\prime}(z)|^{2}dxdy.$
-Now the Koebe distortion theorem [1] implies that
+Now, the Koebe distortion theorem (see [1]) implies that
 $\forall z\in\Omega,\quad\frac{1}{4}\leq\frac{d(\varphi(z),\Gamma^{\prime})}{d(z,\Gamma)|\varphi^{\prime}(z)|}\leq 4,$ (2.1)
 and the result follows.
-Applying this result for $\Omega^{\prime}=\mathbb{D}$, $\partial\mathbb{D}$ being the unit circle $\mathbb{T}$, we see that all the $p$-critical Besov spaces over $\Omega$ are quasi-isometric to $\mathcal{B}_{p,p}^{1/p}(\mathbb{D})$.
+Applying this result to $\Omega^{\prime}=\mathbb{D}$, the inner or outer disk bounded by the unit circle $\mathbb{T}$, we see that all the $p$-critical Besov spaces over $\Omega$ are quasi-isometric to $\mathcal{B}_{p,p}^{1/p}(\mathbb{D})$.
 It is a classical result [44] that the space $\mathcal{C}_{p}(\mathbb{D})$ consisting of functions in $\mathcal{B}_{p,p}^{1/p}(\mathbb{D})$ that have a continuous extension to $\overline{\mathbb{D}}$ is dense in $\mathcal{B}_{p,p}^{1/p}(\mathbb{D})$.
-This can also be derived from that the set of polynomials is dense in the Bergman space $A^{p}_{p-2}$, the set of analytic functions $g$ in $\mathbb{D}$ such that
+This can also be derived from the fact that the set of polynomials is dense in the Bergman space $A^{p}_{p-2}$, the set of analytic functions $g$ in $\mathbb{D}$ such that
 $\iint_{\mathbb{D}}|g(z)|^{p}|1-|z|^{2}|^{p-2}dxdy<\infty.$
-Since a bi-holomorphic homeomorphism between two Jordan domains extends to a homeomorphism between the closures the same property will hold for the analogous spaces $\mathcal{C}_{p}(\Omega)$.
-The next step is to prove that $\mathcal{C}_{p}(\Omega_{i})|_{\Gamma}=\mathcal{C}_{p}(\Omega_{e})|_{\Gamma}$ for any quasicircle $\Gamma$. We first prove the result for $p=2$.
+Since a bi-holomorphic homeomorphism between two Jordan domains extends to a homeomorphism between the closures, the same property will hold for the analogous space $\mathcal{C}_{p}(\Omega)$.
+In this subsection, we prove that $\mathcal{C}_{p}(\Omega_{i})|_{\Gamma}=\mathcal{C}_{p}(\Omega_{e})|_{\Gamma}$ for any quasicircle $\Gamma$. To proceed, we first address the case of $p=2$.
 Let $u_{i}\in\mathcal{C}_{2}(\Omega_{i})$ and $f=u_{i}|_{\Gamma}$.
@@
 Put $v(z)=u_{i}(R(z)),z\in\Omega_{e}$.
-By the bi-Lipschitz property of $R$ around $\Gamma$ we have that
+By the bi-Lipschitz property of $R$ around $\Gamma$ we have the following
 $\iint_{\Omega_{e}}|\nabla v(z)|^{2}dxdy<\infty$
-(stricly speaking, there is a technical problem at $\infty$, which is settled in the appendix). We invoke then the Dirichlet principle which implies that
+(strictly speaking, there is a technical problem at $\infty$, which is settled in the appendix). We then invoke the Dirichlet principle which implies that
 $\iint_{\Omega_{e}}|\nabla u_{e}(z)|^{2}dxdy\leq\iint_{\Omega_{e}}|\nabla v(z)|^{2}dxdy<+\infty,$
-where $u_{e}:=P_{e}(f)$ stands for the harmonic extension of $f$ in $\Omega_{e}$. That is, we have that the transmission operator
+where $u_{e}=P_{e}(f)$ stands for the harmonic extension of $f$ in $\Omega_{e}$. That is, we have that the transmission operator
 $\mathcal{C}_{2}(\Omega_{i})\ni u_{i}\mapsto u_{e}\in\mathcal{C}_{2}(\Omega_{e})$
 is bounded. Actually, it is a bounded isomorphism since the boundedness of the inverse operator can be seen by exchanging the roles of $\Omega_{i}$ and $\Omega_{e}$, which proves the result.
-If $p>1$ and $0<s<1$ (except the case of $p=2$ and $s=1/2$) we loose the Dirichlet principle and use instead the almost-Dirichlet principle.
-Recall that the almost-Dirichlet principle holds for $\mathcal{B}_{p,p}^{s}(\Omega)$ if for any $f$ defined on $\Gamma$, $u$ denotes its harmonic extension in $\Omega$ and $v$ is any continuous extension of $f$ to $\Omega$ which is in $C^{1}(\Omega)$, we have
+For the case of a general $p\neq 2$ we loose the Dirichlet principle and use instead the almost-Dirichlet principle.
+Recall that the almost-Dirichlet principle is said to be valid for $\mathcal{B}_{p,p}^{s}(\Omega)$, p>1, 0<s<1p>1,\,0<s<1, if for any $f$ defined on $\Gamma$, $u$ denotes its harmonic extension in $\Omega$ and $v$ is any continuous extension of $f$ to $\Omega$ which is in $C^{1}(\Omega)$, we have
 $\iint_{\Omega}|\nabla u(z)|^{p}d(z,\Gamma)^{(1-s)p-1}dxdy\leq C\iint_{\Omega}|\nabla v(z)|^{p}d(z,\Gamma)^{(1-s)p-1}dxdy$
-for some constant $C$ depending only on $\Omega,\,p,\,s$. It was proved in [30] that the almost-Dirichlet principle holds for $\mathcal{B}_{p,p}^{s}(\mathbb{D}_{i,e})$ for all p>1, 0<s<1p>1,\,0<s<1.
+for some constant $C$ depending only on $\Omega,\,p,\,s$. In [33] it was proved that the almost-Dirichlet principle holds for $\mathcal{B}_{p,p}^{s}(\mathbb{D}_{i,e})$ for all p>1, 0<s<1p>1,\,0<s<1.
 Lemma 2.1.
@@
 As a consequence, if $f$ is a continuous function on $\Gamma$, its harmonic extension to $\Omega_{i}$ is in $\mathcal{C}_{p}(\Omega_{i})$ if and only if its harmonic extension to $\Omega_{e}$ is in $\mathcal{C}_{p}(\Omega_{e})$.
-Then, $C_{p}(\Omega_{i}\!\to\!\Gamma)=C_{p}(\Omega_{e}\!\to\!\Gamma)$ so that we may define $\mathcal{C}_{p}(\Gamma):=\mathcal{C}_{p}(\Omega_{i,e}\!\to\!\Gamma)$.
+Then, $C_{p}(\Omega_{i}\!\to\!\Gamma)=C_{p}(\Omega_{e}\!\to\!\Gamma)$ so that we may define $\mathcal{C}_{p}(\Gamma)=\mathcal{C}_{p}(\Omega_{i,e}\!\to\!\Gamma)$.
 Let $\varphi_{i,e}$ denote the Riemann mappings from $\mathbb{D}_{i,e}$ onto $\Omega_{i,e}$ as before.
-By the conformal invariance, we have that $f\circ\varphi_{i,e}\in\mathcal{C}_{p}(\mathbb{T})$ and $f\circ\varphi_{e}=f\circ\varphi_{i}\circ h$ with $h:=\varphi_{i}^{-1}\circ\varphi_{e}$ being a circle homeomorphism called conformal welding.
-As we have recalled above this homeomorphism is a quasisymmetry: we deduce from this that $V_{h}:g\mapsto g\circ h$ is an isomorphism of $\mathcal{C}_{p}(\mathbb{T})$ which extends to an isomorphism of $\mathcal{B}_{p,p}^{1/p}(\mathbb{T})$ with $V_{h}^{-1}=V_{h^{-1}}$.
-Now if one consider any quasisymmetry $h$ of the unit circle we know [1] that $h$ is the conformal welding of some quasicircle.
+By conformal invariance, we have $f\circ\varphi_{i,e}\in\mathcal{C}_{p}(\mathbb{T})$ and $f\circ\varphi_{e}=f\circ\varphi_{i}\circ h$ with $h=\varphi_{i}^{-1}\circ\varphi_{e}$ being a circle homeomorphism called conformal welding.
+As we have recalled above, this homeomorphism is a quasisymmetry: we deduce from this that $V_{h}:g\mapsto g\circ h$ is an isomorphism of $\mathcal{C}_{p}(\mathbb{T})$ that extends to an isomorphism of $\mathcal{B}_{p,p}^{1/p}(\mathbb{T})$ with $V_{h}^{-1}=V_{h^{-1}}$.
+Now, if one considers any quasisymmetry $h$ of the unit circle, we know [1] that $h$ is the conformal welding of some quasicircle.
 We deduce from this geometric approach that $V_{h}$ is an isomorphism of $\mathcal{B}_{p,p}^{1/p}(\mathbb{T})$ for any quasisymmetry $h$.
-Conversely, Nag-Sullivan [33] for the case $p=2$ and Bourdaud [5] for the general case have proven that if $h$ is a circle homeomorphism and if $V_{h}$ is an isomorphism of $\mathcal{B}_{p,p}^{1/p}(\mathbb{T})$ then $h$ must be a quasisymmetry.
+Conversely, Nag-Sullivan [37] for the case $p=2$ and Bourdaud [5] for general case have proven that if $h$ is a circle homeomorphism and if $V_{h}$ is an isomorphism of $\mathcal{B}_{p,p}^{1/p}(\mathbb{T})$ then $h$ must be a quasisymmetry.
 Vodop’Yanov[48], and later Bourdaud [5] moreover proved that if $s\neq 1/p$ then the circle homeomorphisms $h$ such that $V_{h}$ operates on $\mathcal{B}_{p,p}^{s}(\mathbb{T})$ isomorphically are the bi-Lipschitz ones which form a proper subgroup of the group of quasisymmetries.
-2.1 Boundary Values of Functions $\mathcal{B}_{p,p}^{1/p}(\Omega)$
-When dealing with a function $u$ in $\mathcal{B}_{p,p}^{1/p}(\Omega_{i})$ that is continuous up to the boundary the boundary values are obvious: it is just the restriction to the boundary, and moreover we have just seen that if we take the harmonic extension of this boundary value in $\Omega_{e}$ we get a function $v\in\mathcal{B}_{p,p}^{1/p}(\Omega_{e})$ with a norm equivalent to the one we started with.
+2.2 Boundary values of functions in $\mathcal{B}_{p,p}^{1/p}(\Omega)$
+When dealing with a function $u$ in $\mathcal{B}_{p,p}^{1/p}(\Omega_{i})$ that is continuous up to the boundary, the boundary values are obvious: it is just the restriction to the boundary, and moreover we have just seen that if we take the harmonic extension of this boundary value in $\Omega_{e}$ we get a function $v\in\mathcal{B}_{p,p}^{1/p}(\Omega_{e})$ with a norm equivalent to the one we started with.
 Now let us consider a general function $u\in\mathcal{B}_{p,p}^{1/p}(\Omega_{i})$: as we have seen, this function is the limit of a sequence $(u_{n})\in\mathcal{C}_{p}(\Omega_{i})$ and we put $f_{n}$ to be the boundary value of $u_{n}$.
 The harmonic extension to $\Omega_{e}$ of $f_{n}$ is then a Cauchy sequence in $\mathcal{B}_{p,p}^{1/p}(\Omega_{e})$, thus convergent to some $v\in\mathcal{B}_{p,p}^{1/p}(\Omega_{e})$ and we may regard the couple $(u,v)$ as an abstract version of the boundary value of $u$, and also $v$.
-If we want to be more concrete, we use the Riemann maps $\varphi_{i,e}$: $u\circ\varphi_{i}\in\mathcal{B}_{p,p}^{1/p}(\mathbb{D}_{i})\subset h^{p}(\mathbb{D})$, the classical harmonic Hardy space of the disk, so it has radial boundary values almost everywhere on the circle defining a function $b_{i}\in L^{p}(\mathbb{T})$.
+The set of such boundary functions is denoted by $\mathcal{B}_{p,p}^{1/p}(\Gamma)$.
+If we want to be concrete, we use the Riemann maps $\varphi_{i,e}$: $u\circ\varphi_{i}\in\mathcal{B}_{p,p}^{1/p}(\mathbb{D}_{i})\subset h^{p}(\mathbb{D})$, the classical harmonic Hardy space of the disk, so it has radial boundary values almost everywhere on the circle defining a function $b_{i}\in L^{p}(\mathbb{T})$.
 Notice that this fact translates in $\Omega_{i}$ by saying that $u$ has limits along $\omega_{i}$-almost every internal ray, where $\omega_{i}$ stands for harmonic measure in $\Omega_{i}$.
@@
 When $u\in\mathcal{C}_{p}(\Omega_{i})$ we have seen that $b_{e}=V_{h}(b_{i})$ where $h$ is the conformal welding of $\Gamma$.
-Unfortunately it does not make sense in $L^{p}(\mathbb{T})$ since $h$ need not be absolutely continuous with respect to the Lebesgue measure, a fact which transfers geometrically to the fact that $\omega_{i}$ and $\omega_{e}$ need not be mutually absolutely continuous.
+Unfortunately, it does not make sense in $L^{p}(\mathbb{T})$ since $h$ need not be absolutely continuous with respect to the Lebesgue measure, a fact that geometrically transfers to the fact that $\omega_{i}$ and $\omega_{e}$ need not be mutually absolutely continuous.
 Nevertheless, the relation $b_{e}=V_{h}(b_{i})$ remains true if we interpret $V_{h}$ on $\mathcal{B}_{p,p}^{1/p}(\mathbb{T})$ as the completion of $V_{h}|_{\mathcal{C}_{p}^{1/p}(\mathbb{T})}$.
-In the case of $p=2$, Schippers-Staubach [36] has given an even more concrete description of the boundary values: we outline their argument.
+In the case of $p=2$, Schippers-Staubach ([40], [43]) has given an even more concrete description of the boundary values: we outline their argument.
 On the one hand, for any $u\in\mathcal{B}_{2,2}^{1/2}(\Omega_{i})$ we have $u\circ\varphi_{i}\in\mathcal{B}_{2,2}^{1/2}(\mathbb{D}_{i})$ as before.
-In this case, $u\circ\varphi_{i}$ has radial boundary values everywhere on the circle $\mathbb{T}$ except on a Borel set $F_{1}$ of Logarithmic capacity zero, defining a function $b_{i}\in L^{2}(\mathbb{T})$.
-On the other hand, let $h:=\varphi_{i}^{-1}\circ\varphi_{e}$, the conformal welding with respect to the quasicircle $\Gamma$, that is a quasisymmetry (so is $h^{-1}$) as we pointed out above.
+In this case, $u\circ\varphi_{i}$ has radial boundary values everywhere on the circle $\mathbb{T}$ except for a Borel set $F_{1}$ of Logarithmic capacity zero, defining a function $b_{i}\in L^{2}(\mathbb{T})$.
+On the other hand, let $h=\varphi_{i}^{-1}\circ\varphi_{e}$, the conformal welding with respect to the quasicircle $\Gamma$, that is a quasisymmetry (so is $h^{-1}$) as we pointed out above.
 Since a quasisymmetry takes a Borel set of Logarithmic capacity zero to a Borel set of Logarithmic capacity zero (see [2]) we see that $F_{2}:=h^{-1}(F_{1})$ is also a Borel set of Logarithmic capacity zero.
 Then $b_{e}=b_{i}\circ\varphi_{i}^{-1}\circ\varphi_{e}=V_{h}(b_{i})$ is well-defined on $\mathbb{T}\setminus F_{2}$ and belongs to $L^{2}(\mathbb{T})$.
-It is known that the union $F_{1}\cup F_{2}$ is also of Logarithmic capacity zero, and in particular, Lebesgue measure zero since a set of Logarithmic capacity zero has zero Lebesgue measure.
-Put $P_{e}(b_{e})$ to be the harmonic extension in $\mathbb{D}_{e}$ of $b_{e}$, and $v:=P_{e}(b_{e})\circ\varphi_{e}^{-1}$ is the harmonic extension in $\Omega_{e}$ of $b_{i}\circ\varphi_{i}^{-1}$.
+It is known that the union $F_{1}\cup F_{2}$ is also of Logarithmic capacity zero, and, in particular, Lebesgue measure zero since a set of Logarithmic capacity zero has zero Lebesgue measure.
+Put $P_{e}(b_{e})$ as the harmonic extension in $\mathbb{D}_{e}$ of $b_{e}$, and $v:=P_{e}(b_{e})\circ\varphi_{e}^{-1}$ is the harmonic extension in $\Omega_{e}$ of $b_{i}\circ\varphi_{i}^{-1}$.
 We can see that $u$ and $v$ have the same boundary values along every internal ray on $\Gamma$, except on a set whose harmonic measure is zero with respect to both $\Omega_{i}$ and $\Omega_{e}$.
-Moreover, since the spaces $\mathcal{B}_{p,p}^{1/p}(\Omega_{i,e})$ increase with $p$ the same argument also holds for the case of $1<p<2$.
+Moreover, since spaces $\mathcal{B}_{p,p}^{1/p}(\Omega_{i,e})$ increase with $p$, the same argument holds for the case of $1<p<2$.
 2.3 Plemelj-Calderón problem for $\mathcal{B}_{2,2}^{1/2}(\Gamma)$
 We can now address Plemelj-Calderón problem for $\mathcal{B}_{2,2}^{1/2}(\Gamma)$.
-The argument is broken into two steps.
-In the first we consider the function in the space $D(\Gamma)$, the restriction space to $\Gamma$ of $C_{c}^{\infty}(\mathbb{C})$.
+The main result of this subsection is Theorem 2.3, which has been proven in the literature; see [40], [29].
+However, the proof we present below is also worth recording, and will be used further in later subsections.
+The argument is broken down into two steps.
+In the first step, we consider the function in the space $C_{c}^{\infty}(\mathbb{C})|_{\Gamma}$, the restriction space to $\Gamma$ of $C_{c}^{\infty}(\mathbb{C})$.
 The second step consists in an approximation process.
-For any test function $f$, it holds that
+For any test function $f$, we have the following.
 $f(z)=-\frac{1}{\pi}\iint_{\mathbb{C}}\frac{\bar{\partial}f(\zeta)}{\zeta-z}d\xi d\eta,\qquad z\in\mathbb{C}.$ (2.4)
@@
 $\widehat{F_{e}}(z)=\frac{1}{\pi}\iint_{\Omega_{i}}\frac{\bar{\partial}f(\zeta)}{\zeta-z}d\xi d\eta,\qquad z\in\Omega_{e}.$ (2.6)
-It is easily seen that $\widehat{F_{i}}$ and $\widehat{F_{e}}$ are holomorphic functions in $\Omega_{i}$ and $\Omega_{e}$, respectively with $\widehat{F_{e}}(z)=O(\frac{1}{|z|})$ at $\infty$.
-Furthermore, it can be shown that both $\widehat{F_{i}}$ and $\widehat{F_{e}}$ define continuous functions at each $z\in\mathbb{C}$.
+It is easily seen that $\widehat{F_{i}}$ and $\widehat{F_{e}}$ are holomorphic functions in $\Omega_{i}$ and $\Omega_{e}$ respectively with $\widehat{F_{e}}(z)=O(\frac{1}{|z|})$ at $\infty$.
+Furthermore, it can be shown that both $\widehat{F_{i}}$ and $\widehat{F_{e}}$ define continuous functions in $\mathbb{C}$.
 It follows from (2.4) that on $\Gamma$,
 $f=\widehat{F_{i}}-\widehat{F_{e}}$. (2.7)
-Here, we have used that the quasicircle $\Gamma$ has zero area with respect to the two-dimensional Lebesgue measure. Taking derivative we have that
+Here, we have used the fact that the quasicircle $\Gamma$ has zero area with respect to the two-dimensional Lebesgue measure. Taking the derivative on both sides of (2.5) we get
 $\widehat{F_{i}}^{\prime}(z)=-\frac{1}{\pi}\iint_{\Omega_{e}}\frac{\bar{\partial}f(\zeta)}{(\zeta-z)^{2}}d\xi d\eta=B(\bar{\partial}f\chi_{\Omega_{e}})(z).$
 Here, $\chi_{\Omega_{e}}$ is the characteristic function of the domain $\Omega_{e}$ and $B$ denotes the Beurling operator, i.e., the convolution with $\frac{1}{\pi}\text{p.v.}\frac{1}{z^{2}}$ which is an isometry of $L^{2}(\mathbb{C})$.
-Then,
+Then
 $\|\widehat{F_{i}}\|_{\mathcal{B}_{2,2}^{1/2}(\Omega_{i})}$ $=\|\widehat{F_{i}}^{\prime}\|_{L^{2}(\Omega_{i})}=\|B(\bar{\partial}f\chi_{\Omega_{e}})\|_{L^{2}(\Omega_{i})}$ (2.8)
-$\leq\|B(\bar{\partial}f\chi_{\Omega_{e}})\|_{L^{2}(\mathbb{C})}\leq\|\bar{\partial}f\|_{L^{2}(\Omega_{e})}\leq\|\bar{\partial}f\|_{L^{2}(\mathbb{C})}.$
-Similarly, it follows from (2.6) that
+$\leq\|B(\bar{\partial}f\chi_{\Omega_{e}})\|_{L^{2}(\mathbb{C})}=\|\bar{\partial}f\|_{L^{2}(\Omega_{e})}\leq\|\bar{\partial}f\|_{L^{2}(\mathbb{C})}.$
+Similarly, from (2.6) it follows that
 $\|\widehat{F_{e}}\|_{\mathcal{B}_{2,2}^{1/2}(\Omega_{e})}\leq\|\bar{\partial}f\|_{L^{2}(\mathbb{C})}.$ (2.9)
 Lemma 2.2.
-For any $f\in C_{c}^{\infty}(\mathbb{C})$, the integrals in (2.5) and (2.6) depend only on the boundary values $f$ on $\Gamma$, not on the specific extensions belonging to $B_{2,2}^{1/2}(\Omega_{i,e})$ of $f$ to respectively $\Omega_{i}$ and $\Omega_{e}$.
+For any $f\in C_{c}^{\infty}(\mathbb{C})$, the integrals in (2.5) and (2.6) depend only on the boundary values $f$ on the quasicircle $\Gamma$, not on the specific extensions belonging to $B_{2,2}^{1/2}(\Omega_{i,e})$ of $f$ to $\Omega_{i}$ and $\Omega_{e}$.
 In particular, $\widehat{F_{i}}$ and $\widehat{F_{e}}$ defined in respectively (2.5) and (2.6) equal to
@@
 $f=F_{i}-F_{e}$ (2.12)
-on $\Gamma$. Here, $u_{i}$ and $u_{e}$ are harmonic extensions of $f$ to respectively $\Omega_{i}$ and $\Omega_{e}$.
+on $\Gamma$. Here, $u_{i}$ and $u_{e}$ are harmonic extensions of $f$ to $\Omega_{i}$ and $\Omega_{e}$.
 Proof.
-Note that $u_{i}$ and $u_{e}$ are continuous on $\overline{\Omega}_{i}$ and $\overline{\Omega}_{e}$, respectively since the boundary function $f$ is continuous on $\Gamma$.
+Note that $u_{i}$ and $u_{e}$ are continuous on $\overline{\Omega}_{i}$ and $\overline{\Omega}_{e}$, respectively, since the boundary function $f$ is continuous on $\Gamma$.
 By the Dirichlet principle, $u_{i,e}\in\mathcal{B}_{2,2}^{1/2}(\Omega_{i,e})$.
-Using the similar computation we can see that the holomorphic functions $F_{i}\in\mathcal{B}_{2,2}^{1/2}(\Omega_{i})$ and $F_{e}\in\mathcal{B}_{2,2}^{1/2}(\Omega_{e})$.
-By the theorem of Gol’dshtein-Latfullin-Vodop’yanov [18] (see also [22]) $u_{i}$ can be extended to $U_{i}\in W^{1,2}(\mathbb{C})$, defining $U_{i}=u_{i}\circ R$ on $\Omega_{e}$ where $R$ is a quasiconformal reflection across $\Gamma$.
-There exists a sequence $(U_{i,n})$ of $C_{c}^{\infty}(\mathbb{C})$ converging to $U_{i}$ in $W^{1,2}(\mathbb{C})$ (see e.g.
-Ch.11 in [26]).
+Using a similar computation to that of $\widehat{F_{i}}$ and $\widehat{F_{e}}$, we can see that the holomorphic functions $F_{i}\in\mathcal{B}_{2,2}^{1/2}(\Omega_{i})$ and $F_{e}\in\mathcal{B}_{2,2}^{1/2}(\Omega_{e})$.
+By the theorem of Gol’dshtein-Latfullin-Vodop’yanov [18] (see also [22]) saying that a simply connected domain $\Omega$ satisfies the condition $W^{1,2}(\mathbb{C})|_{\Omega}=W^{1,2}(\Omega)$ if and only if $\partial\Omega$ is a quasicircle, $u_{i}$ can be extended to $U_{i}\in W^{1,2}(\mathbb{C})$ by defining $U_{i}=u_{i}\circ R$ on $\Omega_{e}$ where $R$ is a quasiconformal reflection across $\Gamma$.
+There exists a sequence $(U_{i,n})$ of $C_{c}^{\infty}(\mathbb{C})$ converging to $U_{i}$ in $W^{1,2}(\mathbb{C})$ (see, e.g., Ch.11 in [26]).
 Denote $U_{i,n}|_{\Gamma}$ by $f_{i,n}$.
@@
 $\widehat{F_{e,n}}(z)=\frac{1}{\pi}\iint_{\Omega_{i}}\frac{\bar{\partial}U_{i,n}(\zeta)}{\zeta-z}d\xi d\eta,\qquad z\in\Omega_{e}.$
-Notice that $\widehat{F_{i,n}}$ and $\widehat{F_{e,n}}$ are holomorphic in $\Omega_{i},\,\Omega_{e}$ respectively, continuous in $\mathbb{C}$, and $F_{e,n}(z)=O(\frac{1}{|z|})$ as $z\to\infty$.
-Similar to the reasoning in (2.8) we have that
+Notice that $\widehat{F_{i,n}}$ and $\widehat{F_{e,n}}$ are holomorphic in $\Omega_{i},\,\Omega_{e}$, respectively, continuous in $\mathbb{C}$, and $\widehat{F_{e,n}}(z)=O(\frac{1}{|z|})$ as $z\to\infty$.
+Similarly to the reasoning in (2.8) we get
 $\begin{split}\|\widehat{F_{e,n}}-F_{e}\|_{\mathcal{B}_{2,2}^{1/2}(\Omega_{e})}&=\|(\widehat{F_{e,n}}-F_{e})^{\prime}\|_{L^{2}(\Omega_{e})}\leq\|\bar{\partial}U_{i,n}-\bar{\partial}u_{i})\|_{L^{2}(\Omega_{i})}\\ &\leq\|\nabla(U_{i,n}-u_{i})\|_{L^{2}(\Omega_{i})}\leq\|U_{i,n}-U_{i}\|_{W^{1,2}(\mathbb{C})}\to 0\end{split}$
@@
 $\|\nabla(P_{i}(f_{i,n})-u_{i})\|_{L^{2}(\Omega_{i})}\leq\|\nabla(U_{i,n}-u_{i})\|_{L^{2}(\Omega_{i})}\to 0$
-where $P_{i}(f_{i,n})$ denotes the harmonic extension in $\Omega_{i}$ of $f_{i,n}$. We thus have that $f_{i,n}(z)\to f(z)$ at each point $z\in\Gamma$.
+where $P_{i}(f_{i,n})$ denotes the harmonic extension in $\Omega_{i}$ of $f_{i,n}$. Thus, we have $f_{i,n}(z)\to f(z)$ at each point $z\in\Gamma$.
 […]
@@
 $f=F_{i}-G_{e}.$ (2.14)
-By uniqueness of the decomposition, proved in [27] mainly based on an identity in [39], we see that these two decomposition (2.13),(2.14) are same as (2.7) so that $\widehat{F_{i}}=F_{i}$,  $\widehat{F_{e}}=F_{e}$ and $f=F_{i}-F_{e}$ on $\Gamma$.
+By the uniqueness of the decomposition (see e.g.
+[28]), we see that these two decompositions (2.13),(2.14) are the same as (2.7) so that $\widehat{F_{i}}=F_{i}$,  $\widehat{F_{e}}=F_{e}$ and $f=F_{i}-F_{e}$ on $\Gamma$.
 ∎
-Concerning the above proof, we give two side remarks.
+Concerning the above proof, we make two side remarks.
 Firstly, using a part of the argument in the proof of Lemma 2.2, we can see that $C_{c}^{\infty}(\mathbb{C})|_{\Gamma}$ is a dense subset of $\mathcal{B}_{2,2}^{1/2}(\Gamma)$.
-Precisely, for any $f\in\mathcal{B}_{2,2}^{1/2}(\Gamma)$, its harmonic extension $u_{i}$ in $\Omega_{i}$ can be extended to $U_{i}\in W^{1,2}(\mathbb{C})$.
+Specifically, for any $f\in\mathcal{B}_{2,2}^{1/2}(\Gamma)$, its harmonic extension $u_{i}$ in $\Omega_{i}$ can be extended to $U_{i}\in W^{1,2}(\mathbb{C})$.
 There exists a sequence $(U_{i,n})$ of $C_{c}^{\infty}(\mathbb{C})$ converging to $U_{i}$ in $W^{1,2}(\mathbb{C})$.
 Using the Dirichlet principle, we may conclude the assertion.
-Secondly, we explain that if the Jordan curve $\Gamma$ is rectifiable then the assertion in Lemma 2.2 can be derived easily from the generalized Green’s formula (see p.150 in [28]).
-It says that let $f$ have $L^{1}$-derivatives in the domain $G$ and if $\overline{\Omega}\subset G$ is a Jordan domain with rectifiable boundary $\Gamma=\partial\Omega$ then
+Secondly, we explain that if the quasicircle $\Gamma$ is rectifiable then the assertion in Lemma 2.2 can be derived easily from the generalized Green’s formula (see p.150 in [30]).
+It says that if $f$ has $L^{1}$-derivatives in the domain $G$ and if $\overline{\Omega}\subset G$ is a Jordan domain with rectifiable boundary $\Gamma=\partial\Omega$ then
 $\iint_{\Omega}f_{\bar{z}}dxdy=-\frac{i}{2}\int_{\Gamma}fdz.$
-Let $f\in C_{c}^{\infty}(\mathbb{C})$ and $\Gamma$ be a rectifiable Jordan curve and let $u_{e}$ be the harmonic extension of $f$, restricted on $\Gamma$, to $\Omega_{e}$, that is continuous on $\Omega_{e}\cup\Gamma$.
+Let $f\in C_{c}^{\infty}(\mathbb{C})$ and $\Gamma$ be a rectifiable Jordan curve and let $u_{e}$ be the harmonic extension of $f$, restricted on $\Gamma$, to $\Omega_{e}$, which is continuous in $\Omega_{e}\cup\Gamma$.
 By the Dirichlet principle, $u_{e}\in W^{1,2}(\Omega_{e})$.
-As we noticed before, $u_{e}$ can be extended to $U_{e}\in W^{1,2}(\mathbb{C})$ using the quasiconformal reflection across $\Gamma$.
+As we have noticed before, $u_{e}$ can be extended to $U_{e}\in W^{1,2}(\mathbb{C})$ using the quasiconformal reflection across $\Gamma$.
 Note that $f-U_{e}$ has $L^{2}$-derivatives in $\mathbb{C}$.
@@
 $\iint_{\Omega_{e}}\frac{\bar{\partial}(f-u_{e})(\zeta)}{\zeta-z}d\xi d\eta=-\frac{i}{2}\int_{\Gamma}\frac{(f-u_{e})|_{\Gamma}(\zeta)}{\zeta-z}d\zeta,\quad z\in\Omega_{i}$
-which equals to 0 since $(f-u_{e})|_{\Gamma}=0$ on $\Gamma$. Thus, we see that
+which equals 0 since $(f-u_{e})|_{\Gamma}=0$ on $\Gamma$. Thus, we see that
 $\iint_{\Omega_{e}}\frac{\bar{\partial}f(\zeta)}{\zeta-z}d\xi d\eta=\iint_{\Omega_{e}}\frac{\bar{\partial}u_{e}(\zeta)}{\zeta-z}d\xi d\eta$
@@
 $\begin{split}&\|F_{i}\|_{\mathcal{B}_{2,2}^{1/2}(\Omega_{i})}\leq\|u_{e}\|_{\mathcal{B}_{2,2}^{1/2}(\Omega_{e})}\simeq\|f\|_{\mathcal{B}_{2,2}^{1/2}(\Gamma)},\\ &\|F_{e}\|_{\mathcal{B}_{2,2}^{1/2}(\Omega_{e})}\leq\|u_{i}\|_{\mathcal{B}_{2,2}^{1/2}(\Omega_{i})}\simeq\|f\|_{\mathcal{B}_{2,2}^{1/2}(\Gamma)}.\end{split}$
-Here and in what follows, the notation $A\simeq B$" means there exists a constant $C$ such that $A/C\leq B\leq CA$. In the above the implicit constant depends only on $\Gamma$.
+Here and in what follows, the notation $A\simeq B$" means that there exists a constant $C$ such that $A/C\leq B\leq CA$. In the above, the implicit constant depends only on $\Gamma$.
 Recall that $C_{c}^{\infty}(\mathbb{C})|_{\Gamma}$ is a dense subset of $\mathcal{B}_{2,2}^{1/2}(\Gamma)$.
@@
 The space $\mathcal{B}_{2,2}^{1/2}(\Gamma)$ has two remarkable properties: it is conformally invariant and its definition does not involve the distance to the boundary.
-As a corollary, as we have seen, Plemelj-Calderón property holds for this space for all quasicircles.
-For $p\neq 2$ these properties split: the conformality property remains for the spaces $\mathcal{B}_{p,p}^{1/p}(\Gamma),\,p>1$ but in its definition a distance term $d(z,\Gamma)^{p-2}$ appears.
-This case will be discussed in the next section, together with the general case of $s\in(0,1)$ with a distance term $d(z,\Gamma)^{(1-s)p-1}$.
+As we have seen, Plemelj-Calderón property holds for this space for all quasicircles.
+For $p\neq 2$ these properties split: the conformality property remains for the spaces $\mathcal{B}_{p,p}^{1/p}(\Gamma),\,p>1$, but in its definition a distance term $d(z,\Gamma)^{p-2}$ appears.
+This case will be discussed in the next subsection, together with the general case of $s\in(0,1)$ with a distance term $d(z,\Gamma)^{(1-s)p-1}$.
 The other property transfers to $\mathcal{B}_{p,p}^{1-1/p}(\Gamma)$.
@@
 Let $\Gamma$ be a quasicircle and $p>2$.
-Every $f\in\widetilde{\mathcal{B}}_{p,p}^{1-1/p}(\Gamma)$ admits a unique decomposition $f=F_{i}-F_{e}$ with $F_{i,e}$ being holomorphic, $F_{i,e}\in\mathcal{B}_{p,p}^{1-1/p}(\Omega_{i,e})$ and
+Each $f\in\widetilde{\mathcal{B}}_{p,p}^{1-1/p}(\Gamma)$ admits a unique decomposition $f=F_{i}-F_{e}$ with $F_{i,e}$ being holomorphic, $F_{i,e}\in\mathcal{B}_{p,p}^{1-1/p}(\Omega_{i,e})$ and
 $\|F_{i,e}\|_{\mathcal{B}_{p,p}^{1-1/p}(\Omega_{i,e})}\leq C\|f\|_{\mathcal{B}_{p,p}^{1-1/p}(\Gamma)}$
@@
 This may be seen by two different ways.
-Firstly, s=11/p>1/ps=1-1/p>1/p, we will prove later that functions in $\mathcal{B}_{p,p}^{s}(\Omega_{i,e})$ are $\alpha$-Hölder continuous in the closure of $\Omega_{i,e}$ with $\alpha=s-1/p$.
-The other way to prove the same thing is to appeal to Jones’s theorem: a function in $\mathcal{B}_{p,p}^{1-1/p}(\Omega_{i,e})$ can be extended to a function in $W^{1,p}(\mathbb{C})$ even though clearly one cannot extend in this case by using quasiconformal reflection argument, used in the case of $p=2$ as above (see [22], also see [8]).
+Firstly, s=11/p>1/ps=1-1/p>1/p, we will prove later that the functions in $\mathcal{B}_{p,p}^{s}(\Omega_{i,e})$ are $\alpha$-Hölder continuous in the closure of $\Omega_{i,e}$ with $\alpha=s-1/p$.
+The other way to prove the same thing is to appeal to Jones’ theorem: a function in $\mathcal{B}_{p,p}^{1-1/p}(\Omega_{i,e})$ can be extended to a function in $W^{1,p}(\mathbb{C})$ even though clearly one cannot extend in this case by using quasiconformal reflection argument, used in the case of $p=2$ as above (see [22], also see [8]).
 The Morrey’s inequality states that $W^{1,p}(\mathbb{C})\subset\Lambda^{\alpha}(\mathbb{C})$, $p>2$ (note that this inclusion relation holds in $\mathbb{R}^{n}$ for $p>n$) for $\alpha=1-2/p$, the space of $\alpha$-Hölder continuous functions in $\mathbb{C}$.
-Knowing this fact one may re-define the space $\mathcal{B}_{p,p}^{1-1/p}(\Gamma)$ as the space of $\Lambda^{1-2/p}$-functions on $\Gamma$ whose harmonic extensions to $\Omega_{i,e}$ belong to $\mathcal{B}_{p,p}^{1-1/p}(\Omega_{i,e})$, that is,
+Knowing this fact, one may re-define the space $\mathcal{B}_{p,p}^{1-1/p}(\Gamma)$ as the space of $\Lambda^{1-2/p}$-functions on $\Gamma$ whose harmonic extensions to $\Omega_{i,e}$ belong to $\mathcal{B}_{p,p}^{1-1/p}(\Omega_{i,e})$, that is,
 $\mathcal{B}_{p,p}^{1-1/p}(\Gamma)=\Lambda^{1-2/p}(\Gamma)\cap\mathcal{B}_{p,p}^{1-1/p}(\Omega_{i}\to\Gamma)\cap\mathcal{B}_{p,p}^{1-1/p}(\Omega_{e}\to\Gamma).$
 Proof.
-Thanks to Jones’ theorem the proof is exactly the same as for $p=2$ because the Beurling transform is bounded on $L^{p}(\mathbb{C})$ for all $p>1$.
-Notice that the argument of uniqueness used in the case $p=2$ goes through here because $L^{p}(\Omega_{i})\subset L^{2}(\Omega_{i})$ for $p>2$.
+Thanks to Jones’ theorem, the proof is exactly the same as that for $p=2$ because the Beurling transform is bounded on $L^{p}(\mathbb{C})$ for all $p>1$.
+Notice that the argument of uniqueness used in the case $p=2$ goes through here because $L^{p}(\Omega_{i})\subset L^{2}(\Omega_{i})$ for $p>2$, and $\mathcal{B}_{p,p}^{1-1/p}(\Omega_{e})\subset\mathcal{B}_{2,2}^{1/2}(\Omega_{e})$ by the argument in Appendix.
 ∎
-3 Fractional Sobolev Spaces on Quasidisks
-3.1 Plemelj-Calderón Property for General $\mathcal{B}_{p,p}^{s}(\Gamma)$
-As mentioned in the proceeding section, in order to deal with Plemelj-Calderón property for general $p>1,s\in(0,1)$ we need to deal with boundedness of the Beurling transform on $L^{p}(\omega,\,\mathbb{C})$ where $\omega$ is a weight.
-By a weight we mean a nonnegative locally integrable function on $\mathbb{C}$.
+2.5 Plemelj-Calderón problem for general $\mathcal{B}_{p,p}^{s}(\Gamma)$
+In order to deal with Plemelj-Calderón property for general $p>1,s\in(0,1)$ we need to consider the boundedness of the Beurling transform on $L^{p}(\omega,\,\mathbb{C})$ where $\omega$ is a weight.
+By weight we mean a nonnegative locally integrable function in $\mathbb{C}$.
 Since the Beurling transform is a Calderón-Zygmund operator, we know [9] that this property holds if and only if $\omega$ belongs to the so-called class $A_{p}$ whose definition we now recall.
@@
 $\frac{1}{|D|}\iint_{D}\omega(z)dxdy\left(\frac{1}{|D|}\iint_{D}w(z)^{-\frac{1}{p-1}}dxdy\right)^{p-1}\leq C.$ (2.15)
-For future use let us notice that $A_{p}$ weights on $\mathbb{T}$ are defined similarly.
+For future use, let us note that the $A_{p}$-weights on $\mathbb{T}$ are defined similarly.
 It is known that for all $p>1$, $A_{1}\subset A_{p}\subset A_{\infty}$, $\omega\in A_{p}$ is equivalent to $\omega^{-1/(p-1)}\in A_{p^{\prime}}$ where $1/p+1/p^{\prime}=1$, and $A_{p_{1}}\subset A_{p_{2}}$ for $1<p_{1}<p_{2}$.
-By Hölder inequality the left part of the inequality (3.1) is always $\geq 1$: one can thus interpret this class as verifying some kind of reverse Hölder inequality.
+By Hölder inequality, the left part of the inequality (2.15) is always not less than $1$: thus one can interpret this class as verifying some kind of reverse Hölder inequality.
 If the weight $\omega$ has $A_{p}$ then any Calderon-Zygmund operator is bounded on the weighted space $L^{p}(\omega,\mathbb{C})$ (see [9]).
@@
 $F_{i}(z)=-\frac{1}{\pi}\iint_{\Omega_{e}}\frac{\bar{\partial}f(\zeta)}{\zeta-z}d\xi d\eta,\quad F_{e}(z)=\frac{1}{\pi}\iint_{\Omega_{i}}\frac{\bar{\partial}f(\zeta)}{\zeta-z}d\xi d\eta.$
-Notice $F_{i}^{\prime}(z)=B(\bar{\partial}f\chi_{\Omega_{e}})(z)=B(\bar{\partial}u_{e}\chi_{\Omega_{e}})(z)$, $F_{e}^{\prime}(z)=-B(\bar{\partial}f\chi_{\Omega_{i}})(z)=-B(\bar{\partial}u_{i}\chi_{\Omega_{i}})(z)$ where $u_{i,e}$ are the harmonic extensions of $f|_{\Gamma}$ to $\Omega_{i,e}$, we have that a sufficient condition on $\Gamma$ that implies $F_{i,e}\in\mathcal{B}_{p,p}^{s}(\Omega_{i,e})$ is that the weight $\omega(z)=d(z,\Gamma)^{(1-s)p-1}$ satisfies the $A_{p}$ condition.
-Precisely, under this condition we have
+Notice
+$F_{i}^{\prime}(z)=B(\bar{\partial}f\chi_{\Omega_{e}})(z)=B(\bar{\partial}u_{e}\chi_{\Omega_{e}})(z)$
+and
+$F_{e}^{\prime}(z)=-B(\bar{\partial}f\chi_{\Omega_{i}})(z)=-B(\bar{\partial}u_{i}\chi_{\Omega_{i}})(z)$
+by Lemma 2.2, where $u_{i,e}$ are the harmonic extensions of $f|_{\Gamma}$ to $\Omega_{i,e}$, we see that a sufficient condition on $\Gamma$ which implies $F_{i,e}\in\mathcal{B}_{p,p}^{s}(\Omega_{i,e})$ is that the weight $\omega(z)=d(z,\Gamma)^{(1-s)p-1}$ satisfies the $A_{p}$ condition.
+Precisely, under this condition we have the following.
 $\|F_{i}\|_{\mathcal{B}_{p,p}^{s}(\Omega_{i})}\leq C\|u_{e}\|_{\mathcal{B}_{p,p}^{s}(\Omega_{e})}\leq C\|f\|_{\mathcal{B}_{p,p}^{s}(\Gamma)},$
@@
 where $C$ is an absolute constant.
-The dependence of the condition that $d(z,\Gamma)^{(1-s)p-1}$ having $A_{p}$ on the geometry of the curve has been studied by Astala [3].
-In order to state his results we first need to introduce some notions.
-For a compact set $E\subset\mathbb{C}$ and $0<\delta\leq 2$ let
+The dependence of the condition $d(z,\Gamma)^{(1-s)p-1}$ having $A_{p}$ on the geometry of the curve has been studied by Astala [3].
+In order to state his results, we first need to introduce some notions.
+For a compact set $E\subset\mathbb{C}$ and $0<\delta\leq 2$, let
 $M_{\delta}(E,t)=\frac{|E+D(0,t)|}{t^{2-\delta}}.$
-Here, $E+D(0,t)$, by the definition $\{e+z:\;e\in E,\;z\in D(0,t)\}$, is a set of points at a distance less than $t$ from $E$ which is called Minkowski sausage.
-We then define a kind of Minkowski content by
+Here, $E+D(0,t)$, by the definition $\{e+z:\;e\in E,\;z\in D(0,t)\}$, is a set of points at a distance less than $t$ from $E$, which is called Minkowski sausage.
+(Here and in the sequel we write $|E+D(0,t)|$ for the Lebesgue measure of the set $E+D(0,t)$.) We then define a kind of Minkowski content by
 $h_{\delta}(E)=\sup\limits_{0<t\leq\mathrm{diam}(E)}{M_{\delta}(E,t)}.$
+We mention here that in the definition of $h_{\delta}(E)$, $M_{\delta}(E,t)$ is sometimes replaced in the literature by
+$H_{\delta}(E,t):=\inf\{nt^{\delta}:\;E\subset\cup_{i=1}^{n}D(z_{i},t),\;n\in\mathbb{N}\}.$
+This replacement does not make any difference since there are universal constants $c_{1}$ and $c_{2}$ such that (see [36])
+$c_{1}M_{\delta}(E,t)\leq H_{\delta}(E,t)\leq c_{2}M_{\delta}(E,t).$
 Definition 2.6.
@@
 $h(\Gamma)=\inf\{\delta:\Gamma\,\mathrm{is}\,\delta\text{-}\mathrm{regular}\}\in[1,2)$
-and we may say the degree of regularity of $\Gamma$ is $h(\Gamma)$. For more information on $h(\Gamma)$, and its relation with the Hausdorff dimension $\text{dim}_{H}(\Gamma)$ of $\Gamma$, see [3].
+and we may say that the degree of regularity of $\Gamma$ is $h(\Gamma)$.
+For more information on $h(\Gamma)$, and its relation with the Hausdorff dimension $\text{dim}_{H}(\Gamma)$ of $\Gamma$; see [3].
 Definition 2.7.
@@
 $d(z,\Gamma)^{\alpha-1}\in A_{p}\Leftrightarrow d(z,\Gamma)^{\alpha-1}\in A_{1}\Leftrightarrow\alpha>h(\Gamma)-1.$ (2.16)
-Define $\widetilde{\mathcal{B}}_{p,p}^{s}(\Gamma)$ to be the closure of
+Define $\widetilde{\mathcal{B}}_{p,p}^{s}(\Gamma)$ as the closure of
 $C_{c}^{\infty}(\mathbb{C})|_{\Gamma}\cap\mathcal{B}_{p,p}^{s}(\Omega_{i}\to\Gamma)\cap\mathcal{B}_{p,p}^{s}(\Omega_{e}\to\Gamma)$
-in $\mathcal{B}_{p,p}^{s}(\Gamma).$ Notice that every quasicircle is porous (see [41]). We may now state the principal theorem of this section:
+in $\mathcal{B}_{p,p}^{s}(\Gamma).$ Notice that every quasicircle is porous (see [47]). We may now state the principal theorem of this subsection.
 Figure 1: Domain formed by points $(h(\Gamma),s)$: case p>2 and case p<2
 Theorem 2.9.
-Let $\Gamma$ be a quasicircle and $1<p<\infty,\;0<s<1$. If the point $(h(\Gamma),s)$ locates in the shadowed region (see Figure 1); that is,
-$(h(\Gamma)-1)\frac{p-1}{p}<s<\frac{p+1-h(\Gamma)}{p},$ (3.3)
-then any $f\in\widetilde{\mathcal{B}}_{p,p}^{s}(\Gamma)$ has a decomposition $f=F_{i}-F_{e}$ with $F_{i,e}$ being holomorphic, $F_{i,e}\in\mathcal{B}_{p,p}^{s}(\Omega_{i,e})$ and
+Let $\Gamma$ be a quasicircle and $1<p<\infty,\;0<s<1$. If the point $(h(\Gamma),s)$ is located in the shadowed region (see Figure 1); that is,
+$(h(\Gamma)-1)\frac{p-1}{p}<s<\frac{p+1-h(\Gamma)}{p}$ (2.17)
+(that would be simplified to $0<s<1$ if $h(\Gamma)=1$, i.e.
+$\Gamma$ is a chord-arc curve ), then any $f\in\widetilde{\mathcal{B}}_{p,p}^{s}(\Gamma)$ has a decomposition $f=F_{i}-F_{e}$ with $F_{i,e}$ being holomorphic, $F_{i,e}\in\mathcal{B}_{p,p}^{s}(\Omega_{i,e})$ and
 $\|F_{i,e}\|_{\mathcal{B}_{p,p}^{s}(\Omega_{i,e})}\leq C\|f\|_{\mathcal{B}_{p,p}^{s}(\Gamma)}$ (2.18)
@@
 Proof.
-Following the discussion above, it is clear that (3.4) holds if it holds that the weight $d(z,\Gamma)^{(1-s)p-1}\in A_{p}$.
+Following the discussion above, clearly (2.18) holds if $d(z,\Gamma)^{(1-s)p-1}\in A_{p}$ holds.
 Assume first that $s>\frac{p-1}{p}$ so that (1s)p<1(1-s)p<1.
@@
 Assume now $s<\frac{p-1}{p}$.
-Then, $d(z,\Gamma)^{(1-s)p-1}\in A_{p}\Leftrightarrow d(z,\Gamma)^{\frac{1-(1-s)p}{p-1}}\in A_{p^{\prime}}$.
+Then $d(z,\Gamma)^{(1-s)p-1}\in A_{p}\Leftrightarrow d(z,\Gamma)^{\frac{1-(1-s)p}{p-1}}\in A_{p^{\prime}}$.
 Here, $\frac{1-(1-s)p}{p-1}=\frac{sp}{p-1}-1$ with $\frac{sp}{p-1}<1$.
-By Theorem 3.4 again we have thus that $d(z,\Gamma)^{(1-s)p-1}\in A_{p}\Leftrightarrow s>\frac{p-1}{p}(h-1)$.
-The decomposition can be proved similar to that of Theorem 2.3.
-The case of s=11/ps=1-1/p has been addressed in the last section.
+By Theorem 2.8 again we have thus $d(z,\Gamma)^{(1-s)p-1}\in A_{p}\Leftrightarrow s>\frac{p-1}{p}(h-1)$.
+The decomposition can be shown to be similar to that of Theorem 2.3.
+The case of s=11/ps=1-1/p has been addressed in the last subsection.
 ∎
-We define $\mathcal{A}$ to be the admissible set involved in Theorem 3.5, that is,
+We define $\mathcal{A}$ as the admissible set involved in Theorem 2.9, that is,
 $\mathcal{A}=\{(h(\Gamma),p,s)\in[1,2)\times(1,+\infty)\times(0,1):(h(\Gamma)-1)(p-1)<sp<p+1-h(\Gamma)\},$
@@
 Figure 2: Domain formed by points $(p,s)$: case h>3/2 and case h<3/2
-We end this part with a comment on the triangular domain above the shadowed region in Figure 1.
-Let the curve $\Gamma$ be a self-similar isotropic fractal which satisfies the open set condition.
-Its fractal dimension $d(\Gamma)$ equals to $h(\Gamma)$ in this case (see [32]).
-For instance, the Koch snowflake curve, we mentioned before, is a such quasicircle with $d(\Gamma)=\log 4/\log 3$.
-Denote by $N(\varepsilon)$ the minimal number of disks of radius $\varepsilon$ necessary to cover $\Gamma$.
-Then,
-$d(\Gamma)=\lim_{\varepsilon\to 0}\frac{\log N(\varepsilon)}{\log\frac{1}{\varepsilon}}.$
-Taking $\varepsilon=2^{-(n+1)}$ with sufficiently large $n$ one has the difference between the number $N(\varepsilon)$ and $2^{(n+1)d(\Gamma)}$ approaches 0 as $n$ increases to $\infty$.
-By dividing the domain $\Omega_{i}$ into pieces:
-$\{z\in\Omega_{i}:\;d(z,\Gamma)\geq 2^{-1}\},\quad\{z\in\Omega_{i}:\;2^{-(j+1)}\leq d(z,\Gamma)<2^{-j}\},\quad j=1,2,3,\cdots$
-and then computing directly one can see that the integral $\iint_{\Omega_{i}}d(z,\Gamma)^{(1-s)p-1}dxdy$ converges if and only if the series $\sum 2^{-n(p-sp+1-h(\Gamma))}$ converges, namely, $s<(p+1-h(\Gamma))/p$.
-Consequently, assuming $s\geq(p+1-h(\Gamma))/p$ one can see that in this case even the simplest function $u(z)=z$ is not in the space $\mathcal{B}^{s}_{p,p}(\Omega_{i})$.
-3.2 Plemelj-Calderón Property for $\mathcal{B}_{p,p}^{1/p}(\Gamma)$
+Remark 2.10.
+We end this subsection with a comment on the triangular domain above the shadowed region in Figure 1. Divide the domain $\Omega_{i}$ into pieces:
+$E_{0}$ $=\{z\in\Omega_{i}:\;d(z,\Gamma)\geq 2^{-1}\},$
+$E_{n}$ $=\{z\in\Omega_{i}:\;2^{-(n+1)}\leq d(z,\Gamma)<2^{-n}\},\quad n=1,2,3,\cdots$
+If $t<R$ it follows from basic covering theorems that there exists a constant $C_{1}$ such that (see p.474 in [3])
+$t^{h(\Gamma)-2}|\{z\in D(z_{0},R):\;d(z,\Gamma)<t\}|\leq C_{1}R^{h(\Gamma)}$
+where $D(z_{0},R)$ stands for the open disk centered at $z_{0}$ and of radius $R$.
+From this we see $|E_{n}|\leq C_{2}2^{-n(2-h(\Gamma))}$.
+Then, by direct computation one can see that the integral $\iint_{\Omega_{i}}d(z,\Gamma)^{(1-s)p-1}dxdy$ converges if and only if the series $\sum 2^{-n(p-sp+1-h(\Gamma))}$ converges, namely, $s<(p+1-h(\Gamma))/p$.
+Consequently, assuming $s\geq(p+1-h(\Gamma))/p$ one can see that even the simplest function $u(z)=z$ is not in the space $\mathcal{B}^{s}_{p,p}(\Omega_{i})$.
+2.6 Plemelj-Calderón problem for $\mathcal{B}_{p,p}^{1/p}(\Gamma)$
 Recall that the dense subspace $\mathcal{C}_{p}(\Omega_{i,e})$ of $\mathcal{B}_{p,p}^{1/p}(\Omega_{i,e})$ has well-defined boundary values so that $\mathcal{C}_{p}(\Omega_{i}\!\to\!\Gamma)=\mathcal{C}_{p}(\Omega_{e}\!\to\!\Gamma)$, this space of boundary values called $\mathcal{C}_{p}(\Gamma)$.
@@
 Remark 2.12.
-Under the assumption of Theorem 3.6, when $s<1/p$ the only thing that we can say is that $\mathcal{B}_{p,p}^{s}(\Gamma)$ contains $\mathcal{B}_{p,p}^{1/p}(\Gamma)$ and thus also $C_{c}^{\infty}(\mathbb{C})|_{\Gamma}$.
+Under the assumption of Theorem 2.11, when $s<1/p$ the only thing we can say is that $\mathcal{B}_{p,p}^{s}(\Gamma)$ contains $\mathcal{B}_{p,p}^{1/p}(\Gamma)$ and thus also $C_{c}^{\infty}(\mathbb{C})|_{\Gamma}$.
 Proof of Theorem 2.11.
 The assumption on $p$ is exactly the condition for $(h(\Gamma),p,\frac{1}{p})\in\mathcal{A}$.
-If it is satisfied we then have that $\omega(z)=d(z,\Gamma)^{p-2}$ is an $A_{p}$ weight, in particular, integrable locally in $\mathbb{C}$.
+If it is satisfied we then have that $\omega(z)=d(z,\Gamma)^{p-2}$ is an $A_{p}$-weight, in particular, integrable locally in $\mathbb{C}$.
 Suppose $g\in C_{c}^{\infty}(\mathbb{C})$.
@@
 By that and the almost-Dirichlet principle (Lemma 2.1) one may show $g|_{\Gamma}\in\mathcal{B}_{p,p}^{1/p}(\Gamma)$.
-For a general $f\in\mathcal{B}_{p,p}^{1/p}(\Gamma)$, we will use the argument on the boundary values of $\mathcal{B}_{p,p}^{1/p}(\Omega_{i,e})$ in the first paragraph of section 2.1.
+For a general $f\in\mathcal{B}_{p,p}^{1/p}(\Gamma)$, we will use the argument on the boundary values of $\mathcal{B}_{p,p}^{1/p}(\Omega_{i,e})$ in the first paragraph of subsection 2.1.
 Let $u$ be the harmonic extension to $\Omega_{i}$ of $f$.
@@
 Remark 2.14.
-We already know that this decomposition is unique if $p=2$. Since the spaces $\mathcal{B}_{p,p}^{1/p}(\mathbb{D})$ increase with $p$ we still have uniqueness for $h(\Gamma)<p\leq 2$.
+We already know that this decomposition is unique if $p=2$. Since the spaces $\mathcal{B}_{p,p}^{1/p}(\mathbb{D})$ increase with $p$, we still have uniqueness for $h(\Gamma)<p\leq 2$.
+3 $A_{p}$-Weighted Sobolev spaces on quasidisks
 3.1 Boundary values of functions in $\mathcal{B}_{p,p}^{s}(\Omega)$
 Let $\Gamma$ be a quasicircle and $\Omega$ one of its complementary domains $\Omega_{i,e}$ as above.
-The boundary values of the critical case $s=1/p$ have been treated in the last section.
-As announced in the introduction, we will define now what we mean by boundary values of functions in $\mathcal{B}_{p,p}^{s}(\Omega),\,p>1,\,s\in(0,1)$.
-As we shall see the cases $s<1/p$ and $s>1/p$ will show to be very different.
-We start with the case of $s>1/p$.
+The boundary values of $\mathcal{B}_{p,p}^{s}(\Omega)$-functions in the critical case $s=1/p$ have been treated in the last section.
+As announced in the introduction, we will define now what we mean by the boundary values of the functions in $\mathcal{B}_{p,p}^{s}(\Omega),\,p>1,\,s\in(0,1)$.
+As we shall see, the cases $s<1/p$ and $s>1/p$ will show to be very different.
+We now consider the case of $s>1/p$.
 Proposition 3.1.
-Let $\Omega$ be a (bounded or unbounded) domain bounded by a bounded Jordan curve $\Gamma$. For any $u\in\mathcal{B}_{p,p}^{s}(\Omega)$ with $s\in(1/p,1)$ there exists a constant $C>0$ such that
+Let $\Omega$ be a domain bounded by a bounded Jordan curve $\Gamma$. For any $u\in\mathcal{B}_{p,p}^{s}(\Omega)$ with $s\in(1/p,1)$ there exists a constant $C>0$ such that
 $|\nabla u(z)|\leq Cd(z,\Gamma)^{\frac{-1-(1-s)p}{p}},\qquad z\in\Omega.$
 Proof.
-Let $z\in\Omega$ and $D$ is the disk $D(z,d(z,\Gamma)/2)$. By the mean value property of harmonic functions one may write
+Let $z\in\Omega$ and $D$ be the disk $D(z,d(z,\Gamma)/2)$. By the mean value property of harmonic functions, one may write
 $\nabla u(z)=\frac{1}{|D|}\iint_{D}\nabla u(\zeta)d\xi d\eta.$
-(Here and in the sequel we write $|D|$ for the Lebesgue measure of $D$.) so that, by Hölder inequality,
+so that, by Hölder inequality,
 $|\nabla u(z)|^{p}$ $\leq\frac{4}{\pi d(z,\Gamma)^{2}}\iint_{D}|\nabla u(\zeta)|^{p}d\xi d\eta$
@@
 Proposition 3.2.
-Let $\Omega$ be a (bounded or unbounded) domain bounded by a quasicircle $\Gamma$ and $\alpha\in(0,1]$. The following are equivalent:
+Let $\Omega$ be a domain bounded by a quasicircle $\Gamma$ and $\alpha\in(0,1]$. The following are equivalent:
   1. (1)
@@
 Here, $\Lambda^{\alpha}(\overline{\Omega})$ stands for the space of Hölder functions of order $\alpha$.
-If now $s\in(1/p,1)$ then $\alpha=s-1/p\in(0,1)$ and Gehring-Martio theorem implies that functions in $\mathcal{B}_{p,p}^{s}(\Omega)$ are continuous on $\overline{\Omega}$ , so that boundary values of functions in $\mathcal{B}_{p,p}^{s}(\Omega_{i,e})$ are well-defined and they characterize $u$, as being the unique harmonic extension of this continuous boundary value.
+If now $s\in(1/p,1)$ then $\alpha=s-1/p\in(0,1)$ and Gehring-Martio theorem implies that the functions in $\mathcal{B}_{p,p}^{s}(\Omega)$ are continuous on $\overline{\Omega}$ , so the boundary values of the functions in $\mathcal{B}_{p,p}^{s}(\Omega_{i,e})$ are well-defined and characterize $u$, as being the unique harmonic extension of this continuous boundary value.
 We may now define the space $\mathcal{B}_{p,p}^{s}(\Gamma)$ for this range $s\in(1/p,1)$ as being the space of (s1/p)(s-1/p)-Hölder functions $f$ on $\Gamma$ whose harmonic extension $u_{i,e}$ to $\Omega_{i,e}$ belongs to $\mathcal{B}_{p,p}^{s}(\Omega_{i,e})$, and the space $\mathcal{B}_{p,p}^{s}(\Gamma)$ is assigned the natural norm $\|\cdot\|_{\mathcal{B}_{p,p}^{s}(\Gamma)}$ so that $\|f\|^{p}_{\mathcal{B}_{p,p}^{s}(\Gamma)}=\|u_{i}\|^{p}_{\mathcal{B}_{p,p}^{s}(\Omega_{i})}+\|u_{e}\|^{p}_{\mathcal{B}_{p,p}^{s}(\Omega_{e})}$.
-3.3.1 Trace Mappings
-As we have already seen, in the case of $s>1/p$ the spaces $\mathcal{B}_{p,p}^{s}(\Omega_{i,e})$ have well-defined boundary values and, with the terminology adopted in the introduction, the trace mapping
+3.2 Trace mappings on $\mathcal{B}_{p,p}^{s}(\Omega)$
+As we have already seen, in the case $s>1/p$ the spaces $\mathcal{B}_{p,p}^{s}(\Omega_{i,e})$ have well-defined boundary values and, with the terminology adopted in the introduction, the trace mapping
 $\mathcal{B}_{p,p}^{s}(\Omega_{i,e})\to\mathcal{B}_{p,p}^{s}(\Omega_{i,e}\to\Gamma)$
-is an isomorphism. The first purpose of this part is to determine conditions that imply
+is an isomorphism. The goal of this subsection is to discuss conditions that imply
 $\mathcal{B}_{p,p}^{s}(\Omega_{i}\to\Gamma)=\mathcal{B}_{p,p}^{s}(\Omega_{e}\to\Gamma).$
 We already know the result for $s=1/p$ and $h(\Gamma)<p<h(\Gamma)/(1-h(\Gamma))$ (but in this case the boundary value has a different meaning).
-Now the space $\Lambda^{s}(\Omega)$ introduced above in connection with Gehring-Martio theorem coincides with the Besov space $\mathcal{B}_{\infty,\infty}^{s}(\Omega)$.
-A theorem of Hinkkanen [20] (see also [48]) states that suppose the curve $\Gamma$ is a Jordan curve and if $f\in\Lambda^{s}(\Gamma)$ with $s<1/2$ then the harmonic extension of $f$ to $\Omega_{i,e}$ is in $\mathcal{B}_{\infty,\infty}^{s}(\Omega)$ with norm bounded by that of $f$ and the index $1/2$ is critical.
+Another case is $p=\infty$: the space $\Lambda^{s}(\Omega)$ introduced above in connection with the Gehring-Martio theorem coincides with the Besov space $\mathcal{B}_{\infty,\infty}^{s}(\Omega)$.
+A theorem of Hinkkanen [20] (see also [54]) states that if the curve $\Gamma$ is any Jordan curve and if $f\in\Lambda^{s}(\Gamma)$ with $s<1/2$ then the harmonic extension of $f$ to $\Omega_{i,e}$ is in $\mathcal{B}_{\infty,\infty}^{s}(\Omega)$ with norm bounded by that of $f$ and the index $1/2$ is critical.
 Hinkkanen’s result is true for all Jordan curves, and it has been improved to $s<1/K$ in [54] for quasicircles such that their Riemann mapping has a $K$-quasiconformal extension with $1\leq K\leq 2$.
-We thus define
+Thus, we define
 $K_{*}:=\min{(K,2)}.$
-We may now state the main result of this part.
-Theorem 3.12.
-Let $\Gamma,p,s$ be such that
-$h(\Gamma)<p\leq\infty,\;1/p\leq s\;\mathrm{and}\;s<\frac{p+K_{*}-h(\Gamma)}{K_{*}p}$ (3.6)
-then
-$\mathcal{B}_{p,p}^{s}(\Omega_{i}\to\Gamma)=\mathcal{B}_{p,p}^{s}(\Omega_{e}\to\Gamma).$ (3.7)
-Proof.
-Let $P_{e}$, as before, be the operator which takes a function on $\Gamma$ to its harmonic extension in $\Omega_{e}$. As we have recalled that if
+Let $P_{e}$, as before, be the operator that takes a function on $\Gamma$ to its harmonic extension in $\Omega_{e}$. We have recalled that if
 $h(\Gamma)<q<\frac{h(\Gamma)}{1-h(\Gamma)},\,\;\;0<\sigma<1/K_{*}$
@@
 $P_{e}:\,\mathcal{B}_{\infty,\infty}^{\sigma}(\Omega_{i}\to\Gamma)$ $\to\mathcal{B}_{\infty,\infty}^{\sigma}(\Omega_{e}),$
-are both bounded. We want to interpolate these results. Using complex interpolation, we have, for $0<\theta<1$,
-$\mathcal{B}_{q,q}^{1/q}(\Omega),\mathcal{B}_{\infty,\infty}^{\sigma}(\Omega)]_{\theta}=\mathcal{B}_{p,p}^{s}(\Omega),$
+are both bounded. We want to interpolate these results. It is known that in the sense of complex interpolation we have in the case of unit disk $\mathbb{D}$, for $0<\theta<1$,
+$\mathcal{B}_{q,q}^{1/q}(\mathbb{D}),\mathcal{B}_{\infty,\infty}^{\sigma}(\mathbb{D})]_{\theta}=\mathcal{B}_{p,p}^{s}(\mathbb{D}),$
 with
@@
 $\frac{1}{p}$ $=\frac{1-\theta}{q}.$
-We must decide what the couples $(p,s)$ are that may be "caught" in this way in view of the constraints: the relations above are equivalent to
-$\sigma$ $=\frac{s-\frac{1}{p}}{\theta},$
-$q$ $=(1-\theta)p,$
-and the constraints are
-$0<\frac{s-\frac{1}{p}}{\theta}$ $<1/K_{*},$
-$h(\Gamma)<(1-\theta)p$ $<\frac{h(\Gamma)}{h(\Gamma)-1}.$
-That correspond to $\theta$ being in the intersection of two intervals
-$(K_{*}(s-\frac{1}{p}),1),\;(1-\frac{h(\Gamma)}{p(h(\Gamma)-1)},1-\frac{h(\Gamma)}{p}).$
-We see that in order to have a solution we must have $s>1/p,p>h(\Gamma)$, and $K_{*}(s-\frac{1}{p})<1-\frac{h(\Gamma)}{p}$, which is the condition of the statement of the theorem. ∎
-3.3.2 Weighted Sobolev Spaces
-Let $\omega$ be an $A_{p}$ weight defined in $\mathbb{C}$, for some $p>1$.
-The weighted Sobolev space $W^{1,p}(\omega,\mathbb{C})$ is the space of tempered distributions $T$ such that $\nabla T$, taken in the sense of distributions, is a function in $L^{p}(\omega,\mathbb{C})$.
+If the interpolation statement above remained true for quasicircles then, using interpolation for the operator $P_{e}$, we would produce many new examples of equality of Besov-Sobolev spaces from inside and outside $\Gamma$.
+We strongly believe that this interpolation for quasicircles is true but we could not prove it so far, except for the rectifiable case to be treated in Section 4.
+3.3 Trace mappings on weighted Sobolev spaces
+Let $\omega$ be an $A_{p}$-weight defined in $\mathbb{C}$, for some $p>1$.
+The weighted Sobolev space $W^{1,p}(\omega,\mathbb{C})$ is the space of tempered distributions $T$ such that $T$ is locally integrable and $\nabla T$, taken in the sense of distributions, is a function in $L^{p}(\omega,\mathbb{C})$.
 We will use these spaces for $\omega(z)=d(z,\Gamma)^{(1-s)p-1}\in A_{p}$ provided that $(h(\Gamma),p,s)\in\mathcal{A}$.
-For the usual technical reasons we need to know that $C_{c}^{\infty}(\mathbb{C})$-functions are dense in $W^{1,p}(\omega,\mathbb{C})$.
+For the usual technical reasons, we need to know that the $C_{c}^{\infty}(\mathbb{C})$-functions are dense in $W^{1,p}(\omega,\mathbb{C})$. See [8] for a proof of a more general case.
 Lemma 3.3.
@@
 Proof.
-Let us consider $f\in W^{1,p}(\omega,\mathbb{C})$: without loss of generally we may assume that $f$ has compact support.
+Consider $f\in W^{1,p}(\omega,\mathbb{C})$: without loss of generality, we may assume that $f$ has compact support.
 Let then $(\varphi_{\varepsilon})$ be a $C_{c}^{\infty}(\mathbb{C})$-approximation of identity, so that $f_{\varepsilon}=f*\varphi_{\varepsilon}\in C_{c}^{\infty}(\mathbb{C})$.
@@
 which converges to 0 by the Lebesgue dominated convergence theorem.
-Indeed $|\nabla f_{\varepsilon}|$ is controlled by the Hardy-Littlewood maximal function of $\nabla f$ and this maximal function is in $L^{p}(\omega,\mathbb{C})$ from the fact that $\omega\in A_{p}$.
+Indeed $|\nabla f_{\varepsilon}|$ is controlled by the Hardy-Littlewood maximal function of $\nabla f$ and this maximal function is in $L^{p}(\omega,\mathbb{C})$ from the fact that $\omega\in A_{p}$ (see [34]).
 ∎
-In the classical case $\omega\equiv 1$ we recall that if $p>2$ then $W^{1,p}(\mathbb{C})\subset\Lambda^{\alpha}(\mathbb{C})$ with $\alpha=1-2/p$.
-In the weighted case we do not know this embedding result in general but we have two particular analogues.
+In the classical case $\omega\equiv 1$ we recall that if $p>2$ then $W^{1,p}(\mathbb{C})\subset\Lambda^{\alpha}(\mathbb{C})$ with $\alpha=1-2/p$, and moreover, the property being local, a similar inclusion holds for $W^{1,p}(D)$, $D$ being a disk.
 Proposition 3.4.
-The following two embedding assertions hold:
-  • If $\omega\in A_{1}$ then every function in $W^{1,p}(\omega,\mathbb{C})$, $p>2$, is continuous;
-  • If $\omega=d(z,\Gamma)^{(1-s)p-1}$ is only assumed to be in $A_{q}$ for some $q>1$ then the same result is true if $p>\frac{6-2h(\Gamma)}{2-h(\Gamma)}$.
-Remark 3.15.
-When $\omega(z)=d(z,\Gamma)^{(1-s)p-1}$ it is known that for any $q>1$, $\omega\in A_{q}\Leftrightarrow\omega\in A_{1}$ if 0<(1s)p<10<(1-s)p<1, i.e., s>11/ps>1-1/p.
+If $p>2$, the following two embedding assertions hold:
+  • If $\omega\in A_{1}$ then all functions in $W^{1,p}(\omega,\mathbb{C})$ are continuous;
+  • If $\omega(z)=d(z,\Gamma)^{(1-s)p-1}$ is only assumed to be locally integrable, then the same result holds if $s>\frac{h(\Gamma)}{2}-\frac{h(\Gamma)-1}{p}$.
+We recall here that when $\omega(z)=d(z,\Gamma)^{(1-s)p-1}$ it is known that for any $q>1$, $\omega\in A_{q}\Leftrightarrow\omega\in A_{1}$ if $h(\Gamma)-1<(1-s)p<1$, i.e., $1-\frac{h-1}{p}s>1-1/p$.
 Proof of Proposition 3.4.
-Suppose first that $\omega\in A_{1}$. Choose $q$ so that $2<q<p$ and let $B$ be an open ball in $\mathbb{C}$ . By Hölder inequality, we get
-$\int_{B}|\nabla f(x)|^{q}dxdy\leq\left(\int_{B}|\nabla f(x)|^{p}\omega(x)dxdy\right)^{\frac{q}{p}}\left(\int_{B}\omega(x)^{-\frac{q}{p-q}}dxdy\right)^{1-\frac{q}{p}}<\infty$
+Suppose first that $\omega\in A_{1}$. Choose $q$ so that $2<q<p$ and let $D$ be an open disk in $\mathbb{C}$ . By Hölder inequality, we get
+$\int_{D}|\nabla f(x)|^{q}dxdy\leq\left(\int_{D}|\nabla f(x)|^{p}\omega(x)dxdy\right)^{\frac{q}{p}}\left(\int_{D}\omega(x)^{-\frac{q}{p-q}}dxdy\right)^{1-\frac{q}{p}}<\infty$ (3.1)
 because $f\in W^{1,p}(\omega,\mathbb{C})$, and $\omega\in A_{1}\subset A_{p/q}$, which is equivalent to $\omega^{-q/(p-q)}\in A_{p/(p-q)}$, and in particular implies that $\omega^{-q/(p-q)}$ is locally integrable.
-For the second case let us recall first that if $B$ is a ball containing $\Gamma$ then
-$\iint_{B}d(z,\Gamma)^{\beta}dxdy<\infty$
-if $\beta>h(\Gamma)-2$. By the previous computation it follows that a sufficient condition for the validity of the theorem is
-$\frac{2}{p-2}>h(\Gamma)-2$
-which is equivalent to the condition stated. ∎
-We return to Theorem 3.12. It states that under certain hypothesis on $(h(\Gamma),p,s)$ we have
-$\mathcal{B}_{p,p}^{s}(\Omega_{i}\to\Gamma)=\mathcal{B}_{p,p}^{s}(\Omega_{e}\to\Gamma).$
-Let us call this space of boundary values $\mathcal{B}_{p,p}^{s}(\Gamma)$.
+We conclude with the Sobolev embedding theorem mentioned above.
+For the second case, if $D$ is a disk containing $\Gamma$ then by the same computation as that in Remark 2.10 we see
+$\iint_{D}d(z,\Gamma)^{\beta}dxdy<\infty$
+if $\beta>h(\Gamma)-2$.
+By the previous computation it follows that a sufficient condition for the validity of the theorem is that the second integral in the right hand-side of (3.1) is finite, and this will be the case if there exists $q\in(2,p)$ such that
+$\frac{q}{p-q}((1-s)p-1)>h(\Gamma)-2.$
+This condition is equivalent to, with $h=h(\Gamma)$:
+F(q,s,p,h)=q(h1(1s)p)+p(2h)>0.F(q,s,p,h)=q(h-1-(1-s)p)+p(2-h)>0.
+The function $F$ being affine in $q$, the last inequality occurs if and only if
+$\max(F(2,s,p,h),F(p,s,p,h))>0.$
+But
+$F(2,s,p,h)$ $=2p\left(s-\left(\frac{h}{2}-\frac{h-1}{p}\right)\right),$
+$F(p,s,p,h)$ $=p^{2}\left(s-\left(1-\frac{1}{p}\right)\right).$
+The proposition easily follows by noticing that, if $p>2$,
+$\frac{h}{2}-\frac{h-1}{p}<1-\frac{1}{p}.$
+Let us now assume that the quasicircle $\Gamma$ satisfies
+$\mathcal{B}_{p,p}^{s}(\Omega_{i}\to\Gamma)=\mathcal{B}_{p,p}^{s}(\Omega_{e}\to\Gamma),$
+and we then call this space of boundary values $\mathcal{B}_{p,p}^{s}(\Gamma)$.
 Theorem 3.5.
-With the following assumption
-$\frac{1}{p}<1-\frac{1}{p}\leq s<\frac{p+K_{*}-h(\Gamma)}{K_{*}p}$ (3.8)
-we have that the elements of $W^{1,p}(\omega,\mathbb{C})$ are continuous functions in $\mathbb{C}$ and that
+If $p>2,\,s>\frac{h(\Gamma)}{2}-\frac{h(\Gamma)-1}{p}$ and $\omega\in A_{p}$ where $\omega(z)=d(z,\Gamma)^{(1-s)p-1}$ then the elements of $W^{1,p}(\omega,\mathbb{C})$ are continuous functions in $\mathbb{C}$ and
 $\mathcal{B}_{p,p}^{s}(\Gamma)=W^{1,p}(\omega,\mathbb{C})|_{\Gamma},$ (3.2)
-where $\omega(z)=d(z,\Gamma)^{(1-s)p-1}$ and $W^{1,p}(\omega,\mathbb{C})|_{\Gamma}=\{f|_{\Gamma}:\,f\in W^{1,p}(\omega,\mathbb{C})\}.$
-Before we start the proof, let us comment on this statement. First of all, the assumption made implies that $\omega\in A_{p}$ and that
-$\mathcal{B}_{p,p}^{s}(\Omega_{i}\to\Gamma)=\mathcal{B}_{p,p}^{s}(\Omega_{e}\to\Gamma)\;(=\mathcal{B}_{p,p}^{s}(\Gamma)).$
-Moreover the inequality 1/p<11/p1/p<1-1/p is equivalent to $p>2$ and the inequality s>11/ps>1-1/p further implies, as we have noticed above, that $\omega\in A_{1}$, from which it follows that the elements of $W^{1,p}(\omega,\mathbb{C})$ are continuous functions (Proposition 3.14).
-Finally, the inequality
-$1-\frac{1}{p}<\frac{p+K_{*}-h(\Gamma)}{K_{*}p}$
-is equivalent to $p<\frac{2K_{*}-h(\Gamma)}{K_{*}-1}$.
+where $W^{1,p}(\omega,\mathbb{C})|_{\Gamma}=\{f|_{\Gamma}:\,f\in W^{1,p}(\omega,\mathbb{C})\}.$
 Proof of Theorem 3.5.
@@
 Let $f\in\mathcal{B}_{p,p}^{s}(\Gamma)$ and let $u_{i}$ the harmonic extension of $f$ to $\Omega_{i}$ .
-By hypothesis $u_{i}\in W^{1,p}(\omega,\Omega_{i})$ and a theorem of [8] asserts that since $\Gamma$ is a quasicircle and $\omega\in A_{p}$, $u_{i}$ can be extended to a function of $W^{1,p}(\omega,\mathbb{C})$.
+By hypotheses $u_{i}\in W^{1,p}(\omega,\Omega_{i})$ and a theorem of [8] asserts that since $\Gamma$ is a quasicircle and $\omega\in A_{p}$, $u_{i}$ can be extended to a function of $W^{1,p}(\omega,\mathbb{C})$.
 In order to prove that $W^{1,p}(\omega,\mathbb{C})|_{\Gamma}\subset\mathcal{B}_{p,p}^{s}(\Gamma)$, let us first consider $f\in C_{c}^{\infty}(\mathbb{C})$. Recall that
 $F_{i}(z)=-\frac{1}{\pi}\iint_{\Omega_{e}}\frac{\bar{\partial}f(\zeta)}{\zeta-z}d\xi d\eta,\quad F_{e}(z)=\frac{1}{\pi}\iint_{\Omega_{i}}\frac{\bar{\partial}f(\zeta)}{\zeta-z}d\xi d\eta$
-are holomorphic respectively in $\Omega_{i}$ and $\Omega_{e}$, and the fact that $\omega\in A_{p}$ implies that $F_{i,e}\in\mathcal{B}_{p,p}^{s}(\Omega_{i,e}\to\Gamma)$, which is equal to $\mathcal{B}_{p,p}^{s}(\Gamma)$ under the assumption of Theorem 3.12.
+are holomorphic respectively in $\Omega_{i}$ and $\Omega_{e}$, and the fact that $\omega\in A_{p}$ implies that $F_{i,e}\in\mathcal{B}_{p,p}^{s}(\Omega_{i,e}\to\Gamma)$, which is equal to $\mathcal{B}_{p,p}^{s}(\Gamma)$ by hypotheses.
 Using $f|_{\Gamma}=F_{i}-F_{e}$ we may see $f|_{\Gamma}\in\mathcal{B}_{p,p}^{s}(\Gamma)$.
@@
 Corollary 3.6.
-Under the assumption of Theorem 3.16, the space $C_{c}^{\infty}(\mathbb{C})|_{\Gamma}$ is dense in $\mathcal{B}_{p,p}^{s}(\Gamma)$, and thus the conclusions in Theorem 3.5 hold for all $f\in\mathcal{B}_{p,p}^{s}(\Gamma)$.
-We end this section by a question.
-Is it true that within the hypothesis of Theorem 3.16 if $f\in\mathcal{B}_{p,p}^{s}(\Gamma)$ then its extension to the whole plane by the harmonic extension on both sides belongs to $W^{1,p}(\omega,\mathbb{C})$?
-We do not know the answer to this question in general but only for the cases with improved regularity (bigger $s$):
+Under the assumption of Theorem 3.5, the space $C_{c}^{\infty}(\mathbb{C})|_{\Gamma}$ is dense in $\mathcal{B}_{p,p}^{s}(\Gamma)$, and thus the conclusions of Theorem 2.9 hold for all $f\in\mathcal{B}_{p,p}^{s}(\Gamma)$.
+We end this subsection with a question.
+Is it true that within the hypotheses of Theorem 3.5 if $f\in\mathcal{B}_{p,p}^{s}(\Gamma)$ then its extension to the whole plane by the harmonic extension on both sides belongs to $W^{1,p}(\omega,\mathbb{C})$?
+We do not know the answer to this question in general, but only for the cases with improved regularity (larger $s$):
 Proposition 3.7.
-Let $\Gamma$ be a quasicircle and let $f$ be a function which is $\alpha$-Hölder continuous in $\mathbb{C}$ with $\alpha>h(\Gamma)-1$ and of class $C^{1}$ in $\mathbb{C}\backslash\Gamma$.
-Then if the function $F(z)=\nabla f(z),\,z\in\mathbb{C}\backslash\Gamma$ (thus defined almost everywhere) is in $L^{1}_{loc}(\mathbb{C})$, we have that, in the sense of distributions, $\nabla f=F$.
+Let $\Gamma$ be a quasicircle, and let $f$ be a function that is $\alpha$-Hölder continuous in $\mathbb{C}$ with $\alpha>h(\Gamma)-1$ and of class $C^{1}$ in $\mathbb{C}\backslash\Gamma$.
+Then if the function $F(z)=\nabla f(z),\,z\in\mathbb{C}\backslash\Gamma$ (defined almost everywhere) is in $L^{1}_{loc}(\mathbb{C})$, we have, in the sense of distributions, $\nabla f=F$.
 Proof.
 Let $\varphi$ be a $C_{c}^{\infty}$-function in $\mathbb{R}$ with support included in [1,1][-1,1].
-As usual we define the approximation of identity $\varphi_{\varepsilon}(x)=\frac{1}{\varepsilon}\varphi(\frac{x}{\varepsilon})$.
+As usual, we define the approximation of identity $\varphi_{\varepsilon}(x)=\frac{1}{\varepsilon}\varphi(\frac{x}{\varepsilon})$.
 From this we can define the approximation of identity in two variables
 $\psi_{\varepsilon}(z)=\psi_{\varepsilon}(x+iy)=\varphi_{\varepsilon}(x)\varphi_{\varepsilon}(y).$
-We may assume without loss of generality that $f$ has compact support, the function $f*\psi_{\varepsilon}$ is then a $C^{\infty}$-function with compact support and it converges uniformly in $\mathbb{C}$ towards $f$.
+We may assume without loss of generality that $f$ has compact support, and then the function $f*\psi_{\varepsilon}$ is a $C^{\infty}$-function with compact support and converges uniformly in $\mathbb{C}$ towards $f$.
 From this we deduce that
@@
 for any $h\in C_{c}^{\infty}(\mathbb{C})$, and the right-hand side converges to $\iint_{\mathbb{C}}f(z)\nabla h(z)dxdy$ as $\varepsilon$ converges to 0.
-We split the left-hand side into two parts:
-Let $\Gamma_{\varepsilon}$ be the set of points at distance less than $2\varepsilon$ from $\Gamma$. Then
+We divide the left-hand side into two parts:
+Let $\Gamma_{\varepsilon}$ be the set of points at a distance less than $2\varepsilon$ from $\Gamma$. Then
 $\iint_{\mathbb{C}\backslash\Gamma_{\varepsilon}}\nabla(f*\psi_{\varepsilon})(z)h(z)dxdy=\iint_{\mathbb{C}\backslash\Gamma_{\varepsilon}}(\nabla F)*\psi_{\varepsilon}(z)h(z)dxdy$
@@
 Now $|f(x-t_{1},y-t_{2})-f(x,y-t_{2})|\leq C\varepsilon^{\alpha}$ and $|\varphi_{\varepsilon}^{\prime}(t_{1})|\leq C\varepsilon^{-2}$.
-Since we integrate on an interval of length $2\varepsilon$, the absolute value of the whole integral is bounded by $C\varepsilon^{\alpha-1}$.
+Since we integrate on an interval of length $2\varepsilon$, the absolute value of the entire integral is bounded by $C\varepsilon^{\alpha-1}$.
 Finally,
 $\iint_{\Gamma_{\varepsilon}}\left|f*\frac{\partial(\psi_{\varepsilon})}{\partial x}(z)\right|dxdy\leq|\Gamma_{\varepsilon}|\varepsilon^{\alpha-1}.$
-But we know that $|K\Gamma_{\varepsilon}|\varepsilon^{h(\Gamma)-2}\leq C$; we conclude that
+But we know $|K\Gamma_{\varepsilon}|\varepsilon^{h(\Gamma)-2}\leq C$; we conclude that
 $|\Gamma_{\varepsilon}|\varepsilon^{\alpha-1}\leq C\varepsilon^{\alpha+1-h(\Gamma)},$
 and the result follows. ∎
-Based on Proposition 3.18, a sufficient condition for a positive answer to the question is the hypothesis of Theorem 3.16 and $s>\frac{1}{p}+(h(\Gamma)-1)$.
-To be precise, the function $f\in\mathcal{B}_{p,p}^{s}(\Gamma)$, defined to be $u_{i,e}$ in $\Omega_{i,e}$, satisfies the assumption of Proposition 3.18, and thus belongs to $L^{1}_{\text{loc}}(\mathbb{C})$ in the sense of distribution, so that $f\in W^{1,p}(\omega,\mathbb{C})$.
-But for the condition $s>\frac{1}{p}+(h(\Gamma)-1)$ to be compatible with (3.8) some elementary arithmetic left to the reader leads to the following sufficient conditions:
-$p>2,\quad\frac{1}{p}+h(\Gamma)-1<s<\frac{p+K_{*}-h(\Gamma)}{pK_{*}}.$
-Notice that in the case of a chord-arc curve ($h(\Gamma)=1$) there are admissible values of $s$ for all $p>2$. In general, for having amissible values of $s$ one needs
-$h(\Gamma)<\frac{K_{*}+1}{K_{*}+\frac{1}{p}}$
-and for this inequality to hold for any $p>2$ and $K_{*}\leq 2$, $h(\Gamma)<\frac{6}{5}$ is necessary and sufficient.
-3.3.3 Almost-Dirichlet Principle
+Based on Proposition 3.7, a sufficient condition for a positive answer to the question is the hypotheses of Theorem 3.5 and $s>\frac{1}{p}+(h(\Gamma)-1)$.
+Notice that this condition need not be compatible with the assumptions of the theorem, in particular with the fact that $\omega\in A_{p}$: we indeed need
+$\frac{1}{p}+h-1<1+\frac{1-h}{p},$
+which is equivalent to
+$p>\frac{h}{2-h},$
+and this condition is a new constraint if $h>4/3$.
+3.4 Almost-Dirichlet principle
 We start by recalling the classical Dirichlet principle.
-If $\Omega$ is a Jordan domain with $\partial\Omega=\Gamma$ and $f$ is a continuous function on $\Gamma$ having a continuous extension $F:\overline{\Omega}\to\mathbb{C}$ which is in $C^{1}(\Omega)$.
+If $\Omega$ is a Jordan domain with $\partial\Omega=\Gamma$ and $f$ is a continuous function on $\Gamma$ having a continuous extension $F:\overline{\Omega}\to\mathbb{C}$ that is in $C^{1}(\Omega)$.
 then, if $u$ stands for the harmonic extension of $f$ to $\Omega$, we have
 $\iint_{\Omega}|\nabla u(z)|^{2}dxdy\leq\iint_{\Omega}|\nabla F(z)|^{2}dxdy.$
-Now let $\Gamma$ be a quasicircle and $p>1,\,s\in(0,1)$: we will say that $\mathcal{B}_{p,p}^{s}(\Omega)$ satisfies an almost-Dirichlet principle if there exists $C>0$ such that
+Now, let $\Gamma$ be a quasicircle, and $p>1,\,s\in(0,1)$. We will say that $\mathcal{B}_{p,p}^{s}(\Omega)$ satisfies an almost-Dirichlet principle if there exists $C>0$ such that
 $\iint_{\Omega}|\nabla u(z)|^{p}d(z,\Gamma)^{(1-s)p-1}dxdy\leq C\iint_{\Omega}|\nabla F(z)|^{p}d(z,\Gamma)^{(1-s)p-1}dxdy.$
 Maz’ya [32] has proven the almost-Dirichlet principle for the unit disk for all $p>1,s\in(0,1)$ and his proof has been simplified by [33].
-Theorem 3.16 shows that if $\Gamma$ satisfies the hypothesis of this theorem, then the almost-Dirichlet principle holds.
-3.3.4 The Case of pp–Dirichlet Energy
-In this part we highlight the case of $p$-Dirichlet energy corresponding to $\omega\equiv 1$, i.e., s=11/ps=1-1/p.
-Theorem 3.19.
-If $\Gamma$ is a quasicircle then for all $p\in\left(2,\frac{2K_{*}-h(\Gamma)}{K_{*}-1}\right)$, there exists a constant $C=C(p,\Gamma,s)$ such that for $v\in C_{c}^{\infty}(\mathbb{C})$ ,
-$\iint_{\Omega}|\nabla u(z)|^{p}dxdy\leq C\iint_{\Omega}|\nabla v(z)|^{p}dxdy,$ (3.10)
-where $u$ stands for the harmonic extension of $v|_{\Gamma}$ in $\Omega$.
-The proof is just by putting s=11/ps=1-1/p in Theorem 3.16 and taking care of the constraints: details are left to the reader.
-In particular, if $\Gamma$ is chord-arc then (3.10) holds for $p\in[2,3)$, and we have the following extension if $K<2$:
-Corollary 3.20.
-If $\Gamma$ is chord-arc with $K<2$ then the almost-Dirichlet principle holds for $2\leq p<\frac{2K-1}{K-1}.$
-Notice that the corollary implies that for any $p>2$ the almost-Dirichlet principle is valid if $K$ is close enough to $1$.
-4 Fractional Sobolev Spaces on Chord-arc Domains
-4.1 A Review of Classical Facts about Fractional Sobolev Spaces on Plane Domains.
-4.1.1 Fractional Sobolev Spaces on Plane Domains
-We have already dealt with the Sobolev space $W^{1,p}(\mathbb{C})$, $1<p<\infty$ consisting of tempered distributions $f\in\mathcal{S}^{\prime}(\mathbb{C})$ such that $f$ is locally integrable and $\nabla f$, taken in the sense of distributions, is in $L^{p}(\mathbb{C})$.
+Theorem 3.5 shows that if $\Gamma$ satisfies the hypotheses of this theorem, then the almost-Dirichlet principle holds.
+4 Plemelj-Calderón problem for $B_{p,p}^{s}(\Gamma)$ on chord-arc curves: the rectifiable case
+4.1 A review of classical facts about fractional Sobolev spaces in plane domains.
+We have already dealt with the Sobolev space $W^{1,p}(\mathbb{C})$, $1<p<\infty$, consisting of tempered distributions $f$ such that $f$ is locally integrable and $\nabla f$, taken in the sense of distributions, is in $L^{p}(\mathbb{C})$.
 The semi-norm
 $\|f\|_{W^{1,p}(\mathbb{C})}=\left(\iint_{\mathbb{C}}|\nabla f(z)|^{p}dxdy\right)^{\frac{1}{p}}$
-equips $W^{1,p}(\mathbb{C})$ with a structure of Banach space modulo constants called homogeneous Sobolev space in the literature.
+equips $W^{1,p}(\mathbb{C})$ with a structure of Banach space modulo constants called the homogeneous Sobolev space in the literature.
 The inhomogeneous Sobolev space is given by $\mathbb{W}^{1,p}(\mathbb{C})=W^{1,p}(\mathbb{C})\cap L^{p}(\mathbb{C})$.
@@
 $W^{1,p}(\Omega)$ and $\mathbb{W}^{1,p}(\Omega)$ should be defined similarly.
-We point out that if the open set $\Omega$ supports classical Poincaré inequality
+We point out that if the open set $\Omega$ satisfies the classical Poincaré inequality, namely,
 $\|f\|_{L^{p}(\Omega)}\leq C\|f\|_{W^{1,p}(\Omega)}$
-for every $f\in C_{c}^{\infty}(\Omega)$ where the constant $C$ depends only on $\Omega$ and $p$ then both spaces are known to coincide.
-For instance, the bounded chord-arc domain is a such domain.
-The bounded quasidisk also supports Poincaré inequality since it is an extension domain for Sobolev spaces $W^{1,p}(\mathbb{C})$ (see [22, 8]) so that one can apply Poincaré inequality in a bigger sufficiently regular domain which, combined with the extension theorem, leads to the conclusion.
-For $0<s<1$ we define the fractional Besov-Sobolev space $B_{p,p}^{s}(\mathbb{C})$ as the space of $L^{p}(\mathbb{C})$-functions $f:\mathbb{C}\to\mathbb{C}$ such that
+for every $f\in C_{c}^{\infty}(\Omega)$ where the constant $C$ depends only on $\Omega$ and $p$ then both spaces coincide.
+For instance, bounded chord-arc domains are such domains.
+The bounded quasidisks also satisfy Poincaré inequality, since it is an extension domain for Sobolev spaces $W^{1,p}(\mathbb{C})$ (see [22, 8]) so that one can apply Poincaré inequality in a sufficiently regular larger domain, which, combined with the extension theorem, leads to the conclusion.
+For $0<s<1$ we define the fractional Besov-Sobolev space $B_{p,p}^{s}(\mathbb{C})$ in $\mathbb{C}$ as the space of $L^{p}(\mathbb{C})$-functions $f:\mathbb{C}\to\mathbb{C}$ such that
 $\int_{\mathbb{C}}\int_{\mathbb{C}}\frac{|f(z)-f(\zeta)|^{p}}{|z-\zeta|^{2+sp}}dzd\zeta<\infty.$
-We remark that suppose $s\geq 1$ and $\Omega$ is a connected open set in $\mathbb{C}$ then any measurable function $f:\Omega\to\mathbb{C}$ such that the integral of the above integrand $|f(z)-f(\zeta)|^{p}/|z-\zeta|^{2+sp}$ over $\Omega\times\Omega$ converges is actually constant.
-These spaces may be viewed as intermediate spaces between $W^{0,p}(\mathbb{C})$, i.e., $L^{p}(\mathbb{C})$ and $\mathbb{W}^{1,p}(\mathbb{C})$.
+(We remark that if $s\geq 1$ and $\Omega$ is a connected open set in $\mathbb{C}$ then any measurable function $f:\Omega\to\mathbb{C}$ such that the integral of the above integrand $|f(z)-f(\zeta)|^{p}/|z-\zeta|^{2+sp}$ over $\Omega\times\Omega$ converges is actually constant.) These spaces may be viewed as intermediate spaces between $W^{0,p}(\mathbb{C})$, i.e., $L^{p}(\mathbb{C})$ and $\mathbb{W}^{1,p}(\mathbb{C})$.
 More precisely, they may be obtained as real-interpolate spaces between these two spaces: $L^{p}(\mathbb{C}),\mathbb{W}^{1,p}(\mathbb{C})]_{s,p}=B^{s}_{p,p}(\mathbb{C})$, see [46].
-We call them Besov-Sobolev spaces because if one uses complex interpolation instead we obtain the spaces $H_{p}^{s}(\mathbb{C})$ consisting of Bessel potential of $L^{p}$-functions: $L^{p}(\mathbb{C}),\mathbb{W}^{1,p}(\mathbb{C})]_{s}=H^{s}_{p}(\mathbb{C})$.
-and $B_{p,p}^{s}(\mathbb{C})=H_{p}^{s}(\mathbb{C})$ only if $p=2$ (if $p\neq 2$ these two spaces are not even isomorphic).
-For further use we extend the definition of Besov-Sobolev spaces to $s\in(1,2)$ by saying that $f\in B_{p,p}^{s}(\mathbb{C})$ with $s\in(1,2)$ if $f\in B_{p,p}^{s-1}(\mathbb{C})$ and $\nabla f\in B_{p,p}^{s-1}(\mathbb{C})$.
-Similarly, it can be further extended to other non-integer orders $s>0$.
+Notice that they are also special Besov spaces.
+We call them fractional Besov-Sobolev spaces.
+If we use complex interpolation instead, we obtain the spaces $H_{p}^{s}(\mathbb{C})$ consisting of Bessel potentials of $L^{p}$-functions: $L^{p}(\mathbb{C}),\mathbb{W}^{1,p}(\mathbb{C})]_{s}=H^{s}_{p}(\mathbb{C})$.
+It is known that $B_{p,p}^{s}(\mathbb{C})=H_{p}^{s}(\mathbb{C})$ only if $p=2$ (if $p\neq 2$ these two spaces are not even isomorphic).
+For further use, we extend the definition of fractional Besov-Sobolev spaces to the case of $s\in(1,2)$ by saying that $f\in B_{p,p}^{s}(\mathbb{C})$ with $s\in(1,2)$ if $f\in B_{p,p}^{s-1}(\mathbb{C})$ and $\nabla f\in B_{p,p}^{s-1}(\mathbb{C})$.
+Similarly, it can be extended to the case of other non-integer orders $s>0$.
 It is known that the space $C_{c}^{\infty}(\mathbb{C})$ is dense in $B_{p,p}^{s}(\mathbb{C})$ for non-integer orders $s>0$ and $1<p<\infty$.
@@
 The name comes from the fact that a Jordan domain $\Omega$ is chord-arc if and only if $\Gamma=\partial\Omega$ is a chord-arc curve.
-We define $B_{p,p}^{s}(\Omega)$ simply as the space of restrictions to $\Omega$ of functions in $B_{p,p}^{s}(\mathbb{C})$. Such a function is in particular such that
+We define $B_{p,p}^{s}(\Omega)$ simply as the space of restrictions to $\Omega$ of functions in $B_{p,p}^{s}(\mathbb{C})$. Such a function , in particular, satisfies
 $\int_{\Omega}\int_{\Omega}\frac{|f(z)-f(\zeta)|^{p}}{|z-\zeta|^{2+sp}}dzd\zeta<\infty$
 and functions satisfying this last property can be extended to a function in $B_{p,p}^{s}(\mathbb{C})$.
-4.1.2 Fractional Sobolev Spaces on Chord-arc Curves
-If $\Omega$ is a (bounded or unbounded) chord-arc domain with boundary $\Gamma$ we may define for $0<s<1$, $p>1$,
+If $\Omega$ is a chord-arc domain with boundary $\Gamma$ we may define the fractional Besov-Sobolev spaces $B_{p,p}^{s}(\Gamma)$ on $\Gamma$ as follows: $0<s<1$, $p>1$,
 $B_{p,p}^{s}(\Gamma)=\{f\in L^{p}(\Gamma):\,\iint_{\Gamma\times\Gamma}\frac{|f(z)-f(\zeta)|^{p}}{|z-\zeta|^{1+ps}}|dz||d\zeta|<\infty\}.$
-We define the trace operator $\gamma$ on $C_{c}^{\infty}(\mathbb{C})$ by $\gamma(u)=u|_{\Gamma}$.
+We define the trace operator $\gamma$ in $C_{c}^{\infty}(\mathbb{C})$ by $\gamma(u)=u|_{\Gamma}$.
 Theorem 4.1 (p.182 in [24]).
-Let $s\in(1/p,1+1/p)$.
+Let $\Omega$ be a chord-arc domain and let $s\in(1/p,1+1/p)$.
 The trace operator $\gamma$ can be extended to a bounded linear operator from $B_{p,p}^{s}(\Omega)$ onto $B_{p,p}^{s-1/p}(\Gamma)$, the order having $1/p$-loss, whose kernel is the closure of $C_{c}^{\infty}(\Omega)$ in $B_{p,p}^{s}(\Omega)$.
-4.2 Plemelj-Calderón Problem on Chord-arc Domains
+4.2 Plemelj-Calderón problem on chord-arc curves
 We now come back to Plemelj-Calderón problem.
-By a deep theorem of David [11] using the Lipschitz result [10], the Cauchy integral operator $T$ is bounded on $L^{p}(\Gamma)$ (i.e., $B_{p,p}^{0}(\Gamma)$) for every $p\in(1+\infty)$ (see section 1.3).
-Using it we will show the boundedness of the operator $T$ on $B_{p,p}^{1}(\Gamma)$.
+Recall that the Cauchy integral operator $T$ is bounded on $L^{p}(\Gamma)$ (i.e., $B_{p,p}^{0}(\Gamma)$) for every $p\in(1+\infty)$ (see subsection 1.3) if $\Gamma$ is chord-arc.
+Our goal in this subsection is to show that the operator $T$ is bounded on $B_{p,p}^{s}(\Gamma)$ for all $p\in(1,+\infty)$ and $s\in(0,1]$.
+We begin by proofing that the result is valid for $B_{p,p}^{1}(\Gamma)$ using that for $B_{p,p}^{0}(\Gamma)$.
 Theorem 4.2.
-The operator $T$ is bounded on the Sobolev space $B_{p,p}^{1}(\Gamma)$, and moreover, the operator norm on $B_{p,p}^{1}(\Gamma)$ equals to that on $L^{p}(\Gamma)$.
-Before we proceed to the proof of this theorem, one must properly define what we mean by $B^{1}_{p,p}(\Gamma)$ when $\Gamma$ is a chord-arc curve (or $K$-chord-arc curve for specifying its chord-arc constant).
-Without loss of generality we suppose $\text{length}(\Gamma)=2\pi$.
+The operator $T$ is bounded on $B_{p,p}^{1}(\Gamma)$, and moreover, the operator norm on $B_{p,p}^{1}(\Gamma)$ is equal to that on $L^{p}(\Gamma)$.
+Before we proceed to the proof of this theorem, one must properly define what we mean by $B^{1}_{p,p}(\Gamma)$ when $\Gamma$ is a chord-arc curve (or $K$-chord-arc curve to specify its chord-arc constant).
+Without loss of generality, we suppose $\text{length}(\Gamma)=2\pi$.
 Let $\varphi:\,\mathbb{D}\to\Omega$ be the Riemann map from the unit disk $\mathbb{D}$ onto the interior domain $\Omega$ of $\Gamma$, which can be extended to a homeomorphism of the closures such that $\varphi$ restricted on $\mathbb{T}$ is a quasisymmetry (see [39]).
 Let $t\mapsto z(t)$ denote the arc-length parametrization of $\Gamma$ and $\lambda(e^{it})=z(t)$ so that $\varphi=\lambda\circ h$ where $h$ is an absolutely continuous homeomorphism of $\mathbb{T}$ with $|h^{\prime}|=|\varphi^{\prime}|$.
-It is easy to see $\lambda$ is a bi-Lipschitz homeomorphism from $\mathbb{T}$ onto $\Gamma$.
+It is easy to see that $\lambda$ is a bi-Lipschitz homeomorphism from $\mathbb{T}$ onto $\Gamma$.
 Indeed, for any $t_{1},t_{2}\in[0,2\pi)$,
@@
 The space $B_{p,p}^{1}(\Gamma)$ is endowed with the natural norm $\|f\|_{B_{p,p}^{1}(\Gamma)}=\|f^{\prime}\|_{L^{p}(\Gamma)}$.
-To prepare for the proof we introduce the following definition.
-Set $\Gamma_{r}=\tau(|z|=r)$ to be "circular curves", where $\tau$ is a Riemann map that takes $\mathbb{D}$ onto $\Omega$.
-The Hardy space $E^{p}(\Omega)$ of index $p$ on the chord-arc domain $\Omega$ consists of holomorphic functions $F$ on $\Omega$ with a finite norm
+To prepare for the proof, we introduce the following definition.
+Set $\Gamma_{r}=\tau(|z|=r)$ as "circular curves", where $\tau$ is a Riemann map that takes $\mathbb{D}$ to $\Omega$.
+The Hardy space $E^{p}(\Omega)$ of the index $p$ in the chord-arc domain $\Omega$ consists of holomorphic functions $F$ in $\Omega$ with finite norm
 $\|F\|_{p}=\Bigg(\frac{1}{2\pi}\sup_{r}\int_{\Gamma_{r}}|F(w)|^{p}|dw|\Bigg)^{1/p}.$
@@
 on $\Gamma$.
-Here, $G_{i}$ and $G_{e}$ belong to the Hardy spaces of index $p$ on $\Omega_{i}$ and $\Omega_{e}$, respectively, and then $G_{i}$ and $G_{e}$ have non-tangential boundary limits almost everywhere on $\Gamma$ with respect to the arc-length measure, denoted still by $G_{i}$ and $G_{e}$, so that $\|G_{i}\|_{L^{p}(\Gamma)}$ and $\|G_{e}\|_{L^{p}(\Gamma)}$ are both controlled from above by $\|f\|_{B_{p,p}^{1}(\Gamma)}$.
+Here, $G_{i}$ and $G_{e}$ belong to the Hardy spaces of index $p$ in $\Omega_{i}$ and $\Omega_{e}$, respectively, and then $G_{i}$ and $G_{e}$ have non-tangential boundary limits almost everywhere on $\Gamma$ with respect to the arc-length measure, denoted still by $G_{i}$ and $G_{e}$, so that $\|G_{i}\|_{L^{p}(\Gamma)}$ and $\|G_{e}\|_{L^{p}(\Gamma)}$ are both controlled from above by $\|f\|_{B_{p,p}^{1}(\Gamma)}$.
 Recall that $G_{i}\in E^{p}(\Omega_{i})\subset E^{1}(\Omega_{i})$, so that $G_{i}\circ\tau\tau^{\prime}\in E^{1}(\mathbb{D}_{i})$.
-Let now $\Phi_{i}$ be an anti-derivative of $G_{i}$ on $\Omega_{i}$.
+Let now $\Phi_{i}$ be an anti-derivative of $G_{i}$ in $\Omega_{i}$.
 Then we see $(\Phi_{i}\circ\tau)^{\prime}=G_{i}\circ\tau\tau^{\prime}\in E^{1}(\mathbb{D}_{i})$.
@@
 on $\Gamma$ with the norm of $\Phi_{i,e}$ in $B^{1}_{p,p}(\Gamma)$ controlled from above by $\|f\|_{B_{p,p}^{1}(\Gamma)}$.
-By Plemelj formula we conclude that the operator $T$ is bounded on $B_{p,p}^{1}(\Gamma)$ with respect to the norm $\|\cdot\|_{B_{p,p}^{1}(\Gamma)}$, and moreover, we have $\|T\|_{B_{p,p}^{1}(\Gamma)\to B_{p,p}^{1}(\Gamma)}=\|T\|_{L^{p}(\Gamma)\to L^{p}(\Gamma)}$.
+By Plemelj formulae we conclude that the operator $T$ is bounded on $B_{p,p}^{1}(\Gamma)$ with respect to the norm $\|\cdot\|_{B_{p,p}^{1}(\Gamma)}$, and moreover, we have $\|T\|_{B_{p,p}^{1}(\Gamma)\to B_{p,p}^{1}(\Gamma)}=\|T\|_{L^{p}(\Gamma)\to L^{p}(\Gamma)}$.
 ∎
-Now we get ready to show the main result of this section: the operator $T$ is bounded on $B_{p,p}^{s}(\Gamma)$ for $0<s<1$.
-Because of the isomorphism between $B_{p,p}^{s}(\Gamma)$ and $B_{p,p}^{s}(\mathbb{T})$, this boils down to proving that the operator with kernel
+Now we get ready to show the main result of this subsection: the operator $T$ is bounded on $B_{p,p}^{s}(\Gamma)$ for $0<s<1$.
+Due to the isomorphism between $B_{p,p}^{s}(\Gamma)$ and $B_{p,p}^{s}(\mathbb{T})$, this boils down to proving that the operator with kernel
 $\text{p.v.}\frac{\lambda^{\prime}(\zeta)}{\lambda(\zeta)-\lambda(\xi)}$
@@
 The following functorial property of complex interpolation is a basic assertion in interpolation theory.
-Let $\{A_{0},A_{1}\}$ and $\{B_{0},B_{1}\}$ be two interpolation couples of Banach spaces and let $L$ be a linear operator mapping from $A_{0}+A_{1}$ into $B_{0}+B_{1}$ such that its restriction to $A_{j}$ is a linear and bounded operator from $A_{j}$ into $B_{j}$ with norm $M_{j}$, where $j=0,1$.
-Then the restriction of $L$ to $A_{0},A_{1}]_{\theta}$, $0<\theta<1$ is a linear and bounded operator from $A_{0},A_{1}]_{\theta}$ into $B_{0},B_{1}]_{\theta}$ with norm $M_{\theta}$.
-Further, it is known that the complex interpolation method is an exact interpolation functor of exponent $\theta$ (see e.g.
-p.88 in [4]) meaning that
+Let $(A_{0},A_{1})$ and $(B_{0},B_{1})$ be two interpolation couples of Banach spaces, and let $L$ be a linear operator mapping from $A_{0}+A_{1}$ to $B_{0}+B_{1}$ such that its restriction to $A_{j}$ is a linear and bounded operator from $A_{j}$ to $B_{j}$ with the norm $M_{j}$, where $j=0,1$.
+Then the restriction of $L$ to $A_{0},A_{1}]_{\theta}$, $0<\theta<1$, is a linear and bounded operator from $A_{0},A_{1}]_{\theta}$ to $B_{0},B_{1}]_{\theta}$ with the norm $M_{\theta}$.
+Furthermore, it is known that the complex interpolation method is an exact interpolation functor of exponent $\theta$ (see, e.g., p.88 in [4]) meaning that
 $M_{\theta}\leq M_{0}^{1-\theta}M_{1}^{\theta}.$ (4.2)
@@
 Proof.
-By taking $s_{0}=0$ and $s_{1}=1$ in Theorem 4.3, we have that
+Taking $s_{0}=0$ and $s_{1}=1$ in Theorem 4.3, we have
 $B_{p,p}^{s}(\mathbb{T})=[L^{p}(\mathbb{T}),B_{p,p}^{1}(\mathbb{T})]_{s}.$ (4.3)
@@
 ∎
-It follows from Plemelj formula that every function $f\in B_{p,p}^{s}(\Gamma)$ may be written uniquely as $f=\Phi_{i}+\Phi_{e}$ with $\|\Phi_{i,e}\|_{B_{p,p}^{s}(\Gamma)}\leq C\|f\|_{B_{p,p}^{s}(\Gamma)}$ for some constant $C$, $\Phi_{i,e}$ being boundary values of holomorphic functions in $\Omega_{i}$ and $\Omega_{e}$, respectively.
-The uniqueness of decomposition for $0<s\leq 1$ just follows from the case of $s=0$.
-4.2.1 On a Theorem of Murai
+It follows from Plemelj formulae that every function $f\in B_{p,p}^{s}(\Gamma)$ can be written uniquely as $f=\Phi_{i}+\Phi_{e}$ with $\|\Phi_{i,e}\|_{B_{p,p}^{s}(\Gamma)}\leq C\|f\|_{B_{p,p}^{s}(\Gamma)}$ for some constant $C$, $\Phi_{i,e}$ being the boundary values of the holomorphic functions in $\Omega_{i}$ and $\Omega_{e}$, respectively.
+The uniqueness of the decomposition for $0<s\leq 1$ follows from the case of $s=0$.
+4.3 Plemelj-Calderón problem on Lipschitz curves
 Clearly, Theorem 4.4 is also valid for chord-arc curves $\Gamma$ passing through $\infty$, and in particular for Lipschitz curves $\Gamma$ (i.e., there is a Lipschitz function $A:\mathbb{R}\to\mathbb{R}$ whose graph is $\Gamma$).
-A theorem by Murai [31] states that for a Lipschitz curve $\Gamma$ with Lipschitz norm $M$, $\|T\|_{L^{p}(\Gamma)\to L^{p}(\Gamma)}$ is no more than $C(1+M)^{3/2}$ (see also [12]), where $C$ is a universal constant.
-If we plug this information in the preceding theorem we obtain the same bound for $\|T\|_{B_{p,p}^{1-1/p}(\Gamma)\to B_{p,p}^{1-1/p}(\Gamma)}$.
+A theorem by Murai [35] states that for a Lipschitz curve $\Gamma$ with the Lipschitz norm $M$, $\|T\|_{L^{p}(\Gamma)\to L^{p}(\Gamma)}$ is no more than $C(1+M)^{3/2}$ (see also [12]), where $C$ is a universal constant.
+If we plug this information into the preceding theorem, we obtain the same bound for $\|T\|_{B_{p,p}^{1-1/p}(\Gamma)\to B_{p,p}^{1-1/p}(\Gamma)}$.
 But the norm $\|T\|_{B_{p,p}^{1-1/p}(\Gamma)\to B_{p,p}^{1-1/p}(\Gamma)}$ depends only on the $L^{p}(\mathbb{C})$-boundedness of the Beurling transform (see Section 2), so that it is actually independent of $M$.
@@
 Theorem 4.5.
-If $\Gamma$ is a Lipschitz curve with Lipschitz constant $M$ then we have, for $0<s<1$,
+If $\Gamma$ is a Lipschitz curve with Lipschitz constant $M$, then we have, for $0<s<1$,
 $\|T\|_{B_{p,p}^{s}(\Gamma)\to B_{p,p}^{s}(\Gamma)}\leq C(1+M)^{\frac{3}{2}|1-\frac{ps}{p-1}|}$
@@
 5 Douglas versus Littlewood-Paley: the chord-arc case
+Let $\Gamma$ be a chord-arc curve of length $2\pi$ with 0 in its inner domain $\Omega$, and $1<p<\infty$, $0<s<1$.
+Recall that we have two ways of generalizing fractional Besov-Sobolev spaces from the unit circle $\mathbb{T}$ to $\Gamma$.
+The first one can be called the Douglas way: we define $B_{p,p}^{s}(\Gamma)$ as the set of functions $f\in L^{p}(\Gamma)$ such that
+$\|f\|_{B_{p,p}^{s}(\Gamma)}^{p}:=\iint_{\Gamma\times\Gamma}\frac{|f(z)-f(\zeta)|^{p}}{|z-\zeta|^{1+ps}}|dz||d\zeta|<\infty,$
+which are intermediate spaces between $B_{p,p}^{0}(\Gamma)$ (i.e., $L^{p}(\Gamma)$) and $B_{p,p}^{1}(\Gamma)$.
+Let $B_{p,p}^{s}(\Omega)$, $0\leq s\leq 1$, be the set of harmonic extensions $u$ in $\Omega$ of all $f\in B_{p,p}^{s}(\Gamma)$.
+Then $B_{p,p}^{0}(\Omega)=E^{p}(\Omega)\oplus\overline{E^{p}(\Omega)}$.
+More precisely, there exists a unique pair of holomorphic functions $\varphi$ and $\psi$ in $E^{p}(\Omega)$ with $\varphi(0)=u(0)$ and $\psi(0)=0$ such that $u=\varphi+\overline{\psi}$.
+Since $\Gamma$ is rectifiable, it makes sense to speak of a tangential direction almost everywhere, and each function $\varphi$ in $E^{p}(\Omega)$ has a non-tangential limit almost everywhere on $\Gamma$, and conversely, $\varphi$ is the harmonic extension of its boundary limit function.
+Furthermore, since $\Omega$ is a chord-arc domain, in particular a Smirnov domain, $E^{p}(\Omega)$ coincides with the $L^{p}(\Gamma)$ closure of polynomials.
+Based on this, $E^{p}(\Omega)$ can be identified with the set of its boundary functions, so it is reasonable to identify $B_{p,p}^{0}(\Omega)$ and $B_{p,p}^{0}(\Gamma)$.
+Finally, for any $0\leq s\leq 1$, since the spaces $B_{p,p}^{s}(\Omega)$ decrease with $s$, we can identify $B_{p,p}^{s}(\Omega)$ and $B_{p,p}^{s}(\Gamma)$ for convenience and switch between the two freely.
+The second one is defined by Littlewood-Paley theory: consider the set of functions $f\in L^{p}(\Gamma)$ such that its harmonic extension $u$ in $\Omega$ satisfies
+$\|u\|_{\mathcal{B}_{p,p}^{s}(\Omega)}^{p}:=\iint_{\Omega}|\nabla u(z)|^{p}d(z,\Gamma)^{(1-s)p-1}dxdy<\infty.$
+The set of such functions $u$ is denoted by $\mathcal{B}_{p,p}^{s}(\Omega)$.
+In this section, our main goal is to prove $B_{p,p}^{s}(\Gamma)=\mathcal{B}_{p,p}^{s}(\Omega)$ under some condition on $\Gamma$, or, more precisely, $\|f\|_{B_{p,p}^{s}(\Gamma)}\simeq\|u\|_{\mathcal{B}_{p,p}^{s}(\Omega)}$.
+Here, the implicit constant depends only on $\Gamma$, $s$ and $p$.
 5.1 The conjugate operator
-Let $\Gamma$ be a chord-arc curve of length $2\pi$, and $1<p<\infty$, $0<s<1$.
-We have seen two ways of generalizing the fractional Sobolev spaces from $\mathbb{R}$ (or from $\mathbb{T}$) to $\Gamma$.
-The first one, which can be called the Douglas way is
-$B_{p,p}^{s}(\Gamma)=\bigg\{f\in L^{p}(\Gamma):\iint_{\Gamma\times\Gamma}\frac{|f(z)-f(\zeta)|^{p}}{|z-\zeta|^{1+ps}}|dz||d\zeta|<\infty\bigg\}.$
-The second one, via Littlewood-Paley theory, is
-$\mathcal{B}_{p,p}^{s}(\Gamma)=\bigg\{f\in L^{p}(\Gamma):\iint_{\Omega_{i,e}}|\nabla u_{i,e}(z)|^{p}d(z,\Gamma)^{(1-s)p-1}dxdy<\infty\bigg\}.$
-Here, $u_{i,e}$ stands for the harmonic extension of $f$ to $\Omega_{i,e}$.
-We have seen that Plemelj-Calderón problem is solvable in some sense on both spaces $B_{p,p}^{s}(\Gamma)$ and $\mathcal{B}_{p,p}^{s}(\Gamma)$ so that the natural question arises of whether these spaces coincide.
-Assume that $\Omega$ is bounded and $\mathbb{D}$ is the unit disk.
+Assume that $\Omega$ is a bounded chord-arc domain and $\mathbb{D}$ is the unit disk.
 Let $z_{0}$ be a point in $\Omega$.
 If $u$ is harmonic in $\Omega$ it is well known that there exists a unique harmonic function $\tilde{u}$ in $\Omega$ such that $\tilde{u}(z_{0})=0$ and $u+i\tilde{u}$ is holomorphic in $\Omega$.
-In the case of $\Omega=\mathbb{D}$, $\tilde{u}$ is the harmonic (Poisson) extension of $H(f)$ where $H$ is the Hilbert transform:
+In the case of $\Omega=\mathbb{D}$, $\tilde{u}$ is the harmonic extension (Poisson) of $H(f)$ where $H$ is the Hilbert transform:
 $H(f)(e^{i\theta})=\lim_{\epsilon\to 0}\frac{1}{2\pi}\int_{|\theta-\varphi|>\epsilon}\cot\Big(\frac{\theta-\varphi}{2}\Big)f(e^{i\varphi})d\varphi.$
-By Cauchy-Riemann equations, $|\nabla\tilde{u}|=|\nabla u|$, so that $\mathcal{B}_{p,p}^{s}(\Omega)$ is conjugate-invariant.
+By the Cauchy-Riemann equations, $|\nabla\tilde{u}|=|\nabla u|$.
+Thus, we see that $\mathcal{B}_{p,p}^{s}(\Omega)$ is conjugate-invariant.
 We denote the subspace of $\mathcal{B}_{p,p}^{s}(\Omega)$ consisting of holomorphic functions by $\mathcal{HB}_{p,p}^{s}(\Omega)$, i.e., the set of $u+i\tilde{u}$ with $u,\tilde{u}\in\mathcal{B}_{p,p}^{s}(\Omega)$.
-A necessary condition for $B_{p,p}^{s}(\Gamma)=\mathcal{B}_{p,p}^{s}(\Gamma)$ to hold is thus that $B_{p,p}^{s}(\Gamma)$ is stable by conjugation.
-Let $\varphi:\,\mathbb{D}\to\Omega$ be the Riemann map with $\varphi(0)=z_{0}$.
+A necessary condition for $B_{p,p}^{s}(\Gamma)=\mathcal{B}_{p,p}^{s}(\Omega)$ to hold is thus that $B_{p,p}^{s}(\Gamma)$ is stable by conjugation.
+Let $\varphi:\,\mathbb{D}\to\Omega$ be a Riemann map with $\varphi(0)=z_{0}$.
 Recall that $t\mapsto z(t)$ denotes the arc-length parametrization of $\Gamma$ and $\lambda(e^{it})=z(t)$ so that $\varphi=\lambda\circ h$ where $h$ is an absolutely continuous homeomorphism of $\mathbb{T}$ with $|h^{\prime}|=|\varphi^{\prime}|$.
 Recall that $\lambda$ is a bi-Lipschitz homeomorphism from $\mathbb{T}$ onto $\Gamma$.
-A simple computation shows that if $f\in B_{p,p}^{s}(\Gamma)$; that is $g=f\circ\lambda\in B_{p,p}^{s}(\mathbb{T})$, and $u$ is its harmonic extension to $\Omega$, then $\tilde{u}$ is the harmonic extension of $\tilde{f}=\tilde{g}\circ\lambda^{-1}$.
+A simple computation shows that if $f\in B_{p,p}^{s}(\Gamma)$; that is, $g=f\circ\lambda\in B_{p,p}^{s}(\mathbb{T})$ and $u$ is its harmonic extension to $\Omega$, then $\tilde{u}$ is the harmonic extension of $\tilde{f}=\tilde{g}\circ\lambda^{-1}$.
 Here,
@@
 where $V_{h}(g)=g\circ h$.
-It is known (see [9], p.247 in [16]) that $V_{h}^{-1}HV_{h}$ is bounded on $L^{p}(\mathbb{T})$ if and only if $|h^{\prime}|=|\varphi^{\prime}|$ belongs to $A_{p}$ on $\mathbb{T}$.
+It is known (see [9], p.247 in [16]) that $V_{h}^{-1}HV_{h}$ is bounded on $L^{p}(\mathbb{T})$, $1<p<\infty$, if and only if $|h^{\prime}|=|\varphi^{\prime}|$ belongs to $A_{p}$ on $\mathbb{T}$.
 Based on this comment on the limiting case $s=0$, we now state a theorem about the general case $0\leq s\leq 1$.
 Theorem 5.1.
-For $p>1,\,0\leq s\leq 1$, suppose $\Gamma$ is such that
+For $p>1,\,0\leq s\leq 1$, suppose that $\Gamma$ is such that
   • $|\varphi^{\prime}|\in A_{p}$ if $1<p\leq 2$;
   • $|\varphi^{\prime}|\in A_{p^{\prime}}$$(1/p^{\prime}+1/p=1)$ if $p>2$.
-Then, $B_{p,p}^{s}(\Gamma)$ is stable by conjugation, and moreover, the conjugate operator $f\mapsto\tilde{f}$ on $B_{p,p}^{s}(\Gamma)$ is bounded.
-Before going to the proof, let us comment on the hypothesis of Theorem 5.1.
-If $\omega$ is an $A_{\infty}$ weight let us define $\delta(\omega)=\inf\{q>1:\,\omega\in A_{q}\}$.
+Then $B_{p,p}^{s}(\Gamma)$ is stable by conjugation, and, moreover, the conjugate operator $f\mapsto\tilde{f}$ on $B_{p,p}^{s}(\Gamma)$ is bounded.
+By this theorem, we define $\text{HB}_{p,p}^{s}(\Gamma)$ as the subspace of $B_{p,p}^{s}(\Gamma)$ consisting of functions $f+i\tilde{f}$ with $f,\tilde{f}\in B_{p,p}^{s}(\Gamma)$, and $\text{HB}_{p,p}^{s}(\Omega)$ as the analytic subspace of $B_{p,p}^{s}(\Omega)$.
+Before going to the proof, let us comment on the hypotheses of Theorem 5.1.
+If $\omega$ is an $A_{\infty}$-weight, let us define $\delta(\omega)=\inf\{q>1:\,\omega\in A_{q}\}$.
 It is known [9] that if $\omega\in A_{q}$ for some $q>1$ then $\delta(\omega)<q$ (see p.254 in [16]).
-Let us then define $\delta=\delta(|\varphi^{\prime}|)$.
-Noting that the hypothesis imply that $|\varphi^{\prime}|\in A_{2}$, we see that the hypothesis of the theorem are equivalent to
-$\delta<q<\frac{\delta}{\delta-1}.$
-On the other hand, it is known (see [21]) that $\Gamma$ being chord-arc implies $|\varphi^{\prime}|$ having $A_{\infty}$ but there are examples of chord-arc curves such that $|\varphi^{\prime}|\notin A_{2}$ (see [23]), we thus know the chord-arc condition is not sufficient for Theorem 5.1 to hold.
+The above theorem means that given $p>1$ and the class of curves $\Gamma$ for which the claim is true depends on $p$.
+If we reverse the point of view and start with a curve $\Gamma$ satisfying $|\varphi^{\prime}|\in A_{2}$, then by defining $\delta=\delta(|\varphi^{\prime}|)$, the claim is true for
+$\delta<p<\frac{\delta}{\delta-1}.$
+On the other hand, it is known (see [21]) that $\Gamma$ being chord-arc implies $|\varphi^{\prime}|$ having $A_{\infty}$ but there are examples of chord-arc curves such that $|\varphi^{\prime}|\notin A_{2}$ (see [23]), we thus know that the chord-arc condition is not sufficient for Theorem 5.1 to hold.
 Proof.
@@
 Specially, a natural condition on the curve $\Gamma$ implying that $|\varphi^{\prime}|\in A_{2}$ is the radial-Lipschitz condition.
-Indeed, if $\Gamma$ is a radial-Lipschitz curve it can be shown that $|\varphi^{\prime}|$ satisfies the Helson-Szegö condition (see section 5.3 for a specific proof): $\log|\varphi^{\prime}|=u+Hv$ with $u\in L^{\infty},\;v\in L^{\infty}$ and $\|v\|_{\infty}<\pi/2$, which is equivalent to $|\varphi^{\prime}|\in A_{2}$ as follows from the Helson-Szegö theorem ([16]).
+Indeed, if $\Gamma$ is a radial-Lipschitz curve, then it can be shown that $|\varphi^{\prime}|$ satisfies the Helson-Szegö condition (see subsection 5.3 for a specific proof): $\log|\varphi^{\prime}|=u+Hv$ with $u\in L^{\infty},\;v\in L^{\infty}$ and $\|v\|_{\infty}<\pi/2$, which is equivalent to $|\varphi^{\prime}|\in A_{2}$ as follows from the Helson-Szegö theorem ([16]).
 We may now state:
 Corollary 5.2.
-For $0\leq s\leq 1$, if $\Gamma$ is such that $|\varphi^{\prime}|\in A_{2}$ then $B_{2,2}^{s}(\Gamma)$ is stable by conjugation, and moreover, the conjugate operator $f\mapsto\tilde{f}$ on $B_{2,2}^{s}(\Gamma)$ is bounded.
+For $0\leq s\leq 1$, if $\Gamma$ is such that $|\varphi^{\prime}|\in A_{2}$ then $B_{2,2}^{s}(\Gamma)$ is stable by conjugation, and moreover, the conjugate operator $f\mapsto\tilde{f}$ in $B_{2,2}^{s}(\Gamma)$ is bounded.
 The conclusion holds in particular for radial-Lipschitz curves.
-For $0\leq s\leq 1$, $1<p<\infty$, we denote the subspace of $B_{p,p}^{s}(\Gamma)$ consisting of functions $f+i\tilde{f}$ such that $f,\;\tilde{f}\in B_{p,p}^{s}(\Gamma)$ by $\text{HB}_{p,p}^{s}(\Gamma)$.
 5.2 The operator $V_{s,p}$ on $\mathcal{B}^{s}_{p,p}(\Omega)$
-Recall that $\mathcal{B}^{s}_{p,p}(\Omega)$, $0\leq s<1$ and $1<p<\infty$, is the set of harmonic functions $u$ on the domain $\Omega$ bounded by the Jordan curve $\Gamma$ such that its norm $\|u\|_{\mathcal{B}^{s}_{p,p}(\Omega)}<\infty$.
-Suppose $\Gamma$ is chord-arc which in particular is rectifiable.
-It is known that the non-tangential boundary values $f$ of $u$ exist almost everywhere on $\Gamma$ with respect to the arc-length measure and $u$ can be recovered from $f$ using the harmonic extension.
-The space of boundary functions of $\mathcal{B}^{s}_{p,p}(\Omega)$ is denoted by $\mathcal{B}_{p,p}^{s}(\Omega\!\to\!\Gamma)$ so that $\mathcal{B}_{p,p}^{s}(\Gamma)=\mathcal{B}_{p,p}^{s}(\Omega_{i}\!\to\!\Gamma)\cap\mathcal{B}_{p,p}^{s}(\Omega_{e}\!\to\!\Gamma)$.
-We may identify $\mathcal{B}_{p,p}^{s}(\Omega)$ with $\mathcal{B}_{p,p}^{s}(\Omega\!\to\!\Gamma)$ for convenience, and will switch between the two freely.
-Similarly, for $0\leq s\leq 1$ and $1<p<\infty$ we also identify $B_{p,p}^{s}(\Gamma)$ with the set $B_{p,p}^{s}(\Omega)$ of its harmonic extensions in $\Omega$.
 Let $\Omega,\,\Omega^{\prime}$ be two Jordan domains containing 0 and $\varphi$ a holomorphic diffeomorphism from $\Omega$ onto $\Omega^{\prime}$ fixing 0.
 Let $f\in\mathcal{HB}_{p,p}^{s}(\Omega^{\prime})$.
-Using the change of variable $\zeta=\varphi(z)$ we have
+Using the change of variables $\zeta=\varphi(z)$ we have
 $\iint_{\Omega^{\prime}}d(\zeta,\Gamma^{\prime})^{(1-s)p-1}|f^{\prime}(\zeta)|^{p}d\xi d\eta=\iint_{\Omega}d(\varphi(z),\Gamma^{\prime})^{(1-s)p-1}|(f\circ\varphi)^{\prime}(z)|^{p}|\varphi^{\prime}(z)|^{2-p}dxdy.$
-By the Koebe distortion theorem (2.1) there exists a universal constant $C>1$ such that
+By the Koebe distortion theorem (see (2.1)) there exists a universal constant $C>1$ such that
 $C^{-|(1-s)p-1|}\iint_{\Omega}d(z,\Gamma)^{(1-s)p-1}|V_{s,p}(f)^{\prime}(z)|^{p}dxdy$
@@
 $\|V_{s,p}\|,\,\|V_{s,p}^{-1}\|\leq C^{|(1-s)-1/p|}.$
-Notice that for s=11/ps=1-1/p, this operator is nothing but the composition by $\varphi$ and that, in this case, $V_{1-1/p,p}$ is an isometry.
-For convenience of later use, when $p=2$, the operator $V_{s,p}$ will be simplified to be $V_{s}$.
-5.2.1 The Operator $V_{s}$ on $\mathcal{B}^{s}_{2,2}(\Omega)$
-In order to understand better the operator $V_{s}$ let us rewrite it by using an integration by parts:
+Notice that for s=11/ps=1-1/p, this operator is nothing more than the composition by $\varphi$ and that, in this case, $V_{1-1/p,p}$ is an isometry.
+For convenience of later use, when $p=2$, the operator $V_{s,p}$ will be simplified to $V_{s}$.
+In order to better understand the operator $V_{s}$ let us rewrite it by using an integration by parts:
 $V_{s}(f)(z)$ $=f\circ\varphi(z)\varphi^{\prime}(z)^{1/2-s}-(1/2-s)\int_{0}^{z}f\circ\varphi(u)\varphi^{\prime}(u)^{1/2-s}\frac{\varphi^{\prime\prime}(u)}{\varphi^{\prime}(u)}du$
@@
 where the supremum is taken over all sub-arcs $I$ of $\mathbb{T}$ and $g_{I}$ denotes the integral mean of $g$ over $I$.
-It is known that $\log\varphi^{\prime}\in{\rm BMOA}(\mathbb{D})$ if $\varphi(\mathbb{D})$ is chord-arc, but not vice versa.
+It is known that $\log\varphi^{\prime}\in{\rm BMOA}(\mathbb{D})$ if the domain $\varphi(\mathbb{D})$ is chord-arc, but not vise versa.
 By Fefferman-Stein, $\log\varphi^{\prime}\in{\rm BMOA}(\mathbb{D})$ if and only if
 $d\mu=\bigg|\frac{\varphi^{\prime\prime}(z)}{\varphi^{\prime}(z)}\bigg|^{2}(1-|z|)dxdy$
-is a Carleson measure in $\mathbb{D}$ (see e.g. Ch.VI in [16]).
+is a Carleson measure in $\mathbb{D}$ (see, e.g., Ch.VI in [16]).
 Theorem 5.3.
@@
 Suppose $\log\varphi^{\prime}\in{\rm BMOA}(\mathbb{D})$; that is, $\mu$ is a Carleson measure in $\mathbb{D}$.
-Let $f\in E^{2}(\Omega)$ so that $T_{0}f\in E^{2}(\mathbb{D})$.
-Then by Carleson (see e.g.
-Theorem 3.9 in [16])
+Let $f\in E^{2}(\Omega)$ be such that $T_{0}f\in E^{2}(\mathbb{D})$.
+Then by Carleson (see, e.g., Theorem 3.9 in [16])
 $\iint_{\mathbb{D}}|T_{0}(f)|^{2}d\mu\leq C\|T_{0}f\|_{E^{2}(\mathbb{D})}^{2}$
@@
 $<\infty.$
-Combined with (5.2), this completes the proof of statement $(1)$.
+Combined with (5.2), this completes the proof of the statement $(1)$.
 For the proof of statement $(2)$ we will need the definition of the (Littlewood-Paley) $\mathsf{g}$-function of a holomorphic function $f$ in $\mathbb{D}$:
@@
 $\iint_{\mathbb{D}}(1-|u|)|(f\circ\varphi)^{\prime}(u)|^{2}|\varphi^{\prime}(u)|dudv<\infty,$
-or, in other words, that $\mathsf{g}(f\circ\varphi)\in L^{2}(\mathbb{T},|\varphi^{\prime}|d\theta)$.
+or, in other words, $\mathsf{g}(f\circ\varphi)\in L^{2}(\mathbb{T},|\varphi^{\prime}|d\theta)$.
 Since $\Gamma$ is assumed to be chord-arc, we have that $|\varphi^{\prime}|$ has the weight $A_{\infty}$ on $\mathbb{T}$.
 We can then apply a theorem of Gundy-Wheeden [19] (see also [21]) which implies that $\mathsf{g}(f\circ\varphi)\in L^{2}(\mathbb{T},|\varphi^{\prime}|d\theta)$ if and only if the non-tangential maximal function $\mathsf{n}(f\circ\varphi)$ of $f\circ\varphi$ is in $L^{2}(\mathbb{T},|\varphi^{\prime}|d\theta)$.
-On the other hand, for any "circular curves" $\Gamma_{r}$ it holds that (see p.233 in [21])
+On the other hand, for any "circular curves" $\Gamma_{r}$ we have that (see p.233 in [21])
 $\int_{\Gamma_{r}}|f(\zeta)|^{2}d\sigma(\zeta)\leq C\|\mathsf{n}(f\circ\varphi)\|^{2}_{L^{2}(\mathbb{T},|\varphi^{\prime}|d\theta)},$
 which implies $f\in E^{2}(\Omega)$. ∎
-Notice that for this theorem we do not need $|\varphi^{\prime}|$ having $A_{2}$ on $\mathbb{T}$.
+Notice that for this theorem, we do not need $|\varphi^{\prime}|$ having $A_{2}$ on $\mathbb{T}$.
 This condition is nevertheless necessary for the following corollary to hold by Corollary 5.2 since if it is not attached then the (real) space of real parts of $E^{2}(\Omega\!\to\!\Gamma)$-functions is a proper subspace of the real space $L_{\mathbb{R}}^{2}(\Gamma,d\sigma)$
 Corollary 5.4.
-Let $\Omega$ be a chord-arc domain bounded by $\Gamma$. If $|\varphi^{\prime}|\in A_{2}$ then $B_{2,2}^{0}(\Gamma)=\mathcal{B}_{2,2}^{0}(\Omega\!\to\!\Gamma)$.
-5.3 The Equality $\mathcal{B}_{2,2}^{s}(\Omega\!\to\!\Gamma)=B_{2,2}^{s}(\Gamma)$
-The aim of this part is to prove the following theorem that is one of the main theorems of this paper.
+Let $\Omega$ be a chord-arc domain bounded by $\Gamma$. If $|\varphi^{\prime}|\in A_{2}$ then $B_{2,2}^{0}(\Gamma)=\mathcal{B}_{2,2}^{0}(\Omega)$.
+5.3 The equality $\mathcal{B}_{2,2}^{s}(\Omega)=B_{2,2}^{s}(\Gamma)$
+In this subsection, our goal is to prove the following theorem, which is one of the main theorems in this paper.
 Theorem 5.5.
@@
 $\log|\varphi^{\prime}|=b_{1}+Hb_{2},\qquad b_{1},\;Hb_{1},\;b_{2}\in L^{\infty}(\mathbb{T})\;\text{and}\;\|b_{2}\|_{\infty}<\frac{\pi}{2},$ (5.3)
-then, for $0\leq s\leq 1$, $\mathcal{B}_{2,2}^{s}(\Omega\!\to\!\Gamma)=B_{2,2}^{s}(\Gamma)$ with comparable norms. The conclusions hold in particular for radial-Lipschitz domains.
-Before going further, let us proceed from understanding why the conditions (5.3) are assumed.
+then, for $0\leq s\leq 1$, $\mathcal{B}_{2,2}^{s}(\Omega)=B_{2,2}^{s}(\Gamma)$ with comparable norms. The conclusions hold in particular for the radial-Lipschitz domains.
+Before going further, let us start by understanding why conditions (5.3) are assumed.
 The Helson-Szegö theorem (see Ch.IV in [16]) says that $\log|\varphi^{\prime}|=b_{1}+Hb_{2}$ with $b_{1},b_{2}\in L^{\infty}(\mathbb{T})$ and $\|b_{2}\|_{\infty}<\pi/2$ if and only if $|\varphi^{\prime}|\in A_{2}$ on $\mathbb{T}$, which is only used to guarantee that $B_{2,2}^{s}(\Gamma)$ is stable by conjugation for all $0\leq s\leq 1$ (see Corollary 5.2).
-For the proof of the corresponding analytic case: $\mathcal{HB}_{2,2}^{s}(\Omega\!\to\!\Gamma)=\text{HB}_{2,2}^{s}(\Gamma)$, this condition can thus be ignored.
 Note that $\text{Arg}\varphi^{\prime}=H\log|\varphi^{\prime}|=Hb_{1}-b_{2}$.
 The extra assumption that $Hb_{1}\in L^{\infty}(\mathbb{T})$ is used to derive that $\text{Arg}\varphi^{\prime}$ is bounded, geometrically which means that the boundary curve cannot spiral too much.
-For instance, for the double spiral curve, $\text{Arg}\varphi^{\prime}$ keeps increasing and goes to infinity.
+For example, for the double spiral curve, $\text{Arg}\varphi^{\prime}$ keeps increasing and goes to infinity.
 Let $r(\theta):\mathbb{R}\to\mathbb{(}0,+\infty)$ be a $2\pi$-periodic continuous function.
 Then the curve $\Gamma=\{r(\theta)e^{i\theta}:\;\theta\in[0,2\pi)\}$ is called a starlike Jordan curve with respect to 0 and the domain $\Omega$ bounded by $\Gamma$ is called a starlike domain.
-Recall that if moreover $r(\theta)$ is a Lipschitz function, i.e., there exists a positive constant $M$ such that $\|r^{\prime}\|_{\infty}\leq M$ then the curve $\Gamma$ is called a radial-Lipschitz curve and $\Omega$ is called a radial-Lipschitz domain.
-Suppose $\varphi$ is a Riemann mapping from $\mathbb{D}$ onto $\Omega$ fixing 0.
+Recall that if, moreover, $r(\theta)$ is a Lipschitz function, i.e., there exists a positive constant $C(\Gamma)$ such that $\|r^{\prime}\|_{\infty}\leq C(\Gamma)$ then the curve $\Gamma$ is called a radial-Lipschitz curve and $\Omega$ is called a radial-Lipschitz domain.
+Suppose $\varphi$ is a Riemann mapping from $\mathbb{D}$ to $\Omega$ that fixes 0.
 It is known that $\Omega$ is starlike with respect to 0 if and only if
@@
 The condition (5.4), combined with this, shows that $\text{Arg}\varphi^{\prime}\in L^{\infty}$.
-A simple computation shows that $\Omega$ is a radial-Lipschitz domain if and only if the condition (5.4) holds where the constant $\alpha<1$ depends on $M$, from which it follows that $\Omega$ is a chord-arc domain (see p.172 in [35]).
+A simple computation shows that $\Omega$ is a radial-Lipschitz domain if and only if condition (5.4) holds where the constant $\alpha<1$ depends on $C(\Gamma)$, from which it follows that $\Omega$ is a chord-arc domain (see p.172 in [39]).
 Notice that
@@
 $H\left(\log\left|\frac{\varphi(z)}{z}\right|\right)=\text{Arg}\frac{\varphi(z)}{z}.$
-We have that a radial-Lipschitz curve $\Gamma$ satisfies the condition (5.3). Consequently, the radial-Lipschitz domain has all properties of Theorem 5.5.
-The following abstract interpolation Theorem of Voigt [43] is the main tool in our proof of Theorem 5.5:
+We have that a radial-Lipschitz curve $\Gamma$ satisfies the condition (5.3). Consequently, the radial-Lipschitz domain has all the properties of Theorem 5.5.
+The following abstract interpolation Theorem of Voigt [49] is the main tool in our proof of Theorem 5.5 and subsequent Theorem 5.7:
 Theorem 5.6.
@@
 We now start the proof of Theorem 5.5.
-Clearly, we only need to show $\mathcal{HB}_{2,2}^{s}(\Omega\!\to\!\Gamma)=\text{HB}_{2,2}^{s}(\Gamma)$ with comparable norms.
-Recall that Bergman spaces with standard weights $A^{p}_{\alpha}$ are defined as the sets of holomorphic functions $g$ on $\mathbb{D}$ such that $g\in L^{p}(\mathbb{D},(1-|z|^{2})^{\alpha}dxdy)$, i.e.,
+Clearly, we only need to show $\mathcal{HB}_{2,2}^{s}(\Omega)=\text{HB}_{2,2}^{s}(\Omega)$ with comparable norms.
+Recall that Bergman spaces with standard weights $A^{p}_{\alpha}$ are defined as the sets of holomorphic functions $g$ in the unit disk $\mathbb{D}$ such that $g\in L^{p}(\mathbb{D},(1-|z|^{2})^{\alpha}dxdy)$, i.e.,
 $\|g\|_{A^{p}_{\alpha}}^{p}=\iint_{\mathbb{D}}|g(z)|^{p}(1-|z|^{2})^{\alpha}dxdy<\infty$
@@
 Here, the assumption that $\alpha>-1$ is essential because the space $L^{p}(\mathbb{D},(1-|z|^{2})^{\alpha}dxdy)$ does not contain any holomorphic function other than 0 when $\alpha\leq-1$.
-The definition of $A^{p}_{\alpha}$ is extended to the case where $\alpha$ is any real number; that is consistent with the traditional definition when $\alpha>-1$ (see Theorem 13 in [47]).
-From the definition it can be seen that if $\alpha<\beta$ then the strict inclusion $A^{p}_{\alpha}\subset A^{p}_{\beta}$ holds.
-For our purpose we in particular mention that the space $A^{2}_{-1}$ is the classical Hardy space $E^{2}(\mathbb{D})$, i.e., that is the set of holomorphic functions $g$ in $\mathbb{D}$ such that
+The definition of $A^{p}_{\alpha}$ is extended to the case where $\alpha$ is any real number; that is consistent with the traditional definition of $\alpha>-1$ (see Theorem 13 in [53]).
+From the definition it can be seen that if $\alpha<\beta$ then the strict inclusion $A^{p}_{\alpha}\subset A^{p}_{\beta}$ is satisfied.
+For our purpose, we in particular mention that the space $A^{2}_{-1}$ is the classical Hardy space $E^{2}(\mathbb{D})$, that is, the set of holomorphic functions $g$ in $\mathbb{D}$ such that
 $\int_{\mathbb{D}}|g^{\prime}(z)|^{2}(1-|z|^{2})dxdy<\infty.$
@@
 with equivalent norms.
-Suppose $\varphi$ is a conformal map from $\mathbb{D}$ onto $\Omega$ fixing 0. For any $f\in\mathcal{HB}_{2,2}^{s}(\Omega)$, $0\leq s<1$, we have seen from (5.2) that
+Suppose that $\varphi$ is a conformal map from $\mathbb{D}$ onto $\Omega$ fixing 0. For any $f\in\mathcal{HB}_{2,2}^{s}(\Omega)$, $0\leq s<1$, we have seen from (5.2) that
 $\iint_{\mathbb{D}}|V_{s}(f)^{\prime}(z)|^{2}(1-|z|)^{1-2s}dxdy\simeq\iint_{\Omega}|f^{\prime}(\zeta)|^{2}d(\zeta,\Gamma)^{1-2s}d\xi d\eta$
@@
 with comparable norms. Note that any $g\in A^{2}_{1-2s}$ can be written in the form $V_{s}(f)^{\prime}$ by picking $f(z)=\int_{0}^{z}g\circ\varphi^{-1}(u)(\varphi^{-1})^{\prime}(u)^{3/2-s}du$.
-Motivated by equivalence (5.6), we define $\mathcal{HB}_{2,2}^{1}(\Omega)$ as the set of analytic functions $f$ on $\Omega$ satisfying $V_{1}(f)^{\prime}\in A^{2}_{-1}$.
+Motivated by equivalence (5.6), we define $\mathcal{HB}_{2,2}^{1}(\Omega)$ as the set of analytic functions $f$ in $\Omega$ satisfying $V_{1}(f)^{\prime}\in A^{2}_{-1}$.
 It is easy to see that $f$ is an anti-derivative of a function in $E^{2}(\Omega)$.
@@
 $\int_{\Gamma}|f^{\prime}(\zeta)|^{2}|d\zeta|=\int_{\mathbb{T}}|f^{\prime}\circ\varphi(z)\varphi^{\prime}(z)^{1/2}|^{2}|dz|=\int_{\mathbb{T}}|V_{1}(f)^{\prime}(z)|^{2}|dz|<\infty.$
-By this definition, $\mathcal{HB}_{2,2}^{1}(\Omega\!\to\!\Gamma)=\text{HB}_{2,2}^{1}(\Gamma)$, and we also have $\mathcal{HB}_{2,2}^{0}(\Omega\!\to\!\Gamma)=\text{HB}_{2,2}^{0}(\Gamma)$ with equivalent norms by Theorem 5.3.
-Combining this with (5.6), we can see that this boils down to prove
+By this definition, $\mathcal{HB}_{2,2}^{1}(\Omega)=\text{HB}_{2,2}^{1}(\Omega)$.
+Recall that we also have $\mathcal{HB}_{2,2}^{0}(\Omega)=\text{HB}_{2,2}^{0}(\Omega)$ with equivalent norms by Theorem 5.3.
+Combining this with (5.6), we can see that this boils down to proving
 $f\in\text{HB}_{2,2}^{s}(\Omega)\;\Leftrightarrow\;V_{s}(f)^{\prime}\in A^{2}_{1-2s}$ (5.7)
 with comparable norms for $0<s<1$. For this purpose, we will use Theorem 5.6 in the following.
-Recall that $(A^{2}_{1},A^{2}_{-1})$ is an interpolation pair and $A^{2}_{1},A^{2}_{-1}]_{s}=A^{2}_{1-2s},\;0<s<1$.
-Notice that the boundary curve $\Gamma$ of $\Omega$ is a chord-arc curve which is the image of $\mathbb{T}$ under the bi-Lipschitz mapping $\lambda$, we then conclude by (4.3) that $(\text{HB}_{2,2}^{0}(\Gamma),\text{HB}_{2,2}^{1}(\Gamma))$ is an interpolation pair and $\text{HB}_{2,2}^{0}(\Gamma),\text{HB}_{2,2}^{1}(\Gamma)]_{s}=\text{HB}_{2,2}^{s}(\Gamma)$.
+Recall that $(A^{2}_{1},A^{2}_{-1})$ is an interpolation pair, and $A^{2}_{1},A^{2}_{-1}]_{s}=A^{2}_{1-2s},\;0<s<1$.
+Notice that the boundary curve $\Gamma$ of $\Omega$ is a chord-arc curve which is the image of $\mathbb{T}$ under the bi-Lipschitz mapping $\lambda$, we then conclude by (4.3) that $(\text{HB}_{2,2}^{0}(\Omega),\text{HB}_{2,2}^{1}(\Omega))$ is an interpolation pair and $\text{HB}_{2,2}^{0}(\Omega),\text{HB}_{2,2}^{1}(\Omega)]_{s}=\text{HB}_{2,2}^{s}(\Omega)$.
 Note that $\Omega$ is a chord-arc domain, in particular a Smirnov domain, that is equivalent to that the Hardy space $E^{2}(\Omega)$ (i.e., $\text{HB}_{2,2}^{0}(\Omega)$) coincides with the $L^{2}(\Gamma)$ closure of the set $\check{B}$ of polynomials on $\Gamma$ (see Theorem 10.6 in [15]).
-Since $\text{HB}_{2,2}^{1}(\Gamma)$ is the set of functions $f\in E^{2}(\Omega)$ such that $f^{\prime}\in E^{2}(\Omega)$, the set $\check{B}$ is also a dense subspace of $\text{HB}_{2,2}^{1}(\Gamma)$.
-By the inclusion relation $\text{HB}_{2,2}^{1}(\Gamma)\subset\text{HB}_{2,2}^{0}(\Gamma)$, we have that $\text{HB}_{2,2}^{0}(\Gamma)\cap\text{HB}_{2,2}^{1}(\Gamma)=\text{HB}_{2,2}^{1}(\Gamma)$.
+Since $\text{HB}_{2,2}^{1}(\Omega)$ is the set of functions $f\in E^{2}(\Omega)$ such that $f^{\prime}\in E^{2}(\Omega)$, the set $\check{B}$ is also a dense subspace of $\text{HB}_{2,2}^{1}(\Omega)$.
+By the inclusion relation $\text{HB}_{2,2}^{1}(\Omega)\subset\text{HB}_{2,2}^{0}(\Omega)$, we have $\text{HB}_{2,2}^{0}(\Omega)\cap\text{HB}_{2,2}^{1}(\Omega)=\text{HB}_{2,2}^{1}(\Omega)$.
 Define a family of linear mappings $(T_{z};\;z\in S)$ on $\check{B}$ by
@@
 Here, the constant $C_{1}$ depends on $f$ and $\Omega$.
-Then,
+Then
 $\iint_{\mathbb{D}}|T_{z}(f)(\zeta)|^{2}(1-|\zeta|)d\xi d\eta\leq C_{1}e^{2c}\left(e^{-ct^{2}}e^{t\|\text{Arg}\varphi^{\prime}\|_{\infty}}\right)^{2}\iint_{\mathbb{D}}|\varphi^{\prime}|^{2(1-s)}|\varphi^{\prime}(\zeta)|(1-|\zeta|)d\xi d\eta$
 Set $A:=\|\text{Arg}\varphi^{\prime}\|_{\infty}$ and pick $c\geq A^{2}/4$ we then have that $e^{-ct^{2}}e^{tA}\leq e$.
-By the Koebe distortion theorem (2.1), $|\varphi^{\prime}(\zeta)|(1-|\zeta|)$ is comparable to $d(\varphi(\zeta),\;\Gamma)$, that is bounded due to the boundedness of the diameter of $\Omega$.
+By the Koebe distortion theorem (see (2.1)), $|\varphi^{\prime}(\zeta)|(1-|\zeta|)$ is comparable to $d(\varphi(\zeta),\;\Gamma)$, which is bounded due to the boundedness of the diameter of $\Omega$.
 By Hölder inequality,
 $\iint_{\mathbb{D}}|\varphi^{\prime}|^{2(1-s)}d\xi d\eta\leq\pi^{s}\left(\iint_{\mathbb{D}}|\varphi^{\prime}|^{2}d\xi d\eta\right)^{1-s}<\infty.$
-We can then conclude that the integral (5.8) converges, and moreover it has an upper-bound depending only on $\Omega$ and given $f$.
+We can then conclude that the integral (5.8) converges, and, moreover, it has an upper-bound depending only on $\Omega$ and given $f$.
 That implies that the linear mapping $T_{z}$ sends $\check{B}$ to $A^{2}_{1}$, and for any $f\in\check{B}$ the function $T_{(\cdot)}f:\;S\to A^{2}_{1}$ is bounded.
@@
 we then have
-$\int_{\mathbb{T}}|T_{1+it}(f)(\zeta)|^{2}|d\zeta|\leq e^{2(1+c)}\int_{\Gamma}|f^{\prime}(\zeta)|^{2}|d\zeta|\leq e^{2(1+c)}$
+$\int_{\mathbb{T}}|T_{1+it}(f)(\zeta)|^{2}|d\zeta|\leq e^{2(1+c)}\int_{\Gamma}|f^{\prime}(\zeta)|^{2}|d\zeta|\leq e^{2(1+c)}.$
 Thus,
 $M_{1}:=\sup\{\|T_{1+it}f\|_{A^{2}_{-1}};\;t\in\mathbb{R},f\in\check{B},\|f\|_{B_{2,2}^{1}(\Omega)}\leq 1\}<\infty.$
-By the continuity of parametrized integral, all functions $S\ni z\mapsto T_{z}f\in A^{2}_{1}$,   $\mathbb{R}\ni t\mapsto T_{it}f\in A^{2}_{1}$, and $\mathbb{R}\ni t\mapsto T_{1+it}f\in A^{2}_{-1}$ are continuous.
+By the continuity of the parametrized integral, all functions $S\ni z\mapsto T_{z}f\in A^{2}_{1}$,   $\mathbb{R}\ni t\mapsto T_{it}f\in A^{2}_{1}$, and $\mathbb{R}\ni t\mapsto T_{1+it}f\in A^{2}_{-1}$ are continuous.
 Finally, to prove that for any $f\in\check{B}$, $\stackrel{{\scriptstyle\circ}}{{S}}\ni z\mapsto T_{z}f\in A^{2}_{1}$ is a holomorphic map, we use a general result about the infinite dimensional holomorphy (see p.206 in [25]).
@@
 For each $\zeta\in\mathbb{D}$, define $l_{\zeta}(g)=g(\zeta)$ for $g\in A^{2}_{1}$.
-By taking $p=2,\alpha=1$ in (5.9), we see that $\|l_{\zeta}\|\leq C(1-|\zeta|^{2})^{-3/2}$, which implies that $l_{\zeta}\in(A^{2}_{1})^{*}$.
+Taking $p=2,\alpha=1$ in (5.9), we see that $\|l_{\zeta}\|\leq C(1-|\zeta|^{2})^{-3/2}$, which implies that $l_{\zeta}\in(A^{2}_{1})^{*}$.
 Set $A:=\{l_{\zeta};\;\zeta\in\mathbb{D}\}$.
@@
 $\|T_{s}(f)\|_{A^{2}_{1-2s}}\leq M_{0}^{1-s}M_{1}^{s}\|f\|_{B_{2,2}^{s}(\Omega)},\qquad\text{for\; all}\;f\in\check{B}.$ (5.10)
-Since $\check{B}$ is dense in $\text{HB}_{2,2}^{0}(\Gamma)\cap\text{HB}_{2,2}^{1}(\Gamma)$ and $\text{HB}_{2,2}^{0}(\Gamma)\cap\text{HB}_{2,2}^{1}(\Gamma)$ is dense in $\text{HB}_{2,2}^{s}(\Gamma)$ for all $s\in[0,1]$ (see [43]), it follows immediately that $\check{B}$ is dense in $\text{HB}_{2,2}^{s}(\Gamma)$.
-Thus, (5.10) holds true for all $f\in\text{HB}_{2,2}^{s}(\Gamma)$.
-The bounded inverse theorem implies the inverse of the operator $T_{s}$ is also bounded.
+Since $\check{B}$ is dense in $\text{HB}_{2,2}^{0}(\Omega)\cap\text{HB}_{2,2}^{1}(\Omega)$ and $\text{HB}_{2,2}^{0}(\Omega)\cap\text{HB}_{2,2}^{1}(\Omega)$ is dense in $\text{HB}_{2,2}^{s}(\Omega)$ for all $s\in[0,1]$ (see [49]), it immediately follows that $\check{B}$ is dense in $\text{HB}_{2,2}^{s}(\Omega)$.
+Thus, (5.10) holds for all $f\in\text{HB}_{2,2}^{s}(\Omega)$.
+The bounded inverse theorem implies that the inverse of the operator $T_{s}$ is also bounded.
 This completes the proof of Theorem 5.5.
-5.4 The Equality $\mathcal{B}_{p,p}^{s}(\Omega\!\to\!\Gamma)=B_{p,p}^{s}(\Gamma)$
-The aim of this part is to prove the following
+5.4 The equality $\mathcal{B}_{p,p}^{s}(\Omega)=B_{p,p}^{s}(\Gamma)$
+In this subsection, our goal is to prove the following
 Theorem 5.7.
@@
   • when $p<2$, $|\varphi^{\prime}|\in A_{p}$
-then it holds that $\mathcal{B}_{p,p}^{s}(\Omega\!\to\!\Gamma)=B_{p,p}^{s}(\Gamma)$ with comparable norms.
-We recall that if $\delta$ stands for the infinimum of the set of $q$ such that $|\varphi^{\prime}|\in A_{q}$ then the conditions in the theorem are equivalent to
-$\delta<q<\frac{\delta}{\delta-1}.$
+then it holds that $\mathcal{B}_{p,p}^{s}(\Omega)=B_{p,p}^{s}(\Gamma)$ with comparable norms.
+We recall that if $\delta=\inf\{q>1:\,|\varphi^{\prime}|\in A_{q}\}$ then the theorem is true for
+$\delta<p<\frac{\delta}{\delta-1}.$
 In particular, $\delta<2$ so that $|\varphi^{\prime}|\in A_{2}$.
 That implies that $B_{p,p}^{s}(\Gamma)$ is stable by conjugation by Theorem 5.1.
-Together with $\text{Arg}\,\varphi^{\prime}\in L^{\infty}$, that also implies that $\mathcal{B}_{2,2}^{s}(\Omega\!\to\!\Gamma)=B_{2,2}^{s}(\Gamma)$, $0\leq s\leq 1$, with comparable norms by Theorem 5.5.
+Together with $\text{Arg}\,\varphi^{\prime}\in L^{\infty}$, this also implies that $\mathcal{B}_{2,2}^{s}(\Omega)=B_{2,2}^{s}(\Gamma)$, $0\leq s\leq 1$, with comparable norms by Theorem 5.5.
 These two results will be used in the proof of Theorem 5.7.
-On the other hand, for the proof of Theorem 5.7 we also need to use the claim: $\mathcal{B}_{p,p}^{1/p}(\Omega\!\to\!\Gamma)=B_{p,p}^{1/p}(\Gamma)$ with comparable norms for $1<p<\infty$ provided that the curve $\Gamma$ is chord-arc.
-That was proved in our recent paper [45] using the conformal invariance of the space $\mathcal{B}_{p,p}^{1/p}(\Omega)$.
-We outline its proof here: suppose $f\in B_{p,p}^{1/p}(\Gamma)$ and its harmonic extension in $\Omega$ is still denoted by $f$ then the assertion follows from the following reasoning:
+On the other hand, for the proof of Theorem 5.7 we also need to use the claim: $\mathcal{B}_{p,p}^{1/p}(\Omega)=B_{p,p}^{1/p}(\Gamma)$ with comparable norms for $1<p<\infty$ provided that the curve $\Gamma$ is chord-arc (see Theorem 1.1).
+For convenience we sketch its proof here: suppose that $f\in B_{p,p}^{1/p}(\Gamma)$ and its harmonic extension in $\Omega$ is still denoted by $f$ then the assertion follows from the following reasoning:
 $\|f\|_{B_{p,p}^{1/p}(\Gamma)}\simeq\|f\circ\lambda\|_{B_{p,p}^{1/p}(\mathbb{T})}\simeq\|f\circ\varphi\|_{B_{p,p}^{1/p}(\mathbb{T})}\simeq\|f\circ\varphi\|_{\mathcal{B}_{p,p}^{1/p}(\mathbb{D})}=\|f\|_{\mathcal{B}_{p,p}^{1/p}(\Omega)},$
 where the implicit constants depend only on $\Omega$ and $p$.
-The above steps in order are due to $(1)$ $\lambda$ is a bi-Lipschitz homeomorphism of $\mathbb{T}$ onto $\Gamma$; $(2)$ $\lambda^{-1}\circ\varphi$ is a quasisymmetry; $(3)$ when $\Gamma=\mathbb{T}$ this is well-known as we mentioned before; $(4)$ the space $\mathcal{B}_{p,p}^{1/p}(\Omega)$ is conformal invariant.
+The above steps are in order due to $(1)$ $\lambda$ is a bi-Lipschitz homeomorphism of $\mathbb{T}$ onto $\Gamma$; $(2)$ $\lambda^{-1}\circ\varphi$ is a quasisymmetry; $(3)$ when $\Gamma=\mathbb{T}$ this is well-known, as we mentioned before; $(4)$ the space $\mathcal{B}_{p,p}^{1/p}(\Omega)$ is conformal invariant.
 Recall that $\mathcal{B}_{p,p}^{s}(\Omega)$ is conjugate invariant.
-We have seen, under the assumption of Theorem 5.7, that $B_{p,p}^{s}(\Gamma)$ is also conjugate invariant, in order to prove Theorem 5.7 we thus only need to prove the holomorphic case:
+We have seen, under the assumption of Theorem 5.7, that $B_{p,p}^{s}(\Gamma)$ is also conjugate invariant (see Theorem 5.1), in order to prove Theorem 5.7 we thus only need to prove the following
 Theorem 5.8.
@@
 If $\Gamma$ is such that $\text{Arg}\,\varphi^{\prime}\in L^{\infty}$ then it holds that
-$\mathcal{HB}_{p,p}^{s}(\Omega\!\to\!\Gamma)=\text{HB}_{p,p}^{s}(\Gamma)$ (5.11)
-with comparable norms. The conclusion holds in particular for radial-Lipschitz domains.
+$\mathcal{HB}_{p,p}^{s}(\Omega)=\text{HB}_{p,p}^{s}(\Gamma)$ (5.11)
+with comparable norms. The conclusion holds in particular for the radial-Lipschitz domains.
 Proof.
-For any $f\in\mathcal{HB}_{p,p}^{s}(\Omega)$, $0<s<1$, $1<p<\infty$ we have seen from (5.2) that
+For any $f\in\mathcal{HB}_{p,p}^{s}(\Omega)$, $0<s<1$, $1<p<\infty$, we have seen from (5.2) that
 $\iint_{\mathbb{D}}|V_{s,p}(f)^{\prime}(z)|^{p}(1-|z|)^{(1-s)p-1}dxdy\simeq\iint_{\Omega}|f^{\prime}(\zeta)|^{p}d(\zeta,\Gamma)^{(1-s)p-1}d\xi d\eta$
@@
 with comparable norms.
-Noting that any $g\in A^{p}_{(1-s)p-1}$ can be written in the form $V_{s,p}(f)^{\prime}$ by choosing $f(z)=\int_{0}^{z}g\circ\varphi^{-1}(u)(\varphi^{-1})^{\prime}(u)^{1+1/p-s}du$, we have that the operator $\mathcal{HB}_{p,p}^{s}(\Omega)\ni f\mapsto V_{s,p}(f)^{\prime}\in A^{p}_{(1-s)p-1}$ is a bounded isomorphism.
-This boils down to prove
+Noting that any $g\in A^{p}_{(1-s)p-1}$ can be written in the form $V_{s,p}(f)^{\prime}$ by choosing $f(z)=\int_{0}^{z}g\circ\varphi^{-1}(u)(\varphi^{-1})^{\prime}(u)^{1+1/p-s}du$, we find that the operator $\mathcal{HB}_{p,p}^{s}(\Omega)\ni f\mapsto V_{s,p}(f)^{\prime}\in A^{p}_{(1-s)p-1}$ is a bounded isomorphism.
+This boils down to proving the following:
 $f\in\text{HB}_{p,p}^{s}(\Omega)\;\Leftrightarrow\;V_{s,p}(f)^{\prime}\in A^{p}_{(1-s)p-1}$ (5.13)
 with comparable norms. For this purpose, we will use Theorem 5.6 in the following.
-We have proved that for any $1<p<\infty$, $\mathcal{B}_{p,p}^{1/p}(\Omega\!\to\!\Gamma)=B_{p,p}^{1/p}(\Gamma)$ if $\Gamma$ is a chord-arc curve; and for any $0<s<1$, $\mathcal{B}_{2,2}^{s}(\Omega\!\to\!\Gamma)=B_{2,2}^{s}(\Gamma)$ if moreover $\Gamma$ satisfies that $|\varphi^{\prime}|\in A_{2}$ and $\text{Arg}\varphi^{\prime}\in L^{\infty}$, thus the equivalence (5.13) has already been known to hold for $s=1/p$, $1<p<\infty$, and also holds for $0<s<1$, $p=2$.
-Recall that $(A^{p}_{p-2},A^{2}_{1-2s})$ is an interpolation pair and for $0\leq\theta\leq 1$,
+We have proved that for any $1<p<\infty$, $\mathcal{HB}_{p,p}^{1/p}(\Omega)=\text{HB}_{p,p}^{1/p}(\Omega)$ if $\Gamma$ is a chord-arc curve; and for any $0<s<1$, $\mathcal{HB}_{2,2}^{s}(\Omega)=\text{HB}_{2,2}^{s}(\Omega)$ if, moreover, $\Gamma$ satisfies $\text{Arg}\varphi^{\prime}\in L^{\infty}$, thus the equivalence (5.13) is already known for $s=1/p$, $1<p<\infty$, and also for $0<s<1$, $p=2$.
+Recall that $(A^{p}_{p-2},A^{2}_{1-2s})$ is an interpolation pair, and for $0\leq\theta\leq 1$,
 $A^{p}_{p-2},A^{2}_{1-2s}]_{\theta}=A^{q}_{(1-\alpha)q-1}.$ (5.14)
-We have also seen that $(\text{HB}_{p,p}^{1/p}(\Gamma),\text{HB}_{2,2}^{s}(\Gamma))$ is an interpolation pair and
-$\text{HB}_{p,p}^{1/p}(\Gamma),\text{HB}_{2,2}^{s}(\Gamma)]_{\theta}=\text{HB}_{q,q}^{\alpha}(\Gamma)$ (5.15)
+We have also seen that $(\text{HB}_{p,p}^{1/p}(\Omega),\text{HB}_{2,2}^{s}(\Omega))$ is an interpolation pair and
+$\text{HB}_{p,p}^{1/p}(\Omega),\text{HB}_{2,2}^{s}(\Omega)]_{\theta}=\text{HB}_{q,q}^{\alpha}(\Omega)$ (5.15)
 provided that $\Gamma$ is a chord-arc curve. Here,
 $\begin{split}\alpha=\alpha(\theta)&:=(1-\theta)/p+\theta s,\\ q=q(\theta)&:=2p/(2(1-\theta)+\theta p).\end{split}$ (5.16)
-Some elementary arithmetic leads to that for any given $\alpha\in(0,1)$ and $q\in(1,\infty)$, there exist $p\in(1,\infty)$, $s\in(0,\,\min(1,5/2-2/p))$ and $\theta\in[0,1]$ such that (5.16) holds.
-Recall that in the proof of Theorem 5.5 we have shown that $\check{B}$, the set of polynomials on $\Gamma$, is dense in $\text{HB}_{2,2}^{s}(\Gamma)$.
-By using the similar argument, we can see that $\check{B}$ is also dense in $\text{HB}_{p,p}^{1/p}(\Gamma)$ and thus $\check{B}$ is a dense subspace of the intersection space $\text{HB}_{p,p}^{1/p}(\Gamma)\cap\text{HB}_{2,2}^{s}(\Gamma)$.
-Define a family of linear mappings $(\widetilde{T}_{z};\;z=\theta+it\in S)$, $S$ standing for a vertical strip with $0\leq\theta\leq 1$ as before, on $\check{B}$ by
+Some elementary arithmetic leads to the conclusion that for any given $\alpha\in(0,1)$ and $q\in(1,\infty)$, there exist $p\in(1,\infty)$, $s\in(0,\,\min(1,5/2-2/p))$, and $\theta\in[0,1]$ such that (5.16) holds.
+Recall that in the proof of Theorem 5.5 we have shown that $\check{B}$, the set of polynomials on $\Gamma$, is dense in $\text{HB}_{2,2}^{s}(\Omega)$.
+Using a similar argument, we can see that $\check{B}$ is also dense in $\text{HB}_{p,p}^{1/p}(\Omega)$ and thus $\check{B}$ is a dense subspace of the intersection space $\text{HB}_{p,p}^{1/p}(\Omega)\cap\text{HB}_{2,2}^{s}(\Omega)$.
+Define a family of linear mappings $(\widetilde{T}_{z};\;z=\theta+it\in S)$, $S$ representing a vertical strip with $0\leq\theta\leq 1$ as before, on $\check{B}$ by
 $\widetilde{T}_{z}(f):=e^{cz^{2}}V_{\alpha(z),q(z)}(f)^{\prime}=e^{cz^{2}}f^{\prime}\circ\varphi(\varphi^{\prime})^{\frac{1}{q(z)}+1-\alpha(z)}=e^{cz^{2}}f^{\prime}\circ\varphi(\varphi^{\prime})^{(\frac{1}{2}-s)z+1}.$
@@
 We will prove that for each $z=\theta+it\in S$, the linear mapping $\widetilde{T}_{z}$ sends $\check{B}$ to $A^{p}_{p-2}+A^{2}_{1-2s}$, and the family $(\widetilde{T}_{z})$ satisfies the properties (i) and (ii) of Theorem 5.6.
-We first estimate $|\widetilde{T}_{z}(f)(\zeta)|^{p}$ for each $z=\theta+it\in S$ and each $f\in\check{B}$. It follows from simple computation that
+We first estimate $|\widetilde{T}_{z}(f)(\zeta)|^{p}$ for each $z=\theta+it\in S$ and each $f\in\check{B}$. From a simple computation, it follows that
 $|e^{cz^{2}}|^{p}=e^{pc(\theta^{2}-t^{2})}\leq e^{pc}e^{-ct^{2}p},$
@@
 $|(\varphi^{\prime})^{(1/2-s)z+1}|^{p}=e^{t(s-1/2)p\text{Arg}\varphi^{\prime}}\left(|\varphi^{\prime}|^{1+(1/2-s)\theta}\right)^{p}\leq e^{|t|pA}\left(|\varphi^{\prime}|^{1+(1/2-s)\theta}\right)^{p}.$
-There exists a constant $C_{1}$ such that $|f^{\prime}\circ\varphi|^{p}<C_{1}$ since the diameter of the domain $\Omega$ with 0 as an interior point is finite. Then,
+There exists a constant $C_{1}$ such that $|f^{\prime}\circ\varphi|^{p}<C_{1}$ since the diameter of the domain $\Omega$, with 0 as an interior point, is finite. Then
 $|\widetilde{T}_{z}(f)(\zeta)|^{p}\leq C_{1}e^{pc}\left(e^{-ct^{2}+|t|A}\right)^{p}\left(|\varphi^{\prime}|^{1+(1/2-s)\theta}\right)^{p}\leq C_{1}e^{pc}e^{p}\left(|\varphi^{\prime}|^{1+(1/2-s)\theta}\right)^{p}.$ (5.17)
-To show the linear mapping $\widetilde{T}_{z}$ sends $\check{B}$ to $A^{p}_{p-2}+A^{2}_{1-2s}$, we split into the following three cases:
+To show that the linear mapping $\widetilde{T}_{z}$ sends $\check{B}$ to $A^{p}_{p-2}+A^{2}_{1-2s}$, we split into the following three cases:
 Case $1$. when $s\leq 1/2$
@@
 Case $3$. When $1/2<s<\min(1,5/2-2/p)$ and $1<p<2$, it suffices to show that for all $\theta\in(0,1)$,
-$\iint_{\mathbb{D}}|\varphi^{\prime}(z)|^{2+p\theta(\frac{1}{2}-s)}d(\varphi(z),\Gamma)^{p-2}dxdy<\infty.$
-First of all, since $s>1/2$ and $p<2$ we have that
-$1<2+p\theta(\frac{1}{2}-s)<2.$
+$\iint_{\mathbb{D}}|\varphi^{\prime}(z)|^{2+p\theta\left(\frac{1}{2}-s\right)}d(\varphi(z),\Gamma)^{p-2}dxdy<\infty.$
+First of all, since $s>1/2$ and $p<2$ we have
+$1<2+p\theta\left(\frac{1}{2}-s\right)<2.$
 By Hölder inequality,
 $\begin{split}&\frac{1}{\pi}\iint_{\mathbb{D}}|\varphi^{\prime}(z)|^{2+p\theta(\frac{1}{2}-s))}d(\varphi(z),\Gamma)^{p-2}dxdy\\ \leq&\left(\frac{1}{\pi}\iint_{\mathbb{D}}|\varphi^{\prime}(z)|^{2}d(\varphi(z),\Gamma)^{\frac{2(p-2)}{2+p\theta(\frac{1}{2}-s)}}dxdy\right)^{\frac{2+p\theta(\frac{1}{2}-s)}{2}},\end{split}$
-and the last integral is equal, by the change of variable $\zeta=\varphi(z)$, to
+and the last integral is equal, by the change of variables $\zeta=\varphi(z)$, to
 $\iint_{\Omega}d(\zeta,\Gamma)^{\frac{2(p-2)}{2+p\theta(\frac{1}{2}-s)}}d\xi d\eta.$
@@
 $0<\theta<2\frac{1-\frac{1}{p}}{s-\frac{1}{2}}.$
-A sufficient condition for the finiteness will thus be that
+Thus, a sufficient condition for finiteness will be that
 $2\frac{1-\frac{1}{p}}{s-\frac{1}{2}}>1\Leftrightarrow s<\frac{5}{2}-\frac{2}{p}.$
-Consequently, when $1<p<\infty,\,0<s<\min(1,5/2-2/p)$, the linear ma??pping $\widetilde{T}_{z}$ sends $\check{B}$ to $A^{p}_{p-2}+A^{2}_{1-2s}$, and for all $f\in\check{B}$, the mapping $\widetilde{T}_{(\cdot)}(f):S\to A^{p}_{p-2}+A^{2}_{1-2s}$ is bounded.
+Consequently, when $1<p<\infty,\,0<s<\min(1,5/2-2/p)$, the linear mapping $\widetilde{T}_{z}$ sends $\check{B}$ to $A^{p}_{p-2}+A^{2}_{1-2s}$, and for all $f\in\check{B}$, the mapping $\widetilde{T}_{(\cdot)}(f):S\to A^{p}_{p-2}+A^{2}_{1-2s}$ is bounded.
 Taking $\theta=0$ in (5.17) we can see
@@
 $M_{1}$ $:=\sup\{\|\widetilde{T}_{1+it}f\|_{A^{2}_{1-2s}};\;t\in\mathbb{R},f\in\check{B},\|f\|_{B_{2,2}^{s}(\Gamma)}\leq 1\}<\infty.$
-By the continuity of parametrized integral, all functions $S\ni z\mapsto\widetilde{T}_{z}f\in A^{p}_{p-2}+A^{2}_{1-2s}$,   $\mathbb{R}\ni t\mapsto\widetilde{T}_{it}f\in A^{p}_{p-2}$, and $\mathbb{R}\ni t\mapsto\widetilde{T}_{1+it}f\in A^{2}_{1-2s}$ are continuous.
+By the continuity of the parametrized integral, all functions $S\ni z\mapsto\widetilde{T}_{z}f\in A^{p}_{p-2}+A^{2}_{1-2s}$,   $\mathbb{R}\ni t\mapsto\widetilde{T}_{it}f\in A^{p}_{p-2}$, and $\mathbb{R}\ni t\mapsto\widetilde{T}_{1+it}f\in A^{2}_{1-2s}$ are continuous.
 Finally, for any $f\in\check{B}$, the proof of holomorphy of the mapping $\stackrel{{\scriptstyle\circ}}{{S}}\ni z\mapsto\widetilde{T}_{z}f\in A^{p}_{p-2}+A^{2}_{1-2s}$ is similar to that in the proof of Theorem 5.5.
@@
 ∎
-5.5 Some Final Remarks
-In this part, we let $0<s<1$ and $1<p<\infty$, and suppose $\Gamma$ satisfies the assumption of Theorem 5.7, namely $\Gamma$ is a bounded chord-arc curve with 0 in its interior domain $\Omega$, and $\Gamma$ is such that $\text{Arg}\,\varphi^{\prime}\in L^{\infty}$ and $|\varphi^{\prime}|\in A_{q}$ with $q=\min{(p,p^{\prime})}$.
-Here, $1/p+1/p^{\prime}=1$ and $\varphi_{i}^{\prime}$ is a Riemann mapping from the unit disk $\mathbb{D}$ onto $\Omega$ keeping 0 fixed.
-Recall that when $p=2$ we have shown that the radial-Lipschitz curve satisfies this assumption.
-Since $h(\Gamma)=1$ one may see that $(h(\Gamma),p,s)$ is always in the admissible set $\mathcal{A}$.
-It follows from Theorem 6.2 and Theorem 5.7 (or Theorem 5.5 for the case of $p=2$) that in this case $\mathcal{B}_{p,p}^{s}(\Omega_{i}\!\to\!\Gamma)=\mathcal{B}_{p,p}^{s}(\Omega_{e}\!\to\!\Gamma)$ with comparable norms.
-Recall that we explored earlier that this equality also holds under a different assumption where there are some restrictions on $p$ and $s$ in Theorem 3.12.
-If $p>2$, $s\geq 1-1/p$ we then have that
-  • (i)
-    $\mathcal{B}_{p,p}^{s}(\Gamma)=W^{1,p}(\omega,\mathbb{C})|_{\Gamma}$ where $\omega(z)=d(z,\Gamma)^{(1-s)p-1}$, i.e., the space $\mathcal{B}_{p,p}^{s}(\Gamma)$ is the trace space of $W^{1,p}(\omega,\mathbb{C})$ on $\Gamma$;  (by Theorem 3.16)
-  • (ii)
-    The space $C_{c}^{\infty}(\mathbb{C})|_{\Gamma}$ is dense in $\mathcal{B}_{p,p}^{s}(\Gamma)$, and thus Plemelj-Calderón property holds for any function in $\mathcal{B}_{p,p}^{s}(\Gamma)$.
-(by Corollary 3.17 and Theorem 3.5)
-  • (iii)
-    for any $f\in\mathcal{B}_{p,p}^{s}(\Gamma)$, its extension to $\mathbb{C}$ by the harmonic extensions on both connected components $\Omega_{i,e}$ of $\overline{\mathbb{C}}\setminus\Gamma$ belongs to $W^{1,p}(\omega,\mathbb{C})$;  (by section 3.3.2)
-  • (iv)
-    Any function in $\mathcal{B}_{p,p}^{s}(\Omega_{i,e})$ satisfies the almost-Dirichlet principle.  (by section 3.2.3)
-The above statements are claims: given $p>2$ and thus the class of curves $\Gamma$ for which claims are true depends on $p$ (i,e., $|\varphi^{\prime}|\in A_{p^{\prime}}$).
-If we reverse the point of view and start with a curve $\Gamma$ satisfying $\text{Arg}\,\varphi^{\prime}\in L^{\infty}$ and $|\varphi^{\prime}|\in A_{2}$ then the claims are true for $p\in(\delta,\frac{\delta}{\delta-1})$ where $\delta$ is the infinimum of the set of $p^{\prime}s$ such that $|\varphi^{\prime}|\in A_{p}$.
-By the theory of $A_{\infty}$ weights, $\delta<2$.
+Remark 5.9.
+Recall that in Theorems 5.5 and 5.7 we only focused on the case of bounded domains.
+The case of unbounded domains can actually be transformed to the bounded case; see Appendix for details, and similar assertions also hold for the unbounded case.
+Let $\Gamma$ be a radial Lipschitz curve and $\Omega_{i,e}$ be the inner domain and the outer domain of $\Gamma$.
+Let $\varphi_{i,e}$ be a Riemann map for $\Omega_{i,e}$.
+By Helson-Szegö theorem, we see $|\varphi_{i,e}^{\prime}|\in A_{p}$, $p<2$ on $\mathbb{T}$.
+Set
+$p_{0}^{i,e}=\inf\{p>1:\;|\varphi_{i,e}^{\prime}|\in A_{p}\}\;\;\mathrm{and}\;\;p_{0}=\max(p_{0}^{i},p_{0}^{e}).$
+We end this subsection with an estimate of $p_{0}$ in terms of the geometry of the curve. Notice that the two functions
+$\text{Arg}\left(\frac{z\varphi^{\prime}_{i,e}(z)}{\varphi_{i,e}(z)}\right)$
+have the same $L^{\infty}$-norm that is less than $\pi/2$.
+We define $M$ as being the $\tan$ of this norm and call this number the Lipschitz-norm of $\Gamma$.
+We now use the same argument as that below Theorem 5.5:
+$\log|\varphi_{i,e}^{\prime}(z)|^{p}=\log\left|\frac{\varphi_{i,e}(z)}{z}\right|^{p}+\log\left|\frac{z\varphi_{i,e}^{\prime}(z)}{\varphi_{i,e}(z)}\right|=\log\left|\frac{\varphi_{i,e}(z)}{z}\right|^{p}-H\left(p\text{Arg}\frac{z\varphi_{i,e}^{\prime}(z)}{\varphi_{i,e}(z)}\right),$
+we thus conclude by Helson-Szegö theorem that $|\varphi_{i,e}^{\prime}|^{p}$ is an $A_{2}$-weight on the unit circle $\mathbb{T}$ if $p<\frac{\pi}{2\arctan M}$, and this holds in particular for $p=1$.
+On the other hand, let $\omega$ be an $A_{2}$-weight and $\tilde{p}$ the supremum of the set of $p$ such that $\omega^{p}\in A_{2}$.
+It is not difficult to see that $\delta(\omega)(:=\inf\{q>1:\,\omega\in A_{q}\})\leq 1+\frac{1}{\tilde{p}}$.
+Then we get an estimate value of $p_{0}$:
+$1+\frac{1}{\frac{\pi}{2\arctan M}}=1+\frac{2\arctan M}{\pi}.$
+We may then state the following corollary of Theorem 5.7.
+Corollary 5.10.
+If $\Gamma$ is a $M$-radial-Lipschitz curve and $p\in\left(1+\frac{2\arctan M}{\pi},1+\frac{\pi}{2\arctan M}\right)$, $s\in(0,1)$ then
+$\mathcal{B}_{p,p}^{s}(\Omega_{i}\!\to\!\Gamma)=\mathcal{B}_{p,p}^{s}(\Omega_{e}\!\to\!\Gamma)=B_{p,p}^{s}(\Gamma).$ (5.18)
+Notice that this interval of $p$-value converges to $(1,\infty)$ as $M\to 0$ and to a singleton $\{2\}$ as $M\to\infty$.
+5.5 Almost-Dirichlet principle revisited.
+In Section 3 we have proven Theorem 3.5 which states that given a quasicircle $\Gamma,\,p>2,\,$ $\frac{h(\Gamma)}{2}-\frac{h(\Gamma)-1}{p}<s<1-\frac{h(\Gamma)-1}{p}$ (notice that when $\Gamma$ is chord-arc $h(\Gamma)=1$ and thus $1/2<s<1$) then if
+$\mathcal{B}_{p,p}^{s}(\Omega_{i}\to\Gamma)=\mathcal{B}_{p,p}^{s}(\Omega_{e}\to\Gamma)$
+we then have that this space, which we have denoted by $\mathcal{B}_{p,p}^{s}(\Gamma)$, is equal to $W^{1,p}(\mathbb{C},\omega)|_{\Gamma}$.
+Moreover, we have shown that the inclusion $W^{1,p}(\mathbb{C},\omega)|_{\Gamma}\subset\mathcal{B}_{p,p}^{s}(\Gamma)$ implies that the spaces $\mathcal{B}_{p,p}^{s}(\Omega_{i,e})$ obey the almost-Dirichlet principle, which means that there exists $C>0$ such that for any continuous function $F$ on $\overline{\Omega}$ which is a $C^{1}$-function in $\Omega$, if we call $u$ the harmonic extension of $F|_{\Gamma}$ in $\Omega$ then
+$\|u\|_{\mathcal{B}_{p,p}^{s}(\Omega)}^{p}\leq C\iint_{\Omega}|\nabla F(z)|^{p}d(z,\Gamma)^{(1-s)p-1}dxdy.$
+As before, let $\varphi_{i,e}$ be a Riemann map for $\Omega_{i,e}$, and let
+$p_{0}^{i,e}=\inf\{p>1:\;|\varphi_{i,e}^{\prime}|\in A_{p}\}\;\;\mathrm{and}\;\;p_{0}=\max(p_{0}^{i},p_{0}^{e}).$
+Theorem 5.11.
+If $\Gamma$ is a radial-Lipschitz curve then $\mathcal{B}_{p,p}^{s}(\Omega_{i})$ has the almost-Dirichlet property if $2<p<\frac{p_{0}}{p_{0}-1}$ and $s>1/2$.
+Proof.
+It follows immediately from Theorems 3.5, 5.1, 5.7. ∎
+Using the estimate value of $p_{0}$ in the last subsection, we have the following corollary.
+Corollary 5.12.
+If $\Gamma$ is a radial-Lipschitz curve with norm $M>0$ then $\mathcal{B}_{p,p}^{s}(\Omega_{i})$ has the almost-Dirichlet property if
+$2<p<1+\frac{\pi}{2\arctan M},\quad\frac{1}{2}<s<1.$
+This interval of the $p$-value converges to $(2,\infty)$ as $M\to 0$ and to a singleton $\{2\}$ as $M\to\infty$.
 6 Appendix
 Let $\Gamma$ be a Jordan curve with $\infty\notin\Gamma$, and $\Omega_{i,e}$ the connected components of $\overline{\mathbb{C}}\setminus\Gamma$ as above.
-In all this work we have treated $\Omega_{i}$ and $\Omega_{e}$ similarly: in doing so we have overlooked the fact that $\Omega_{e}$ is unbounded, thus requiring an extra argument.
+In all this work, we have treated $\Omega_{i}$ and $\Omega_{e}$ similarly: in doing so we have overlooked the fact that $\Omega_{e}$ is unbounded, thus requiring an extra argument.
 The purpose of this appendix is to fill this gap.
 Without loss of generality, suppose that $\Gamma$ is included in the unit disk $\mathbb{D}$ with 0 being an interior point.
-The map $z\mapsto 1/z$ transforms $\Omega_{e}$ into a bounded domain $\Omega^{\prime}_{i}$ with boundary $\Gamma^{\prime}=1/\Gamma$; $\Omega_{i}^{\prime}$ contains 0 which corresponds to $\infty$ in $\Omega_{e}$.
+The map $z\mapsto 1/z$ transforms $\Omega_{e}$ into a bounded domain $\Omega^{\prime}_{i}$ with boundary $\Gamma^{\prime}=1/\Gamma$; $\Omega_{i}^{\prime}$ contains 0 that corresponds to $\infty$ in $\Omega_{e}$.
 We say $u$ is harmonic in the domain $\Omega_{e}$ with $\infty$ being an interior point if $u(1/z)$ is harmonic in the bounded domain $\Omega_{i}^{\prime}$.
-Let $p>1$ and $s\in(0,1)$ as usual and consider $u\in\mathcal{B}_{p,p}^{s}(\Omega_{e})$.
-Let $v:\Omega_{i}^{\prime}\to\mathbb{C}$ being defined by $v(z)=u(1/z)$.
-By a change of the variable $z=1/w$ we conclude by the Koebe distortion theorem (2.1) that
+Let $p>1$ and $s\in(0,1)$ be as usual and consider $u\in\mathcal{B}_{p,p}^{s}(\Omega_{e})$.
+Let $v:\Omega_{i}^{\prime}\to\mathbb{C}$ be defined by $v(z)=u(1/z)$.
+By a change of variables $z=1/w$ we conclude by the Koebe distortion theorem (2.1) that
 $\iint_{\Omega_{e}}|\nabla u(w)|^{p}d(w,\Gamma)^{(1-s)p-1}dudv\simeq\iint_{\Omega_{i}^{\prime}}|\nabla v(z)|^{p}d(z,\Gamma^{\prime})^{(1-s)p-1}|z|^{2(sp-1)}dxdy$ (6.1)
-where the implicit constant is a universal one. We have the following
+where the implicit constant is universal. We have the following
 Proposition 6.1.
@@
 For any $\zeta$ such that $|\zeta|<r$ the annulus $A_{\zeta}=\{R_{1}<|z-\zeta|<R_{2}\}$ is included in the annulus $\{\frac{1}{2}<|z|<1\}$.
-By the mean value property for harmonic functions we have, for $|\zeta|<r$,
-$|\nabla v(\zeta)|=\frac{1}{\pi R_{i}^{2}}\iint_{|z-\zeta|<R_{i}}\nabla v(z)dxdy,\quad i=1,2$
+By the mean value property for harmonic functions, we have, for $|\zeta|<r$,
+$\nabla v(\zeta)=\frac{1}{\pi R_{i}^{2}}\iint_{|z-\zeta|<R_{i}}\nabla v(z)dxdy,\quad i=1,2$
 so that
 $|\nabla v(\zeta)|\leq\frac{1}{\pi(R_{2}^{2}-R_{1}^{2})}\iint_{\{1/2<|z|<1\}}|\nabla v(z)|dxdy.$
-Noting (6.1) we can conclude by using Hölder inequality that $|\nabla v|$ is uniformly bounded in the disk $|\zeta|<r$ by a constant depending only on $\Gamma$ multiplied by the norm of $u$ in $\mathcal{B}_{p,p}^{s}(\Omega_{e})$.
+Noting (6.1) we can conclude using Hölder inequality that $|\nabla v|$ is uniformly bounded in the disk $|\zeta|<r$ by a constant depending only on $\Gamma$ multiplied by the norm of $u$ in $\mathcal{B}_{p,p}^{s}(\Omega_{e})$.
 This estimate implies that
 $\iint_{|z|<r}|\nabla v(z)|^{p}d(z,\Gamma^{\prime})^{(1-s)p-1}dxdy<\infty.$ (6.2)
-In order to control the rest of the integral giving the norm of $v$:
-$\iint_{\Omega^{\prime}_{i}\setminus(|z|<r)}|\nabla v(z)|^{p}d(z,\Gamma^{\prime})^{(1-s)p-1}dxdy,$ (6.3)
+In order to control the rest of the integral, giving the norm of $v$:
+$\iint_{\Omega^{\prime}_{i}\setminus\{|z|\leq r\}}|\nabla v(z)|^{p}d(z,\Gamma^{\prime})^{(1-s)p-1}dxdy,$ (6.3)
 we first observe that the map $w=1/z$ is bi-Lipschitz in the domain $\Omega^{\prime}_{i}\setminus(|z|<r)$ and then that there exists a constant $C$ depending only on $\Gamma$ such that
@@
 The details are left to the reader. ∎
-Theorem 6.2.
-If $\mathcal{B}_{p,p}^{s}(\Omega_{i}\!\to\!\Gamma)=B_{p,p}^{s}(\Gamma)$, and the identity operator from $\mathcal{B}_{p,p}^{s}(\Omega_{i}\!\to\!\Gamma)$ onto $B_{p,p}^{s}(\Gamma)$ is a bounded isomorphism, then the same assertions hold for the unbounded component $\Omega_{e}$ of $\overline{\mathbb{C}}\setminus\Gamma$, and in particular, the transmission operator from $\mathcal{B}_{p,p}^{s}(\Omega_{i})$ onto $\mathcal{B}_{p,p}^{s}(\Omega_{e})$,
-$\mathcal{B}_{p,p}^{s}(\Omega_{i})\to\mathcal{B}_{p,p}^{s}(\Omega_{i}\!\to\!\Gamma)\to\mathcal{B}_{p,p}^{s}(\Omega_{e}),$
-is a bounded isomorphism.
-Proof.
-Let us use the above proposition to prove that $\mathcal{B}_{p,p}^{s}(\Omega_{e}\!\to\!\Gamma)=B_{p,p}^{s}(\Gamma)$.
-If $u\in\mathcal{B}_{p,p}^{s}(\Omega_{e})$ then $v\in\mathcal{B}_{p,p}^{s}(\Omega^{\prime}_{i})$ so that $v\in\mathcal{B}_{p,p}^{s}(\Omega_{i}^{\prime}\!\to\!\Gamma^{\prime})=B_{p,p}^{s}(\Gamma^{\prime})$ from which it follows easily that $u\in B_{p,p}^{s}(\Gamma)$.
-∎
-Statements and Declarations. The authors declare that there is no conflict of interest, and there is no data associated to this work.
+Statements and Declarations. The authors declare that there are no conflicts of interest and that there are no data associated with this work.
 Acknowledgments.
-This work is supported by the National Natural Science Foundation of China (Grant No.
+This work is supported by the National Natural Science Foundation of China (Grant Nos.
 12271218 and 12571083) and the University of Orléans.